### v2

# Optimal Multi-Distribution Learning

Abstract
Multi-distribution learning (MDL), which seeks to learn a shared model that minimizes the worst-case risk across $k$ distinct data distributions, has emerged as a unified framework in response to the evolving demand for robustness, fairness, multi-group collaboration, etc. Achieving data-efficient MDL necessitates adaptive sampling, also called on-demand sampling, throughout the learning process. However, there exist substantial gaps between the state-of-the-art upper and lower bounds on the optimal sample complexity. Focusing on a hypothesis class of Vapnik–Chervonenkis (VC) dimension $d$, we propose a novel algorithm that yields an $\varepsilon$-optimal randomized hypothesis with a sample complexity on the order of $\frac{d+k}{\varepsilon^{2}}$ (modulo some logarithmic factor), matching the best-known lower bound. Our algorithmic ideas and theory have been further extended to accommodate Rademacher classes. The proposed algorithms are oracle-efficient, which access the hypothesis class solely through an empirical risk minimization oracle. Additionally, we establish the necessity of randomization, unveiling a large sample size barrier when only deterministic hypotheses are permitted. These findings successfully resolve three open problems presented in COLT 2023 (i.e., Awasthi et al., (2023, Problems 1, 3 and 4)).
Keywords: multi-distribution learning; on-demand sampling; game dynamics; VC classes; Rademacher classes; oracle efficiency
1 Introduction
Driven by the growing need of robustness, fairness and multi-group collaboration in machine learning practice, the multi-distribution learning (MDL) framework has emerged as a unified solution in response to these evolving demands (Blum et al.,, 2017; Haghtalab et al.,, 2022; Mohri et al.,, 2019; Awasthi et al.,, 2023). Setting the stage, imagine that we are interested in a collection of $k$ unknown data distributions $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{k}$ supported on $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ (resp. $\mathcal{Y}$) stands for the instance (resp. label) space. Given a hypothesis class $\mathcal{H}$ and a prescribed loss function [^1] For example, for each hypothesis $h\in\mathcal{H}$ and each datapoint $(x,y)\in\mathcal{X}\times\mathcal{Y}$, we employ $\ell(h,(x,y))$ to measure the risk of using hypothesis $h$ to predict $y$ based on $x$. :×$\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$, we are asked to identify a (possibly randomized) hypothesis $\widehat{h}$ achieving near-optimal worst-case loss across these data distributions, namely, [^2] Here, the expectation on the left-hand side of (1) is taken over the randomness of both the datapoints $(x,y)$ and the (randomized) hypothesis $\widehat{h}$.
$\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},\widehat{h}}\big{[}\ell\big{(}\widehat{h},(x,y)\big{)}\big{]}\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}+\varepsilon$ (1)
with $\varepsilon\in(0,1]$ a target accuracy level. In light of the unknown nature of these data distributions, the learning process involves data collection, allowing the learner to sample from $\{\mathcal{D}_{i}\}_{i=1}^{k}$. The performance of a learning algorithm is then gauged by its sample complexity — the number of samples required to fulfil (1). Our objective is to design a learning paradigm that achieves optimal sample complexity.
The MDL framework described above, which can viewed as an extension of agnostic learning (Valiant,, 1984; Blumer et al.,, 1989) tailored to multiple data distributions, has found a wealth of applications across multiple domains. Here, we highlight a few representative examples, and refer the interested reader to Haghtalab et al., (2022) for more in-depth discussions.
  • Collaborative and agnostic federated learning. In the realm of collaborative and agnostic federated learning (Blum et al.,, 2017; Nguyen and Zakynthinou,, 2018; Chen et al.,, 2018; Mohri et al.,, 2019; Blum et al., 2021a, ; Du et al.,, 2021; Deng et al.,, 2020; Blum et al., 2021b, ), a group of $k$ agents, each having access to distinct data sources as characterized by different data distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$, aim to learn a shared prediction model that ideally would achieve low risk for each of their respective data sources. A sample-efficient MDL paradigm would help unleash the potential of collaboration and information sharing in jointly learning a complicated task.
  • Min-max fairness in learning. The MDL framework is well-suited to scenarios requiring fairness across multiple groups (Dwork et al.,, 2021; Rothblum and Yona,, 2021; Du et al.,, 2021). For instance, in situations where multiple subpopulations with distinct data distributions exist, a prevailing objective is to ensure that the learned model does not adversely impact any of these subpopulations. Such a criterion, known as “min-max fairness” in the literature (Mohri et al.,, 2019; Abernethy et al.,, 2022), plays a pivotal role in mitigating the worst disadvantage experienced by any particular subpopulation.
  • Distributionally robust optimization/learning. Another context where MDL naturally finds applications is group distributionally robust optimization and learning (DRO/DRL). Group DRO and DRL aim to develop algorithms that offer robust performance guarantees across a finite number of possible distributional models (Sagawa et al.,, 2019, 2020; Hashimoto et al.,, 2018; Hu et al.,, 2018; Xiong et al.,, 2023; Zhang et al.,, 2020; Wang et al.,, 2023; Deng et al.,, 2020), and have garnered substantial attention recently due to the pervasive need for robustness in modern decision-making applications (Carmon and Hausler,, 2022; Asi et al.,, 2021; Haghtalab et al.,, 2022; Kar et al.,, 2019). When applying MDL to the context of group DRO/DRL, the resultant sample complexity reflects the price that needs to be paid for learning a robust solution.
The MDL framework is also closely related to other topics like multi-source domain adaptation, maximum aggregation, to name just a few (Mansour et al.,, 2008; Zhao et al.,, 2020; Bühlmann and Meinshausen,, 2015; Guo,, 2023).
In stark contrast to single-distribution learning, achieving data-efficient MDL necessitates adaptive sampling throughout the learning process, also known as on-demand sampling (Haghtalab et al.,, 2022). More specifically, pre-determining a sample-size budget for each distribution beforehand and sampling non-adaptively could result in a loss of sample efficiency, as we lack knowledge about the complexity of learning each distribution before the learning process begins. The question then boils down to how to optimally adapt the online sampling strategy to effectively tackle diverse data distributions.
Inadequacy of prior results.
The sample complexity of MDL has been explored in a strand of recent works under various settings. Consider first the case where the hypothesis class $\mathcal{H}$ comprises a finite number of hypotheses. If we sample non-adaptively and draw the same number of samples from each individual distribution $\mathcal{D}_{i}$, then this results in a total sample size exceeding the order of $\frac{k\log(|\mathcal{H}|)}{\varepsilon^{2}}$ (given that learning each distribution requires a sample size at least on the order of $\frac{\log(|\mathcal{H}|)}{\varepsilon^{2}}$). Fortunately, this sample size budget can be significantly reduced with the aid of adaptive sampling. In particular, the state-of-the-art approach, proposed by Haghtalab et al., (2022), accomplishes the objective (1) with probability $1-\delta$ using $O\big{(}\frac{\log(|\mathcal{H}|)+k\log(k/\delta)}{\varepsilon^{2}}\big{)}$ samples. In comparison to agnostic learning on a single distribution, it only incurs an extra additive cost of $k\log(k/\delta)/\varepsilon^{2}$ as opposed to a multiplicative factor in $k$, thus underscoring the importance of adaptive sampling.
A more challenging scenario arises when $\mathcal{H}$ has a finite Vapnik–Chervonenkis (VC) dimension $d$. The sample complexity for VC classes has only been settled for the reliazable case (Blum et al.,, 2017; Chen et al.,, 2018; Nguyen and Zakynthinou,, 2018), a special scenario where the loss function takes the form of $\ell\big{(}h,(x,y)\big{)}=\mathds{1}\{h(x)\neq y\}$ and it is feasible to achieve zero mean loss. For the general non-realizable case, the best-known lower bound for such VC classes is (Haghtalab et al.,, 2022) [^3] Here and throughout, $\widetilde{O}(\cdot)$ and $\widetilde{\Omega}(\cdot)$ hide all logarithimic factors in $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$.
$\widetilde{\Omega}\left(\frac{d+k}{\varepsilon^{2}}\right),$ (2)
which serves as a theoretical benchmark. By first collecting $\widetilde{O}\left(\frac{dk}{\varepsilon}\right)$ samples to help construct a cover of $\mathcal{H}$ with reasonable resolution, Haghtalab et al., (2022) established a sample complexity upper bound of
[…]
Nevertheless, the term $dk/\varepsilon$ in (3a) fails to match the lower bound (2); put another way, this term might result in a potentially large burn-in cost, as the optimality of this approach is only guaranteed (up to log factors) when the total sample size already exceeds an enormous threshold on the order of $\frac{d^{2}k^{2}}{d+k}$. In an effort to alleviate this $dk/\varepsilon$ factor, Awasthi et al., (2023) put forward an alternative algorithm — which utilizes an oracle to learn on a single distribution and obliviates the need for computing an epsilon-net of $\mathcal{H}$ — yielding a sample complexity of
[…]
However, this result (3b) might fall short of optimality as well, given that the scaling $d/\varepsilon^{4}$ is off by a factor of $1/\varepsilon^{2}$ compared with the lower bound (2). A more comprehensive list of past results can be found in Table 1.
Given the apparent gap between the state-of-the-art lower bound (2) and achievability bounds (1), a natural question arises:
  • Question: Is it plausible to design a multi-distribution learning algorithm with a sample complexity of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ for VC classes, thereby matching the established lower bound (2)?
Notably, this question has been posed as an open problem during the Annual Conference on Learning Theory (COLT) 2023; see Awasthi et al., (2023, Problem 1).
Paper Sample complexity bound
Haghtalab et al., (2022) $\frac{\log(|\mathcal{H}|)+k}{\varepsilon^{2}}$
Haghtalab et al., (2022) $\frac{d+k}{\varepsilon^{2}}+\frac{dk}{\varepsilon}$
Awasthi et al., (2023) $\frac{d}{\varepsilon^{4}}+\frac{k}{\varepsilon^{2}}$
Peng, (2023) $\frac{d+k}{\varepsilon^{2}}\left(\frac{k}{\epsilon}\right)^{o(1)}$
our work (Theorem 1) $\frac{d+k}{\varepsilon^{2}}$
lower bound: Haghtalab et al., (2022) $\frac{d+k}{\varepsilon^{2}}$
Table 1: Sample complexity bounds of MDL with $k$ data distributions and a hypothesis class of VC dimension $d$. Here, we only report the polynomial depedency and hide all logarithmic dependency on $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$.
A glimpse of our main contributions.
The present paper delivers some encouraging news: we come up with a new MDL algorithm that successfully resolves the aforementioned open problem in the affirmative. Specifically, focusing on a hypothesis class with VC dimension $d$ and a collection of $k$ data distributions, our main findings can be summarized as follows.
Theorem 1.
There exists an algorithm (see Algorithm 1 for details) such that: with probability exceeding $1-\delta$, the randomized hypothesis $h^{\sf final}$ returned by this algorithm achieves
$\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},h^{\mathsf{final}}}\left[\ell\big{(}h^{\mathsf{final}},(x,y)\big{)}\right]\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}+\varepsilon,$
provided that the total sample size exceeds
$\frac{d+k}{\varepsilon^{2}}\mathrm{poly}\log\Big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\Big{)}.$ (4)
The polylog factor in (4) will be specified momentarily. In a nutshell, we develop the first algorithm that provably achieves a sample complexity matching the lower bound (2) modulo logarithmic factors. Following the game dynamics template adopted in previous methods — namely, viewing MDL as a game between the learner (who selects the best hypothesis) and the adversary (who chooses the most challenging mixture of distributions) — our algorithm is built upon a novel and meticulously designed sampling scheme that deviates significantly from previous methods. Further, we extend our algorithm and theory to accommodate Rademacher classes, establishing a similar sample complexity bound when the weighted Rademacher complexity of the hypothesis class on $n$ points is upper bounded by $O\big{(}\sqrt{\frac{d\log(n)}{n}}\big{)}$.
Additionally, we solve two other open problems posed by Awasthi et al., (2023):
  • Oracle-efficient solutions. An algorithm is said to be oracle-efficient if it only accesses $\mathcal{H}$ through an empirical risk minimization (ERM) oracle (Dudík et al.,, 2020). Awasthi et al., (2023, Problem 4) then asks what the sample complexity of MDL is when confined to oracle-efficient paradigms. Encouragingly, our algorithm (i.e., Algorithm 1) adheres to the oracle-efficient criterion, thus uncovering that the sample complexity of MDL remains unchanged when restricted to oracle-efficient algorithms.
  • Necessity of randomization. Both our algorithm and the most sample-efficient methods preceding our work produce randomized hypotheses. As discussed around Awasthi et al., (2023, Problem 3), a natural question concerns characterization of the sample complexity when restricting the final output to deterministic hypotheses from $\mathcal{H}$. Our result (see Theorem 18) delivers a negative message: under mild conditions, for any MDL algorithm, there exists a hard problem instance such that it requires at least $\Omega(dk/\varepsilon^{2})$ samples to find a deterministic hypothesis $h\in\mathcal{H}$ that attains $\varepsilon$-accuracy. This constitutes a sample size barrier that lies far above what is achievable under randomized hypotheses.
Concurrent work.
While finalizing our paper, we became aware of a concurrent work Peng, (2023), which also studied the MDL problem and significantly improved upon the prior results. More specifically, Peng, (2023) established a sample complexity of $O\left(\frac{(d+k)\log(d/\delta)}{\varepsilon^{2}}\left(\frac{k}{\varepsilon}\right)^{o(1)}\right)$, which is optimal up to some sub-polynomial factor in $k/\varepsilon$; in comparison, our sample complexity is optimal only up to some logarithmic factor. Additionally, it is worth noting that the algorithm therein relies upon a certain recursive structure to eliminate the non-optimal hypothesis, thus incurring exponential computational cost even when an ERM oracle is available.
Notation.
Throughout this paper, we denote $[N]\coloneqq\{1,\ldots,N\}$ for any positive integer $N$. Let $\mathsf{conv}(\mathcal{A})$ represent the convex hull of a set $\mathcal{A}$, and denote by $\Delta(n)$ the $n$-dimensional simplex for any positive integer $n$. For two vectors $v=[v_{i}]_{1\leq i\leq n}$ and $v^{\prime}=[v^{\prime}_{i}]_{1\leq i\leq n}$ with the same dimension, we overload the notation by using $\max\{v,v^{\prime}\}=\big{[}\max\{v_{i},v^{\prime}_{i}\}\big{]}_{1\leq i\leq n}$ to denote the coordinate-wise maximum of $v$ and $v^{\prime}$. Also we say $v\leq v^{\prime}$ iff $v_{i}\leq v^{\prime}_{i}$ for all $i\in[n]$. For any random variable $X$, we use $\mathbb{V}[X]$ to denote its variance, i.e., $\mathbb{V}[X]=\mathbb{E}\left[(X-\mathbb{E}[X])^{2}\right]$. For any two distributions $P$ and $Q$ supported on $\mathcal{X}$, the Kullback-Leibler (KL) divergence from $Q$ to $P$ is defined and denoted by
$\mathsf{KL}(P\,\|\,Q)\coloneqq\mathbb{E}_{Q}\bigg{[}\frac{\mathrm{d}P}{\mathrm{d}Q}\log\frac{\mathrm{d}P}{\mathrm{d}Q}\bigg{]}.$ (5)
2 Problem formulation
This section formulates the multi-distribution learning problem. We assume throughout that each datapoint takes the form of $(x,y)\in\mathcal{X}\times\mathcal{Y}$, with $\mathcal{X}$ (resp. $\mathcal{Y}$) representing the instance space (resp. label space).
Learning from multiple distributions.
The problem setting encompasses several elements as follows.
  • Hypothesis class. Suppose that we are interested in a hypothesis class $\mathcal{H}$, consisting of a set of candidate functions from the instance space $\mathcal{X}$ to the label space $\mathcal{Y}$. Overloading the notation, we use $h_{\pi}$ to represent a randomized hypothesis associated with a probability distribution $\pi\in\Delta(\mathcal{H})$, meaning that a hypothesis $h$ from $\mathcal{H}$ is randomly selected according to distribution $\pi$. Additionally, the Vapnik–Chervonenkis (VC) dimension (Vapnik et al.,, 1994) of $\mathcal{H}$ is assumed to be
    $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})=d;$ (6)
    see Shalev-Shwartz and Ben-David, (2014) for an introduction of the VC dimension.
  • Loss function. Suppose we are given a loss function :×$\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$, so that $\ell(h,(x,y))$ quantifies the risk of using hypothesis $h\in\mathcal{H}$ to make prediction on a datapoint $(x,y)\in\mathcal{X}\times\mathcal{Y}$ (i.e., predicting $y$ based on $x$). One example is the 0-1 loss function $\ell(h,(x,y))=\mathds{1}\{h(x)\neq y\}$, which is often used to measure the misclassification error.
  • (Multiple) data distributions. Suppose that there are $k$ data distributions of interest supported on $\mathcal{X}\times\mathcal{Y}$, denoted by $\mathcal{D}=\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$. We are permitted to draw independent samples from each of these data distributions.
Given a target accuracy level $\varepsilon\in(0,1)$, our objective is to identify a (randomized) hypothesis, represented by $h_{\pi}$ with $\pi\in\Delta(\mathcal{H})$, such that
$\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},h_{\pi}\sim\pi}\big{[}\ell\big{(}h_{\pi},(x,y)\big{)}\big{]}\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}+\varepsilon.$ (7)
Sampling and learning processes.
In order to achieve the aforementioned goal (7), we need to draw samples from the available data distributions in $\mathcal{D}$, and the current paper focuses on sampling in an online fashion. More precisely, the learning process proceeds as follows: in each step $\tau$,
  • the learner selects $i_{\tau}\in[k]$ based on the previous samples;
  • the learner draws an independent sample $(x_{\tau},y_{\tau})$ from the data distribution $\mathcal{D}_{i_{\tau}}$.
The sample complexity of a learning algorithm thus refers to the total number of samples drawn from $\mathcal{D}$ throughout the learning process. A desirable learning algorithm would yield an $\varepsilon$-optimal (randomized) hypothesis (i.e., a hypothesis that achieves (7)) using as few samples as possible.
3 Algorithm
In this section, we present our proposed algorithm. Before proceeding, we find it convenient to introduce some notation concerning the loss under mixed distributions. Specifically, for any distribution $w=[w_{i}]_{1\leq i\leq k}\in\Delta(k)$ and any hypothesis $h\in\mathcal{H}$, the risk over the mixture $\sum_{i\in[k]}w_{i}\mathcal{D}_{i}$ of data distributions is denoted by:
$L(h,w)\coloneqq\sum_{i=1}^{k}w_{i}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]};$ (8a)
similarly, the risk of a randomized hypothesis $h_{\pi}$ (associated with $\pi\in\Delta(\mathcal{H})$) over $\sum_{i\in[k]}w_{i}\mathcal{D}_{i}$ is given by
$L(h_{\pi},w)\coloneqq\sum_{i=1}^{k}w_{i}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},h\sim\pi}\big{[}\ell\big{(}h_{\pi},(x,y)\big{)}\big{]}=\mathop{\mathbb{E}}\limits_{h\sim\pi}\big{[}L(h,w)\big{]}.$ (8b)
1 input: $k$ data distributions $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, hypothesis class $\mathcal{H}$, target accuracy level $\varepsilon$, target success rate $1-\delta$.
2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}:=\frac{400\left(k\log(\frac{k}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}.$
3 initialization: for all $i\in[k]$, set $W^{1}_{i}=1$, $\widehat{w}^{0}_{i}=0$ and $n_{i}^{0}=0$; $\mathcal{S}=\emptyset$.
4 for $t=1,2,\ldots,T$ do
5       set $w^{t}=[w^{t}_{i}]_{1\leq i\leq k}$ and $\widehat{w}^{t}=[\widehat{w}^{t}_{i}]_{1\leq i\leq k}$, with $w^{t}_{i}\leftarrow\frac{W^{t}_{i}}{\sum_{j}W^{t}_{j}}$ and $\widehat{w}^{t}_{i}\leftarrow\widehat{w}^{t-1}_{i}$ for all $i\in[k]$.
       /* recompute $\widehat{w}^{t}$ draw new samples for $\mathcal{S}_{\mathrm{w}}$ only if $w^{t}$ changes sufficiently. */
6       if there exists $j\in[k]$ such that $w_{j}^{t}\geq 2\widehat{w}_{j}^{t-1}$  then
7             $\widehat{w}^{t}_{i}\leftarrow\max\{w^{t}_{i},\widehat{w}^{t-1}_{i}\}$ for all $i\in[k]$;
8             for $i=1,\ldots,k$ do
9                   $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
10                   draw $n_{i}^{t}-n_{i}^{t-1}$ independent samples from $\mathcal{D}_{i}$, and add these samples to $\mathcal{S}$.
11
      /* estimate the near-optimal hypothesis for weighted data distributions. */
12       compute $h^{t}\leftarrow\arg\min_{h\in\mathcal{H}}\widehat{L}(h,w^{t})$, where
$\widehat{L}^{t}(h,w^{t}):=\sum_{i=1}^{k}\frac{w^{t}_{i}}{n_{i}^{t}}\cdot\sum_{j=1}^{n_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (9)
with $(x_{i,j},y_{i,j})$ being the $j$-th datapoint from $\mathcal{D}_{i}$ in $\mathcal{S}$.
       /* estimate the loss vector and execute weighted updates. */
13       $\overline{w}^{t}_{i}\leftarrow\max_{1\leq\tau\leq t}w^{\tau}_{i}$ for all $i\in[k]$.
14       for $i=1,\ldots,k$ do
             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
$\widehat{r}^{t}_{i}=\frac{1}{\lceil k\overline{w}^{t}_{i}\rceil}\sum_{j=1}^{\lceil k\overline{w}^{t}_{i}\rceil}\ell\big{(}h^{t},(x_{i,j}^{t},y_{i,j}^{t})\big{)};$
update the weight as $W^{t+1}_{i}=W^{t}_{i}\exp(\eta\widehat{r}_{i}^{t})$. // Hedge updates.
15
16
output: a randomized hypothesis $h^{\mathsf{final}}$ as a uniform distribution over $\{h^{t}\}_{t=1}^{T}$.
Algorithm 1 Hedge for Multi-distribution Learning ($\mathtt{MDL-Hedge}$)
Following the game dynamics proposed in previous works (Awasthi et al.,, 2023; Haghtalab et al.,, 2022), our algorithm alternates between computing the most favorable hypothesis (performed by the learner) and estimating the most challenging mixture of data distributions (performed by the adversary), with the aid of no-regret learning algorithms (Roughgarden,, 2016; Shalev-Shwartz,, 2012). More specifically, in each round $t$, our algorithm performs the following two steps:
  • (a)
    Given a mixture of data distributions $\mathcal{D}^{(t)}=\sum_{i\in[k]}w_{i}^{t}\mathcal{D}_{i}$ (with $w^{t}=[w_{i}^{t}]_{i\in[k]}\in\Delta(k)$), we construct a dataset to compute a hypothesis $h^{t}$ that nearly minimizes the loss under $\mathcal{D}^{(t)}$, namely,
    $h^{t}\approx\arg\min_{h\in\mathcal{H}}L(h,w^{t}).$ (10)
    This is accomplished by calling an empirical risk minimization oracle.
  • (b)
    Given hypothesis $h^{t}$, we compute an updated weight vector $w^{t+1}\in\Delta(k)$ — and hence an updated mixed distribution $\mathcal{D}^{(t+1)}=\sum_{i\in[k]}w_{i}^{t+1}\mathcal{D}_{i}$. The weight updates are carried out using the celebrated Hedge algorithm (Freund and Schapire,, 1997) designed for online adversarial learning, [^4] Note that the Hedge algorithm is closely related to Exponentiated Gradient Descent, Multiplicative Weights Update, Online Mirror Descent, etc (Arora et al.,, 2012; Shalev-Shwartz,, 2012; Hazan,, 2022). in an attempt to achieve low regret even when the loss vectors are adversarially chosen. More precisely, we run
    $w^{t+1}_{i}~{}\propto~{}w^{t}_{i}\exp\big{(}\eta\widehat{r}_{i}^{t}\big{)},\qquad i\in[k],$ (11)
    where the loss vector $\widehat{r}^{t}=[\widehat{r}_{i}^{t}]_{i\in[k]}$ contains the empirical loss of $h^{t}$ under each data distribution, i.e.,
    $\widehat{r}_{i}^{t}\approx\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h^{t},(x,y)\big{)}\big{]},\qquad i\in[k],$
    computed over another set of data samples.
At the end of the algorithm, we output a randomized hypothesis $h^{\sf final}$ that is uniformly distributed over the hypothesis iterates $\{h^{t}\}_{1\leq t\leq T}$ over all $T$ rounds, following common practice in online adversarial learning.
While the above paradigm has been adopted in past works (Awasthi et al.,, 2023; Haghtalab et al.,, 2022), the resulting sample complexity depends heavily upon how data samples are collected and utilized throughout the learning process. For instance, Awasthi et al., (2023, Algorithm 1) draws fresh data at each step of every round, in order to ensure reliable estimation of the loss function of interest through elementary concentration inequalities. This strategy, however, becomes wasteful over time, constituting the main source of its sample sub-optimality.
In order to make the best use of data, we propose the following key strategies.
  • Sample reuse in Step (a). In stark contrast to Awasthi et al., (2023, Algorithm 1) that draws new samples for estimating each $h^{t}$, we propose to reuse all samples collected in Step (a) up to the $t$-th round to assist in computing $h^{t}$. As will be made precise in lines 1-1 of Algorithm 1, we shall maintain a growing dataset $\mathcal{S}$ for conducting Step (a) throughout, ensuring that there are $n_{i}^{t}$ samples drawn from distribution $\mathcal{D}_{i}$ in the $t$-th round. These datapoints are employed to construct an empirical loss estimator $\widehat{L}^{t}(h,w^{t})$ for each $h\in\mathcal{H}$ in each round $t$, with the aim of achieving uniform convergence $|\widehat{L}^{t}(h,w^{t})-L(h,w^{t})|\leq O(\varepsilon)$ over all $h\in\mathcal{H}$. More detailed explanations are provided in Section 4.1.
  • Weighted sampling for Step (b). As shown in line 1 of Algorithm 1, in each round $t$, we sample each $\mathcal{D}_{i}$ a couple of times to compute the empirical estimator for $\mathbb{E}_{(x,y)\in\mathcal{D}_{i}}\big{[}\ell(h^{t},(x,y))\big{]}$, where the number of samples depends upon the running weights $\{w_{i}^{\tau}\}$. More precisely, we collect $\left\lceil k\overline{w}_{i}^{t}\right\rceil$ fresh samples from each $\mathcal{D}_{i}$, where $\overline{w}_{i}^{t}:=\max_{1\leq\tau\leq t}w_{i}^{\tau}$ is the maximum weight assigned to $\mathcal{D}_{i}$ up to now. Informally speaking, this strategy ensures reduced variance of the estimators and ultimately allows for an improved bound for $\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$. The interested reader is referred to Section 4.2 and Lemma 29 for more detailed explanations.
The whole procedure can be found in Algorithm 1.
4 Technique overview
In this section, we introduce the high-level ideas of the algorithm and analysis. Since Haghtalab et al., (2022) provided a nearly matching upper bound when $k=O(1/\varepsilon)$ (See the second row in Table 1), we assume that $k\geq\frac{100}{\varepsilon}$ throughout this work. We discuss the two primary technical novelties of the uniform convergence of the importance sampling estimator that allows for reusing samples (Section 4.1), and the upper bound of the coordinate-wise maximal $\ell_{1}$ norm of the $\{w^{t}\}$ iterates (Section 4.2).
4.1 Uniformly-convergent estimators
Let $\mathcal{X}$ be a proper discretization of $\Delta(k)$ with size $\exp(O(k\log(kd/\varepsilon)))$. We consider to study the approximation of the natural importance sampling estimator. Fix $n=\{n_{i}\}_{i=1}^{k}$ where each $n_{i}$ is a positive integer. Suppose that we have $n_{i}$ i.i.d. samples from $\mathcal{D}_{i}$ for each $i\in[k]$, denoted by $\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}$. For $w=[w_{i}]_{1\leq i\leq k}\in\mathcal{X}$, we let
$\widehat{L}_{n}(h,w)\coloneqq\sum_{i=1}^{k}w_{i}\cdot\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}.$ (12)
In view of the concentration property of VC classes (Mohri et al.,, 2018), we have
$\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\exp\left(\lambda\cdot\max_{h\in\mathcal{H}}\left|\widehat{L}_{n}(h,w)-L(h,w)\right|\right)\right]$
$\leq 2\cdot(2\|n\|_{1})^{d+1}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right)$ (13)
for all $0\leq\lambda\leq\min_{i}\frac{n_{i}}{w_{i}}$. As a result, for any $\kappa\geq 0$ and $0\leq\lambda\leq\kappa$ one has
$\sum_{w\in\mathcal{X}}\mathds{1}\left\{\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\exp\left(\lambda\cdot\max_{h\in\mathcal{H}}\left|\widehat{L}_{n}(h,w)-L(h,w)\right|\right)\right]$
$\leq\sum_{w\in\mathcal{X}}\mathds{1}\left\{\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}\cdot 2\cdot(2\|n\|_{1})^{d+1}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right)$
$\leq\sum_{w\in\mathcal{X}}\mathds{1}\left\{\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}\cdot 2\cdot(2\|n\|_{1})^{d+1}\exp\left(\frac{2\lambda^{2}}{\kappa}\right)$
$\leq 2\cdot|\mathcal{X}|\cdot(2\|n\|_{1})^{d+1}\exp\left(\frac{2\lambda^{2}}{\kappa}\right).$ (14)
We can then readily see that: for any $\varepsilon^{\prime}\in(0,1)$,
$\mathbb{P}\left\{\exists w\in\mathcal{X},\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa,\max_{h\in\mathcal{H}}\left|\widehat{L}_{n}(h,w)-L(h,w)\right|\geq\varepsilon^{\prime}\right\}$
$\leq 2\cdot|\mathcal{X}|\cdot(2\|n\|_{1})^{d+1}\min_{0\leq\lambda\leq\kappa}\exp\left(-\lambda\varepsilon^{\prime}+\frac{2\lambda^{2}}{\kappa}\right)$
$\leq 2\cdot|\mathcal{X}|\cdot(2\|n\|_{1})^{d+1}\exp\left(-\frac{\kappa(\varepsilon^{\prime})^{2}}{4}\right).$ (15)
Choose $\varepsilon^{\prime}=c\varepsilon$ for some proper constant $c$. By virtue of (15), for any $\delta\in(0,1)$, we can choose $\kappa\geq N_{0}(\varepsilon,\delta):=100\frac{k\log(kd/\varepsilon)+d\log(\|n\|_{1})+\log(1/\delta)}{\varepsilon^{2}}$ to ensure (15) is bounded by $\delta$. In other words, with probability at least $1-\delta$, $|\widehat{L}_{n}(h,w)-L(h,w)|\leq c\varepsilon$ holds uniformly for all $w\in\mathcal{X}$ such that $\max_{i}\frac{w_{i}}{n_{i}}\leq\frac{1}{N_{0}(\varepsilon,\delta)}$.
In the analysis above, we fix $n$ and show that uniform convergence holds with high probability. In fact, with similar arguments we can obtain uniform convergence over all proper choices of $(n,w)$. Let
[…]
$\mathcal{L}(N,\kappa)$$\coloneqq\left\{(w,n)\mid w\in\mathcal{X},n\in\mathcal{N}(N),\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}.$ (16b)
It can then be shown that for any $\varepsilon^{\prime}\in(0,1)$ and $\kappa>0$,
$\mathbb{P}\left[\exists(w,n)\in\mathcal{L}(N,\kappa),\left|\widehat{L}_{h}(h,w)-L(h,w)\right|\geq\varepsilon^{\prime}\right]$
$\leq 2\cdot|\mathcal{N}(N)|\cdot|\mathcal{X}|\cdot(2N)^{d+1}\exp\left(-\frac{\kappa(\varepsilon^{\prime})^{2}}{4}\right).$
Since $|\mathcal{N}(N)|\leq N^{k}$, it suffices to choose $\kappa=N^{*}\coloneqq O\left(\frac{(k+d)\mathrm{polylog}(k,d,1/\varepsilon)+\log(1/\delta)}{\varepsilon^{2}}\right)$ to ensure that the $O(\varepsilon)$-uniform convergence holds with probability at least $1-\delta$.
With the arguments above, with probability at least $1-\delta$, whenever we collect $n=\{n_{i}\}_{i=1}^{k}$ samples in the learning process, we could obtain $\varepsilon$-approximation $\widehat{L}_{n}(h,w)$ (see (12)) of $L(h,w)$ for all $h\in\mathcal{H}$ and $w\in\Delta(k)$ such that $\min_{i}\frac{n_{i}}{w_{i}}\geq N^{*}$. Most importantly, this result allows us to reuse samples in the learning process, instead of querying new samples to estimate $L(h,w^{t})$ in each round $t$.
Recall that $\overline{w}^{t}_{i}=\max_{1\leq\tau\leq t}w_{i}^{\tau}$. To ensure the $O(\varepsilon)$-uniform convergence with respect to $w^{t}$ in each round $t$, it suffices to choose $n_{i}^{t}\geq N^{*}\overline{w}_{i}^{t}$ for all $i\in[k]$.
4.2 Measuring the trajectory of Hedge
Our second technical point is to bound the maximal point of the whole trajectory $\{w^{t}\}_{t=1}^{T}$ of the Hedge algorithm. That is, we aim to control
$\|\overline{w}^{T}\|_{1}=\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}.$
To explain the high-level idea, we consider the learning process in Algorithm 2 with full knowledge of the loss vector.
1 Input: $\mathcal{Y}\subseteq[-1,1]^{k}$, target accuracy level $\varepsilon\in(0,1)$.
2 Initialization: $T=\frac{100\log(k)}{\varepsilon^{2}}$, $\eta=\frac{1}{10}\varepsilon$, $W^{1}_{i}=1$ for all $1\leq i\leq k$.
3 for $t=1,2,\ldots,T$ do
4       $w_{i}^{t}\leftarrow\frac{W_{i}^{t}}{\sum_{j}W_{j}^{t}}$ for all $1\leq i\leq k$.
5       $y^{t}\leftarrow\arg\min_{y\in\mathcal{Y}}(w^{t})^{\top}y$.
6       Update $W_{i}^{t+1}\leftarrow W_{i}^{t}\cdot\exp(\eta y_{i}^{t})$.
Algorithm 2 The Hedge algorithm for bilinear games.
As presented in Algorithm 2, there is a set of loss vectors $\mathcal{Y}$, and we want to solve the minimax problem
$\min_{y\in\mathsf{conv}(\mathcal{Y})}\max_{w\in\Delta(k)}w^{\top}y\qquad(\text{or equivalently, }\max_{w\in\Delta(k)}\min_{y\in\mathsf{conv}(\mathcal{Y})}w^{\top}y),$ (17)
where the equivalence arises from von Neumann’s minimax theorem (v. Neumann,, 1928). By Hedge algorithm, we can obtain an $O(\varepsilon)$-minimax solution in $\frac{100\log(k)}{\varepsilon^{2}}$ rounds. The question is that:
  • Can we bound $\|\overline{w}^{T}\|_{1}=\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ by polylogarithmic terms?
To explain our two key techniques, we assume that
$\min_{y\in\mathsf{conv}(\mathcal{Y})}\max_{w\in\Delta(k)}w^{\top}y=\max_{w\in\Delta(k)}\min_{y\in\mathsf{conv}(\mathcal{Y})}w^{\top}y=0,$
Given this assumption, we then have $(w^{t})^{\top}y^{t}=\min_{y\in\mathcal{Y}}(w^{t})^{\top}y\leq 0,\forall t\in[T]$ and $\frac{1}{T}\sum_{t=1}^{T}(w^{t})^{\top}y^{t}\geq-C\epsilon$ for some constant $C$. In the following analysis, we assume $-(w^{t})^{\top}y^{t}=O(\epsilon)$ holds for any $t\in[T]$. While this assumption is not true in most cases, we can divide $[T]$ into $j_{\mathrm{max}}:=O\left(\log_{2}(1/\epsilon)\right)$ disjoint subsets as $\{\{t\in[T]:-(w^{t})^{\top}y^{t}\in(2^{-(j+1)},2^{j}]\}\}_{j=0}^{j_{\mathrm{max}}-2}$ and $\{t\in[T]:-(w^{t})^{\top}y^{t}\leq 2^{-(j_{\mathrm{max}}-1)}=O(\epsilon)\}$ and then work on each subset. As a trade-off, this operation leads to some additional logarithmic factors (see Lemma 29).
4.2.1 Doubling $w_{j}$ needs $\widetilde{\Omega}(1/\varepsilon^{2})$ steps
Instead of bounding $\|\overline{w}^{T}\|_{1}$ directly, our first attempt is to show that there are at most $\widetilde{O}(1)$ dimensions $i$ such that $\max_{1\leq t\leq T}w^{t}_{i}\geq c$, with $c>0$ some universal constant. For simplicity, we take $c=1/4$.
Let
$\mathcal{W}\coloneqq\Big{\{}i\in[k]\mid\max_{1\leq t\leq T}w_{i}^{t}\geq 1/4\big{\}}.$ (18)
Given that the stepsize $\eta$ is small enough, one can find, for each $i\in[k]$, an interval $[s_{i},e_{i}]$ such that
$\frac{1}{16}\leq w_{i}^{s_{i}}\leq\frac{1}{8},\qquad w_{i}^{e_{i}}\geq\frac{1}{4}\qquad\text{and}\qquad w_{i}^{t}\geq\frac{1}{8}$ (19)
for any $t$ obeying $s_{i}<t\leq e_{i}$. Then we would like to show that $e_{i}-s_{i}\geq\widetilde{\Omega}(1/\varepsilon^{2})$.
Since the stepsize $\eta=O(\varepsilon)$, we can easily get that $e_{i}-s_{i}=\widetilde{\Omega}(1/\varepsilon)$. Meanwhile, when running the Hedge algorithm for adversarial bandits, the adversarial could choose the loss properly so that $e_{i}-s_{i}=\widetilde{\Theta}(1/\varepsilon)$. Fortunately, we can focus on a bilinear game where the opponent plays best response. By definition, for any $1\leq t\leq\tau\leq T$,
$(w^{t})^{\top}y^{t}\leq 0,\quad(w^{t})^{\top}y^{\tau}\geq(w^{t})^{\top}y^{t}.$
For any $1\leq t_{1}\leq t_{2}\leq T$, we then have
$(w^{t_{1}})^{\top}\sum_{\tau=t_{1}}^{t_{2}-1}y^{\tau}\geq(t_{2}-t_{1})(w^{t_{1}})^{\top}y^{t_{1}}.$ (20)
By observing that $\sum_{\tau=t_{1}}^{t_{2}-1}y_{i}^{\tau}=\frac{1}{\eta}\log\left(\frac{W^{t_{2}}_{i}}{W^{t_{1}}_{i}}\right)$, we have
$\sum_{i=1}^{k}w_{i}^{t_{1}}\log\left(\frac{W_{i}^{t_{2}}}{W_{i}^{t_{1}}}\right)\geq\eta(t_{2}-t_{1})(w^{t_{1}})^{\top}y^{t_{1}}.$ (21)
On the other hand, since the optimal minimax value is 00, we can show that
$\log\left(\frac{\sum_{i=1}^{k}W_{i}^{t_{2}}}{\sum_{i=1}^{k}W_{i}^{t_{1}}}\right)=\sum_{\tau=t_{1}}^{t_{2}-1}\log\left(\frac{\sum_{i=1}^{k}W_{i}^{\tau+1}}{\sum_{i=1}^{k}W_{i}^{\tau}}\right)\leq O\big{(}(t_{2}-t_{1})\eta^{2}\big{)}.$ (22)
Putting (21) and (22) together, we arrive at
$-\sum_{i=1}^{k}w_{i}^{t_{1}}\log\left(\frac{w_{i}^{t_{2}}}{w_{i}^{t_{1}}}\right)\leq O\left(\eta^{2}(t_{2}-t_{1})-\eta(t_{2}-t_{1})(w^{t_{1}})^{\top}y^{t_{1}}\right)=O(\eta^{2}(t_{2}-t_{1})),$ (23)
which is exactly
$\mathsf{KL}\big{(}w^{t_{1}}\,\|\,w^{t_{2}}\big{)}\leq O\big{(}\eta^{2}(t_{2}-t_{1})\big{)}.$ (24)
Finally, we obtain that
$t_{2}-t_{1}\geq\Omega\left(\frac{\mathsf{KL}\big{(}w^{t_{1}}\,\|\,w^{t_{2}}\big{)}}{\varepsilon^{2}}\right).$ (25)
By setting $t_{1}=s_{i}$ and $t_{2}=e_{i}$, we can deduce that
$e_{i}-s_{i}\geq\Omega\left(\frac{\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}}{\varepsilon^{2}}\right)=\Omega\left(\frac{1}{\varepsilon^{2}}\right).$ (26)
Also noting that for any $t$, there are at most $8$ dimensions $i$ such that $s_{i}\leq t\leq e_{i}$, it holds that
$8T\geq\sum_{i\in\mathcal{W}}(e_{i}-s_{i})\geq|\mathcal{W}|\cdot\Omega\left(\frac{1}{\varepsilon^{2}}\right),$ (27)
which implies that $|\mathcal{W}|$ is at most $O(T\varepsilon^{2})=O(\log(k))$.
In the analysis above, we analyze the property of the trajectory of Hedge algorithm with exact loss vector. However, in Algorithm 1, we need to sample to estimate the loss vector. In this case, (24) would be replaced by
$\mathsf{KL}(w^{t_{1}}\,\|\,w^{t_{2}})\leq O(\eta^{2}(t_{2}-t_{1}))+O\left(\eta\sum_{\tau=t_{1}}^{t_{2}-1}(w^{t_{1}}+w^{\tau})^{\top}\Delta^{\tau}\right),$ (28)
where $\Delta^{\tau}:=\left(\widehat{y}^{\tau}-y^{\tau}\right)$ with $\widehat{y}^{\tau}$ as the stochastic estimation of $y^{\tau}$. A naive bound for the second term in right hand side of (28) would be $\widetilde{O}\left(\eta\sqrt{(t_{2}-t_{1})}\right)$ by Azuma’s inequality. However, this bound is not sufficient because when $(t_{2}-t_{1})\varepsilon^{2}\ll 1$, the term $\widetilde{O}\left(\eta\sqrt{(t_{2}-t_{1})}\right)$would be the leading term. To address this problem, we consider weighted sampling to construct the empirical loss vector. As presented in Line 1 Algorithm 1, we manage to reduce the variance of $\Delta^{\tau}$ properly so that $\left(\eta\sum_{\tau=t_{1}}^{t_{2}-1}(w^{t_{1}}+w^{\tau})^{\top}\Delta^{\tau}\right)$ is roughly bounded by $O(\eta^{2}(t_{2}-t_{1}))$.
4.2.2 Dealing with the segments
In Section 4.2.1, we show that for any $1\leq t_{1}\leq t_{2}\leq T$
$t_{2}-t_{1}\geq\Omega\left(\frac{\mathsf{KL}\big{(}w^{t_{1}}\,\|\,w^{t_{2}}\big{)}}{\varepsilon^{2}}\right).$ (29)
By this inequality, we bound the size of $\mathcal{W}$ by $O(\log(k))$. A natural idea is to generalize the arguments to bound the size of
$\mathcal{W}(p):=\left\{i\in[k]|\max_{1\leq t\leq T}w_{i}^{t}\in[2p,4p]\right\}$
for $p\in[0,1]$. However, the arguments above can not provide a desired bound for $\mathcal{W}(p)$ when $p$ is small. For each $i\in\mathcal{W}(p)$, let $[s_{i},e_{i}]$ be the interval such that $\frac{p}{2}\leq w_{i}^{s_{i}}\leq p$, $w_{i}^{e_{i}}\geq 2p$ and $w_{i}^{t}\geq p$ for any $s_{i}<t\leq e_{i}$. By (29), we learn that
$e_{i}-s_{i}\geq\Omega\left(\frac{\mathsf{KL}\big{(}\mathrm{Ber}(p)\,\|\,\mathrm{Ber}(2p)\big{)}}{\varepsilon^{2}}\right)\geq\Omega\left(\frac{p}{\varepsilon^{2}}\right),$ (30)
where $\mathrm{Ber}(q)$ denote the Bernoulli distribution with mean $q$. Following the arguments in (27), we hope to bound
$\sum_{i\in\mathcal{W}(p)}(e_{i}-s_{i})=O(T).$ (31)
However, this might be not true because for each $t\in[T]$, there are possibly more than $\frac{1}{2p}$ coordinates $i\in\mathcal{W}(p)$ such that $t\in[s_{i},e_{i}]$, where we can at most derive $|\mathcal{W}(p)|\leq O(T\varepsilon^{2}/p^{2})$. As a result, the contribution of coordinates in $\mathcal{W}(p)$ to the target $\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ is $\widetilde{O}(1/p)$, which could be much larger than $\widetilde{O}(1)$ when $p$ is small.
To address this problem, we note that if there are some coordinates $i\in\mathcal{W}(p)$ sharing similar $[s_{i},e_{i}]$, we can obtain a better bound. For example, suppose that for any $i\in\overline{\mathcal{W}}(p)\subset\mathcal{W}(p)$, $s_{i}=s$ and $e_{i}=e$, we then have that
$e-s\geq\Omega\left(\frac{|\overline{\mathcal{W}}(p)|p}{\varepsilon^{2}}\right),$ (32)
which is much better compared to the original bound $\Omega\left(\frac{p}{\varepsilon^{2}}\right)$ if $|\overline{\mathcal{W}}(p)|$ is large.
By this observation, we consider to merge the dimensions in $\mathcal{W}(p)$ with similar $[s_{i},e_{i}]$. To facilitate analysis, we introduce the notation of segment.
Definition 2 (Segment).
For $p,x>0$ and $i\in[k]$, we say $(t_{1},t_{2})$ is a $(p,q,x)$-segment iff there exists $\mathcal{I}\subset[K]$ such that $\sum_{i\in\mathcal{I}}w^{t_{1}}_{i}\in\left[\frac{p}{2},p\right]$, $\sum_{i\in\mathcal{I}}w^{t_{2}}_{i}\geq p\exp(x)$ and $\sum_{i\in\mathcal{I}}w_{i}^{t}\geq q$ for any $t_{1}\leq t\leq t_{2}$. We name $t_{1}$ as the starting point and $t_{2}$ as the ending point. And we also call $\mathcal{I}$ as the index set.
By choosing $p/4\leq q\leq p/2$, a $(p,q,x)$-segment captures the variation of probabilities from $p$ to $p\exp(x)$. We say two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$ are disjoint iff $s_{1}<e_{1}\leq s_{2}<e_{2}$ or $s_{2}<e_{2}\leq s_{1}<e_{1}$. Using (29), it is easy to derive $t_{2}-t_{1}\geq\Omega(px^{2}/\varepsilon^{2})$ for a $(p,p/2,x)$-segment $(t_{1},t_{2})$.
For each $i\in\mathcal{W}(p)$, there exists a $\left(\frac{p}{4},\frac{p}{8},\log(2)\right)$-segment $(s_{i},e_{i})$ with index set =$\mathcal{I}=\{i\}$. It then follows that $e_{i}-s_{i}\geq\Omega\left(\frac{p}{\varepsilon^{2}}\right)$ for each $i\in\mathcal{W}(p)$. As a result, if the segments $(s_{i},e_{i})$ are mutually disjoint, $|\mathcal{W}(p)|$ is at most $O(\log(k)/p)$. More generally, if we can divide the $|\mathcal{W}(p)|$ segments into disjoint blocks such that the segments in the same block share the same starting and ending points, we can also derive that $|\mathcal{W}(p)|\leq O(\log(k)/p)$. Taking Figure 1 as an example. Suppose we have $\ell$ blocks and the $i$-th block (with starting-ending points as $(\widetilde{s}^{i},\widetilde{e}^{i})$) contains $m_{i}$ coordinates satisfying that $\tilde{s}^{1}<\tilde{e}^{1}\leq\tilde{s}^{2}<\tilde{e}^{2}\leq\dots\leq\tilde{s}^{\ell}<\tilde{e}^{\ell}$. Then $(\widetilde{s}^{i},\widetilde{e}^{i})$ constructs a $\left(m_{i}\cdot\frac{p}{4},m_{i}\cdot\frac{p}{8},\log(2)\right)$-segment with index set as the $m_{i}$ coordinates in the $i$-th block, which means that
$\widetilde{e}^{i}-\widetilde{s}^{i}\geq\Omega\left(\frac{m_{i}p}{\varepsilon^{2}}\right).$
Taking sum over $1\leq i\leq\ell$, we obtain
$|\mathcal{W}(p)|\leq\sum_{i=1}^{\ell}m_{i}\leq O\left(\sum_{i=1}^{\ell}\frac{(\widetilde{e}^{i}-\widetilde{s}^{i})\varepsilon^{2}}{p}\right)\leq O\left(\frac{T\varepsilon^{2}}{p}\right)=O\left(\frac{\log(k)}{p}\right),$ (33)
which implies that $\sum_{i\in\mathcal{W}(p)}\max_{1\leq t\leq T}w_{i}^{t}\leq O(\log(k))$. With standard doubling arguments, it follows that $\sum_{i\in[k]}\max_{1\leq t\leq T}w_{i}^{t}=O(\log^{2}(k))$.
By this observation, we introduce the concept of a regular configuration (see an example in Figure 1).
Definition 3 (Configuration).
A configuration $\mathsf{Conf}$ is a set of intervals $\mathsf{Conf}=\{[a_{i},b_{i}]\}_{i=1}^{m}$ with $b_{i}>a_{i}$ for each $i\in[m]$ (repeatable elements are allowed). We say the configuration $\mathsf{Conf}$ is regular iff for any $i,j\in[m]$, one of the three following claims holds.
  • $a_{i}<b_{i}\leq a_{j}<b_{j}$;
  • $a_{j}<b_{j}\leq a_{i}<b_{i}$;
  • $a_{i}=a_{j}$, $b_{i}=b_{j}$.
In words, (33) states that if $\{[s_{i},e_{i}]\}_{i\in\mathcal{W}(p)}$ forms a regular configuration, we then have $|\mathcal{W}(p)|=O\left(\frac{\log(k)}{p}\right)$ However, the general configuration of the segments might be irregular because it is possible that two segments are intersecting (see Figure 2). That is, for two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$, it holds that $s_{1}<s_{2}<e_{1}<e_{2}$. In this case, the approach above can at most bound $e_{2}-s_{1}$ by $\max\{e_{2}-s_{2},e_{1}-s_{1}\}$ instead of the desired bound $(e_{1}-s_{1})+(e_{2}-s_{2})$. To address this problem, we consider to construct a regular configuration with sub-segments [^5] A sub-segment is equivalent to a sub-interval of a segment. We refer to Figure 6 for more details. of the original segments.
Our first step is to align one side of the segments. As presented in Figure 3, we divide the whole learning process into disjoint blocks such that the segments in each block have a common inner point. In the meantime, a segment is discarded if it intersects with more than one blocks. We show that there exists a regular configuration of the blocks such that at most $(1-\frac{1}{3(\log_{2}(T)+1)})\cdot|\mathcal{W}(p)|$ segments are discarded. In other words, at least $\frac{1}{3(\log_{2}(T)+1)}\cdot|\mathcal{W}(p)|$ segments are contained by these blocks. Since the blocks are disjoint to each other, it suffices to operate on one block. Without loss of generality, we assume that there is only one block, which also means that all segments have a common inner point $t_{\mathrm{mid}}$. Then we divide this block into two parts according to $t_{\mathrm{mid}}$ (see Figure 4). Since
$\log\left(\frac{w_{i}^{e_{i}}}{w_{i}^{t_{\mathrm{mid}}}}\right)+\log\left(\frac{w_{i}^{t_{\mathrm{mid}}}}{w_{i}^{s_{i}}}\right)\geq\log(2),$
it then holds either $(s_{i},t_{\mathrm{mid}})$ is a $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment with index ${i}$ or $(t_{\mathrm{mid}},e_{i})$ is a $\left(w_{i}^{t_{\mathrm{mid}}},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment with index set $\{i\}$. Since $w_{i}^{t_{\mathrm{mid}}}\in[p/4,p/2]$, we can roughly view $(t_{\mathrm{mid}},e_{i})$ as a $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment in the latter case. Therefore, we can find at least $\frac{|\mathcal{W}(p)|}{3(\log_{2}(T)+1)}\cdot\frac{1}{2}$ different $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$ segments with the same starting (ending) points.
Without loss of generality, we assume that these $\frac{|\mathcal{W}(p)|}{6(\log_{2}(T)+1)}\cdot$ $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segments share a common starting point. As presented in Figure 5, we denote the common starting point as $e_{0}$, and re-order the dimensions so that $e_{i}$ is non-decreasing in $i$. Then we construct $O(\log_{2}(k))$ regular configurations using a recursion (see Lemma 28) , and show that at least one of these configurations contains $\widetilde{O}\left(|\mathcal{W}(p)|\right)$ $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2(\log_{2}(k)+2)}\right)$-segments.
In Figure 6, we present a toy example in the case $k=8$. In this toy example We in total have $5$ configurations with 5 different colors. For each $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment in Figure 5, it contributes an $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2(\log_{2}(k)+2)}\right)$-segment to at least one of these configurations. By the pigeonhole princeton, there must be one regular configuration with $\frac{|\mathcal{W}(p)|}{6(\log_{2}(T)+1)(\log_{2}(k)+2)}$ $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2(\log_{2}(k)+2)}\right)$-segments (see Lemma 28).
Figure 1: Regular configuration.
Figure 2: General irregular configuration.
Figure 3: Partition of blocks.
Figure 4: Aligning one side of the configuration. The unfilled part of the segments mean that the variation over $w_{i}^{s_{i}}\to w_{i}^{t_{\mathrm{mid}}}$ (i.e., $\log\left(w_{i}^{t_{\mathrm{mid}}}/w_{i}^{s_{i}}\right)$) is not significant enough.
Figure 5: A group of segments with common starting point.
Figure 6: Construction of the regular configurations. Each segment is cut into at most $\log_{2}(k)+1$ sub-segments with different colors.
5 Analysis (proof of Theorem 1)
Theorem 1 follows by combining Lemma 4, Lemma 5 and Lemma 6.
Recall that $\varepsilon_{1}=\frac{\varepsilon}{100}$. In Lemma 4, we show that, with high-probability,
$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (34)
holds for any $1\leq t\leq T$. Assuming (34) holds for any $1\leq t\leq T$, we continue by proving that $h^{\mathsf{final}}$ is $\varepsilon$-optimal in Lemma 5, and providing a bound for the sample complexity in Lemma 6.
Lemma 4.
With probability at least $41-\delta/$ 4,
$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (35)
holds for any $1\leq t\leq T$.
Lemma 5.
Assume lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability exceeding $41-\delta/$ 4, the output policy $h^{\mathsf{final}}$ is $\varepsilon$-optimal.
Lemma 6.
Assume lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability at least $21-\delta/$ 2, the sample complexity of Algorithm 1 is bounded by
$O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\left(\log^{5}(k)\log(1/\varepsilon)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)\right).$
5.1 Optimality (proof of Lemma 5)
Let $\mathrm{OPT}:=\min_{h\in\mathcal{H}}\max_{i}L(h,\mathbf{e}_{i})$. Let $v^{t}=L(h^{t},w^{t})-\mathrm{OPT}$ and $f^{t}=\min_{h\in\mathcal{H}}L(h,w^{t})-\mathrm{OPT}\leq 0$. By the assumption that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$, we have $v^{t}\leq f^{t}+\varepsilon_{1}\leq\varepsilon_{1}$ for any $1\leq t\leq T$.
Let $\delta^{\prime}=\frac{\delta}{T+k+1}$. Note that $\eta\leq 1$ and $|\hat{r}_{i}^{t}|\leq 1$ for any proper $(i,t)$. Direct computation gives that
$\log\left(\frac{\sum_{i=1}^{k}W_{i}^{t+1}}{\sum_{i=1}^{k}W_{i}^{t}}\right)$ $=\log\left(\sum_{i=1}^{k}w_{i}^{t}\exp(\eta\hat{r}^{t}_{i})\right)\leq\log\left(\sum_{i=1}^{k}w_{i}^{t}(1+\eta\hat{r}_{i}^{t}+\eta^{2}(\hat{r}_{i}^{t})^{2})\right)\leq\eta\sum_{i=1}^{k}w_{i}^{t}\hat{r}_{i}^{t}+\eta^{2}.$ (36)
As a result, we have that
$\eta\sum_{t=1}^{T}w^{t}\cdot\hat{r}^{t}$ $\geq\log\left(\sum_{i=1}^{k}W_{i}^{T+1}\right)-T\eta^{2}-\log(k)$
$\geq\max_{i}\log(W_{i}^{T})-T\eta^{2}-\log(k)$
$\geq\eta\max_{i}\sum_{t=1}^{T}\hat{r}_{i}^{t}-T\eta^{2}-\log(k).$ (37)
By dividing both side with $\eta$, we have
$\sum_{t=1}^{T}w^{t}\cdot\hat{r}^{t}\geq\max_{i}\sum_{t=1}^{T}\hat{r}^{t}_{i}-\left(\frac{\log(k)}{\eta}+\eta T\right).$ (38)
By Azuma’s inequality, with probability $1-(k+1)\delta^{\prime}$, we have that
$\left|\sum_{t=1}^{T}w^{t}\cdot\hat{r}^{t}-\sum_{t=1}^{T}L(h^{t},w^{t})\right|\leq 2\sqrt{T\log(1/\delta^{\prime})};$
$\left|\sum_{t=1}^{T}\hat{r}_{i}^{t}-\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})\right|\leq 2\sqrt{T\log(1/\delta^{\prime})}.$
As a result, with probability $1-(k+1)\delta^{\prime}$, it holds that
$\sum_{t=1}^{T}L(h^{t},w^{t})$ $\geq\max_{i}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})-\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)$
$\geq-\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right).$ (39)
By the fact that $\varepsilon_{1}=\eta=\frac{1}{100}\varepsilon$ and $T=\frac{20000\log(\frac{k}{\delta^{\prime}\varepsilon})}{\varepsilon^{2}}$, we have that
$\max_{i}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})$ $\leq T\mathrm{OPT}+\sum_{t=1}^{T}v^{t}+\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)$
$\leq T\mathrm{OPT}+T\varepsilon_{1}+\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)$
$\leq T\mathrm{OPT}+T\varepsilon.$ (40)
As a result, we obtain that
$\max_{i}L(h^{\mathsf{final}},\mathbf{e}_{i})=\max_{i}\frac{1}{T}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})\leq\mathrm{OPT}+\varepsilon.$ (41)
The proof is finished by noting that $\delta^{\prime}=\frac{\delta}{(T+k+1)}$.
5.2 Sample complexity (proof of Lemma 6)
Recall that $\{w^{t}\}_{t=1}^{T+1}$ is the trajectory by running Algorithm 1. Also recall that $\overline{w}^{T}_{i}=\max_{1\leq t\leq T}w_{i}^{t}$ for all $i\in[k]$. We start with the lemma below to bound the sample complexity by $\|\overline{w}^{T}\|_{1}$.
Lemma 7.
The sample complexity of Algorithm 1 is bounded by $O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\|\overline{w}^{T}\|_{1}\right)$.
Proof.
Recall the definition of $\check{w}^{t}$ and $\overline{w}^{t}$ in Algorithm 1. It is easy to observe that $\check{w}^{t}\leq\overline{w}^{t}.$ Then the total number of samples is bounded by
$O\left(T_{1}\cdot\|\overline{w}^{T}\|_{1}+T(k\|\overline{w}^{T}\|_{1}+k)\right)=O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\|\overline{w}^{T}\|_{1}\right).$
On the other hand, we have the lemma below to bound $\|\overline{w}^{T}\|_{1}$.
Lemma 8.
Assume Line 1-1 in Algorithm 1 is replaced by some oracle which returns a hypothesis $h^{t}$ satisfies that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability $41-\delta/$ 4, the sample complexity of Algorithm 1 is bounded by
$O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\left(\log^{5}(k)\log(1/\varepsilon)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)\right).$
The proof of Lemma 8 is rather involved. We postpone the proof to Appendix C.1 due to space limitation.
By Lemma 7 and Lemma 8, with probability $21-\delta/$ 2, the sample complexity of Algorithm 1 is bounded by
$O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\left(\log^{5}(k)\log(1/\varepsilon)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)\right).$
The proof is completed.
6 Extension: learning Rademacher classes
In this section, we study how to adapt our algorithm and theory to accommodate MDL for Rademacher classes.
1 input: $k$ data distributions $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, hypothesis class $\mathcal{H}$, target accuracy level $\varepsilon$, target success rate $1-\delta$, the constant $\{C_{n}\}_{n\geq 1}$ in Assumption 10.
2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}=\min\left\{t\geq\frac{400\left(k\log(\frac{k}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}\Big{|}C_{t}\leq\frac{\epsilon_{1}}{1200}\right\}$.
3 initialization: for all $i\in[k]$, set $W^{1}_{i}=1$, $\widehat{w}^{0}_{i}=0$ and $n_{i}^{0}=0$; $\mathcal{S}=\emptyset$.
4 draw $\left\lceil 12\log(2k)\right\rceil$ samples from $\mathcal{D}_{i}$ for each $i$, and add these samples to $\mathcal{S}$.
5 for $t=1,2,\ldots,T$ do
6       set $w^{t}=[w^{t}_{i}]_{1\leq i\leq k}$ and $\widehat{w}^{t}=[\widehat{w}^{t}_{i}]_{1\leq i\leq k}$, where $w^{t}_{i}\leftarrow\frac{W^{t}_{i}}{\sum_{j}W^{t}_{j}}$ and $\widehat{w}^{t}_{i}\leftarrow\widehat{w}^{t-1}_{i}$ for all $i\in[k]$.
       /* recompute $\widehat{w}^{t}$ draw new samples for $\mathcal{S}_{\mathrm{w}}$ only if $w^{t}$ changes sufficiently. */
7       if there exists $j\in[k]$ such that $w_{j}^{t}\geq 2\widehat{w}_{j}^{t-1}$  then
8             $\widehat{w}^{t}_{i}\leftarrow\max\{w^{t}_{i},\widehat{w}^{t-1}_{i}\}$ for all $i\in[k]$;
9             for $i=1,\ldots,k$ do
10                   $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
11                   draw $n_{i}^{t}-n_{i}^{t-1}$ independent samples from $\mathcal{D}_{i}$, and add these samples to $\mathcal{S}$.
12
      /* estimate the near-optimal hypothesis for weighted data distributions. */
13       compute $h^{t}\leftarrow\arg\min_{h\in\mathcal{H}}\widehat{L}(h,w^{t})$, where
$\widehat{L}^{t}(h,w^{t}):=\sum_{i=1}^{k}\frac{w^{t}_{i}}{\check{n}_{i}^{t}}\cdot\sum_{j=1}^{\check{n}_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (42)
with $\check{n}_{i}^{t}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$, and $(x_{i,j},y_{i,j})$ being the $j$-th datapoint from $\mathcal{D}_{i}$ in $\mathcal{S}$.
       /* estimate the loss vector and execute weighted updates. */
14       $\overline{w}^{t}_{i}\leftarrow\max_{1\leq\tau\leq t}w^{\tau}_{i}$ for all $i\in[k]$.
15       for $i=1,\ldots,k$ do
             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
$\widehat{r}^{t}_{i}=\frac{1}{\lceil k\overline{w}^{t}_{i}\rceil}\sum_{j=1}^{\lceil k\overline{w}^{t}_{i}\rceil}\ell\big{(}h^{t},(x_{i,j}^{t},y_{i,j}^{t})\big{)};$
update the weight as $W^{t+1}_{i}=W^{t}_{i}\exp(\eta\widehat{r}_{i}^{t})$. // Hedge updates.
16
17
output: a randomized hypothesis $h^{\mathsf{final}}$ as a uniform distribution over $\{h^{t}\}_{t=1}^{T}$.
Algorithm 3 Hedge for Multi-distribution Learning on Rademacher Class ($\mathtt{MDL-Hedge-Rademacher}$)
6.1 Preliminaries: Rademacher complexity
Let us first introduce the formal definition of the Rademacher complexity; more detailed introduction can be found in Shalev-Shwartz and Ben-David, (2014).
Definition 9 (Rademacher complexity).
Given a distribution $\mathcal{D}$ supported on $\mathcal{Z}\coloneqq\mathcal{X}\times\mathcal{Y}$ and a positive integer $n$, the (average) Rademacher complexity is defined as
$\mathsf{Rad}_{n}(\mathcal{D})\coloneqq\mathop{\mathbb{E}}_{\{z_{i}\}_{i=1}^{n}}\left[\mathop{\mathbb{E}}_{\{\sigma_{i}\}_{i=1}^{n}}\left[\max_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\ell(h,z_{i})\right]\right],$ (43)
where $\{z_{i}\}_{i=1}^{n}$ are drawn independently from $\mathcal{D}$, and $\{\sigma_{i}\}_{i=1}^{n}$ are i.i.d. Rademacher random variables obeying $\mathbb{P}\{\sigma_{i}=1\}=\mathbb{P}\{\sigma_{i}=-1\}=1/2$ for each $1\leq i\leq n$.
Next, we would like to make an assumption on the Rademacher complexity of mixtures of distributions. Denoting by $\mathcal{D}(w)$ the mixed distribution
$\mathcal{D}(w)\coloneqq\sum_{i=1}^{k}w_{i}\mathcal{D}_{i}$ (44)
for any probability vector $w\in\Delta(k)$, we can state our assumption as follows.
Assumption 10.
For each $n\geq 1$, there exists a universal constant $C_{n}>0$ such that
$\mathsf{Rad}_{n}\big{(}\mathcal{D}(w)\big{)}\leq C_{n}$ (45)
holds for all $w\in\Delta^{k}$.
Remark 1.
One might raise a question about Assumption 10: Can we only assume $\mathrm{Rad}_{n}(\mathcal{D}_{i})\leq C_{n}$ for $i\in[k]$ without incuring a worse sample complexity? The answer is negative. Actually, the Rademacher complexity $\mathsf{Rad}_{n}(\mathcal{D}(w))$ is not convex in $w$, where we fail to use $\max_{i}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ to bound $\max_{w\in\Delta(k)}\mathsf{Rad}_{n}(\mathcal{D}(w))$. We refer to Appendix D.4 for more details.
It is well known that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ implies Assumption 10 holds with $C_{n}=\sqrt{\frac{2d\log(en/d)}{n}}$ (Mohri et al.,, 2018).
To facilitate our analysis in this section, we find it helpful to introduce the notion of the weighted Rademacher complexity as follows.
Definition 11 (Weighted Rademacher complexity).
Given a collection of distributions $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a set of positive integers $\{n_{i}\}_{i=1}^{k}$, the weighted (average) Rademacher complexity is defined as
$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\mathcal{D})\coloneqq\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right],$ (46)
where $\big{\{}\{z_{i}^{j}\}_{j=1}^{n_{k}}\big{\}}_{i=1}^{k}$ are independently generated with each $z_{i}^{j}$ drawn from $\mathcal{D}_{i}$, and $\big{\{}\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\big{\}}_{i=1}^{k}$ are independent Rademacher random variables obeying $\mathbb{P}\{\sigma_{i}^{j}=1\}=\mathbb{P}\{\sigma_{i}^{j}=-1\}=1/2$. Throughout the rest of this paper, we shall often abbreviate $\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}=\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\mathcal{D})$.
The weighted Rademacher complexity defined above satisfies an important property below.
Lemma 12.
For any two groups of positive integers $\{n_{i}\}_{i=1}^{k}$ and $\{m_{i}\}_{i=1}^{k}$, it holds that
$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$ $\leq\left(\sum_{i=1}^{k}(m_{i}+n_{i})\right)\widetilde{\mathsf{Rad}}_{\{m_{i}+n_{i}\}_{i=1}^{k}}$
$\leq\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$ (47)
In addition, the following allows us to bound the weighted Rademacher complexity under Assumption 10.
Lemma 13.
Consider any $\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for each $i\in[k]$. By taking $w\in\Delta^{k}$ with $w_{i}=\frac{n_{i}}{\sum_{l=1}^{k}n_{l}}$, one has
$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}\leq 72\mathsf{Rad}_{\sum_{i=1}^{k}n_{i}}\big{(}\mathcal{D}(w)\big{)}.$
6.2 Algorithm and sample complexity
We present in Algorithm 3 to learn a Rademacher class under multi-distribution tasks. Algorithm 3 is also a Hedge-like algorithm to learn a convex (concave) game. The major difference from Algorithm 1 lies in the subroutine to learn $h^{t}$ (See lines 3-3 in Algorithm 3). More precisely, to compute the estimator $\widehat{L}^{t}(h,w^{t})$ for $L(h,w^{t})$, instead of using the first $n_{i}^{t}$ samples from $\mathcal{D}_{i}$ for each $i\in[k]$, we choose to use the first $\check{n}_{i}^{t}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ samples from $\mathcal{D}_{i}$ for each $i$.
Formally, we have the result below.
Theorem 14.
Suppose Assumption 10 holds. By running Algorithm 3, with probability $1-\delta$, the output $h^{\mathsf{final}}$ satisfies that
$\max_{i}L(h^{\mathsf{final}},\mathbf{e}_{i})=\max_{i}\frac{1}{T}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})\leq\min_{h\in\mathcal{H}}\max_{i}L(h,\mathbf{e}_{i})+\varepsilon.$
Meanwhile, the sample complexity of Algorithm 3 is bounded by
$O\left(T_{\epsilon}\cdot\log^{5}(k)\log\left(\frac{1}{\varepsilon}\right)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)$
where $T_{\epsilon}$ is defined as
$T_{\epsilon}=\min\left\{t\geq\frac{400\left(k\log(\frac{k}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}\Big{|}C_{t}\leq\frac{\epsilon_{1}}{1200}\right\}.$
In the case $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$, we have $C_{n}\leq\sqrt{\frac{2d\log(en/d)}{n}}$, which implies that $T_{\epsilon}=\tilde{O}\left(\frac{d+k}{\epsilon^{2}}\right)$ and a sample complexity bound of $\tilde{O}\left(\frac{d+k}{\epsilon^{2}}\right)$.
Proof of Theorem 14.
By Lemma 5 and Lemma 6, it suffices to show that by running Algorithm 3 with probability $41-\delta/$ 4, $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ for any $1\leq t\leq T$. Formally, we have the lemma below.
Lemma 15.
Suppose Assumption 10 holds. By running Algorithm 3, with probability $41-\delta/$ 4, it holds that
$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (48)
for any $1\leq t\leq T$.
The proof of Lemma 15 is postponed to Appendix D
7 Extension: oracle-efficient multi-group learning
Given a hypothesis set $\mathcal{H}$ where each $h\in\mathcal{H}:\mathcal{X}\to\mathcal{Y}$, a set of groups $\mathcal{G}$ such that each $g\in\mathcal{G}$ is a subset of $\mathcal{X}$, and a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ and a distribution $\mathcal{D}$ with support $\mathcal{X}\times\mathcal{Y}$, define $L_{\mathcal{D}}(h|g):=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(h(x),y)|x\in g]$. Define $P_{g}=\mathrm{Pr}_{(x,y)\sim\mathcal{D}}[x\in g]$. Let $\gamma=\min_{g\in\mathcal{G}}P_{g}$. The goal of the learning algorithm is to find a (possibly randomized) hypothesis $h$ to minimize $L_{\mathcal{D}}(h|g)$ for all $g\in\mathcal{G}$ up to some threshold $\epsilon\in(0,1]$. That is,
$L_{\mathcal{D}}(h|g)\leq\min_{h^{\prime}\in\mathcal{H}}L_{\mathcal{D}}(h^{\prime}|g)+\epsilon,\forall g\in\mathcal{G}.$
We continue with the compatibility assumption as below, which ensures that such near-optimal hypothesis $h$ exists.
Assumption 16.
There exists $h^{*}\in\mathcal{H}$, such that
$L_{\mathcal{D}}(h^{*}|g)\leq\min_{h\in\mathcal{G}}L_{\mathcal{D}}(h|g)+\frac{\epsilon}{4},\quad\forall g\in\mathcal{G}.$
The learning algorithm
As presented in Algorithm 4, we first sample $N=O\left(\frac{\log(|\mathcal{G}|/\delta)+d\log(d/\epsilon)}{\gamma\epsilon^{2}}\right)$ datapoints from $\mathcal{D}$ and then estimate the optimal value $f_{g}$ for each group $g\in\mathcal{G}$. By taking $L_{\mathcal{D}}(h|g)-f_{g}$ as the loss function, we then invoke a Hedge algorithm over $\mathcal{G}$ to learn the problem
$\min_{h\in\Delta(\mathcal{H})}\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h|g)-f_{g}).$ (49)
By Assumption 16, it suffices to find some (possibly randomized) $h$ such that
$\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h|g)-f_{g}).\leq\min_{h^{\prime}\in\mathcal{H}}\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h^{\prime}|g)-f_{g})\leq\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h^{*}|g)-f_{g})=O(\epsilon).$
Formally, we have the following theorem.
Theorem 17.
Suppose Assumption 16 holds. By running Algorithm 4, with probability, the output $h^{\mathsf{final}}$ satisfies that
$L_{\mathcal{D}}(h^{\mathsf{final}}|g)=\frac{1}{T}\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)\leq\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)+\epsilon$ (50)
for any $g\in\mathcal{G}$. Meanwhile, the sample complexity of Algorithm 4 is bounded by
$O\left(\frac{d\log(d/\epsilon)+\log(|\mathcal{G}|/\delta)}{\gamma\epsilon^{2}}\right).$
Moreover, Algorithm 4 only access $\mathcal{H}$ with an $\mathrm{ERM}$ oracle.
Under Assumption 16, Theorem 17 recovers the result in Tosh and Hsu, (2022) without direct access to the hypothesis set $\mathcal{H}$. In high-level idea, the proof of Theorem 17 is based on the uniform convergence argument, which is a natural extension of our main results. We postpone the proof to Appendix E due to space limitation.
8 Necessity of randomization
Given that the best-known sample complexities prior to our work were derived for algorithms that either output randomized hypotheses or invoke majority votes, Awasthi et al., (2023) raised the question about how the sample complexity is impacted if only deterministic (or “proper”) hypotheses are permitted as the output of the learning algorithms. As it turns out, the restriction to deterministic hypotheses substantially worsens the sample efficiency, as revealed by the following theorem.
Theorem 18.
Assume that $d\geq 2\log(8k)$. Consider any $\varepsilon\in(0,1/100)$, and let $N_{0}=\frac{2^{d}-1}{k}$. One can find
  • a hypothesis class $\mathcal{H}$ containing at most $kN_{0}+1$ hypothesis,
  • a collection of $k$ distributions $\mathcal{D}\coloneqq\{\mathcal{D}_{i}\}_{i=1}^{k}$,
  • a loss function :×$\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$,
such that it takes at least $\frac{dk}{240000\varepsilon^{2}}$ samples to find $h\in\mathcal{H}$ obeying
$\max_{1\leq i\leq k}\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}\leq\min_{h^{\prime}\in\mathcal{H}}\max_{i\in[k]}\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h^{\prime},(x,y)\big{)}\big{]}+\varepsilon$ (51)
with probability exceeding $3/4$.
Let us breifly desribe the high-level strategy for our construction of the hard instance: for each $i\in[k]$, we build a hypothesis set $\mathcal{H}_{i}$ that performs poorly solely on $\mathcal{D}_{i}$. To discriminate the optimal hypothesis — denoted by $h^{\star}$ — from some $\mathcal{H}_{i}$, the learner has to call $\Omega(\log(|\mathcal{H}_{i}|)/\varepsilon^{2})=\widetilde{\Omega}(d/\varepsilon^{2})$ times to $\mathrm{Query}(\mathcal{D}_{i})$ in expectation. The result follows by taking sum over $i\in[k]$.
Proof of Theorem 18.
Note that $N_{0}=\frac{2^{d}-1}{k}$. Set $N=kN_{0}+1=2^{d}$. Set $\mathcal{X}=\{-1,0,1\}^{kN}$. We set $\mathcal{Y}=\{1\}$ to be a set with only one element. Without loss of generality we write $l(h,(x,y))=l(h,x)$.
We now introduce our construction. There are $N$ hypotheses in $\mathcal{H}$, where each hypothesis corresponds to $k$ dimensions of the ground set $\mathcal{X}$. Without loss of generality, for $h\in\mathcal{H}$, we let $\mathcal{I}_{h}=\{j_{h,i}\}_{i=1}^{k}$ be the $k$ dimensions related to $h$. Note that $\mathcal{I}_{h}\cap\mathcal{I}_{h^{\prime}}=\emptyset$ for $h\neq h^{\prime}$. We construct $\mathcal{H}$ as $\mathcal{H}=(\cup_{i=1}^{k}\mathcal{H}_{i})\cup\{h^{\star}\}$. Then we define $l(h,x)$ as $l(h,x)=\min_{i\in\mathcal{I}_{h},x_{i}\neq 0}x_{i}$. Now we design the $k$ distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$. Fix $i\in[k]$, we let $\mathrm{Pr}_{\mathcal{D}_{i}}[x]=\Pi_{\ell=1}^{kN}\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]$, where $\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]=\mathbb{I}[x_{\ell}=0]$ for $\ell\notin\{j_{h,i}|h\in\mathcal{H}\}$, $\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]=\frac{1}{2}\mathbb{I}[x_{\ell}=1]+\frac{1}{2}\mathbb{I}[x_{\ell}=-1]$ for $\ell\in\{j_{h,i}|h\notin\mathcal{H}_{i}\}$, and $\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]=\left(\frac{1}{2}+4\varepsilon\right)\mathbb{I}[x_{\ell}=1]+\left(\frac{1}{2}-4\varepsilon\right)\mathbb{I}[x_{\ell}=-1]$ for $\ell\in\{j_{h,i}|h\in\mathcal{H}_{i}\}$. In words, for each $i\in[k]$, we draw $x\sim\mathcal{D}_{i}$ by independently draw each dimension of $x$, and there are $|\mathcal{H}|=N$ non-zero dimensions. Furthermore, for $x\sim\mathcal{D}_{i}$ and $h\in\mathcal{H}$, we have $l(h,x)=x_{j_{h,i}}$. By this construction, we have the following properties:
  1. (i).
    $\ell(h,x)\in[-1,1]$ for any $h\in\mathcal{H}$ and $x\in\mathcal{X}$;
  2. (ii).
    $\mathbb{E}_{x\sim\mathcal{D}_{i}}[\ell(h,x)]=\mathbb{E}_{x\sim\mathcal{D}_{i,j_{h,i}}}[x_{j_{h,i}}]=2\varepsilon\cdot\mathbb{I}[h\in\mathcal{H}_{i}]$ for any $i\in[k]$ and $h\in\mathcal{H}$;
  3. (iii).
    the only $\varepsilon$-optimal hypothesis is $h^{\star}$ because for any $h\neq h^{\star}$, there exists some $i$ such that $h\in\mathcal{H}_{i}$;
  4. (iv).
    $|\mathcal{H}|=N=kN_{0}+1=2^{d}$, which implies $\mathrm{VC}(\mathcal{H})\leq\log_{2}(N)\leq d$.
For each call to $\mathrm{Query}(\mathcal{D}_{i})$, we can get independent observations $\{x_{j_{h,i}}\}_{h\in\mathcal{H}}$ where $x_{j_{h,i}}\sim D_{i,j_{h,i}}$ for each $h\in\mathcal{H}$.
Now we let the number of calls to $\mathrm{Query}(\mathcal{D}_{i})$ be $M_{i}$ for $i\in[k]$. Our target is to show to distinguish $h^{\star}$ from $\mathcal{H}_{i}$, $M_{i}$ has to be at least $\Omega(d/\varepsilon^{2})$.
Now suppose there is an algorithm $\mathcal{G}$ with numbers of samples $\{M_{i}\}_{i=1}^{k}$ such that the output is $h^{\star}$ with probability $\frac{3}{4}$. Let $\mathrm{Pr}_{\mathcal{G}}[\cdot]$ and $\mathbb{E}_{\mathcal{G}}[\cdot]$ denote respectively the probability and expectation under running the algorithm $\mathcal{G}$. Let $h_{\mathrm{out}}$ be the output hypothesis. It then follows that
$\mathrm{Pr}_{\mathcal{G}}[h_{\mathrm{out}}=h^{\star}]\geq\frac{3}{4}.$
Let $\Pi_{\mathcal{H}}$ be the set of permutations over $\mathcal{H}$. Let $\mathrm{U}(\Pi_{\mathcal{H}})$ be the uniform distribution over $\Pi(\mathcal{H})$. With a slight abuse of notations, for $x\in\{-1,0,1\}^{kN}$ and $\sigma\in\Pi_{\mathcal{H}}$, we define $\sigma(x)$ to be the vector $y$ such that $y_{j_{h,i}}=x_{j_{\sigma(h),i}}$ for all $h\in\mathcal{H}$ and $i\in[k]$.
Let $\mathcal{G}^{\prime}$ be the algorithm with $\mathcal{H}$ replaced by $\sigma(\mathcal{H})$ in the input where $\sigma\sim\mathrm{U}(\Pi_{\mathcal{H}})$. Because $\mathcal{G}$ returns the optimal hypothesis with probability at least $3/4$ for all problem instances, we have that
$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star}]\geq\frac{3}{4}.$
We note that $\mathcal{G}^{\prime}$ is a symmetric algorithm with respect to the hypothesis set. Formally, we have the lemma below to bound the probability of outputting a non-optimal hypothesis.
Lemma 19.
Fix $m\geq 0$ and $\widetilde{i}\in[k]$. Suppose $\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{i}\leq m]\geq\frac{1}{2}$. It then holds that for any $h\in\mathcal{H}_{\widetilde{i}}$
$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m]\geq\frac{1}{2}\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m]\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}).$
Moreover, it holds that $m\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}$.
By Lemma 19, we have that
$\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right]\leq\frac{1}{2}.$ (52)
Noting that $\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star}]\geq\frac{3}{4}$, we learn that
$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}]\geq\frac{1}{4},$ (53)
which implies that $\mathbb{E}_{\mathcal{G}^{\prime}}[M_{\widetilde{i}}]\geq\frac{\log(N_{0}/4)}{120000\varepsilon^{2}}\geq\frac{d-\log_{2}(8k)}{120000\varepsilon^{2}}\geq\frac{d}{240000\varepsilon^{2}}$. Taking sum over $i\in[k]$, we have that
$\mathbb{E}_{\mathcal{G}^{\prime}}\left[\sum_{i=1}^{k}M_{\widetilde{i}}\right]\geq\frac{dk}{240000\varepsilon^{2}}.$ (54)
The proof is completed.
9 Discussion
In this paper, we have settled the problem of achieving optimal sample complexity in multi-distribution learning, assuming the availability of adaptive (or on-demand) sampling. We have put forward a novel oracle-efficient algorithm that provably attains a sample complexity of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ for VC classes, which matches the best-known lower bound up to some logarithmic factor. From the technical perspective, the key novelty of our analysis lies in carefully bounding the trajectory of the Hedge algorithm on a convex (concave) optimization problem. We have further unveiled the necessesity of randomization, revealing that a considerably larger sample size is necessary if the learning algorithm is constrained to return deterministic hypotheses. Notably, our work manages to solve three open problems presented in COLT 2023 (namely, Awasthi et al., (2023, Problems 1, 3 and 4)).
Our work not only addresses existing challenges but also opens up several directions for future exploration. To begin with, while our sample complexity results are optimal up to logarithmic factors, further studies are needed in order to sharpen the logarithmic depdency. Additionally, the current paper assumes a flexible sampling protocol that allows the learner to take samples arbitrarily from any of the $k$ distributions; how will the sample complexity be impacted under additional constraints imposed on the sampling process? Furthermore, can we extend our current analysis (which bounds the dynamics of the Hedge algorithm) to control the trajectory of more general first-order/second-order algorithms, in the context of robust online learning? Another venue for exploration is the extension of our multi-distribution learning framework to tackle other related tasks like multi-calibration (Hébert-Johnson et al.,, 2018; Haghtalab et al.,, 2023). We believe that our algorithmic and analysis framework can shed light on making progress in all of these directions.
Acknowledgements
We thank Eric Zhao for answering numerous questions about the open problems. YC is supported in part by the Alfred P. Sloan Research Fellowship, the NSF grants CCF-1907661, DMS-2014279, IIS-2218713 and IIS-2218773. JDL acknowledges support of the ARO under MURI Award W911NF-11-1-0304, the Sloan Research Fellowship, NSF CCF 2002272, NSF IIS 2107304, NSF CIF 2212262, ONR Young Investigator Award, and NSF CAREER Award 2144994.
Appendix A Auxiliary lemmas
In this section, we introduce several technical lemmas that are used multiple times in our analysis.
We begin by introducing three handy concentrations inequalities. The first result is the well-renowned Freedman inequality (Freedman,, 1975), which assists in deriving variance-aware concentration inequalities for martingales.
Lemma 20 (Freedman’s inequality (Freedman,, 1975)).
Let $(M_{n})_{n\geq 0}$ be a martingale obeying $M_{0}=0$. Define $V_{n}\coloneqq\sum_{k=1}^{n}\mathbb{E}[(M_{k}-M_{k-1})^{2}\,|\,\mathcal{F}_{k-1}]$ for each $n\geq 0$, where $\mathcal{F}_{k}$ denotes the $\sigma$-algebra generated by $(M_{1},M_{2},\dots,M_{k})$. Suppose that $M_{k}-M_{k-1}\leq 1$ for all $k\geq 1$. Then for any $x>0$ and $y>0$, one has
$\mathbb{P}\big{(}M_{n}\geq nx,V_{n}\leq ny\big{)}\leq\exp\left(-\frac{nx^{2}}{2(y+\frac{1}{3}x)}\right).$ (55)
The second concentration result bounds the difference between the sum of a sequence of random variables and the sum of their respective conditional means (w.r.t. the associated $\sigma$-algebra).
Lemma 21 (Lemma 10 in Zhang et al., (2022)).
Let $X_{1},X_{2},\ldots$ be a sequence of random variables taking value in the interval $[0,l]$. For any $k\geq 1$, let $\mathcal{F}_{k}$ be the $\sigma$-algebra generated by $(X_{1},X_{2},\ldots,X_{k})$, and define $Y_{k}:=\mathbb{E}[X_{k}\mid\mathcal{F}_{k-1}]$. Then for any $\delta>0$, we have
$\mathbb{P}\left\{\exists n\in\mathbb{N},\sum_{k=1}^{n}X_{k}\geq 3\sum_{k=1}^{n}Y_{k}+l\log\frac{1}{\delta}\right\}\leq\delta,$
$\mathbb{P}\left\{\exists n\in\mathbb{N},\sum_{k=1}^{n}Y_{k}\geq 3\sum_{k=1}^{n}X_{k}+l\log\frac{1}{\delta}\right\}\leq\delta.$
The third concentration result is the Mcdiarmid inequality, a celebrated inequality widely used to control the flucutaion of multivariate functions when the input variables are independently generated.
Lemma 22 (Mcdiarmid’s inequality).
Let $X_{1},X_{2},\ldots,X_{n}$ be a sequence of independent random variables, with $X_{i}$ supported on $\mathcal{X}_{i}$. Let $f:\mathcal{X}_{1}\times\mathcal{X}_{2}\times\dots\times\mathcal{X}_{n}\to\mathbb{R}$ be a function such that: for any $i\in[n]$ and any $\{x_{1},\ldots,x_{n}\}\in\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$,
$\sup_{x_{i}^{\prime}\in\mathcal{X}_{i}}\big{|}f(x_{1},\cdots,x_{i},\cdots,x_{n})-f(x_{1},\cdots,x^{\prime}_{i},\cdots,x_{n})\big{|}\leq c$
holds for some quantity $c>0$. It then holds that
$\mathbb{P}\Big{\{}\big{|}f(X_{1},X_{2},\cdots,X_{n})-\mathbb{E}\big{[}f(X_{1},X_{2},\cdots,X_{n})\big{]}\big{|}\geq\varepsilon\Big{\}}\leq 2\exp\left(-\frac{2\varepsilon^{2}}{nc^{2}}\right).$
Additionally, the following lemma presents a sort of the data processing inequality w.r.t. the Kullback-Leibler (KL) divergence, which is a classical result from information theory.
Lemma 23.
Let $\mathcal{X}$ and $\mathcal{Y}$ be two sets, and consider any function $f:\mathcal{X}\to\mathcal{Y}$. For any two random variables $X_{1}$ and $X_{2}$ supported on $\mathcal{X}$, it holds that
$\mathsf{KL}\big{(}\mu(X_{1})\,\|\,\mu(X_{2})\big{)}\geq\mathsf{KL}\big{(}\mu\big{(}f(X_{1})\big{)}\,\|\,\mu\big{(}f(X_{2})\big{)}\big{)},$ (56)
where we use $\mu(Z)$ to denote the distribution of a random variable $Z$.
Lastly, let us make note of an elementary bound regarding the KL divergence between two Bernoulli distributions.
Lemma 24.
Consider any $q>0$ and $x\in[0,\log(2)]$. Also, consider any $y,y^{\prime}\in(0,1)$ obeying $y\geq q$ and $y^{\prime}\geq\exp(x)y$. It then holds that
$\mathsf{KL}\left(\mathsf{Ber}(y)\,\|\,\mathsf{Ber}(y^{\prime})\right)\geq\frac{qx^{2}}{4},$
where $\mathsf{Ber}(z)$ denotes the Bernoulli distribution with mean $z$.
Proof.
To begin with, the function defined below satisfies
$f(a,b)\coloneqq\mathsf{KL}\big{(}\mathsf{Ber}(a)\,\|\,\mathsf{Ber}(b)\big{)}$ $=a\log\left(\frac{a}{b}\right)+(1-a)\log\left(\frac{1-a}{1-b}\right).$
For any $0<a\leq b\leq 1$, it is readily seen that
$\frac{\partial f(a,b)}{\partial b}=-\frac{a}{b}+\frac{1-a}{1-b}=\frac{b-a}{b(1-b)}\geq 0.$
It follows from our assumptions $y\geq q$ and $y^{\prime}\geq\exp(x)y$ that
$\mathsf{KL}\left(\mathsf{Ber}(y)\,\|\,\mathsf{Ber}(y^{\prime})\right)$ $=f(y,y^{\prime})=f(y,y)+\int_{y}^{y^{\prime}}\frac{\partial f(y,z)}{\partial z}\mathrm{d}z=\int_{y}^{y^{\prime}}\frac{z-y}{z(1-z)}\mathrm{d}z$
$\geq\frac{1}{y^{\prime}}\int_{y}^{y^{\prime}}(z-y)\mathrm{d}z\geq\frac{(y^{\prime}-y)^{2}}{2y^{\prime}}$
$\geq\frac{(y^{\prime}-y)(1-\exp(-x))}{2}$
$\geq\frac{y(\exp(x)-1)^{2}}{4}\geq\frac{qx^{2}}{4},$
where the penultimate inequality uses $x\in[0,\log(2)]$, and the last inequality holds since $y\geq q$. ∎
Appendix B Missing proofs for optimality
B.1 Proof of Lemma 4
For ease of presentation, suppose there exists a dataset $\widetilde{\mathcal{S}}$ containing $T_{1}$ independent samples drawn from each distribution $\mathcal{D}_{i}$ ($1\leq i\leq k$), so that in total it contains $kT_{1}$ samples. We find it helpful to introduce the following notation.
  • For each $i\in[k]$ and $j\in[n_{i}]$, denote by $(x_{i,j},y_{i,j})$ the $j$-th sample in $\widetilde{\mathcal{S}}$ that is drawn from $\mathcal{D}_{i}$.
  • For each set of integers $n=\{n_{i}\}_{i=1}^{k}\in\mathbb{N}^{k}$, we define $\widetilde{\mathcal{S}}(n)$ to be the dataset containing $\big{\{}(x_{i,j},y_{i,j})\big{\}}_{1\leq j\leq n_{i}}$ for all $i\in[k]$; namely, it comprises, for each $i\in[k]$, the first $n_{i}$ samples in $\widetilde{\mathcal{S}}$ that are drawn from $\mathcal{D}_{i}$.
  • We shall also let $\widetilde{\mathcal{S}}^{+}(n)=\big{\{}\big{\{}\big{(}x_{i,j}^{+},y_{i,j}^{+}\big{)}\big{\}}_{j=1}^{n_{i}}\big{\}}_{i=1}^{k}$ be an independent copy of $\widetilde{\mathcal{S}}(n)$, where for each $i\in[k]$, $\big{\{}\big{(}x_{i,j}^{+},y_{i,j}^{+}\big{)}\big{\}}$ are independent samples drawn from $\mathcal{D}_{i}$.
Equipped with the above notation, we are ready to present our proof.
Step 1: concentration bounds for any fixed $n=\{n_{i}\}_{i=1}^{k}$ and $w\in\Delta(k)$.
Consider any fixed $n=\{n_{i}\}_{i=1}^{k}$ obeying $0\leq n_{i}\leq T_{1}$ for all $i\in[k]$, and any fixed $w\in\Delta(k)$. For any quantity $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$, if we take
$E(\lambda)\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right\}\right)\right]$ (57)
with the expectation taken over the randomness of $\widetilde{\mathcal{S}}(n)$, then we can perform a standard “symmetrization” trick to bound $E(\lambda)$ as follows:
$E(\lambda)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right\}\right)\right]$
$=\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}^{+}(n)}\left[\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\right]\right\}\right)\right]$
$\leq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}^{+}(n)}\left[\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\right]\right]$
$\leq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n),\widetilde{\mathcal{S}}^{+}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\right],$ (58)
where the last two inequalities follow from Jensen’s inequality.
Next, let $\sigma(n)\coloneqq\big{\{}\{\sigma_{i,j}\}_{j=1}^{n_{i}}\big{\}}_{i=1}^{k}$ be a collection of i.i.d. Rademacher random variables obeying $\mathbb{P}(\sigma_{i,j}=1)=\mathbb{P}(\sigma_{i,j}=-1)=1/2$. Denoting $\mathcal{C}=\big{\{}(x_{i,j},y_{i,j})\big{\}}\bigcup\big{\{}(x_{i,j}^{+},y_{i,j}^{+})\big{\}}$, we obtain
$\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n),\widetilde{\mathcal{S}}^{+}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\right]$
$=\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n),\widetilde{\mathcal{S}}^{+}(n)}\left[\mathop{\mathbb{E}}\limits_{\sigma(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i,j}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\,\Big{|}\,\mathcal{C}\right]\right].$ (59)
Note that for any dataset $\mathcal{C}$ with cardinality $|\mathcal{C}|$, the Sauer–Shelah lemma (Wainwright,, 2019, Proposition 4.18) together with our assumption that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ tells us that the cardinality of the following set obeys
$\big{|}\mathcal{H}(\mathcal{C})\big{|}\leq(|\mathcal{C}|+1)^{d}\leq(2kT_{1}+1)^{d},$ (60)
where $\mathcal{H}(\mathcal{C})$ denotes the set obtained by applying all $h\in\mathcal{H}$ to the data points in $\mathcal{C}$, namely,
$\mathcal{H}(\mathcal{C})\coloneqq\Big{\{}\big{(}h(x_{1,1}),h(x_{1,1}^{+}),h(x_{1,2}),h(x_{1,2}^{+}),\cdots\big{)}\mid h\in\mathcal{H}\Big{\}}.$ (61)
We shall also define $\mathcal{H}_{\min,\mathcal{C}}\subseteq\mathcal{H}$ to be the minimum-cardinality subset of $\mathcal{H}$ that results in the same outcome as $\mathcal{H}$ when applied to $\mathcal{C}$, namely,
$\mathcal{H}_{\min,\mathcal{C}}(\mathcal{C})=\mathcal{H}(\mathcal{C})\qquad\text{and}\qquad\big{|}\mathcal{H}_{\min,\mathcal{C}}\big{|}=\big{|}\mathcal{H}(\mathcal{C})\big{|}.$
With these in place, we can demonstrate that
$\mathop{\mathbb{E}}\limits_{\sigma(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i,j}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\,\Big{|}\,\mathcal{C}\right]$
$=\mathop{\mathbb{E}}\limits_{\sigma(n)}\left[\max_{h\in\mathcal{H}_{\min,\mathcal{C}}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i,j}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\,\Big{|}\,\mathcal{C}\right]$
$\leq\mathop{\mathbb{E}}\limits_{\sigma(n)}\left[\sum_{h\in\mathcal{H}_{\min,\mathcal{C}}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i,j}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\,\Big{|}\,\mathcal{C}\right]$
$\leq\big{|}\mathcal{H}_{\min,\mathcal{C}}\big{|}\max_{h\in\mathcal{H}_{\min,\mathcal{C}}}\mathop{\mathbb{E}}\limits_{\sigma(n)}\left[\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i,j}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\,\Big{|}\,\mathcal{C}\right]$
$\leq\big{(}2kT_{1}+1\big{)}^{d}\max_{h\in\mathcal{H}}\prod_{i=1}^{k}\prod_{j=1}^{n_{i}}\mathop{\mathbb{E}}\limits_{\sigma_{i,j}}\left[\exp\left(\lambda\left\{\frac{w_{i}}{n_{i}}\sigma_{i,j}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\,\Big{|}\,\mathcal{C}\right]$
$\leq\big{(}2kT_{1}+1\big{)}^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right).$ (62)
Here, the last inequality makes use of fact $\big{|}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}|\leq 2$ as well as the following elementary inequality
$\mathop{\mathbb{E}}\limits_{\sigma_{i,j}}\big{[}\exp(\sigma_{i,j}x)\big{]}=\frac{1}{2}\big{(}\exp(x)+\exp(-x)\big{)}\leq\exp\big{(}0.5x^{2}\big{)}.$
Taking (58) and (62) together and applying the Markov inequality reveal that, for any $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$,
$E(\lambda)\leq(2\|n\|_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right).$ (63)
Repeating the same arguments also yields
$\overline{E}(\lambda)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}w_{i}L(h,e_{i})-\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\right\}\right)\right]$
$\leq(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right)$
for any $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$. Taking the above two inequalities reveals that, for any $0<\varepsilon^{\prime}\leq 1$,
$\mathbb{P}\left(\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}\right)$
$\qquad\leq\min_{0\leq\lambda\leq\min_{i}\frac{n_{i}}{w_{i}}}\frac{E(\lambda)+\overline{E}(\lambda)}{\exp(\lambda\varepsilon^{\prime})}$
$\qquad\leq\min_{0\leq\lambda\leq\min_{i}\frac{n_{i}}{w_{i}}}2\cdot(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}-\lambda\varepsilon^{\prime}\right).$ (64)
Step 2: uniform concentration bounds over all $n$ and $w$.
Next, we move on to extend the above result to uniform concentration bounds. Towards this, let us first introduce a couple of notation.
  • Let us use $\Delta_{\varepsilon_{2}}(k)\subseteq\Delta(k)$ to denote an $\varepsilon_{2}$-net of $\Delta(k)$ — namely, for any $x\in\Delta(k)$, there exists a vector $x_{0}\in\Delta_{\varepsilon_{2}}(k)$ obeying $\|x-x_{0}\|_{\infty}\leq\varepsilon_{2}$. We shall choose $\Delta_{\varepsilon_{2}}(k)$ properly so that
    $|\Delta_{\varepsilon_{2}}(k)|\leq(1/\varepsilon_{2})^{k}.$
  • Define the following set
    =$\mathcal{B}=\Big{\{}n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\mid T_{1}w_{i}\leq 2n_{i},0\leq n_{i}\leq T_{1},\forall i\in[k],w\in\Delta_{\varepsilon_{1}/(8k)}(k)\Big{\}}.$
Applying the union bound yields
$\mathbb{P}\left(\exists(n,w)\in\mathcal{B},\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}\right)$
$\leq\sum_{(n,w)\in\mathcal{B}}\min_{0\leq\lambda\leq\min_{i}\frac{n_{i}}{w_{i}}}2\cdot(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}-\lambda\varepsilon^{\prime}\right)$
$\leq\sum_{(n,w)\in\mathcal{B}}\min_{0\leq\lambda\leq\frac{T_{1}}{2}}2\cdot(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\frac{2}{T_{1}}-\lambda\varepsilon^{\prime}\right)$
$\leq\sum_{(n,w)\in\mathcal{B}}2\cdot(2kT_{1}+1)^{d}\exp\left(-\frac{T_{1}(\varepsilon^{\prime})^{2}}{16}\right)$
$\leq|\mathcal{B}|\cdot 2\cdot(2kT_{1}+1)^{d}\exp\left(\frac{-T_{1}(\varepsilon^{\prime})^{2}}{16}\right)$
$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(\frac{-T_{1}(\varepsilon^{\prime})^{2}}{16}\right),$
where the second inequality holds since $\sum_{i=1}^{k}\frac{w_{i}^{2}}{n_{i}}\leq\frac{2}{T_{1}}\sum_{i=1}^{k}w_{i}=\frac{2}{T_{1}}$ (according to the definition of $\mathcal{B}$).
Let $\mathcal{S}^{t}$ denote the value of $\mathcal{S}$ after line 1 of Algorithm 1 in the $t$-th round. Recall that $n^{t}=[n^{t}_{i}]_{1\leq i\leq k}$ denotes the number of samples in $\mathcal{S}^{t}$. Consequently, we can take $\mathcal{S}^{t}=\widetilde{\mathcal{S}}(n^{t})$. In view of the update rule in Algorithm 1, one can always find $(n^{t},\widetilde{w}^{t})\in\mathcal{B}$ satisfying $\|\widetilde{w}^{t}-w^{t}\|_{1}\leq k\|\widetilde{w}^{t}-w^{t}\|_{\infty}\leq\varepsilon_{1}/8$. As a result, for any $0<\varepsilon^{\prime}\leq 1$, we can deduce that
$\mathbb{P}\left(\exists t\in[T],\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w^{t}_{i}\frac{1}{n^{t}_{i}}\sum_{i=1}^{n^{t}_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w^{t}_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}+\frac{\varepsilon_{1}}{4}\right)$
$\qquad\leq\mathbb{P}\left(\exists t\in[T],\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\widetilde{w}^{t}_{i}\frac{1}{n^{t}_{i}}\sum_{i=1}^{n^{t}_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}\widetilde{w}^{t}_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}\right)$
$\qquad\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(\frac{-T_{1}(\varepsilon^{\prime})^{2}}{16}\right),$ (65)
where the second inequality arises from the fact that $\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\in[-1,1]$ and $L(h,e_{i})\in[-1,1]$. Substituting $\varepsilon^{\prime}=\frac{\varepsilon_{1}}{4}$ and $T_{1}=\frac{400\left(k\log(k/\varepsilon_{1})+d\log(kd/\varepsilon_{1})+\log(1/\delta)\right)}{\varepsilon_{1}^{2}}$ into (65), we can obtain
$\mathbb{P}\left(\exists t\in[T],\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w^{t}_{i}\cdot\frac{1}{n^{t}_{i}}\sum_{i=1}^{n^{t}_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w^{t}_{i}L(h,e_{i})\right|\geq\frac{\varepsilon_{1}}{2}\right)$
$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(\frac{-T_{1}\varepsilon_{1}^{2}}{16}\right)$
$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\Big{(}-10\big{(}k\log(k/\varepsilon_{1})+d\log(kd/\varepsilon_{1})+\log(1/\delta)\big{)}\Big{)}$
$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot(k/\varepsilon_{1})^{-10k}\cdot(kd/\varepsilon_{1})^{-10d}\cdot\delta$
$\leq\delta/4.$ (66)
Step 3: putting all this together.
Recalling that
$\widehat{L}^{t}(h,w^{t})=\sum_{i=1}^{k}w_{i}^{t}\cdot\frac{1}{n_{i}^{t}}\sum_{i=1}^{n_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)},$
one can see from (66) that, with probability exceeding $41-\delta/$ 4,
$\left|\widehat{L}^{t}(h,w^{t})-L(h,w^{t})\right|\leq\frac{\varepsilon_{1}}{2}$ (67)
holds simultaneously for all $t\in[T]$ and all $h\in\mathcal{H}$. Additionally, observing that
$h^{t}=\arg\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t}),$ (68)
we can immediately deduce that
$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}.$ (69)
This concludes the proof of Lemma 4.
Appendix C Missing proofs for sample complexity
We let $\delta^{\prime}=\delta/(32T^{4}k^{2})$ throughout this section.
C.1 Proof of Lemma 8
Define $\mathcal{W}_{j}=\{i\in[k]|\max_{1\leq t\leq T}w_{i}^{t}\in(2^{-j},2^{-(j-1)}]\}$ for $11\leq j\leq\left\lfloor\log_{2}(k)\right\rfloor+$ 1 and $\overline{\mathcal{W}}=[K]/\cup_{j}\mathcal{W}_{j}$. We consider to bound the size of each $\mathcal{W}_{j}$.
Lemma 25.
Assume the conditions in Lemma 8 hold. For any $21\leq j\leq\log_{2}(k)-$ 2, with probability $1-8T^{4}k\delta^{\prime}$,
$|\mathcal{W}_{j}|\leq 8\cdot 10^{7}\cdot\left((\log_{2}(k)+1)^{4}\left(\log(k)+\log(1/\delta^{\prime})\right)^{2}(\log_{2}(T)+1)\right)\cdot 2^{j}.$
Now we prove Lemma 8. With probability $41-8T^{4}k^{2}\delta^{\prime}=1-\delta/$ 4, we have
$\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ $\leq k\cdot\frac{16}{k}+\sum_{j=1}^{\left\lfloor\log_{2}(k)]\right\rfloor-2}|\mathcal{W}_{j}|2^{-j+1}$
$\leq 2\cdot 10^{8}\cdot\left((\log_{2}(k)+1)^{5}\left(\log(Tk)+\log(1/\delta)\right)^{2}(\log_{2}(T)+1)\right).$
The proof of Lemma 8 is finished.
C.2 Proof of Lemma 25
Fix $j\in[1,\left\lfloor\log_{2}(k)\right\rfloor]-2$. Before step into the proof, we first introduce the high-level idea. In Lemma 29, we provide a lower bound of the length of a $(p,q,x)$ segment. We then want to prove that if $|\mathcal{W}_{j}|$ is large, there will be many disjoint segments, which forces the total length $T$ to be large enough to contain these segments.
As discussed in Section 4.2, we will extract a regular configuration from a general configuration. In high-level intuition, we have the following steps.
  1. 1.
    Prove the existence of segments for all $i\in\mathcal{W}_{j}$ (See Lemma 26);
  2. 2.
    Find some disjoint blocks such that (See Lemma 27 and Figure 3 )
    • The segments in the same blocks have a common inner point;
    • The number of segments in these blocks is at least $\frac{1}{3(\log_{2}(T)+1)}$ of the number of all the segments, i.e., $\frac{1}{3(\log_{2}(T)+1)}|\mathcal{W}_{j}|$ ;
    • Continue the analysis on a single block on benefit of the fact that these blocks are disjoint;
  3. 3.
    For each single block, use the common inner point to align either the starting points or the ending points of at least half of these segments (See Lemma 28 and Figure 4 );
  4. 4.
    Design a group of regular configurations (at most $O(\log_{2}(k))$ groups) so that at least one of the regular configurations contains enough segments with significant variation (See Lemma 28 and Figure 6 ).
C.2.1 The existence of segments
For $i\in\mathcal{W}_{j}$, by noting that $w^{1}_{i}=\frac{1}{k}\leq\frac{1}{2^{j+1}}$ and the fact that the trajectory $\{w^{t}_{i}\}$ changes slowly in $t$, we have the lemma below to show that for each $i\in\mathcal{W}_{j}$, there is an $(\frac{1}{2^{j+1}},\frac{1}{2^{j+2}},\log(2))$-segment $(s_{i},e_{i})$ for $i$.
Lemma 26.
For each $i\in\mathcal{W}^{j}$, there exists $1\leq s_{i}<e_{i}\leq T$, such that $\frac{1}{2^{j+2}}<w^{s_{i}}_{i}\leq\frac{1}{2^{j+1}}$, $\frac{1}{2^{j}}<w^{e_{i}}_{i}$ and $w^{t}_{i}>2^{-(j+2)}$ for any $s_{i}\leq t\leq e_{i}$. In words, there exists a $\left(\frac{1}{2^{j+1}},\frac{1}{2^{j+2}},\log(2)\right)$-segment $(s_{i},e_{i})$ with the index set as $\{i\}$.
Proof.
By definition, it is easy to find $e_{i}$. To find $s_{i}$, define $\tau=\max\{t\leq e_{i}|w^{t}_{i}\leq 2^{-(j+2)}\}$. $\tau$ is properly defined because $w^{1}_{i}=\frac{1}{k}\leq 2^{-(j+2)}$. By definition, we have that for any $\tau+1\leq t\leq e_{i}$, $w^{t}_{i}>2^{-(j+2)}$. Noting that $\log_{2}(w^{t+1}_{i}/w^{t}_{i})\leq\frac{\eta-\log_{2}(Z_{t+1}/Z_{t})}{\log(2)}\leq\frac{2\eta}{\log(2)}\leq\frac{1}{4}$, we have that $w^{\tau+1}_{i}\leq\frac{1}{2^{j+2}}\cdot 2^{\frac{1}{4}}\leq\frac{1}{2^{j+1}}$. So it suffices to choose $s_{i}=\tau+1$. ∎
C.2.2 Extracting regular segments from irregular segments
Lemma 27.
Recall the definition of $\mathcal{W}_{j}$ and $(s_{i},e_{i})$ for $i\in\mathcal{W}_{j}$ in Lemma 26. Then there exists a group of subsets $\{\mathcal{W}_{j}^{p}\}_{p=1}^{P}$ of $\mathcal{W}_{j}$ such that
  1. (i).
    $\mathcal{W}_{j}^{p}\subset\mathcal{W}_{j}$, $\mathcal{W}_{j}^{p}\cap\mathcal{W}_{j}^{p^{\prime}}=\emptyset$, $\forall p\neq p^{\prime}$;
  2. (ii).
    $\sum_{p=1}^{P}|\mathcal{W}_{j}^{p}|\geq\frac{|\mathcal{W}_{j}|}{3(\log_{2}(T)+1)}$;
  3. (iii).
    There exists $1\leq\tilde{s}_{1}<\tilde{e}_{1}\leq\tilde{s}_{2}<\tilde{e}_{2}\leq\dots\leq\tilde{s}_{P}<\tilde{e}_{P}\leq T$, such that for each $1\leq p\leq P$, $\tilde{s}_{p}=\min_{i\in\mathcal{W}_{j}^{p}}s_{i},\tilde{e}_{p}=\max_{i\in\mathcal{W}_{j}^{p}}e_{i}$ and $\max_{i\in\mathcal{W}_{j}^{p}}s_{i}\leq\min_{i\in\mathcal{W}_{j}^{p}}e_{i}$.
Proof.
For $11\leq x\leq\log_{2}(T)+$ 1, we define $\mathcal{W}_{j}(x):=\{i|2^{x-1}\leq e_{i}-s_{i}\leq 2^{x}\}$ and let $x^{*}=\arg\max_{1\leq x\leq\log_{2}(T)+1}|\mathcal{W}_{j}(x^{*})|$. Then we have that $|\mathcal{W}_{j}(x^{*})|\geq\frac{|\mathcal{W}_{j}|}{\log_{2}(T)+1}$. Then we consider to choose the subsets $\{\mathcal{W}_{j}^{m}\}_{m=1}^{M}$ from $\mathcal{W}_{j}(x^{*})$.
Let $\kappa_{1}=\min_{i\in\mathcal{W}_{j}(x^{*})}e_{i}$. We also define $\mathcal{U}_{j}^{1}:=\{i\in\mathcal{W}_{j}(x^{*})|s_{i}\leq\kappa_{1}\}$. For $o\geq 1$, we let
$\kappa_{o+1}=\min_{i\in\mathcal{W}_{j}(x^{*})/\cup_{o^{\prime}=1}^{o}\mathcal{U}_{j}^{o^{\prime}}}e_{i}$
$\mathcal{U}_{j}^{o+1}:=\{i\in\mathcal{W}_{j}(x^{*})/\cup_{o^{\prime}=1}^{o}\mathcal{U}_{j}^{o^{\prime}}|s_{i}\leq\kappa_{o+1}\}.$
And we stop until $\cup_{o\geq 1}\mathcal{U}_{j}^{o}=\mathcal{W}_{j}(x^{*})$.
Clearly, for each $o$ and $i_{1},i_{2}\in\mathcal{U}_{j}^{o}$, we have that $e_{i_{1}}\geq\kappa_{o}\geq s_{i_{2}}$. That is, $\max_{i\in\mathcal{U}_{j}^{o}}s_{i}\leq\min_{i\in\mathcal{U}_{j}^{o}}e_{i}$. We then consider the three groups of subsets of $\mathcal{W}_{j}(x^{*})$: $\{\mathcal{U}_{j}^{3o-2}\}_{o\geq 1}$, $\{\mathcal{U}_{j}^{3o-1}\}_{o\geq 1}$ and $\{\mathcal{U}_{j}^{3o}\}_{o\geq 1}$. Then there exists $i\in\{0,1,2\}$ such that $\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-i}|\geq\frac{1}{3}\sum_{o\geq 1}\mathcal{U}_{j}^{o}$. Without loss of generality, we assume that $\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-2}|\geq\frac{1}{3}\sum_{o\geq 1}\mathcal{U}_{j}^{o}=|\mathcal{W}_{j}(x^{*})|$.
Now we verify that $\{\mathcal{U}_{j}^{3o-2}\}_{o\geq 1}$ is the desired group of sets. The first condition holds because of the definition of $\{\mathcal{U}_{j}^{o}\}_{o\geq 1}$. As for the second condition, we have that
$\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-2}|\geq\frac{1}{3}\sum_{o\geq 1}\mathcal{U}_{j}^{o}=\frac{1}{3}|\mathcal{W}_{j}(x^{*})|\geq\frac{|\mathcal{W}_{j}|}{3(\log_{2}(T)+1)}.$
As for the last condition, it suffices to verify that $\max_{i\in\mathcal{U}_{j}^{3o-2}}e_{i}\leq\min_{i\in\mathcal{U}_{j}^{3o+1}}s_{i}$ for any proper $o$. By definition, for each proper $o\geq 1$, there exists $i\in\mathcal{W}_{j}(x^{*})$ such that $s_{i}\geq\kappa_{o}$ and $\kappa_{o+1}=e_{i}$. Then we have that $\kappa_{o+1}=e_{i}\geq s_{i}+2^{x^{*}-1}\geq\kappa_{o}+2^{x^{*}-1}$. It then follows that, for any $i\in\mathcal{U}_{j}^{3o+1}$, $s_{i}\geq\kappa_{3o}\geq\kappa_{3o-2}+2^{x^{*}}$. On the other hand, for any $i\in\mathcal{U}_{j}^{3o-2}$, we have that $e_{i}\leq s_{i}+2^{x^{*}}\leq\kappa_{3o-2}+2^{x^{*}}$. Putting all together, we have that
$\max_{i\in\mathcal{U}_{j}^{3o-2}}e_{i}\leq\kappa_{3o-2}+2^{x^{*}}\leq\min_{i\in\mathcal{U}_{j}^{3o+1}}s_{i}.$
The proof is finished.
Lemma 28.
Recall the definition of $\mathcal{W}_{j}$ and $(s_{i},e_{i})$ for $i\in\mathcal{W}_{j}$. Then there exists a group of subsets $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ such that the conditions below hold
  1. (i).
    $\mathcal{V}_{j}^{n}\subset\mathcal{W}_{j}$, $\mathcal{V}_{j}^{n}\cap\mathcal{V}_{j}^{n^{\prime}}=\emptyset$, $\forall n\neq n^{\prime}$;
  2. (ii).
    $\sum_{n=1}^{N}|\mathcal{V}_{j}^{n}|\geq\frac{|\mathcal{W}_{j}|}{24\log_{2}(k)(\log_{2}(T)+1)}$;
  3. (iii).
    There exists $1\leq\hat{s}_{1}<\hat{e}_{1}\leq\hat{s}_{2}<\hat{e}_{2}\leq\dots\leq\hat{s}_{N}<\hat{e}_{N}\leq T$, and $\{g_{n}\}_{n=1}^{N}\in\left[1,\infty\right)^{N}$ such that for each $1\leq n\leq N$, $(\hat{s}_{n},\hat{e}_{n})$ is a $\left(2^{-(j+1)}g_{n}|\mathcal{V}_{j}^{n}|,2^{-(j+2)}|\mathcal{V}_{j}^{n}|,\frac{\log(2)}{2\log_{2}(k)}\right)$-segment with index set as $\mathcal{V}_{j}^{n}$. That is, the following hold for each $1\leq n\leq N$:
    • a.
      $\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j+2}}\leq\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{\hat{s}_{n}}\leq\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j+1}}$ ;
    • b.
      $\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j}}\cdot\exp\left(\frac{\log(2)}{2\log_{2}(k)}\right)\leq\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{\hat{e}_{n}}$;
    • c.
      $\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{t}\geq\frac{|\mathcal{V}_{j}^{n}|}{2^{j+2}}$ for any $\hat{s}_{n}\leq t\leq\hat{e}_{n}$.
Proof.
Let $\{\mathcal{W}_{j}^{p}\}_{p=1}^{P}$ and $\{(\tilde{s}_{p},\tilde{e}_{p})\}_{p=1}^{P}$ be the construction in Lemma 27. Let $x=\log(2)$ be the parameter for the segment.
Fix $1\leq p\leq P$. Let $t_{\mathrm{mid}}^{p}=\min_{i\in\mathcal{W}_{j}^{p}}e_{i}$. For each $i\in\mathcal{W}_{j}^{p}$, we have
$\max\left\{\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right),\log\left(\frac{w^{t_{\mathrm{mid}}^{p}}_{i}}{w^{s_{i}}_{i}}\right)\right\}\geq\frac{1}{2}\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)+\frac{1}{2}\log\left(\frac{w^{t_{\mathrm{mid}}^{p}}_{i}}{w^{s_{i}}_{i}}\right)\geq\frac{x}{2}.$
Then we have
$\sum_{i\in\mathcal{W}_{j}^{p}}\left(\mathbb{I}\left[\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)\geq\frac{x}{2}\right]+\mathbb{I}\left[\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)\geq\frac{x}{2}\right]\right)\geq|\mathcal{W}_{j}^{p}|.$
Without loss of generality, we assume that
$\sum_{i\in\mathcal{W}_{j}^{p}}\left(\mathbb{I}\left[\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)\geq\frac{x}{2}\right]\right)\geq\frac{|\mathcal{W}_{j}^{p}|}{2},$ (70)
Let $\widehat{\mathcal{W}}_{j}^{p}:=\{i\in\mathcal{W}_{j}^{p}|\log(w^{e_{i}}_{i}/w^{t_{\mathrm{mid}}^{p}}_{i})\geq\frac{x}{2}\}$. Then $|\widehat{\mathcal{W}}_{j}^{p}|\geq\frac{|\mathcal{W}_{j}^{p}|}{2}$. Let $Q$ denote $|\widehat{\mathcal{W}}_{j}^{p}|$. Without loss of generality, we assume $\widehat{\mathcal{W}}_{j}^{p}=\{1,2,\ldots,Q\}$ and $e_{1}\leq e_{2}\leq\ldots\leq e_{Q}$. Let ~$\tilde{\ell}=\max\{\ell\geq 0|2^{\ell}\leq Q\}$ and $\tilde{Q}=2^{\tilde{\ell}}$.  [^6] We assume $\tilde{Q}\geq 2$ without loss of generality. In the case $\tilde{Q}=1$, we simply choose arbitrary element in $\widehat{\mathcal{W}}_{j}^{p}$ as a single subset. In this way, we can collect at least $\frac{1}{4}|\widehat{\mathcal{W}_{j}^{p}}|$ segments.
Define $e_{0}=t_{\mathrm{mid}}^{p}$. We use $[e_{0},e_{k}]\oplus a$ as a shorthand of $[e_{0+a},e_{k+a}]$. Then we define a group of disjoint intervals of $[T]$ as below.
$\mathcal{K}_{1}=\left\{[e_{0},e_{2^{\tilde{\ell}-1}}]\right\};$
$\mathcal{K}_{2}=\left\{[e_{0},e_{2^{\tilde{\ell}-2}}],[e_{0},e_{2^{\tilde{\ell}-2}}]\oplus 2^{\tilde{\ell}-1}\right\};$
$\mathcal{K}_{3}=\left\{[e_{0},e_{2^{\tilde{\ell}-3}}],[e_{0},e_{2^{\tilde{\ell}-3}}]\oplus 2^{\tilde{\ell}-2},[e_{0},e_{2^{\tilde{\ell}-3}}]\oplus 2\cdot 2^{\tilde{\ell}-2},[e_{0},e_{2^{\tilde{\ell}-3}}]\oplus 3\cdot 2^{\tilde{\ell}-2}\right\};$
$\dots$
$\mathcal{K}_{\ell}=\left\{[e_{0},e_{2^{\tilde{\ell}-\ell}}],[e_{0},e_{2^{\tilde{\ell}-\ell}}]\oplus 2^{\tilde{\ell}-\ell+1},[e_{0},e_{2^{\tilde{\ell}-\ell}}]\oplus 2\cdot 2^{\tilde{\ell}-\ell+1},\dots,[e_{0},e_{2^{\tilde{\ell}-\ell}}]\oplus(2^{\ell-1}-1)2^{\tilde{\ell}-\ell+1}\right\};$
$\dots$
$\mathcal{K}_{\tilde{\ell}}=\left\{[e_{2i},e_{2i+1}]|i=0,1,2,\ldots,2^{\tilde{\ell}-1}-1\right\}$
$\mathcal{K}_{\tilde{\ell}+1}=\left\{[e_{2i+1},e_{2i+2}]|i=0,1,2,\ldots,2^{\tilde{\ell}-1}-1\right\}.$
For each $i\in[\tilde{Q}-1]$ with binary form $\{i_{\ell}\}_{\ell=1}^{\tilde{\ell}}$ and $0\leq\ell\leq\tilde{\ell}$, we define $\mathrm{cut}(i,\ell)$ to be the number with binary form $\{i_{1},i_{2},\ldots,i_{\ell},0,0,\ldots,0\}$. For example, $\mathrm{cut}(i,0)=0$ and $\mathrm{cut}(i,\tilde{\ell})=i$.
$\mathcal{J}(i)=\left\{\left[e_{\mathrm{cut}(i,\ell-1)},e_{\mathrm{cut}(i,\ell)}\right]\right\}_{\ell=1}^{\tilde{\ell}}.$
Noting that for each $i\in[\tilde{Q}-1]$, we have that
$\frac{x}{2}\leq\log\left(\frac{w_{i}^{e_{i}}}{w_{i}^{e_{0}}}\right)=\sum_{\ell=1}^{\tilde{\ell}}\log\left(\frac{w_{i}^{e_{\mathrm{cut}(i,\ell)}}}{w_{i}^{e_{\mathrm{cut}(i,\ell-1)}}}\right)=\sum_{\ell=1}^{\tilde{\ell}}\log\left(\frac{w_{i}^{e_{\mathrm{cut}(i,\ell)}}}{w_{i}^{e_{\mathrm{cut}(i,\ell-1)}}}\right)\mathbb{I}\left[e_{\mathrm{cut}(i,\ell)}\neq e_{\mathrm{cut}(i,\ell-1)}\right].$ (71)
Then
$\max_{1\leq\ell\leq\tilde{\ell}}\log\left(\frac{w_{i}^{e_{\mathrm{cut}(i,\ell)}}}{w_{i}^{e_{\mathrm{cut}(i,\ell-1)}}}\right)\geq\frac{x}{2\tilde{\ell}}.$ (72)
By defining $\widetilde{\mathcal{W}}_{j}^{p}(\ell):=\left\{i\in\widehat{\mathcal{W}}_{j}^{p}:\ell=\arg\max_{1\leq\ell^{\prime}\leq\tilde{\ell}}\log\left(\frac{w_{i}^{e_{\mathrm{cut}(i,\ell^{\prime})}}}{w_{i}^{e_{\mathrm{cut}(i,\ell^{\prime}-1)}}}\right)\right\}$ for each $1\leq\ell\leq\tilde{\ell}$ [^7] Without loss of generality, we assume the $\arg\max$ function is a single value function., we then have that
$\sum_{\ell=1}^{\tilde{\ell}}\left|\widetilde{\mathcal{W}}_{j}^{p}(\ell)\right|\geq\tilde{Q}-1.$ (73)
As a result, there exists some $1\leq\ell^{*}\leq\tilde{\ell}$ such that $|\widetilde{W}_{j}^{p}(\ell^{*})|\geq\frac{\tilde{Q}-1}{\tilde{\ell}}\geq\frac{\tilde{Q}}{2\tilde{\ell}}$.
By definition, for any $i$, if $\mathrm{cut}(i,\ell^{*})\neq\mathrm{cut}(i,\ell^{*}-1)$, then $[e_{\mathrm{cut}(i,\ell^{*}-1)},e_{\mathrm{cut}(i,\ell^{*})}]\in\mathcal{K}_{\ell^{*}}$. On the other hand, by definition of $\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*})$, we know that $\mathrm{cut}(i,\ell^{*})\neq\mathrm{cut}(i,\ell^{*}-1)$ for any $i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*})$.
For each $1\leq o\leq 2^{\ell^{*}-1}$, we define
$\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*},o):=\left\{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*})|[e_{\mathrm{cut}(i,\ell^{*}-1)},e_{\mathrm{cut}(i,\ell^{*})}]=[e_{0},e_{2^{\tilde{\ell}-\ell^{*}}}]\oplus(o-1)2^{\tilde{\ell}-\ell^{*}+1}\right\}.$ (74)
Since we fix $p$ in the analysis above, we use *$\ell^{*}(p)$ and ~$\tilde{\ell}(p)$ to denote $\ell^{*}$ and $\tilde{\ell}$ respectively in the following analysis.
Fix $1\leq p\leq P$ and $1\leq o\leq 2^{\ell^{*}(p)-1}$. We let $\hat{s}(p,o)=e_{(o-1)2^{\tilde{\ell}(p)-\ell^{*}(p)+1}}$ and $\hat{e}(p,o)=e_{2^{\tilde{\ell}(p)-\ell^{*}(p)}+(o-1)2^{\tilde{\ell}(p)-\ell^{*}(p)+1}}$. In words, $[\hat{s}(p,o),\hat{e}(p,o)]$ is the $o$-th interval in the set $\mathcal{K}_{\ell^{*}(p)}$.
Then we claim that $\left\{\{\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)\}_{o=1}^{2^{\ell^{*}(p)-1}}\right\}_{p=1}^{P}$ with intervals $\left\{\{\hat{s}(p,o),\hat{e}(p,o)\}_{o=1}^{2^{\ell^{*}(p)-1}}\right\}_{p=1}^{P}$ is the desired sub-partition for $\mathcal{W}_{j}^{p}$. With a slight abuse of notations, we use $(p,o)$ as the index of the segments instead of $n$. We then verify the three conditions as below.
The condition (i).
It is clear that $\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)\subset\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p))\subset\widehat{\mathcal{W}}_{j}^{p}\subset\mathcal{W}_{j}^{p}\subset\mathcal{W}_{j}$. On the other hand, $\tilde{W}_{j}^{p_{1}}(\ell^{*}(p_{1}),o_{1})\cap\tilde{W}_{j}^{p_{2}}(\ell^{*}(p_{2}),o_{2})\neq\emptyset$ implies that $\mathcal{W}_{j}^{p_{2}}\cap\mathcal{W}_{j}^{p_{2}}\neq\emptyset$. Then we have $p_{1}=p_{2}$. Again by definition in (74), we have $o_{1}=o_{2}$.
The condition (ii).
Because
$\sum_{o=1}^{2^{\ell^{*}(p)-1}}|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|=|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p))|\geq\frac{|\widehat{\mathcal{W}}_{j}^{p}|}{4\log_{2}(|\widehat{\mathcal{W}}_{j}^{p}|)}\geq\frac{|\mathcal{W}_{j}^{p}|}{8\log_{2}(|\mathcal{W}_{j}^{p}|)}.$ (75)
Taking sum over $p$, and using Lemma 27, we have that
$\sum_{p=1}^{P}\sum_{o=1}^{2^{\ell^{*}(p)-1}}|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|=|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p))|\geq\sum_{p=1}^{P}\frac{|\mathcal{W}_{j}^{p}|}{8\log_{2}(k)}\geq\frac{|\mathcal{W}_{j}|}{24\log_{2}(k)(\log_{2}(T)+1)}.$ (76)
The condition (iii).
We set the parameter $\left\{\{g(p,o)\}_{o=1}^{2^{\ell^{*}(p)}}\right\}_{p=1}^{P}$ as $g(p,o)=\frac{\sum_{i\in\widetilde{\mathcal{W}}_{j}(\ell^{*}(p),o)}w_{i}^{\hat{s}(p,o)}}{2^{-(j+2)}\cdot|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|}\geq 1$. Then we can verify $(a)$ as
$\sum_{i\in\widetilde{\mathcal{W}}_{j}(\ell^{*}(p),o)}w_{i}^{\hat{s}(p,o)}=\frac{g(p,o)|\widetilde{\mathcal{W}}_{j}(\ell^{*}(p),o)|}{2^{j+2}}.$
Next we verify $(b)$. For any $i\in\widehat{\mathcal{W}}_{j}^{p}$, we have that $s_{i}\leq e_{\mathrm{cut}(i,\ell-1)}\leq e_{i}$ for any $1\leq\ell\leq\tilde{\ell}(p)$. It then holds that $s_{i}\leq\hat{s}(p,o)\leq e_{i}$ for any $i\in\widehat{\mathcal{W}}^{p}_{j}$.By definition of $(s_{i},e_{i})$, we have that $w_{i}^{\hat{s}(p,o)}\geq 2^{-(j+2)}$.
By definition, for any $i\in\tilde{\mathcal{w}}_{j}^{p}(\ell^{*}(p),o)$, it then holds that
$\log\left(\frac{w_{i}^{\hat{e}(p,o)}}{w_{i}^{\hat{s}(p,o)}}\right)\geq\frac{x}{2\tilde{\ell}(p)};$
$w_{i}^{\hat{s}(p,o)}\geq 2^{-(j+2)}.$
Recall that $x=\log(2)$. It then follow that
$\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{s}(p,o)}_{i}\geq 2^{-(j+2)}\cdot|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|$
$\log\left(\frac{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{e}(p,o)}_{i}}{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{s}(p,o)}_{i}}\right)\geq\log\left(\frac{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{s}(p,o)}_{i}\cdot\exp\left(\frac{x}{2\tilde{\ell}(p)}\right)}{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{s}(p,o)}_{i}}\right)=\frac{x}{2\tilde{\ell}(p)}\geq\frac{\log(2)}{2\log_{2}(k)}.$
As for $(c)$, we note that for any $\hat{s}(p,o)\leq t\leq\hat{e}(p,o)$, and $i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)$, it holds that $s_{i}\leq t\leq e_{i}$. As a result, we have that
$\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{t}_{i}\geq|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|\cdot 2^{-(j+2)}.$
Finally, we need to verify that $[\hat{s}(p_{1},o_{1}),\hat{o}(p_{1},o_{1})]$ and $[\hat{s}(p_{2},o_{2}),\hat{e}(p_{2},o_{2})]$ have no common inner point. In the case $p_{1}\neq p_{2}$ (assuming $p_{1}<p_{2}$), we have that $\tilde{s}_{p_{1}}\leq\hat{s}(p_{1},o_{1})<\hat{e}(p_{1},o_{1})\leq\tilde{e}_{p_{1}}\leq\tilde{s}_{p_{2}}\leq\hat{s}(p_{2},o_{2})<\hat{e}(p_{2},o_{2})$. In the case $p_{1}=p_{2}$ and $o_{1}\neq o_{2}$ (assuming $o_{1}<o_{2}$), we have that $\hat{s}(p_{1},o_{1})<\hat{e}(p_{1},o_{1})<\hat{s}(p_{1},o_{2})<\hat{e}(p_{1},o_{2})$. The proof is finished.
C.2.3 Putting all this together
Let $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ and $\{(\hat{s}_{n},\hat{e}_{n})\}_{n=1}^{N}$ be the construction in Lemma 28. We then divide $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ into two parts. The first part consists of $\mathcal{V}_{j}^{n}$ such that
$2^{-(j+2)}|\mathcal{V}_{j}^{n}|\cdot\frac{\log^{2}(2)}{50(\log_{2}(k)+1)^{2}\log^{2}_{2}(k)}\geq\max\left\{2\log(1/\delta^{\prime})\eta,1/k\right\},$ (77)
which we denote as $\{\bar{\mathcal{V}}_{j}^{n}\}_{n=1}^{\bar{N}}$. And the left part consists of the left sets, which we denote as $\{\underline{\mathcal{V}}_{j}^{n}\}_{n=1}^{\underline{N}}$.
By Lemma 29, for each $1\leq n\leq\bar{N}$, there exists $11\leq j^{\prime}\leq 2\log_{2}(k)+$ 1 such that
$\sum_{\tau=\hat{s}_{n}}^{\hat{e}_{n}-1}\mathbb{I}[-v^{\tau}\geq 2^{-j^{\prime}}]\geq\frac{2^{-(j+2)}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)\cdot 2^{j^{\prime}-1}}{200\log^{2}_{2}(k)(2\log_{2}(k)+1)^{2}\eta}$
or $\hat{e}_{n}-\hat{s}_{n}\geq\frac{2^{-(j+2)}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)}{200\log^{2}_{2}(k)(\log_{2}(k)+1)^{2}\eta^{2}}$.
As a result, for each $1\leq n\leq\bar{N}$, we have that
$\sum_{j^{\prime}=1}^{2\log_{2}(k)+1}\sum_{\tau=\hat{s}_{n}}^{\hat{e}_{n}-1}\mathbb{I}[-v^{\tau}\geq 2^{-j^{\prime}}]2^{-(j^{\prime}-1)}+(\hat{e}_{n}-\hat{s}_{n})\eta\geq\frac{2^{-(j+2)}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)\cdot}{200\log^{2}_{2}(k)(2\log_{2}(k)+1)^{2}\eta}.$ (78)
By noting that $\sum_{j^{\prime}=1}^{2\log_{2}(k)+1}\mathbb{I}[x\geq 2^{-j^{\prime}}]\cdot 2^{-(j^{\prime}-1)}\leq 4x$ for any $x\geq 0$, we have that
$\sum_{\tau=\hat{s}_{n}}^{\hat{e}_{n}-1}4\cdot(-v^{\tau})+2(\hat{e}_{n}-\hat{s}_{n})\eta\geq\frac{2^{-(j+2)}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)\cdot}{200\log^{2}_{2}(k)(2\log_{2}(k)+1)^{2}\eta}.$ (79)
Taking sum over $n$ we have that
$T\eta+4\sum_{t=1}^{T}(-v^{t})\geq\sum_{n=1}^{\bar{n}}\left(\sum_{\tau=\hat{s}_{n}}^{\hat{e}_{n}-1}4\cdot(-v^{\tau})+(\hat{e}_{n}-\hat{s}_{n})\eta\right)\geq\frac{2^{-(j+2)}\sum_{n=1}^{\bar{N}}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)\cdot}{200\log^{2}_{2}(k)(2\log_{2}(k)+1)^{2}\eta}.$ (80)
On the other hand, by (39)
$\sum_{t=1}^{T}(-v^{t})\leq 100\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right),$ (81)
we have that
$\sum_{n=1}^{\bar{N}}|\bar{\mathcal{V}}_{j}^{n}|$ $\leq\frac{3200(\log_{2}(k)+1)^{4}\eta\cdot 2^{j+2}}{\log^{2}(2)}\cdot\left(100\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)\right)+$ (82)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{800T(\log_{2}(k)+1)^{4}\cdot 2^{j+2}\eta^{2}}{\log^{2}(2)}.$
On the other hand, by the first part of Lemma 29, we have that $\hat{e}_{n}-\hat{s}_{n}\geq\frac{\log(2)}{4\eta\log_{2}(k)}$. As a result, we have that
$\sum_{n=1}^{\underline{N}}|\underline{\mathcal{V}}_{j}^{n}|$ $\leq\frac{4T\eta\log_{2}(k)}{\log(2)}\cdot 2^{j+2}\frac{2\log_{2}(k)}{\log(2)}\cdot\max\{2\log(1/\delta^{\prime})\eta,1/k\}$
$\leq\frac{16\cdot 2^{j}T\eta^{2}\log_{2}^{2}(k)\log(1/\delta^{\prime})}{\log^{2}(2)}+\frac{82^{j}T\eta\log_{2}^{2}(k)}{k\log^{2}(2)}.$
Note that $2\log(1/\delta^{\prime})\eta\geq\frac{1}{k}$ by the parameter configuration. We obtain that
$\sum_{n=1}^{N}|\mathcal{V}_{j}^{n}|\leq 3200000\left((\log_{2}(k)+1)^{3}(\log(k)+\log(1/\delta^{\prime})^{2}\right)\cdot 2^{j}.$ (83)
It then follows that
$|\mathcal{W}_{j}|\leq 8\cdot 10^{7}\cdot\left((\log_{2}(k)+1)^{4}\left(\log(k)+\log(1/\delta^{\prime})\right)^{2}(\log_{2}(T)+1)\right)\cdot 2^{j}.$
The proof is thus completed.
C.3 Statement and proof of Lemma 29
Recall the definition of segment in Definition 2. We have the following lemma to bound the length of segments.
Lemma 29.
Assume the conditions in Lemma 8 hold. Suppose $(t_{1},t_{2})$ is a $\left(p,q,x\right)$-segment such that $p\geq 2q$. Then $t_{2}-t_{1}\geq\frac{x}{2\eta}$. Moreover, if we assume that $\frac{qx^{2}}{50(\log_{2}(k)+1)^{2}}\geq\max\{2\log(1/\delta^{\prime})\eta,1/k\}$, it then holds that: with probability $1-6T^{4}k\delta^{\prime}$, at least one of the two claims holds: (1) there exists $11\leq j\leq 2\log_{2}(k)+$ 1 such that $\sum_{\tau=t_{1}}^{t_{2}-1}\mathbb{I}[-v^{\tau}\geq 2^{-j}]\geq\frac{qx^{2}\cdot 2^{\tilde{j}-1}}{100(2\log_{2}(k)+1)^{2}\eta}$; (2) $t_{2}-t_{1}\geq\frac{qx^{2}}{100(\log_{2}(k)+1)^{2}\eta^{2}}$.
Proof.
Let $Z^{t}=\sum_{i=1}^{k}W_{i}^{t}$. We start by proving the first argument. By definition, there exists $i\in[k]$ such that
$\log\left(\frac{w^{t_{2}}_{i}}{w^{t_{1}}_{i}}\right)\geq x,$
which corresponds to
$\eta\sum_{\tau=t_{1}}^{t_{2}-1}\hat{r}_{i}^{\tau}-\log(Z^{t_{2}}/Z^{t_{1}})\geq x.$
Because $\log(Z^{t_{2}}/Z^{t_{1}})\geq-\eta(t_{2}-t_{1})$ and $\hat{r}_{i}^{\tau}\leq 1$ for any $1\leq\tau\leq T$, we have that
$x\leq 2(t_{2}-t_{1})\eta.$ (84)
Then the conclusion follows.
With a slight abuse of notation, we also use $h$ to denote the vector $\{L(h,\mathbf{e}_{i})-\mathrm{OPT}\}_{i\in[k]}$. By the conditions about the $\varepsilon$-optimility of $h^{t}$ in Lemma 8, we have that for any $1\leq\tau,t\leq T$, it holds that
$h^{\tau}\cdot w^{t}\geq h^{t}\cdot w^{t}-\varepsilon_{1}=v^{t}-\varepsilon_{1}.$ (85)
Taking sum over $\tau$, we have that
$\sum_{\tau=t_{1}}^{t_{2}-1}h^{\tau}\cdot w^{t}\geq(t_{2}-t_{1})(v^{t}-\varepsilon_{1}).$
On the other hand, we note that $W^{t}_{i}=\exp(\eta\sum_{\tau=1}^{t}\left(h^{\tau}_{i}+\mathrm{OPT}+\xi^{\tau}_{i}\right))$, where $\xi_{i}^{\tau}=\hat{r}_{i}^{\tau}-h^{\tau}_{i}-\mathrm{OPT}$ is a zero-mean noise.
Then we learn that
$\sum_{\tau=t_{1}}^{t_{2}-1}h^{\tau}\cdot w^{t}=w^{t}\cdot\left(\frac{1}{\eta}\log\left(\frac{W^{t_{2}}}{W^{t_{1}}}\right)-\Delta_{t_{1},t_{2}}\right)-(t_{2}-t_{1})\mathrm{OPT}\geq(t_{2}-t_{1})(v^{t}-\varepsilon_{1}),$ (86)
where $\log(x/y)$ denotes the vector $\{\log(x_{i}/y_{i})\}_{i\in[k]}$ for two $k$-dimensional vectors $(x,y)$, and $\Delta_{t_{1},t_{2}}=\sum_{\tau=t_{1}}^{t_{2}-1}\xi^{\tau}$.
Recall that $Z^{t}=\sum_{i=1}^{k}W^{t}_{i}$ and $w^{t}_{i}=\frac{W^{t}_{i}}{Z^{t}}$. By choose $t_{1}=t$, we then have that
$w^{t}\cdot\left(\log\left(\frac{w^{t_{2}}}{w^{t}}\right)-\eta\Delta_{t,t_{2}}\right)+\log\left(\frac{Z^{t_{2}}}{Z^{t}}\right)-\eta(t_{2}-t)\mathrm{OPT}\geq\eta(t_{2}-t)(v^{t}-\varepsilon_{1}).$
Reversing the inequality, we have that
$\mathsf{KL}(w^{t}\,\|\,w^{t_{2}})$
$:=w^{t}\cdot\left(\log\left(\frac{w^{t}}{w^{t_{2}}}\right)\right)$
$\leq\log(Z^{t_{2}}/Z^{t})-\eta(t_{2}-t)\mathrm{OPT}-\eta w^{t}\cdot\Delta_{t,t_{2}}+(t_{2}-t)\eta(\varepsilon_{1}-v^{t}).$ (87)
(87) provides an upper bound for the KL divergence between $w^{t}$ and $w^{t_{2}}$. We then carefully deal with the RHS of (87).
We start with bounding $\log(Z^{t_{2}}/Z^{t})$. With probability $1-2T^{2}k\delta^{\prime}$, it holds that
$\log(Z^{t_{2}}/Z^{t})$ $:=\sum_{\tau=t}^{t_{2}-1}\log(Z^{\tau+1}/Z^{\tau})$
$=\sum_{\tau=t}^{t_{2}-1}\log\left(\sum_{i=1}^{k}w^{\tau}_{i}\exp(\eta\hat{r}_{i}^{\tau})\right)$
$\leq\sum_{\tau=t}^{t_{2}-1}\left(\eta\sum_{i=1}^{k}w_{i}^{\tau}\hat{r}_{i}^{\tau}+2\eta^{2}\right)$
$\leq\eta\sum_{\tau=t}^{t_{2}-1}v^{\tau}+\eta(t_{2}-t)\mathrm{OPT}+\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}(w_{i}^{\tau}(\hat{r}_{i}^{\tau}-h_{i}^{\tau}-\mathrm{OPT}))+2(t_{2}-t)\eta^{2}$ (88)
$\leq\eta(t_{2}-t)\varepsilon_{1}+\eta(t_{2}-t)\mathrm{OPT}+\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}(w_{i}^{\tau}(\hat{r}_{i}^{\tau}-h_{i}^{\tau}-\mathrm{OPT}))+2(t_{2}-t)\eta^{2}.$ (89)
(88) holds by the fact that $v^{t}=w^{t}\cdot h^{t}$, and (89) is by the fact that $v^{\tau}\leq f^{\tau}+\varepsilon_{1}\leq\varepsilon_{1}$.
We then consider to bound the two following noise terms.
$\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}(w_{i}^{\tau}(\hat{r}_{i}^{\tau}-h_{i}^{\tau}-\mathrm{OPT}))$ $=\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{\tau}\xi_{i}^{\tau},$
$-\eta w^{t}\Delta_{t,t_{2}}$ $=-\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{t}\xi_{i}^{\tau}.$ (90)
Let $\mathcal{F}^{\tau}$ denote the event field before the $\tau$-th round in Algorithm 1. Note that the variance of $\xi_{i}^{\tau}$ is at most $O\left(\frac{1}{k\bar{w}^{\tau}_{i}}\right)$ by the update rule (See Line 1 in Algorithm 1), and $\{\xi_{i}^{\tau}\}$ are independent conditioned on $\mathcal{F}^{\tau}$. Let $q^{\tau}=\sum_{i=1}^{k}w_{i}^{t}\xi_{i}^{\tau}$. Then we have that $|q^{\tau}|\leq 1$, $\mathbb{E}[q^{\tau}|\mathcal{F}^{\tau}]=0$ and $\mathbb{V}[q^{\tau}|\mathcal{F}^{\tau}]\leq\sum_{i=1}^{k}\frac{(w_{i}^{t})^{2}}{k\bar{w}_{i}^{\tau}}\leq\sum_{i=1}^{k}\frac{w_{i}^{t}}{k}=\frac{1}{k}$. By Freedman’s inequality (Lemma 20), with probability $1-\delta^{\prime}$,
$\left|\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{t}\xi_{i}^{\tau}\right|\leq 2\sqrt{\frac{(t_{2}-t)}{k}\log(2/\delta^{\prime}))}+2\log(2/\delta^{\prime});$ (91)
For the other term, by letting $\hat{q}^{\tau}=\sum_{i=1}^{k}w_{i}^{\tau}\xi_{i}^{\tau}$, we then have $|\hat{q}^{\tau}|\leq 1$, $\mathbb{E}[\hat{q}^{\tau}|\mathcal{F}^{\tau}]=0$ and $\mathbb{V}[\hat{q}^{\tau}|\mathcal{F}^{\tau}]\leq\sum_{i=1}^{k}\frac{(w_{i}^{\tau})^{2}}{k\bar{w}_{i}^{\tau}}\leq\sum_{i=1}^{k}\frac{w_{i}^{\tau}}{k}=\frac{1}{k}$. Again, by Freedman’s inequality (Lemma 20), with probability $1-\delta^{\prime}$
$\left|\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{\tau}\xi_{i}^{\tau}\right|\leq 2\sqrt{\frac{(t_{2}-t)}{k}\log(2/\delta^{\prime}))}+2\log(2/\delta^{\prime}).$ (92)
By (87), (89), (91) and (92), and using union bounds over $(t,t_{2})$, with probability $1-6T^{4}k\delta^{\prime}$, we have that
$\mathsf{KL}(w^{t}\,\|\,w^{t_{2}})$ $\leq 2(t_{2}-t)\eta\varepsilon_{1}-(t_{2}-t)\eta v^{t}$
$\quad\quad\quad+2\eta\sqrt{\frac{(t_{2}-t)\log(1/\delta^{\prime})}{k}}+2(t_{2}-t)\eta^{2}+2\log(1/\delta^{\prime})\eta$ (93)
holds for any $1\leq t<t_{2}\leq T$.
[…]
For $1\leq j\leq j_{\mathrm{max}}:=\left\lfloor 2\log_{2}(k)+1\right\rfloor$, we let
$\tau_{j}:=\min\{\tau|t_{1}\leq\tau\leq t_{2}-1,\,-v^{\tau}\leq 2^{-(j-1)}\}.$
We also define $\tau_{j_{\mathrm{max}+1}}=t_{2}$. By definition, we know that $\tau_{1}=t_{1}$ and $\tau_{j_{2}}\geq\tau_{j_{1}}$ for $j_{2}\geq j_{1}$. Let $\mathcal{I}$ be the index set of this segment. Let $y_{j}:=\sum_{i\in\mathcal{I}}w_{i}^{\tau_{j}}$.
We then claim that there exists $1\leq\tilde{j}\leq j_{\mathrm{max}}$ such that
$\log\left(\frac{y_{\tilde{j}+1}}{y_{\tilde{j}}}\right)$ $\geq\frac{x}{2\log_{2}(k)+1}.$ (94)
Suppose that none of $1\leq\tilde{j}\leq j_{\mathrm{max}}$ satisfies (94). Then for any $j$, it holds that $\log\left(\frac{y_{j+1}}{y_{j}}\right)\leq\frac{x}{2\log_{2}(k)+1}$, which implies that $y_{j}\geq y_{j+1}\exp\left(\frac{x}{2\log_{2}(k)+1}\right)$. As a result, we have that $y_{1}\geq y_{j_{\mathrm{max}}+1}\cdot\exp\left(-j_{\mathrm{max}}\cdot\frac{x}{2\log_{2}(k)+1}\right)\geq p$, which leads to contradiction.
Now we assume $\tilde{j}$ satisfies (94). By the definition, we have that $y_{\tilde{j}}\geq q$.
By (93), we have that
$\mathsf{KL}(w^{\tilde{j}}\,\|\,w^{\tau_{\tilde{j}+1}})$ $\leq 2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta\varepsilon_{1}+(\tau_{{\tilde{j}}+1}-\tau_{\tilde{j}})\eta 2^{-(\tilde{j}-1)}$
$\quad+2\eta\sqrt{\frac{(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\log(1/\delta^{\prime})}{k}}+2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta^{2}+2\log(1/\delta^{\prime})\eta.$ (95)
Since $\log\left(\frac{y_{\tilde{j}+1}}{y_{\tilde{j}}}\right)\geq\frac{x}{2\log_{2}(k)+1}$ and $y_{\tilde{j}}\geq q$, by Lemma 23 and Lemma 24, we have that
$\mathsf{KL}(w^{\tau_{\tilde{j}}}\,\|\,w^{\tau_{\tilde{j}+1}})\geq\mathsf{KL}\left(\mathrm{Ber}\left(y_{\tilde{j}}\right)\,\|\,\mathrm{Ber}\left(y_{\tilde{j}+1}\right)\right)\geq\frac{qx^{2}}{4(2\log_{2}(k)+1)^{2}},$
where $\mathrm{Ber}(x)$ denote the Bernoulli distribution with mean $x\in[0,1]$. As a result, we obtain that
$\frac{qx^{2}}{4(2\log_{2}(k)+1)^{2}}$ $\leq 2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta\varepsilon_{1}+(\tau_{{\tilde{j}}+1}-\tau_{\tilde{j}})\eta 2^{-({\tilde{j}}-1)}$
$\quad+2\eta\sqrt{\frac{(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\log(1/\delta^{\prime})}{k}}+2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta^{2}+2\log(1/\delta^{\prime})\eta.$ (96)
By noting that $\frac{qx^{2}}{50(2\log_{2}(k)+1)^{2}}\geq\max\{2\log(1/\delta^{\prime})\eta,1/k\}$ and the fact that $\frac{2^{j_{\mathrm{max}}}}{\eta}\geq\frac{1}{\eta^{2}}$, we have that
$\tau_{\tilde{j}+1}-\tau_{\tilde{j}}$ $\geq\min\left\{\frac{qx^{2}}{100(2\log_{2}(k)+1)^{2}}\min\left\{\frac{1}{\eta\varepsilon_{1}},\frac{2^{\tilde{j}-1}}{\eta},\frac{1}{\eta^{2}}\right\},\frac{kx^{2}2^{-2l_{1}}}{10000\eta^{2}\log(1/\delta^{\prime})(\log_{2}(k)+1)^{4}}\right\}$
$\geq\frac{qx^{2}\cdot 2^{\tilde{j}-1}}{100(2\log_{2}(k)+1)^{2}\eta}.$ (97)
By definition, we learn that $\sum_{\tau=t_{1}}^{t_{2}-1}\mathbb{I}[-v^{\tau}\geq 2^{-j}]\geq\tau_{j+1}-t_{1}\geq\tau_{j+1}-\tau_{j}$ for any $11\leq j\leq 2\log_{2}(k)+$ 1. As a result we have that $\sum_{\tau=t_{1}}^{t_{2}-1}\mathbb{I}[-v^{\tau}\geq 2^{-\tilde{j}}]\geq\frac{qx^{2}\cdot 2^{\tilde{j}-1}}{100(2\log_{2}(k)+1)^{2}\eta}$.
The proof is completed by noting that $\tau_{j_{\mathrm{max}}+1}=t_{2}$.
Appendix D Missing proofs for learning Rademacher classes
D.1 Restatement and Proof of Lemma 15
Lemma 15. [Restatement] By running Algorithm 3, with probability $41-\delta/$ 4, it holds that
$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (98)
for any $1\leq t\leq T$.
Proof.
We will follow the notation in the proof of Lemma 4. Fix $n=\{n_{i}\}_{i=1}^{k}$ such that $n_{i}\geq 12\log(2k)$ for all $i\in[k]$ and $w\in\Delta(k)$. Let $\kappa=\min_{i}\frac{n_{i}}{w_{i}}$. Recall that $(x_{i,j},y_{i,j})$ is the $j$-th sample from $D_{i}$.
Define
$F(n,w):=\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right)\right].$
Let $\{\{(\tilde{x}_{i,j},\tilde{y}_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}$ be a group of ghost samples for $\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}$. We then have that
$F(n,w)$ (99)
$\leq\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{(\tilde{x}_{i,j},\tilde{y}_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\left(\ell(h,(x_{i,j},y_{i,j}))-\ell(h,(\tilde{x}_{i,j},\tilde{y}_{i,j}))\right)\right)\right]$
$=\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{(\tilde{x}_{i,j},\tilde{y}_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\left(\ell(h,(x_{i,j},y_{i,j}))-\ell(h,(\tilde{x}_{i,j},\tilde{y}_{i,j}))\right)\right)\right]$
$\leq 2\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{1}{\kappa}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,(x_{i,j},y_{i,j}))\right)\right]$ (100)
$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}.$ (101)
Here (100) is by Lemma 30 as below.
Lemma 30.
Let $\mathcal{L}$ be a subset of $\mathbb{R}^{n}$. Let $w^{1},w^{2}\in\mathbb{R}^{n}$ be such that $|w^{1}_{i}|\leq|w_{i}^{2}|$ for all $i\in[n]$. Then it holds
$\mathbb{E}_{\{\sigma_{i}\}_{i=1}^{n}\overset{\mathrm{iid}}{\sim}\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{1}_{i}f_{i}\right]\leq\mathbb{E}_{\{\sigma\}_{i=1}^{n}\overset{\mathrm{iid}}{\sim}\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{2}_{i}f_{i}\right].$ (102)
Proof.
It suffices to prove (102) under the case where $w_{i}^{1}=w^{2}_{i}$ for $11\leq i\leq n-$ 1, and $|w^{1}_{n}|\leq|w^{2}_{n}|$. Fix $\sigma_{i}$ for $11\leq i\leq n-$ 1.
$\mathbb{E}_{\sigma_{n}\sim\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w_{i}^{1}f_{i}\right]$
$=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{1}f_{i}+w_{n}^{1}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{1}f_{i}-w_{n}^{1}f_{n}\right)$
$=\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{1}(f_{i}^{1}+f_{i}^{2})+w_{n}^{1}(f_{n}^{1}-f_{n}^{2})\right)$
$=\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}(f_{i}^{1}+f_{i}^{2})+w_{n}^{1}(f_{n}^{1}-f_{n}^{2})\right)$
$\leq\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}(f_{i}^{1}+f_{i}^{2})+|w_{n}^{2}(f_{n}^{1}-f_{n}^{2})|\right)$
$=\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}(f_{i}^{1}+f_{i}^{2})+w_{n}^{2}(f_{n}^{1}-f_{n}^{2})\right)$
$=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}f_{i}+w_{n}^{2}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}f_{i}-w_{n}^{2}f_{n}\right)$
$=\mathbb{E}_{\sigma_{n}\sim\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w_{i}^{2}f_{i}\right].$
The proof is completed by taking expectation over $\{\sigma_{i}\}_{i=1}^{n-1}$. ∎
By Lemma 22 with $c=\frac{1}{\kappa}$, we have that
$\mathrm{Pr}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\left|\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right)-F(n,w)\right|\geq\varepsilon\right]$
$\leq$ $2\exp\left(-\frac{2\kappa^{2}\varepsilon^{2}}{\sum_{i=1}^{k}n_{i}}\right).$ (103)
By (101) and (103), for any $\delta^{\prime}\in(0,1]$, with probability $1-\delta^{\prime}$, it holds that
$\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right)$
$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{d\log(2\sum_{i=1}^{k}n_{i})+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (104)
By replacing $(\ell,L)$ with $(-\ell,-L)$, using similar arguments we obtain that, with probability $1-2\delta^{\prime}$,
$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right|$
$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{d\log(2\sum_{i=1}^{k}n_{i})+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (105)
Now we fix $\kappa\geq 0$. Define that
~=$\tilde{\mathcal{L}}=\left\{n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\in\Delta_{\varepsilon_{1}/(8k)}(k)\mid T_{1}w_{i}\leq 2n_{i},12\log(2k)\leq n_{i}\leq T_{1},\forall i\in[k],\sum_{i=1}^{k}n_{i}\leq 2T_{1}\right\}.$
By (105), for any $\delta^{\prime}>0$, with probability $1-\delta^{\prime}$, for any $\{n,w\}\in\tilde{\mathcal{L}}$, we have that
$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right|$
$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{T_{1}}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{T_{1}}\sqrt{\frac{d\log(2\sum_{i=1}^{k}n_{i})+\log(|\tilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}$
$\leq 4\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{d\log(4T_{1})+\log(2|\tilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{T_{1}}}$
$\leq 4\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{d\log(4T_{1})+2k\log(16kT_{1}/\epsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}$
$\leq 600C_{T_{1}}+4\sqrt{\frac{d\log(4T_{1})+2k\log(16kT_{1}/\epsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}.$
Here the last inequality is by Assumption 10, Lemma 12, Lemma 13 and the fact that $\sum_{i=1}^{k}n_{i}\geq\frac{T_{1}}{2}$.
Note that in Algorithm 3,
$\hat{L}^{t}(h,w^{t})=\sum_{i=1}^{k}\frac{w_{i}^{t}}{\check{n}_{i}^{t}}\cdot\sum_{j=1}^{\check{n}_{i}^{t}}\ell(h,(x_{i,j},y_{i,j})).$
By the definition that $\check{n}_{i}^{t}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ for $i\in[k]$, we have that $T_{1}w_{i}^{t}\leq\check{n}_{i}^{t}-1$ and $12\log(2k)\leq\check{n}_{i}^{t}\leq T_{1}$ for all $i\in[k]$. On the other hand, $\sum_{1=1}^{k}\check{n}_{i}^{t}\leq T_{1}+k+12k\log(2k)\leq 2T_{1}-2$. Therefore, there exists some $\tilde{w}^{t}\in\Delta(k)$ such that $\{\{\check{n}_{i}^{t}\}_{i=1}^{k},\tilde{w}^{t}\}\in\tilde{\mathcal{L}}$ and $\|w^{t}-\tilde{w}^{t}\|_{1}\leq\frac{\varepsilon_{1}}{8k}$ for each $1\leq t\leq T$. Choose $\delta^{\prime}=\delta/4$. We then obtain that: with probability $41-\delta/$ 4, it holds that
$\max_{h\in\mathcal{H}}|\hat{L}^{t}(h,w^{t})-L(h,w^{t})|$
$\leq 600C_{T_{1}}+4\sqrt{\frac{d\log(4T_{1})+2k\log(16kT_{1}/\epsilon_{1})+\log(8/\delta)}{T_{1}}}$
$\leq\frac{\varepsilon_{1}}{2}.$ (106)
for any $1\leq t\leq T$. Here the last inequality is by definition of $T_{1}$.
Finally, by noting that $h^{t}=\arg\min_{h\in\mathcal{H}}\hat{L}^{t}(h,w^{t})$, we have that
$L(h^{t},w^{t})\leq\hat{L}^{t}(h^{t},w^{t})+\frac{\varepsilon_{1}}{2}=\min_{h\in\mathcal{H}}\hat{L}^{t}(h,w^{t})+\frac{\varepsilon_{1}}{2}\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}.$
The proof is completed ∎
D.2 Restatement and Proof of Lemma 12
Lemma 12.[Restatement] For any two groups of positive integers $\{n_{i}\}_{i=1}^{k}$ and $\{m_{i}\}_{i=1}^{k}$, it holds that
$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$ $\leq\left(\sum_{i=1}^{k}(m_{i}+n_{i})\right)\widetilde{\mathsf{Rad}}_{\{m_{i}+n_{i}\}_{i=1}^{k}}$
$\leq\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$ (107)
Proof.
In what follows, assume that each $z_{i}^{j}$ obeys $z_{i}^{j}\sim\mathcal{D}_{i}$, and each $\sigma_{i}^{j}$ is a zero-mean Rademacher random variable. Direct computation then gives
$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$
$=\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
=$\overset{\mathrm{(i)}}{=}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]\right]$
$\overset{\mathrm{(ii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
$=\left(\sum_{i=1}^{k}(n_{i}+m_{i})\right)\widetilde{\mathsf{Rad}}_{\{n_{i}+m_{i}\}_{i=1}^{k}}$
$\overset{\mathrm{(iii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
$\qquad\qquad\qquad\qquad+\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=n_{i}+1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
$=\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$
Here, (i) is valid due to the zero-mean property of $\{\sigma_{i}^{j}\}$, (ii) comes from Jensen’s inequality, and (iii) holds since $\max_{x}(f_{1}(x)+f_{2}(x))\leq\max_{x}f_{1}(x)+\max_{x}f_{2}(x)$. ∎
D.3 Restatement and Proof of Lemma 13
Lemma 13.[Restatement] Consider any $\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for each $i\in[k]$. By taking $w\in\Delta^{k}$ with $w_{i}=\frac{n_{i}}{\sum_{l=1}^{k}n_{l}}$, one has
$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}\leq 72\mathsf{Rad}_{\sum_{i=1}^{k}n_{i}}\big{(}\mathcal{D}(w)\big{)}.$
Proof.
Set $\widetilde{n}=\sum_{i=1}^{k}n_{i}$. Let $\{X_{j}\}_{j=1}^{n}$ be $n$ i.i.d. multinomial random variables with parameter $\{w_{i}\}_{i=1}^{k}$, and take $\widehat{n}_{i}=\sum_{j=1}^{n}\mathds{1}\{X_{j}=i\}$ for each $i\in[k]$. From (44) and Definition 9, it is easily seen that
$\mathsf{Rad}_{\widetilde{n}}\big{(}D(w)\big{)}$ $=\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{\widetilde{n}}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{\widetilde{n}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right]\right],$
where each $z_{i}^{j}$ is independently drawn from $\mathcal{D}_{i}$, and each $\sigma_{i}^{j}$ is an independent Rademacher random variable.
In addition, Lemma 21 tells us that: for any $i\in[k]$, one has
$\widehat{n}_{i}\geq\frac{1}{3}n_{i}-2\log(2k)\geq\frac{1}{6}n_{i}\qquad\Longrightarrow\qquad\widehat{n}_{i}\geq\left\lceil\frac{1}{6}n_{i}\right\rceil\eqqcolon\widetilde{n}_{i}$ (108)
with probability exceeding $1-\frac{1}{2k}$. By defining $\mathcal{E}$ to be the event where $\widehat{n}_{i}\geq n_{i}/6$ for all $i\in[k]$, we can immediately see from the union bound that
$\mathbb{P}(\mathcal{E})\geq 1/2.$
Consequently, we can derive
$\mathsf{Rad}_{\widetilde{n}}\big{(}\mathcal{D}(w)\big{)}$ $\geq\mathbb{P}(\mathcal{E})\cdot\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{\widetilde{n}}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{\widetilde{n}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\mid\mathcal{E}\right]\right]\right]$
$\geq\frac{1}{2}\cdot\frac{\sum_{i=1}^{k}\widetilde{n}_{i}}{\widetilde{n}}\widetilde{\mathsf{Rad}}_{\{\widetilde{n}_{i}\}_{i=1}^{k}}$
$\geq\frac{1}{12}\cdot\frac{1}{6}\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}},$ (109)
where the last two inequalities both follow from Lemma 39. This concludes the proof. ∎
D.4 Necessity of Assumption 10
In this section, we will discuss the necessity of Assumption 10 in comparison to the following weaker assumption, where we only assume that the Rademacher complexity on each $\mathcal{D}_{i}$ is well bounded.
Assumption 31.
For each $n\geq 1$, there exists a universal constant $C_{n}\geq 0$ such that
$\mathsf{Rad}_{n}(\mathcal{D}_{i})\leq C_{n}$ (110)
for all $i\in[k]$.
Formally, we have the following results.
Lemma 32.
Let $w^{0}=[1/k,1/k,\ldots,1/k]^{\top}$. There exists a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that
$\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))\geq\Omega\left(\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})\right)$ (111)
for $n\geq 12k\log(k)$.
Proof.
Without loss of generality, we set $\mathcal{Y}=\{0\}$ and $\ell(h,(x,y))=h(x)-y=h(x)$. We can then regard $\mathcal{D}_{i}$ as a distribution over $\mathcal{X}_{i}$ because there is only one element in $\mathcal{Y}$.
We pick $k$ subsets of $\mathcal{X}$ as $\{\mathcal{X}_{i}\}_{i=1}^{k}$. For each $i\in[k]$, we choose the distribution $\mathcal{D}_{i}$ to be an arbitrary distribution supported by $\mathcal{X}_{i}$. On the other hand, we also define $\mathcal{H}_{i}$ to be a set of hypothesis $h(x)=0$ for all $x\notin\mathcal{X}_{i}$ for each $i\in[k]$. For $\{h_{i}\}_{i=1}^{k}$ such that $h_{i}\in\mathcal{H}_{i}$, we define $\mathsf{joint}(\{h_{i}\}_{i=1}^{k})$ to be the hypothesis $h$ such that $h(x)=h_{i}(x),\forall x\in\mathcal{X}_{i},i\in[k]$ and $h(x)=0$ for $x\notin\cup_{i}\mathcal{X}_{i}$. Now we construct the hypothesis set $\mathcal{H}$ as
=$\mathcal{H}=\{\mathsf{joint}(\{h_{i}\}_{i=1}^{k})|h_{i}\in\mathcal{H}_{i},\forall i\in[k]\}.$ (112)
Recall the definition of $\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\{\mathcal{D}_{i}\}_{i=1}^{k})$. Then we have that
$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\{\mathcal{D}_{i}\}_{i=1}^{k})$ $=\mathop{\mathbb{E}}\limits_{\{x_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h(x_{i}^{j})\right]\right]$
$=\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}\max_{h_{i}\in\mathcal{H}_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h_{i}(x_{i}^{j})\right]\right]$ (113)
$=\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}n_{i}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}}\left[\frac{1}{n_{i}}\max_{h_{i}\in\mathcal{H}_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h_{i}(x_{i}^{j})\right]\right]$
$=\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}n_{i}\mathsf{Rad}_{n_{i}}(\mathcal{D}_{i}).$
Here (113) is by the definition of $\mathcal{H}$.
By taking $n_{i}=\frac{n}{k}$ for all $i\in[k]$ and Lemma 13, we have
$\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})=\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\{\mathcal{D}_{i}\}_{i=1}^{k})\leq 72\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))$ (114)
By Lemma 32, if we set $C_{n}\sim\sqrt{\frac{d}{n}}$ in Assumption 31, the best upper bound for $\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))\sim\sqrt{\frac{dk}{n}}$, which implies that more samples are needed to learn the mixture distribution $w^{0}$.
Moreover, under the construction in Lemma 32, by further assuming $\min_{h_{i}\in\mathcal{H}_{i}}\mathbb{E}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]=\frac{1}{2}$ for all $i\in[k]$, to find $h$ such that
$\max_{i}\mathbb{E}_{x\sim\mathcal{D}_{i}}[h(x)]\leq\frac{1}{2}+\epsilon,$ (115)
we need to find $h_{i}\in\mathcal{H}_{i}$ such that
$\mathbb{E}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]\leq\frac{1}{2}+\epsilon$ (116)
for all $i\in[k]$.
Following this intuition, we can construct a counter example under Assumption 31 as below.
Theorem 33.
There exists a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that Assumption 31 holds with $C_{n}=O\left(\sqrt{\frac{d}{n}}\right)$, and it costs at least $\tilde{\Omega}\left(\frac{dk}{\epsilon^{2}}\right)$ samples to find some $h\in\mathcal{H}$ such that
$\max_{i\in[k]}L(h,\mathbf{e}_{i})\leq\min_{h^{\prime}\in\mathcal{H}}\max_{i\in[k]}L(h^{\prime},\mathbf{e}_{i})+\epsilon.$
Proof.
With the construction in Lemma 32, it suffices to find some $\mathcal{H}^{\prime}$ and $\mathcal{D}^{\prime}$ such that the following three conditions holds.
  1. 1.
    $\mathsf{Rad}_{n}(\mathcal{D}^{\prime},\mathcal{H}^{\prime}):=\frac{1}{n}\mathbb{E}_{\{x^{j}\}_{j=1}^{n}\overset{\mathrm{iid}}{\sim}\mathcal{D}^{\prime},\{\sigma^{j}\}_{j=1}^{n}\overset{\mathrm{iid}}{\sim}\{\pm 1\}}\left[\max_{h^{\prime}\in\mathcal{H}^{\prime}}\sum_{j=1}^{n}\sigma^{j}h^{\prime}(x^{j})\right]\leq C_{n};$ (117)
  2. 2.
    $\min_{h^{\prime}\in\mathcal{H}}\mathbb{E}_{x\sim\mathcal{D}^{\prime}}[h(x)]=\frac{1}{2}$;
  3. 3.
    It cost at least $\tilde{\Omega}\left(\frac{d}{\epsilon^{2}}\right)$ samples to find some $h$ such that $\mathbb{E}_{x\sim\mathcal{D}^{\prime}}[h(x)]\leq\frac{1}{2}+\epsilon$.
This construction is also straightfoward. We set $N=2^{d}$ and $\mathcal{X}^{\prime}=\{0,1\}^{N}$. Let $\mathcal{D}^{\prime}$ to be the distribution $\mathrm{Pr}_{\mathcal{D}^{\prime}}[x]=\Pi_{n=1}^{N}\mathrm{Pr}_{\mathcal{D}^{\prime}_{n}}[x_{n}]$, where $\mathrm{Pr}_{\mathcal{D}^{\prime}_{n^{*}}}[x_{n^{*}}]=\frac{1}{2}\mathbb{I}[x_{n^{*}}=1]+\frac{1}{2}\mathbb{I}[x_{n^{*}}=0]$ for some $n^{*}$ , $\mathrm{Pr}_{\mathcal{D}^{\prime}_{n}}[x_{n}]=(\frac{1}{2}+2\epsilon)\mathbb{I}[x_{n^{*}}=1]+(\frac{1}{2}-2\epsilon)\mathbb{I}[x_{n^{*}}=0]$ for all $n\neq n^{*}$. Then we set =$\mathcal{H}^{\prime}=\{h^{n}\}_{n=1}^{N}$ with $h^{n}(x)=x_{n}$ for each $n\in[N]$. It is then easy to verify that the first two condition holds. As for the third condition, ollowing the arguments in Theorem 18, we need at least $\tilde{\Omega}\left(\frac{d}{\epsilon^{2}}\right)$ i.i.d. samples from $\mathcal{D}^{\prime}$ to identify $n^{*}$. The proof is finished.
Appendix E Proof of Theorem 17
1 Input: The hypothesis set $\mathcal{H}$, the group set $\mathcal{G}$, target accuracy level $\epsilon$, target success rate $1-\delta$, minimal probability $\gamma$.
2 Initialization: $T=\frac{10000\log(|\mathcal{G}|/\delta)}{\gamma\varepsilon^{2}}$, $\eta=\frac{1}{10}\varepsilon\gamma$, $N=\frac{10000(\log(|\mathcal{G}|/\delta)+d\log(d/\varepsilon))}{\gamma\varepsilon^{2}}$, $W^{1}_{g}=1$ for all $g\in\mathcal{G}$.
3 Call $\mathrm{Query}(\mathcal{D})$ for $N$ times to get $N$ i.i.d. samples from $\mathcal{D}$, denoted by $\{(x_{i},y_{i})\}_{i=1}^{N}$.
4 $f_{g}\leftarrow\min_{h\in\mathcal{L}}\frac{1}{\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i}),\forall g\in\mathcal{G}$,
5 $N_{g}\leftarrow\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g],$ $p_{g}\leftarrow\frac{N_{g}}{N}$, $\forall g\in\mathcal{G}$
6 for $t=1,2,\ldots,T$ do
7       $w_{g}^{t}\leftarrow\frac{W_{g}^{t}}{\sum_{g\in\mathcal{G}}W_{g}^{t}},\forall g\in\mathcal{G}$,
8       $h^{t}\leftarrow\arg\min_{h\in\mathcal{H}}\left(\sum_{g\in\mathcal{G}}w_{g}^{t}\cdot\left(\frac{1}{N_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-f_{g}\right)\right)$,
9       Call $\mathrm{Query}(\mathcal{D})$ to get a data point $(\hat{x}_{t},\hat{y}_{t})$,
10       $\hat{r}^{t}_{g}\leftarrow\frac{1}{p_{g}}\mathbb{I}[\hat{x}_{t}\in g]\ell(h^{t}(\hat{x}_{t}),\hat{y}_{t})-f_{g}$, $\forall g\in\mathcal{G}$,
11       $W_{g}^{t+1}\leftarrow W_{g}^{t}\cdot\exp(\eta\hat{r}^{t}_{g})$, $\forall g\in\mathcal{G}$;
Return: A randomized hypothesis $h^{\mathsf{final}}$ as a uniform distribution over $\{h^{t}\}_{t=1}^{T}$.
Algorithm 4 Hedge for Multi-group Learning
Throughout this section, we let $\mathcal{F}_{t}$ be the event field before the $t$-th round. Clearly, the number of samples used in Algorithm 4 is bounded by $T+N=O\left(\frac{(\log(|\mathcal{G}|/\delta))+d\log(d/\epsilon)}{\gamma\epsilon^{2}}\right)$. So it suffices to prove the optimality of $h^{\mathsf{final}}$. Formally, we have the lemma below.
Lemma 34.
With probability $1-\delta$, it holds that
$\frac{1}{T}\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)\leq\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)+\epsilon$ (118)
for any $g\in\mathcal{G}$.
Proof.
Let $W^{t}=\sum_{g\in\mathcal{G}}W^{t}_{g}$. By definition and the fact that $\eta\hat{r}_{g}^{t}\leq\frac{\eta}{\gamma}\leq 1$ for any proper $(g,t)$ pair,
$\log(W_{g}^{T+1})=\eta\sum_{t=1}^{T}\hat{r}_{g}^{t},$
$\log\left(\frac{W^{t+1}}{W^{t}}\right)=\log\left(\sum_{g\in\mathcal{G}}w_{g}^{t}\exp(\eta\hat{r}_{g}^{t})\right)\leq\log\left(\sum_{g\in\mathcal{G}}w_{g}^{t}(1+\eta\hat{r}_{g}^{t}+\eta^{2}(\hat{r}_{g}^{t})^{2})\right)\leq\sum_{g\in\mathcal{G}}w_{g}^{t}(\eta\hat{r}_{g}^{t}+\eta^{2}(\hat{r}_{g}^{t})^{2}).$
Noting that
$\log(W^{T+1}_{g})\log(W^{T+1})=\sum_{t=1}^{T}\log(W^{t+1}/W^{t})+\log(|\mathcal{G}|),$
for any $\tilde{g}\in\mathcal{G}$,
$\eta\sum_{t=1}^{T}\hat{r}_{\tilde{g}}^{t}\leq\eta\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}+\eta^{2}\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}+\log(|\mathcal{G}|).$
Dividing both side by $\eta$, we learn that
$\sum_{t=1}^{T}\hat{r}_{\tilde{g}}^{t}-\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}\hat{r}_{g}^{t}\leq\frac{\log(|\mathcal{G}|)}{\eta}+\eta\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}.$ (119)
By Lemma 36, with probability $1-\frac{\delta}{16}$,
$\sum_{t=1}^{T}\hat{r}_{\tilde{g}}^{t}-\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}\leq\frac{\log(|\mathcal{G}|)}{\eta}+\frac{3\eta T}{\gamma}\leq\frac{T\epsilon}{32}$ (120)
for any $\tilde{g}\in\mathcal{G}$. By (120) and Lemma 37, with probability $1-\frac{3\delta}{4}$,
$\sum_{t=1}^{T}\left(L_{\mathcal{D}}(h^{t}|g)-f_{g}\right)\leq\sum_{t=1}^{T}\hat{r}_{g}^{t}+\frac{T\epsilon}{32}\leq\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}+\frac{5T\epsilon}{32}\leq T\epsilon$ (121)
for any $g\in\mathcal{G}$. By Lemma 35, with probability $1-\delta$, we have that
$\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)-T\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)\leq T\epsilon,$
which means that
$\frac{1}{T}\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)-\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)\leq\epsilon.$
The proof is completed.
Lemma 35.
With probability $1-\frac{\delta}{16}$, it holds that
$\left|f_{g}-\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)\right|\leq\frac{\epsilon}{8}$
for any $g\in\mathcal{G}$.
Proof.
Fix $g\in\mathcal{G}$. Let $\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}$ to be the ghost samples, another group of $N$ i.i.d. datapoints from $\mathcal{D}$ which is independent of $\{(x_{i},y_{i})\}_{i=1}^{N}$. Let $\lambda\in(0,1/2]$ be a positive real number.
Then we have that
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{N}}\left[\exp\left(\lambda\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-N\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathbb{I}[x\in g]\ell(h(x),y)]\right)\right)\right]$
$\leq\mathbb{E}_{\{(x_{i},y_{i}),(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}}\left[\exp\left(\lambda\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{N}\left(\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-\mathbb{I}[x^{\prime}_{i}\in g]\ell(h(x^{\prime}_{i}),y^{\prime}_{i})\right)\right)\right)\right]$ (122)
$=\mathbb{E}_{\{(x_{i},y_{i}),(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}}\mathbb{E}_{\{\sigma_{i}\}_{i=1}^{N}\overset{\mathrm{i.i.d.}}{\sim}\{\pm 1\}}\left[\exp\left(\lambda\max_{h\in\mathcal{H}}\sum_{i=1}^{N}\sigma_{i}\left(\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-\mathbb{I}[x^{\prime}_{i}\in g]\ell(h(x^{\prime}_{i}),y^{\prime}_{i})\right)\right)\right]$
$\leq\mathbb{E}_{\{(x_{i},y_{i}),(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}}\left[\exp\left(\sum_{i=1}^{N}2\lambda^{2}(\mathbb{I}[x_{i}\in g]+\mathbb{I}[x^{\prime}_{i}\in g])^{2}\right)\right]$
$\leq\mathbb{E}_{\{(x_{i},y_{i}),(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}}\left[\exp\left(\sum_{i=1}^{N}4\lambda^{2}(\mathbb{I}[x_{i}\in g]+\mathbb{I}[x^{\prime}_{i}\in g])\right)\right]$
$\leq\exp\left(16N\lambda^{2}P_{g}\right).$ (123)
Here (122) is by Jenson’s inequality.
As a result, we obtain that
$\mathrm{Pr}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-N\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathbb{I}[x\in g]\ell(h(x),y)]\right)>P_{g}\epsilon/32\right]$
$\leq\min_{\lambda\in(0,1/2]}\exp\left(4N\lambda^{2}P_{g}-\lambda P_{g}\epsilon/32\right)$
$\leq\exp\left(-\frac{P_{g}\epsilon^{2}}{8192N}\right)$ (124)
$\leq\exp\left(-\frac{\gamma\epsilon^{2}}{8192N}\right)$
$\leq\frac{\delta}{64|\mathcal{G}|}.$ (125)
Using similar arguments, we can obtain that
$\mathrm{Pr}\left[\max_{h\in\mathcal{H}}\left(-\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})+N\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathbb{I}[x\in g]\ell(h(x),y)]\right)>P_{g}\epsilon/32\right]\leq\frac{\delta}{64|\mathcal{G}|}$ (126)
Therefore,
$\mathrm{Pr}\left[\left|\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-L_{\mathcal{D}}(h|g)\right|\leq\epsilon/32\right]\leq\frac{\delta}{64|\mathcal{G}|}.$ (127)
Recall that $p_{g}=\frac{\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]}{N}$. With Chernoff’s inequality, we have that
$\mathrm{Pr}\left[\left|NP_{g}-Np_{g}\right|\geq 2\sqrt{NP_{g}\log(128|\mathcal{G}|/\delta)}\right]\leq\frac{\delta}{64|\mathcal{G}|},$ (128)
which means that
$\mathrm{Pr}\left[\left|\frac{1}{NP_{g}}-\frac{1}{Np_{g}}\right|\geq\frac{\epsilon}{32}\right]\leq\frac{\delta}{64|\mathcal{G}|}.$ (129)
By (127) and (129), and taking a union bound over $\mathcal{G}$, we obtain that with probability $1-\frac{\delta}{16}$,
$\left|\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-L_{\mathcal{D}}(h|g)\right|\leq\frac{\epsilon}{16}$ (130)
for any $h\in\mathcal{H},g\in\mathcal{G}$. By the definition of $h_{g}$, we have that
$\left|f_{g}-\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)\right|\leq 2\max_{h\in\mathcal{H}}\left|\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-L_{\mathcal{D}}(h|g)\right|\leq\frac{\epsilon}{8}.$
The proof is completed. ∎
Lemma 36.
With probability $1-\frac{\delta}{16}$, it holds that
$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}\leq\frac{3T}{\gamma}.$ (131)
Proof.
By definition, we have that
$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}\leq\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\frac{\mathbb{I}[x_{t}\in g]}{p_{g}^{2}}.$ (132)
Let $X_{t}=\sum_{g\in\mathcal{G}}w_{g}^{t}\frac{\mathbb{I}[x_{t}\in g]}{p^{2}_{g}}$ and recall $\mathcal{F}_{t}$ is the event field before the $t$-th round. Then $X_{t}\leq\frac{1}{\lambda^{2}}$, and $\mathbb{E}[X_{t}^{2}|\mathcal{F}_{t}]=\sum_{g,g^{\prime}}w_{g}^{t}w_{g^{\prime}}^{t}\frac{\mathbb{E}[\mathbb{I}[\hat{x}_{t}\in g,\hat{x}_{t}\in g^{\prime}]]}{p_{g}^{2}p_{g^{\prime}}^{2}}\leq\sum_{g,g^{\prime}}w_{g}^{t}w_{g^{\prime}}^{t}\frac{1}{(p_{g}p_{g^{\prime}})^{3/2}}\leq\frac{1}{\gamma^{3}}$. Using Freedman’s inequality (Lemma 20), with probability $1-\frac{\delta}{16}$,
$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}\leq\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\frac{1}{p_{g}}+2\sqrt{\frac{T\log(16/\delta)}{\gamma^{3}}}\leq\frac{3T}{\gamma},$ (133)
where the last inequality is by the fact that $T\geq\frac{\log(16/\delta)}{\gamma}$.
Lemma 37.
Assume the events in Lemma 35 hold. With probability $1-\frac{\delta}{4}$, it holds that
$\left|\sum_{t=1}^{T}\hat{r}_{g}^{t}-\sum_{t=1}^{T}(L_{\mathcal{D}}(h^{t}|g)-f_{g})\right|\leq\frac{T\epsilon}{8},\forall g\in\mathcal{G};$ (134)
$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}\leq\frac{13T\epsilon}{16}.$ (135)
Proof.
We start with the first inequality. By definition,
$\sum_{t=1}^{T}\hat{r}_{g}^{t}-\sum_{t=1}^{T}(L_{\mathcal{D}}(h^{t}|g)-f_{g})$
$=\sum_{t=1}^{T}\left(\frac{1}{p_{g}}\mathbb{I}[\hat{x}_{t}\in g]\ell(h(\hat{x}_{t}),\hat{y}_{t})-L_{\mathcal{D}}(h^{t}|g)\right).$ (136)
Defining $\hat{X}_{t}=\frac{1}{p_{g}}\mathbb{I}[\hat{x}_{t}\in g]\ell(h(\hat{x}_{t}),\hat{y}_{t})$. We then have that $\mathbb{E}[\hat{X}_{t}|\mathcal{F}_{t}]=\frac{P_{g}}{p_{g}}L_{\mathcal{D}}(h^{t}|g)$ and $\mathbb{E}[\hat{X}_{t}^{2}|\mathcal{F}_{t}]\leq\frac{P_{g}}{p_{g}^{2}}$. With Freedman’s inequality (Lemma 20) , with probability $1-\frac{\delta}{64}$,
$\left|\sum_{t=1}^{T}\hat{r}^{g}_{t}-\frac{P_{g}}{p_{g}}\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)\right|\leq\frac{2}{p_{g}}\sqrt{TP_{g}\log(128|\mathcal{G}|/\delta)}\leq\frac{T\epsilon}{16}$ (137)
for any $g\in\mathcal{G}$. Combining (137) with (129), we learn that: with probability $1-\frac{\delta}{32}$,
$\left|\sum_{t=1}^{T}\hat{r}_{g}^{t}-\sum_{t=1}^{T}(L_{\mathcal{D}}(h^{t}|g)-f_{g})\right|\leq\frac{T\epsilon}{8},$ (138)
holds for any $g\in\mathcal{G}$.
As for the second inequality, we define $\tilde{X}_{t}=\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{p_{g}}\mathbb{I}[\hat{x}_{t}\in g]\ell(h^{t}(\hat{x}_{t}),\hat{y}_{t})\right)$. It then follows that $\mathbb{E}[\tilde{X}_{t}|\mathcal{F}_{t}]=\sum_{g\in\mathcal{G}}w_{g}^{t}L_{\mathcal{D}}(h^{t}|g)\leq\frac{1}{\gamma}$ and
$\mathbb{E}[\tilde{X}_{t}^{2}|\mathcal{F}_{t}]\leq\sum_{g^{\prime},g^{\prime}}w_{g}^{t}w_{g^{\prime}}^{t}\frac{\mathbb{E}[\mathbb{I}[\hat{x}_{t}\in g,\hat{x}_{t}\in g^{\prime}]]}{p_{g}p_{g^{\prime}}}\leq\sum_{g^{\prime},g^{\prime}}w_{g}^{t}w_{g^{\prime}}^{t}\frac{1}{\sqrt{p_{g}p_{g^{\prime}}}}\leq\frac{1}{\gamma}.$
Using Freedman’s inequality (Lemma 20), we learn that, with probability $1-\frac{\delta}{64}$
$\sum_{t=1}^{T}\left(\tilde{X}_{t}-\sum_{g\in\mathcal{G}}w_{g}^{t}L_{\mathcal{D}}(h^{t}|g)\right)\leq 2\sqrt{\frac{T\log(128/\delta)}{\gamma}}+2\frac{\log(128/\delta)}{\gamma}\leq\frac{T\epsilon}{16}.$ (139)
As a result,
$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}\leq\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(L_{\mathcal{D}}(h^{t}|g)-f_{g})+\frac{T\epsilon}{16}.$ (140)
Recall the definition that
$h^{t}=\arg\min_{h\in\mathcal{H}}\left(\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{N_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-f_{g}\right)\right).$ (141)
By Lemma 35,
$\max_{h\in\mathcal{H}}\left|\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-L_{\mathcal{D}}(h|g)\right|\leq\frac{\epsilon}{16}$ (142)
for any $g\in\mathcal{G}$.
By the definition that $Np_{g}=N_{g}$, we have
$\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h^{t}(x_{i}),y_{i})-f_{g}\right)$ $\leq\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h^{*}(x_{i}),y_{i})-f_{g}\right)$
$\leq\sum_{g\in\mathcal{G}}w_{g}^{t}(L_{\mathcal{D}}(h^{*}|g)-f_{g})+\frac{\epsilon}{16}.$ (143)
By (143), Assumption 16 and Lemma 35, we have that
$\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h^{t}(x_{i}),y_{i})-f_{g}\right)\leq\frac{1}{2}\epsilon.$ (144)
On the other hand, we also have
$\sum_{g\in\mathcal{G}}w_{g}^{t}\left(L_{\mathcal{D}}(h^{t}|g)-f_{g}\right)\leq\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h^{t}(x_{i}),y_{i})-f_{g}\right)+\frac{\epsilon}{16},$
which implies that
$\sum_{g\in\mathcal{G}}w_{g}^{t}\left(L_{\mathcal{D}}(h^{t}|g)-f_{g}\right)\leq\frac{13\epsilon}{16}.$ (145)
The conclusion follows by (140) and (145).
Appendix F Missing proofs in Section 8
F.1 Restatement and Proof of Lemma 19
Lemma 19. [Restatement] Fix $m\geq 0$ and $\widetilde{i}\in[k]$. Suppose $\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{i}\leq m]\geq\frac{1}{2}$. It then holds that for any $h\in\mathcal{H}_{\widetilde{i}}$
$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m]\geq\frac{1}{2}\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m]\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}).$ (146)
Moreover, it holds that $m\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}$.
Proof.
For $v\in\{-1,1\}^{m}$, we let $n^{+}(v)=|\{p:v_{p}=1\}|$ be the number of $1$’s in all coordiantes of $v$.
Let $\mathcal{V}$ be a subset of $\{-1,1\}^{2m}$ defined as $\mathcal{V}:=\{v^{1},v^{2}\in\{-1,1\}^{m}|n^{+}(v^{1})-n^{+}(v^{2})\leq 4\sqrt{m}+2m\epsilon\}$. Let $\mathrm{Pr}_{\mathcal{C}}[\cdot]$ denote the probability distribution of $\left\{\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}\right\}$ and $\mathrm{Pr}_{\mathcal{C}^{\prime}}[\cdot]$ denote the probability distribution of $\left\{\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}\right\}$.
Recall the definition of $\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}$ and $\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}$. By Hoeffding’s inequality, we have that $\mathrm{Pr}_{\mathcal{C}}[\mathcal{V}]\geq\frac{3}{4}.$ Also noting that the distribution of $\mathcal{V}$ is independent of the algorithm $\mathcal{G}^{\prime}$, we have that
$\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\left\{\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}\right\}\in\mathcal{V}\right]\geq\frac{3}{4}-(1-\frac{3}{4})=\frac{1}{2}.$ (147)
That is,
$\sum_{v=\{v^{1},v^{2}\}\in\mathcal{V}}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\geq\frac{1}{2}.$
On the other hand, for any $v=\{v^{1},v^{2}\}\in\mathcal{V}$, we have
$\mathrm{Pr}_{\mathcal{C}^{\prime}}[v]$ $=\mathrm{Pr}_{\mathcal{C}}[v]\cdot\left(1-8\epsilon\right)^{n^{+}(v^{1})-n^{+}(v^{2})}(1+8\epsilon)^{n^{+}(v^{2})-n^{+}(v^{1})}$
$=\mathrm{Pr}_{\mathcal{C}}[v]\left(\frac{1-8\epsilon}{1+8\epsilon}\right)^{n^{+}(v^{1})-n^{+}(v^{2})}$
$\geq\mathrm{Pr}_{\mathcal{C}}[v]\exp(-20(n^{+}(v^{1})-n^{+}(v^{2}))\epsilon)$
$\geq\mathrm{Pr}_{\mathcal{C}}[v]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2}).$
As a result, we have that
$\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m,\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\mathrm{Pr}_{\mathcal{C^{\prime}}}[v]$
$\geq\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\mathrm{Pr}_{\mathcal{C}}[v]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2})$
$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\mathrm{Pr}_{\mathcal{C}}[v]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2})$ (148)
$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2}).$
Here (148) is by Lemma 38. We present the detailed computation as follows. Let $v^{o}_{\ell}$ be the $\ell$-th dimension of $v^{1}$ for $1\leq\ell\leq m$ and $o=1,2$.
$\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
$=\sum_{m^{\prime}=1}^{m}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}=m^{\prime}|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
$=\sum_{m^{\prime}=1}^{m}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}=m^{\prime}|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m^{\prime}}=\{v^{1}_{\ell}\}_{\ell=1}^{m^{\prime}},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m^{\prime}}=\{v^{2}_{\ell}\}_{\ell=1}^{m^{\prime}}\right]$ (149)
$=\sum_{m^{\prime}=1}^{m}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}=m^{\prime}|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m^{\prime}}=\{v^{1}_{\ell}\}_{\ell=1}^{m^{\prime}},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m^{\prime}}=\{v^{2}_{\ell}\}_{\ell=1}^{m^{\prime}}\right]$
$=\sum_{m^{\prime}=1}^{m}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}=m^{\prime}|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$ (150)
$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right].$
Here (149) and (150) hold because for $h^{\prime}=h,h^{*}$ the event $\{h_{\mathrm{out}}=h^{\prime},M_{\tilde{i}}=m^{\prime}\}$ is independent of $\{x^{\ell}(\tilde{i})\}_{\ell\geq m^{\prime}+1}$.
Taking sum over $v=\{v^{1},v^{2}\}\in\mathcal{V}$, we have that
$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m]$
$\geq\sum_{v\in\mathcal{V}}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m,\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
$\geq\sum_{v\in\mathcal{V}}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2})$
$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\left\{\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}\right\}\in\mathcal{V}\right]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2})$
$\geq\frac{1}{2}\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2}).$
(146) is proven.
Taking sum over all $h\in\mathcal{H}_{\tilde{i}}$, we have that
$1\geq\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}\in\mathcal{H}_{\tilde{i}},M_{\tilde{i}}\leq m]\geq\frac{N_{0}}{2}\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m]\exp(-40m\epsilon^{2}-80\sqrt{m}\epsilon)$ (151)
which implies that
$\frac{1}{2}\leq\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m\right]\leq\frac{2}{N_{0}}\exp(40m\epsilon^{2}+80\sqrt{m}\epsilon).$ (152)
Therefore, we have that $40m\epsilon^{2}+80\sqrt{m}\epsilon\geq\log(N_{0}/4)$, which implies that $m\geq\min\{\frac{\log(N_{0}/4)}{80\epsilon^{2}},\frac{\log^{2}(N_{0}/4)}{30000\epsilon^{2}}\}\geq\frac{\log(N_{0}/4)}{30000\epsilon^{2}}$
F.2 Statement and Proof of Lemma 38
Lemma 38.
Fix $\tilde{i}\in[k]$, and let $x^{\ell}(i)$ be the $\ell$-sample from $D_{i}$ for any $i\in[k]$ and $\ell\geq 1$. Fix $h\in\mathcal{H}_{\tilde{i}}$, $m>0$ and $v^{1},v^{2}\in\{-1,1\}^{2m}$.
$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{*},M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}]$
$=\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h,M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1}].$
Proof.
Let $\bar{\sigma}$ be the permutation over $\mathcal{H}$ with $\bar{\sigma}(h^{*})=h,\bar{\sigma}(h)=h^{*}$ and $\bar{\sigma}(h^{\prime})=h^{\prime}$ for all $h^{\prime}\notin\{h,h^{*}\}$. It then holds that $\bar{\sigma}^{-1}=\bar{\sigma}$.
Fix $\{m_{i}\}_{i=1}^{k}$ and $X(i)=\{X^{\ell}(i)\}_{\ell=1}^{m_{i}}\in\{-1,0,1\}^{kNm_{i}}$ for $i\in[k]$. Let $x(i)=\{x^{\ell}(i)\}_{\ell=1}^{m_{i}}$ be the datapoints of the first $m_{i}$ calls to $\mathrm{Query}(D_{i})$. With a slight abuse of notations, we let $\sigma(x(i))=\{\sigma(x^{\ell}(i))\}_{\ell=1}^{m_{i}}$ for each $i\in[k]$. We continue with the following lemmas.
By Lemma 39, we have that
$\mathrm{Pr}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right]$
$=\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=\sigma^{-1}(h),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right],$
which means
$\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h^{*},\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\sigma(x(i))=X(i),\forall i\in[k]\right]$
$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\bar{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right]\cdot\frac{\mathrm{Pr}[\sigma(x(i))=X(i),\forall i\in[k]]}{\mathrm{Pr}[\bar{\sigma}\sigma(x(i))=X(i),\forall i\in[k]]}$
$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\bar{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right]\cdot\frac{\mathrm{Pr}[\sigma(x(\tilde{i}))=X(\tilde{i})]}{\mathrm{Pr}[\bar{\sigma}\sigma(x(\tilde{i}))=X(\tilde{i})]}$
$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\bar{\sigma}\sigma(x(i)))=X(i),\forall i\in[k]\right]$
$\quad\qquad\qquad\qquad\qquad\cdot\frac{\mathrm{Pr}\left[\{x_{j_{h,i}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h^{*},\tilde{i}}}^{\ell}({\tilde{i}})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),{\tilde{i}}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\right]}{\mathrm{Pr}\left[\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\right]}.$ (153)
Re-arrange the equation, we have that
$\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\Bigg{[}h_{\mathrm{out}}=h^{*},\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\sigma(x(i))=X(i),\forall i\in[k]$
|$\qquad\qquad\qquad\qquad\Big{|}\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\Bigg{]}$
$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\Bigg{[}h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\bar{\sigma}\sigma(x(i))=X(i),\forall i\in[k]$
|$\qquad\qquad\qquad\quad\Big{|}\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\Bigg{]}.$
Taking sum over all possible choices of $\{X(i)\}_{i\neq\tilde{i}}$, $\left\{\{X^{\ell}_{j_{h^{\prime},i^{\prime}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\right\}_{i^{\prime}\in[k],h^{\prime}\notin\{h,h^{*}\}}$,
$\left\{\{X^{\ell}_{j_{h^{\prime},i^{\prime}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\right\}_{h^{\prime}\in\{h,h^{*}\},i^{\prime}\neq\tilde{i}}$ and $\{m_{i}\}_{i\neq\tilde{i}}$, we obtain that
$\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\Bigg{[}h_{\mathrm{out}}=h^{*},M_{\tilde{i}}=m_{\tilde{i}}$
|$\qquad\qquad\qquad\qquad\Big{|}\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\Bigg{]}$
$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\Bigg{[}h_{\mathrm{out}}=h,M_{\tilde{i}}=m_{\tilde{i}}$
|$\qquad\qquad\qquad\qquad\Big{|}\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\Bigg{]}$
for any $X(\tilde{i})\in\{-1,0,1\}^{kNm_{\tilde{i}}}$.
Fix $m_{\tilde{i}}=m$, and choose $\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=v_{1}$, $\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=v_{2}$. We then have that
$\mathrm{Pr}_{\mathcal{G}^{\prime},\mathcal{H}}\left[h_{\mathrm{out}}=h^{*},M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h^{*},M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=\bar{\sigma}^{-1}(h^{*}),M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1}\right]$
$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h,M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1}\right]$
$=\mathrm{Pr}_{\mathcal{G}^{\prime},\mathcal{H}}\left[h_{\mathrm{out}}=h,M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1}\right].$
The proof is completed. ∎
F.3 Statement and Proof of Lemma 39
Lemma 39.
Fix $\{m_{i}\}_{i\in[k]}$ , $\sigma\in\Pi_{\mathcal{H}}$ and $X\in\{-1,0,1\}^{kN\sum_{i=1}^{k}m_{i}}$. Let $\{X(i)\}_{i\in[k]}$ and $\{x(i)\}$ be defined as Lemma 38. We then have that
$\mathrm{Pr}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right]$
$=\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=\sigma^{-1}(h),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right].$ (154)
Proof.
Let $\mathcal{H}^{\prime}=\sigma(\mathcal{H})$. Let $h_{p}(\cdot)$ denote the $p$-th hypothesis in the hypothesis set.
$\mathrm{Pr}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h_{p}(\mathcal{H}),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right]$
$=\mathrm{Pr}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h_{p}(\mathcal{H}),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\{\{x^{l}_{j_{h_{p^{\prime}}(\mathcal{H}),i}}(i^{\prime})\}_{p^{\prime}=1,i=1}^{|\mathcal{H}|,k}\}_{\ell=1}^{m_{i^{\prime}}}\}_{i^{\prime}=1}^{k}=X\right]$
$=\mathrm{Pr}_{\mathcal{G},\mathcal{H}^{\prime}}\left[h_{\mathrm{out}}=h_{p}(\mathcal{H}^{\prime}),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\{\{x^{l}_{j_{h_{p^{\prime}}(\mathcal{H}^{\prime}),i}}(i^{\prime})\}_{p^{\prime}=1,i=1}^{|\mathcal{H}^{\prime}|,k}\}_{\ell=1}^{m_{i^{\prime}}}\}_{i^{\prime}=1}^{k}=X\right]$ (155)
$=\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h_{p}(\sigma(\mathcal{H})),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right]$
$=\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=\sigma^{-1}(h_{p}(\mathcal{H})),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right].$ (156)
[…]

### v3

@@
 Focusing on a hypothesis class of Vapnik–Chervonenkis (VC) dimension $d$, we propose a novel algorithm that yields an $\varepsilon$-optimal randomized hypothesis with a sample complexity on the order of $\frac{d+k}{\varepsilon^{2}}$ (modulo some logarithmic factor), matching the best-known lower bound.
-Our algorithmic ideas and theory have been further extended to accommodate Rademacher classes.
+Our algorithmic ideas and theory are further extended to accommodate Rademacher classes.
 The proposed algorithms are oracle-efficient, which access the hypothesis class solely through an empirical risk minimization oracle.
-Additionally, we establish the necessity of randomization, unveiling a large sample size barrier when only deterministic hypotheses are permitted.
-These findings successfully resolve three open problems presented in COLT 2023 (i.e., Awasthi et al., (2023, Problems 1, 3 and 4)).
+Additionally, we establish the necessity of randomization, revealing a large sample size barrier when only deterministic hypotheses are permitted.
+These findings resolve three open problems presented in COLT 2023 (i.e., awasthi2023sample).
 Keywords: multi-distribution learning; on-demand sampling; game dynamics; VC classes; Rademacher classes; oracle efficiency
 1 Introduction
-Driven by the growing need of robustness, fairness and multi-group collaboration in machine learning practice, the multi-distribution learning (MDL) framework has emerged as a unified solution in response to these evolving demands (Blum et al.,, 2017; Haghtalab et al.,, 2022; Mohri et al.,, 2019; Awasthi et al.,, 2023).
+Driven by the growing need of robustness, fairness and multi-group collaboration in machine learning practice, the multi-distribution learning (MDL) framework has emerged as a unified solution in response to these evolving demands (blum2017collaborative; haghtalab2022demand; mohri2019agnostic; awasthi2023sample).
 Setting the stage, imagine that we are interested in a collection of $k$ unknown data distributions $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{k}$ supported on $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ (resp.
@@
 with $\varepsilon\in(0,1]$ a target accuracy level.
-In light of the unknown nature of these data distributions, the learning process involves data collection, allowing the learner to sample from $\{\mathcal{D}_{i}\}_{i=1}^{k}$.
+In light of the unknown nature of these data distributions, the learning process is often coupled with data collection, allowing the learner to sample from $\{\mathcal{D}_{i}\}_{i=1}^{k}$.
 The performance of a learning algorithm is then gauged by its sample complexity — the number of samples required to fulfil (1).
-Our objective is to design a learning paradigm that achieves optimal sample complexity.
-The MDL framework described above, which can viewed as an extension of agnostic learning (Valiant,, 1984; Blumer et al.,, 1989) tailored to multiple data distributions, has found a wealth of applications across multiple domains.
-Here, we highlight a few representative examples, and refer the interested reader to Haghtalab et al., (2022) for more in-depth discussions.
+Our objective is to design a learning paradigm that achieves the optimal sample complexity.
+The MDL framework described above, which can viewed as an extension of agnostic learning (valiant1984theory; blumer1989learnability) tailored to multiple data distributions, has found a wealth of applications across multiple domains.
+Here, we highlight a few representative examples, and refer the interested reader to haghtalab2022demand and the references therein for more extensive discussions.
   • Collaborative and agnostic federated learning.
-In the realm of collaborative and agnostic federated learning (Blum et al.,, 2017; Nguyen and Zakynthinou,, 2018; Chen et al.,, 2018; Mohri et al.,, 2019; Blum et al., 2021a, ; Du et al.,, 2021; Deng et al.,, 2020; Blum et al., 2021b, ), a group of $k$ agents, each having access to distinct data sources as characterized by different data distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$, aim to learn a shared prediction model that ideally would achieve low risk for each of their respective data sources.
+In the realm of collaborative and agnostic federated learning (blum2017collaborative; nguyen2018improved; chen2018tight; mohri2019agnostic; blum2021one; du2021fairness; deng2020distributionally; blum2021communication), a group of $k$ agents, each having access to distinct data sources as characterized by different data distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$, aim to learn a shared prediction model that ideally would achieve low risk for each of their respective data sources.
 A sample-efficient MDL paradigm would help unleash the potential of collaboration and information sharing in jointly learning a complicated task.
   • Min-max fairness in learning.
-The MDL framework is well-suited to scenarios requiring fairness across multiple groups (Dwork et al.,, 2021; Rothblum and Yona,, 2021; Du et al.,, 2021).
+The MDL framework is well-suited to scenarios requiring fairness across multiple groups (dwork2021outcome; rothblum2021multi; du2021fairness).
 For instance, in situations where multiple subpopulations with distinct data distributions exist, a prevailing objective is to ensure that the learned model does not adversely impact any of these subpopulations.
-Such a criterion, known as “min-max fairness” in the literature (Mohri et al.,, 2019; Abernethy et al.,, 2022), plays a pivotal role in mitigating the worst disadvantage experienced by any particular subpopulation.
+One criterion designed to meet this objective, known as “min-max fairness” in the literature (mohri2019agnostic; abernethy2020active), plays a pivotal role in mitigating the worst disadvantage experienced by any particular subpopulation.
   • Distributionally robust optimization/learning.
 Another context where MDL naturally finds applications is group distributionally robust optimization and learning (DRO/DRL).
-Group DRO and DRL aim to develop algorithms that offer robust performance guarantees across a finite number of possible distributional models (Sagawa et al.,, 2019, 2020; Hashimoto et al.,, 2018; Hu et al.,, 2018; Xiong et al.,, 2023; Zhang et al.,, 2020; Wang et al.,, 2023; Deng et al.,, 2020), and have garnered substantial attention recently due to the pervasive need for robustness in modern decision-making applications (Carmon and Hausler,, 2022; Asi et al.,, 2021; Haghtalab et al.,, 2022; Kar et al.,, 2019).
+Group DRO and DRL aim to develop algorithms that offer robust performance guarantees across a finite number of possible distributional models (sagawa2019distributionally; sagawa2020investigation; hashimoto2018fairness; hu2018does; xiong2023distributionally; zhang2020coping; wang2023distributionally; deng2020distributionally), and have garnered substantial attention recently due to the pervasive need for robustness in modern decision-making (carmon2022distributionally; asi2021stochastic; haghtalab2022demand; kar2019meta).
 When applying MDL to the context of group DRO/DRL, the resultant sample complexity reflects the price that needs to be paid for learning a robust solution.
-The MDL framework is also closely related to other topics like multi-source domain adaptation, maximum aggregation, to name just a few (Mansour et al.,, 2008; Zhao et al.,, 2020; Bühlmann and Meinshausen,, 2015; Guo,, 2023).
-In stark contrast to single-distribution learning, achieving data-efficient MDL necessitates adaptive sampling throughout the learning process, also known as on-demand sampling (Haghtalab et al.,, 2022).
-More specifically, pre-determining a sample-size budget for each distribution beforehand and sampling non-adaptively could result in a loss of sample efficiency, as we lack knowledge about the complexity of learning each distribution before the learning process begins.
-The question then boils down to how to optimally adapt the online sampling strategy to effectively tackle diverse data distributions.
+The MDL framework is also closely related to other topics like multi-source domain adaptation, maximum aggregation, to name just a few (mansour2008domain; zhao2020multi; buhlmann2015magging; guo2023statistical).
+In contrast to single-distribution learning, achieving data-efficient MDL necessitates adaptive sampling throughout the learning process, also known as on-demand sampling (haghtalab2022demand).
+More specifically, pre-determining a sample-size budget for each distribution beforehand and sampling non-adaptively could result in loss of sample efficiency, as we lack knowledge about the complexity of learning each distribution before the learning process begins.
+The question then comes down to how to optimally adapt the online sampling strategy to effectively tackle diverse data distributions.
 Inadequacy of prior results.
@@
 Fortunately, this sample size budget can be significantly reduced with the aid of adaptive sampling.
-In particular, the state-of-the-art approach, proposed by Haghtalab et al., (2022), accomplishes the objective (1) with probability $1-\delta$ using $O\big{(}\frac{\log(|\mathcal{H}|)+k\log(k/\delta)}{\varepsilon^{2}}\big{)}$ samples.
+In particular, the state-of-the-art approach, proposed by haghtalab2022demand, accomplishes the objective (1) with probability at least $1-\delta$ using $O\big{(}\frac{\log(|\mathcal{H}|)+k\log(k/\delta)}{\varepsilon^{2}}\big{)}$ samples.
 In comparison to agnostic learning on a single distribution, it only incurs an extra additive cost of $k\log(k/\delta)/\varepsilon^{2}$ as opposed to a multiplicative factor in $k$, thus underscoring the importance of adaptive sampling.
 A more challenging scenario arises when $\mathcal{H}$ has a finite Vapnik–Chervonenkis (VC) dimension $d$.
-The sample complexity for VC classes has only been settled for the reliazable case (Blum et al.,, 2017; Chen et al.,, 2018; Nguyen and Zakynthinou,, 2018), a special scenario where the loss function takes the form of $\ell\big{(}h,(x,y)\big{)}=\mathds{1}\{h(x)\neq y\}$ and it is feasible to achieve zero mean loss.
-For the general non-realizable case, the best-known lower bound for such VC classes is (Haghtalab et al.,, 2022) [^3] Here and throughout, $\widetilde{O}(\cdot)$ and $\widetilde{\Omega}(\cdot)$ hide all logarithimic factors in $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$.
+The sample complexity for VC classes has only been settled for the realizable case (blum2017collaborative; chen2018tight; nguyen2018improved), a special scenario where the loss function takes the form of $\ell\big{(}h,(x,y)\big{)}=\mathds{1}\{h(x)\neq y\}$ and it is feasible to achieve zero mean loss.
+For the general non-realizable case, the best-known lower bound for such VC classes is (haghtalab2022demand) [^3] Let $\mathcal{X}=\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$.
+Here and throughout, the notation $f(\mathcal{X})=O\big{(}g(\mathcal{X})\big{)}$ or $f(\mathcal{X})\lesssim g(\mathcal{X})$ (resp.
+$f(\mathcal{X})=\Omega\big{(}g(\mathcal{X})\big{)}$ or $f(\mathcal{X})\gtrsim g(\mathcal{X})$) mean that there exists some universal constant $C_{1}>0$ (resp.
+$C_{2}>0$) such that $f(\mathcal{X})\leq C_{1}g(\mathcal{X})$ (resp.
+$f(\mathcal{X})\geq C_{2}g(\mathcal{X})$); the notation $f(\mathcal{X})=\Theta\big{(}g(\mathcal{X})\big{)}$ or $f(\mathcal{X})\asymp g(\mathcal{X})$ mean $f(\mathcal{X})=O\big{(}g(\mathcal{X})\big{)}$ and $f(\mathcal{X})=\Omega\big{(}g(\mathcal{X})\big{)}$ hold simultaneously.
+The notation $\widetilde{O}(\cdot)$, $\widetilde{\Theta}(\cdot)$ and $\widetilde{\Omega}(\cdot)$ are defined analogously except that all log factors in $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$ are hidden.
 $\widetilde{\Omega}\left(\frac{d+k}{\varepsilon^{2}}\right),$ (2)
 which serves as a theoretical benchmark.
-By first collecting $\widetilde{O}\left(\frac{dk}{\varepsilon}\right)$ samples to help construct a cover of $\mathcal{H}$ with reasonable resolution, Haghtalab et al., (2022) established a sample complexity upper bound of
+By first collecting $\widetilde{O}\left(\frac{dk}{\varepsilon}\right)$ samples to help construct a cover of $\mathcal{H}$ with reasonable resolution, haghtalab2022demand established a sample complexity upper bound of
 […]
-Nevertheless, the term $dk/\varepsilon$ in (3a) fails to match the lower bound (2); put another way, this term might result in a potentially large burn-in cost, as the optimality of this approach is only guaranteed (up to log factors) when the total sample size already exceeds an enormous threshold on the order of $\frac{d^{2}k^{2}}{d+k}$.
-In an effort to alleviate this $dk/\varepsilon$ factor, Awasthi et al., (2023) put forward an alternative algorithm — which utilizes an oracle to learn on a single distribution and obliviates the need for computing an epsilon-net of $\mathcal{H}$ — yielding a sample complexity of
+Nevertheless, the term $dk/\varepsilon$ in (3a) fails to match the lower bound (2); put another way, this term might result in a potentially large burn-in cost, as the optimality of this approach is only guaranteed (up to log factors) when the total sample size already exceeds a (potentially large) threshold on the order of $\frac{d^{2}k^{2}}{d+k}$.
+In an effort to alleviate this $dk/\varepsilon$ factor, awasthi2023sample put forward an alternative algorithm — which utilizes an oracle to learn on a single distribution and obliviates the need for computing an epsilon-net of $\mathcal{H}$ — yielding a sample complexity of
 […]
@@
   • Question: Is it plausible to design a multi-distribution learning algorithm with a sample complexity of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ for VC classes, thereby matching the established lower bound (2)?
-Notably, this question has been posed as an open problem during the Annual Conference on Learning Theory (COLT) 2023; see Awasthi et al., (2023, Problem 1).
+Notably, this question has been posed as an open problem during the Annual Conference on Learning Theory (COLT) 2023; see awasthi2023sample.
 Paper Sample complexity bound
-Haghtalab et al., (2022) $\frac{\log(|\mathcal{H}|)+k}{\varepsilon^{2}}$
-Haghtalab et al., (2022) $\frac{d+k}{\varepsilon^{2}}+\frac{dk}{\varepsilon}$
-Awasthi et al., (2023) $\frac{d}{\varepsilon^{4}}+\frac{k}{\varepsilon^{2}}$
-Peng, (2023) $\frac{d+k}{\varepsilon^{2}}\left(\frac{k}{\epsilon}\right)^{o(1)}$
+haghtalab2022demand $\frac{\log(|\mathcal{H}|)+k}{\varepsilon^{2}}$
+haghtalab2022demand $\frac{d+k}{\varepsilon^{2}}+\frac{dk}{\varepsilon}$
+awasthi2023sample $\frac{d}{\varepsilon^{4}}+\frac{k}{\varepsilon^{2}}$
+peng2023sample $\frac{d+k}{\varepsilon^{2}}\left(\frac{k}{\varepsilon}\right)^{o(1)}$
 our work (Theorem 1) $\frac{d+k}{\varepsilon^{2}}$
-lower bound: Haghtalab et al., (2022) $\frac{d+k}{\varepsilon^{2}}$
+lower bound: haghtalab2022demand $\frac{d+k}{\varepsilon^{2}}$
 Table 1: Sample complexity bounds of MDL with $k$ data distributions and a hypothesis class of VC dimension $d$.
-Here, we only report the polynomial depedency and hide all logarithmic dependency on $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$.
+Here, we only report the polynomial dependency and hide all logarithmic dependency on $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$.
 A glimpse of our main contributions.
 The present paper delivers some encouraging news: we come up with a new MDL algorithm that successfully resolves the aforementioned open problem in the affirmative.
-Specifically, focusing on a hypothesis class with VC dimension $d$ and a collection of $k$ data distributions, our main findings can be summarized as follows.
+Specifically, focusing on a hypothesis class with VC dimension $d$ and a collection of $k$ data distributions, our main findings can be summarized in the following theorem.
+[^4] Following the definition of the VC dimension, the hypotheses in $\mathcal{H}$ are assumed to be binary-valued for VC classes.
 Theorem 1.
@@
 Following the game dynamics template adopted in previous methods — namely, viewing MDL as a game between the learner (who selects the best hypothesis) and the adversary (who chooses the most challenging mixture of distributions) — our algorithm is built upon a novel and meticulously designed sampling scheme that deviates significantly from previous methods.
-Further, we extend our algorithm and theory to accommodate Rademacher classes, establishing a similar sample complexity bound when the weighted Rademacher complexity of the hypothesis class on $n$ points is upper bounded by $O\big{(}\sqrt{\frac{d\log(n)}{n}}\big{)}$.
-Additionally, we solve two other open problems posed by Awasthi et al., (2023):
+Further, we extend our algorithm and theory to accommodate Rademacher classes, establishing a similar sample complexity bound; see Section 7 for details.
+Additionally, we solve two other open problems posed by awasthi2023sample:
   • Oracle-efficient solutions.
-An algorithm is said to be oracle-efficient if it only accesses $\mathcal{H}$ through an empirical risk minimization (ERM) oracle (Dudík et al.,, 2020).
-Awasthi et al., (2023, Problem 4) then asks what the sample complexity of MDL is when confined to oracle-efficient paradigms.
+An algorithm is said to be oracle-efficient if it only accesses $\mathcal{H}$ through an empirical risk minimization (ERM) oracle (dudik2020oracle).
+awasthi2023sample then asked what the sample complexity of MDL is when confined to oracle-efficient paradigms.
 Encouragingly, our algorithm (i.e., Algorithm 1) adheres to the oracle-efficient criterion, thus uncovering that the sample complexity of MDL remains unchanged when restricted to oracle-efficient algorithms.
@@
 Both our algorithm and the most sample-efficient methods preceding our work produce randomized hypotheses.
-As discussed around Awasthi et al., (2023, Problem 3), a natural question concerns characterization of the sample complexity when restricting the final output to deterministic hypotheses from $\mathcal{H}$.
+As discussed around awasthi2023sample, a natural question concerns characterization of the sample complexity when restricting the final output to deterministic hypotheses from $\mathcal{H}$.
 Our result (see Theorem 2) delivers a negative message: under mild conditions, for any MDL algorithm, there exists a hard problem instance such that it requires at least $\Omega(dk/\varepsilon^{2})$ samples to find a deterministic hypothesis $h\in\mathcal{H}$ that attains $\varepsilon$-accuracy.
-This constitutes a sample size barrier that lies far above what is achievable under randomized hypotheses.
+This constitutes an enormous sample complexity gap between what is achievable under randomized hypotheses and what is achievable using deterministic hypotheses.
 Concurrent work.
-While finalizing our paper, we became aware of a concurrent work Peng, (2023), which also studied the MDL problem and significantly improved upon the prior results.
-More specifically, Peng, (2023) established a sample complexity of $O\left(\frac{(d+k)\log(d/\delta)}{\varepsilon^{2}}\left(\frac{k}{\varepsilon}\right)^{o(1)}\right)$, which is optimal up to some sub-polynomial factor in $k/\varepsilon$; in comparison, our sample complexity is optimal only up to some logarithmic factor.
+We shall mention that a concurrent work peng2023sample, posted around the same time as our work, also studied the MDL problem and significantly improved upon the prior results.
+More specifically, peng2023sample established a sample complexity of $O\left(\frac{(d+k)\log(d/\delta)}{\varepsilon^{2}}\left(\frac{k}{\varepsilon}\right)^{o(1)}\right)$, which is optimal up to some sub-polynomial factor in $k/\varepsilon$; in comparison, our sample complexity is optimal up to polylogarithmic factor.
 Additionally, it is worth noting that the algorithm therein relies upon a certain recursive structure to eliminate the non-optimal hypothesis, thus incurring exponential computational cost even when an ERM oracle is available.
@@
 Also we say $v\leq v^{\prime}$ iff $v_{i}\leq v^{\prime}_{i}$ for all $i\in[n]$.
-For any random variable $X$, we use $\mathbb{V}[X]$ to denote its variance, i.e., $\mathbb{V}[X]=\mathbb{E}\left[(X-\mathbb{E}[X])^{2}\right]$.
+For any random variable $X$, we use $\mathsf{Var}[X]$ to denote its variance, i.e., $\mathsf{Var}[X]=\mathbb{E}\left[(X-\mathbb{E}[X])^{2}\right]$.
+Let $\mathrm{e}^{\mathsf{basis}}_{1},\dots,\mathrm{e}^{\mathsf{basis}}_{k}$ represent the standard basis vectors in $\mathbb{R}^{k}$.
 For any two distributions $P$ and $Q$ supported on $\mathcal{X}$, the Kullback-Leibler (KL) divergence from $Q$ to $P$ is defined and denoted by
@@
 We assume throughout that each datapoint takes the form of $(x,y)\in\mathcal{X}\times\mathcal{Y}$, with $\mathcal{X}$ (resp.
-$\mathcal{Y}$) representing the instance space (resp.
+$\mathcal{Y}$) the instance space (resp.
 label space).
 Learning from multiple distributions.
-The problem setting encompasses several elements as follows.
+The problem setting encompasses several elements below.
   • Hypothesis class.
-Suppose that we are interested in a hypothesis class $\mathcal{H}$, consisting of a set of candidate functions from the instance space $\mathcal{X}$ to the label space $\mathcal{Y}$.
+Suppose we are interested in a hypothesis class $\mathcal{H}$, comprising a set of candidate functions from the instance space $\mathcal{X}$ to the label space $\mathcal{Y}$.
 Overloading the notation, we use $h_{\pi}$ to represent a randomized hypothesis associated with a probability distribution $\pi\in\Delta(\mathcal{H})$, meaning that a hypothesis $h$ from $\mathcal{H}$ is randomly selected according to distribution $\pi$.
-Additionally, the Vapnik–Chervonenkis (VC) dimension (Vapnik et al.,, 1994) of $\mathcal{H}$ is assumed to be
-    $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})=d;$ (6)
-    see Shalev-Shwartz and Ben-David, (2014) for an introduction of the VC dimension.
+When it comes to a VC class, we assume that the hypotheses in $\mathcal{H}$ are binary-valued, and that the VC dimension (vapnik1994measuring) of $\mathcal{H}$ is
+    $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})=d.$ (6)
   • Loss function.
@@
 3 Algorithm
-In this section, we present our proposed algorithm.
+In this section, we present our proposed algorithm for learning VC classes.
 Before proceeding, we find it convenient to introduce some notation concerning the loss under mixed distributions.
@@
 1 input: $k$ data distributions $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, hypothesis class $\mathcal{H}$, target accuracy level $\varepsilon$, target success rate $1-\delta$.
-2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}:=\frac{400\left(k\log(\frac{k}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}.$
+2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}\coloneqq\frac{4000\left(k\log(k/\varepsilon_{1})+d\log(kd/\varepsilon_{1})+\log(1/\delta)\right)}{\varepsilon_{1}^{2}}.$
 3 initialization: for all $i\in[k]$, set $W^{1}_{i}=1$, $\widehat{w}^{0}_{i}=0$ and $n_{i}^{0}=0$; $\mathcal{S}=\emptyset$.
@@
 5       set $w^{t}=[w^{t}_{i}]_{1\leq i\leq k}$ and $\widehat{w}^{t}=[\widehat{w}^{t}_{i}]_{1\leq i\leq k}$, with $w^{t}_{i}\leftarrow\frac{W^{t}_{i}}{\sum_{j}W^{t}_{j}}$ and $\widehat{w}^{t}_{i}\leftarrow\widehat{w}^{t-1}_{i}$ for all $i\in[k]$.
-       /* recompute $\widehat{w}^{t}$ draw new samples for $\mathcal{S}_{\mathrm{w}}$ only if $w^{t}$ changes sufficiently. */
+       /* recompute $\widehat{w}^{t}$ draw new samples for $\mathcal{S}$ only if $w^{t}$ changes sufficiently. */
 6       if there exists $j\in[k]$ such that $w_{j}^{t}\geq 2\widehat{w}_{j}^{t-1}$  then
@@
       /* estimate the near-optimal hypothesis for weighted data distributions. */
-12       compute $h^{t}\leftarrow\arg\min_{h\in\mathcal{H}}\widehat{L}(h,w^{t})$, where
-$\widehat{L}^{t}(h,w^{t}):=\sum_{i=1}^{k}\frac{w^{t}_{i}}{n_{i}^{t}}\cdot\sum_{j=1}^{n_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (9)
+12       compute $h^{t}\leftarrow\arg\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})$, where
+$\widehat{L}^{t}(h,w^{t})\coloneqq\sum_{i=1}^{k}\frac{w^{t}_{i}}{n_{i}^{t}}\cdot\sum_{j=1}^{n_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (9)
 with $(x_{i,j},y_{i,j})$ being the $j$-th datapoint from $\mathcal{D}_{i}$ in $\mathcal{S}$.
@@
 14       for $i=1,\ldots,k$ do
-             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
+15             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
 $\widehat{r}^{t}_{i}=\frac{1}{\lceil k\overline{w}^{t}_{i}\rceil}\sum_{j=1}^{\lceil k\overline{w}^{t}_{i}\rceil}\ell\big{(}h^{t},(x_{i,j}^{t},y_{i,j}^{t})\big{)};$
             update the weight as $W^{t+1}_{i}=W^{t}_{i}\exp(\eta\widehat{r}_{i}^{t})$. // Hedge updates.
-15
 16
-output: a randomized hypothesis $h^{\mathsf{final}}$ as a uniform distribution over $\{h^{t}\}_{t=1}^{T}$.
-Algorithm 1 Hedge for Multi-distribution Learning ($\mathtt{MDL-Hedge}$)
-Following the game dynamics proposed in previous works (Awasthi et al.,, 2023; Haghtalab et al.,, 2022), our algorithm alternates between computing the most favorable hypothesis (performed by the learner) and estimating the most challenging mixture of data distributions (performed by the adversary), with the aid of no-regret learning algorithms (Roughgarden,, 2016; Shalev-Shwartz,, 2012).
+17
+output: a randomized hypothesis $h^{\mathsf{final}}$ uniformly distributed over $\{h^{t}\}_{t=1}^{T}$.
+Algorithm 1 Hedge for multi-distribution learning on VC classes ($\mathtt{MDL}\text{-}\mathtt{Hedge}\text{-}\mathtt{VC}$)
+Remark 1.
+Note that in Algorithm 1, the quantities $\{n_{i}^{t}\}_{1\leq t\leq T}$ are designed to be non-decreasing in $t$.
+Remark 2.
+In line 1 of Algorithm 1, we also allow $h^{t}$ to be $\varepsilon_{1}$-best response that obeys $\widehat{L}^{t}(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})+\varepsilon_{1}$; our theory remains unchanged if we make this modification.
+Following the game dynamics proposed in previous works (awasthi2023sample; haghtalab2022demand), our algorithm alternates between computing the most favorable hypothesis (performed by the learner) and estimating the most challenging mixture of data distributions (performed by the adversary), with the aid of no-regret learning algorithms (roughgarden2016twenty; shalev2012online).
 More specifically, in each round $t$, our algorithm performs the following two steps:
@@
     Given hypothesis $h^{t}$, we compute an updated weight vector $w^{t+1}\in\Delta(k)$ — and hence an updated mixed distribution $\mathcal{D}^{(t+1)}=\sum_{i\in[k]}w_{i}^{t+1}\mathcal{D}_{i}$.
-The weight updates are carried out using the celebrated Hedge algorithm (Freund and Schapire,, 1997) designed for online adversarial learning, [^4] Note that the Hedge algorithm is closely related to Exponentiated Gradient Descent, Multiplicative Weights Update, Online Mirror Descent, etc (Arora et al.,, 2012; Shalev-Shwartz,, 2012; Hazan,, 2022).
+The weight updates are carried out using the celebrated Hedge algorithm (freund1997decision) designed for online adversarial learning, [^5] Note that the Hedge algorithm is closely related to Exponentiated Gradient Descent, Multiplicative Weights Update, Online Mirror Descent, etc (arora2012multiplicative; shalev2012online; hazan2022introduction).
 in an attempt to achieve low regret even when the loss vectors are adversarially chosen.
@@
     computed over another set of data samples.
+In words, the min-player and the max-player in our algorithm follow the best-response dynamics and no-regret dynamics, respectively (note that the Hedge algorithm is known to be a no-regret algorithm (roughgarden2016twenty)).
 At the end of the algorithm, we output a randomized hypothesis $h^{\sf final}$ that is uniformly distributed over the hypothesis iterates $\{h^{t}\}_{1\leq t\leq T}$ over all $T$ rounds, following common practice in online adversarial learning.
-While the above paradigm has been adopted in past works (Awasthi et al.,, 2023; Haghtalab et al.,, 2022), the resulting sample complexity depends heavily upon how data samples are collected and utilized throughout the learning process.
-For instance, Awasthi et al., (2023, Algorithm 1) draws fresh data at each step of every round, in order to ensure reliable estimation of the loss function of interest through elementary concentration inequalities.
-This strategy, however, becomes wasteful over time, constituting the main source of its sample sub-optimality.
+While the above paradigm has been adopted in past works (awasthi2023sample; haghtalab2022demand), the resulting sample complexity depends heavily upon how data samples are collected and utilized throughout the learning process.
+For instance, awasthi2023sample — which also alternates between best response and no-regret algorithm — draws fresh data at each step of every round, in order to ensure reliable estimation of the loss function of interest through elementary concentration inequalities.
+This strategy, however, becomes wasteful over time, constituting one of the main sources of its sample sub-optimality.
 In order to make the best use of data, we propose the following key strategies.
   • Sample reuse in Step (a).
-In stark contrast to Awasthi et al., (2023, Algorithm 1) that draws new samples for estimating each $h^{t}$, we propose to reuse all samples collected in Step (a) up to the $t$-th round to assist in computing $h^{t}$.
+In stark contrast to awasthi2023sample that draws new samples for estimating each $h^{t}$, we propose to reuse all samples collected in Step (a) up to the $t$-th round to assist in computing $h^{t}$.
 As will be made precise in lines 1-1 of Algorithm 1, we shall maintain a growing dataset $\mathcal{S}$ for conducting Step (a) throughout, ensuring that there are $n_{i}^{t}$ samples drawn from distribution $\mathcal{D}_{i}$ in the $t$-th round.
@@
 More precisely, we collect $\left\lceil k\overline{w}_{i}^{t}\right\rceil$ fresh samples from each $\mathcal{D}_{i}$, where $\overline{w}_{i}^{t}:=\max_{1\leq\tau\leq t}w_{i}^{\tau}$ is the maximum weight assigned to $\mathcal{D}_{i}$ up to now.
-Informally speaking, this strategy ensures reduced variance of the estimators and ultimately allows for an improved bound for $\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$.
+Informally speaking, this strategy is chosen carefully to ensure reduced variance of the estimators given a sample size budget.
 The interested reader is referred to Section 4.2 and Lemma 17 for more detailed explanations.
 The whole procedure can be found in Algorithm 1.
-4 Technique overview
-In this section, we introduce the high-level ideas of the algorithm and analysis.
-Since Haghtalab et al., (2022) provided a nearly matching upper bound when $k=O(1/\varepsilon)$ (See the second row in Table 1), we assume that $k\geq\frac{100}{\varepsilon}$ throughout this work.
-We discuss the two primary technical novelties of the uniform convergence of the importance sampling estimator that allows for reusing samples (Section 4.1), and the upper bound of the coordinate-wise maximal $\ell_{1}$ norm of the $\{w^{t}\}$ iterates (Section 4.2).
-4.1 Uniformly-convergent estimators
-Let $\mathcal{X}$ be a proper discretization of $\Delta(k)$ with size $\exp(O(k\log(kd/\varepsilon)))$.
-We consider to study the approximation of the natural importance sampling estimator.
-Fix $n=\{n_{i}\}_{i=1}^{k}$ where each $n_{i}$ is a positive integer.
-Suppose that we have $n_{i}$ i.i.d.
-samples from $\mathcal{D}_{i}$ for each $i\in[k]$, denoted by $\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}$.
-For $w=[w_{i}]_{1\leq i\leq k}\in\mathcal{X}$, we let
-$\widehat{L}_{n}(h,w)\coloneqq\sum_{i=1}^{k}w_{i}\cdot\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}.$ (12)
-In view of the concentration property of VC classes (Mohri et al.,, 2018), we have
-$\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\exp\left(\lambda\cdot\max_{h\in\mathcal{H}}\left|\widehat{L}_{n}(h,w)-L(h,w)\right|\right)\right]$
-$\leq 2\cdot(2\|n\|_{1})^{d+1}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right)$ (13)
-for all $0\leq\lambda\leq\min_{i}\frac{n_{i}}{w_{i}}$. As a result, for any $\kappa\geq 0$ and $0\leq\lambda\leq\kappa$ one has
-$\sum_{w\in\mathcal{X}}\mathds{1}\left\{\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\exp\left(\lambda\cdot\max_{h\in\mathcal{H}}\left|\widehat{L}_{n}(h,w)-L(h,w)\right|\right)\right]$
-$\leq\sum_{w\in\mathcal{X}}\mathds{1}\left\{\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}\cdot 2\cdot(2\|n\|_{1})^{d+1}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right)$
-$\leq\sum_{w\in\mathcal{X}}\mathds{1}\left\{\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}\cdot 2\cdot(2\|n\|_{1})^{d+1}\exp\left(\frac{2\lambda^{2}}{\kappa}\right)$
-$\leq 2\cdot|\mathcal{X}|\cdot(2\|n\|_{1})^{d+1}\exp\left(\frac{2\lambda^{2}}{\kappa}\right).$ (14)
-We can then readily see that: for any $\varepsilon^{\prime}\in(0,1)$,
-$\mathbb{P}\left\{\exists w\in\mathcal{X},\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa,\max_{h\in\mathcal{H}}\left|\widehat{L}_{n}(h,w)-L(h,w)\right|\geq\varepsilon^{\prime}\right\}$
-$\leq 2\cdot|\mathcal{X}|\cdot(2\|n\|_{1})^{d+1}\min_{0\leq\lambda\leq\kappa}\exp\left(-\lambda\varepsilon^{\prime}+\frac{2\lambda^{2}}{\kappa}\right)$
-$\leq 2\cdot|\mathcal{X}|\cdot(2\|n\|_{1})^{d+1}\exp\left(-\frac{\kappa(\varepsilon^{\prime})^{2}}{4}\right).$ (15)
-Choose $\varepsilon^{\prime}=c\varepsilon$ for some proper constant $c$.
-By virtue of (15), for any $\delta\in(0,1)$, we can choose $\kappa\geq N_{0}(\varepsilon,\delta):=100\frac{k\log(kd/\varepsilon)+d\log(\|n\|_{1})+\log(1/\delta)}{\varepsilon^{2}}$ to ensure (15) is bounded by $\delta$.
-In other words, with probability at least $1-\delta$, $|\widehat{L}_{n}(h,w)-L(h,w)|\leq c\varepsilon$ holds uniformly for all $w\in\mathcal{X}$ such that $\max_{i}\frac{w_{i}}{n_{i}}\leq\frac{1}{N_{0}(\varepsilon,\delta)}$.
-In the analysis above, we fix $n$ and show that uniform convergence holds with high probability.
-In fact, with similar arguments we can obtain uniform convergence over all proper choices of $(n,w)$.
-Let
-[…]
-$\mathcal{L}(N,\kappa)$$\coloneqq\left\{(w,n)\mid w\in\mathcal{X},n\in\mathcal{N}(N),\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}.$ (16b)
-It can then be shown that for any $\varepsilon^{\prime}\in(0,1)$ and $\kappa>0$,
-$\mathbb{P}\left[\exists(w,n)\in\mathcal{L}(N,\kappa),\left|\widehat{L}_{h}(h,w)-L(h,w)\right|\geq\varepsilon^{\prime}\right]$
-$\leq 2\cdot|\mathcal{N}(N)|\cdot|\mathcal{X}|\cdot(2N)^{d+1}\exp\left(-\frac{\kappa(\varepsilon^{\prime})^{2}}{4}\right).$
-Since $|\mathcal{N}(N)|\leq N^{k}$, it suffices to choose $\kappa=N^{*}\coloneqq O\left(\frac{(k+d)\mathrm{polylog}(k,d,1/\varepsilon)+\log(1/\delta)}{\varepsilon^{2}}\right)$ to ensure that the $O(\varepsilon)$-uniform convergence holds with probability at least $1-\delta$.
-With the arguments above, with probability at least $1-\delta$, whenever we collect $n=\{n_{i}\}_{i=1}^{k}$ samples in the learning process, we could obtain $\varepsilon$-approximation $\widehat{L}_{n}(h,w)$ (see (12)) of $L(h,w)$ for all $h\in\mathcal{H}$ and $w\in\Delta(k)$ such that $\min_{i}\frac{n_{i}}{w_{i}}\geq N^{*}$.
-Most importantly, this result allows us to reuse samples in the learning process, instead of querying new samples to estimate $L(h,w^{t})$ in each round $t$.
-Recall that $\overline{w}^{t}_{i}=\max_{1\leq\tau\leq t}w_{i}^{\tau}$.
-To ensure the $O(\varepsilon)$-uniform convergence with respect to $w^{t}$ in each round $t$, it suffices to choose $n_{i}^{t}\geq N^{*}\overline{w}_{i}^{t}$ for all $i\in[k]$.
-4.2 Measuring the trajectory of Hedge
-Our second technical point is to bound the maximal point of the whole trajectory $\{w^{t}\}_{t=1}^{T}$ of the Hedge algorithm. That is, we aim to control
-$\|\overline{w}^{T}\|_{1}=\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}.$
-To explain the high-level idea, we consider the learning process in Algorithm 2 with full knowledge of the loss vector.
+4 A glimpse of key technical novelties
+In this section, we highlight two technical novelties that empower our analysis: (i) uniform convergence of the weighted sampling estimator that allows for sample reuse (see Section 4.1), and (ii) tight control of certain $\|\cdot\|_{1,\infty}$ norm of the iterates $\{w^{t}\}_{1\leq t\leq T}$ that dictates the sample efficiency (see Section 4.2).
+4.1 Towards sample reuse: uniform concentration and a key quantity
+Recall that in Algorithm 1, we invoke the empirical risk estimator $\widehat{L}^{t}(h,w^{t})$ as an estimate of the true risk of hypothesis $h$ over the weighted distribution specified by $w^{t}$ (cf.
+(9)).
+In order to facilitate sample reuse when constructing such risk estimators across all iterations, it is desirable to establish uniform concentration results to control the errors of such risk estimators throughout the execution of the algorithm.
+Towards this end, our analysis strategy proceeds as follows.
+Step 1: concentration for any fixed set of parameters.
+Consider any given set of integers $n=\{n_{i}\}_{i=1}^{k}$ and any given vector $w\in\Delta(k)$.
+Suppose, for each $i\in[k]$, we have $n_{i}$ i.i.d.
+samples drawn from $\mathcal{D}_{i}$ — denoted by $\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}$ — and let us look at the empirical risk estimator,
+$\widehat{L}_{n}(h,w)\coloneqq\sum\nolimits_{i=1}^{k}w_{i}\cdot\frac{1}{n_{i}}\sum\nolimits_{j=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)},$ (12)
+which is a sum of independent random variables. Evidently, for a given hypothesis $h$, the variance of $\widehat{L}_{n}(h,w)$ is upper bounded by
+$\mathsf{Var}\big{(}\widehat{L}_{n}(h,w)\big{)}\leq\sum\nolimits_{i=1}^{k}\frac{w_{i}^{2}}{n_{i}}\leq\left(\sum\nolimits_{i=1}^{k}w_{i}\right)\frac{1}{\min_{i}n_{i}/w_{i}}=\frac{1}{\min_{i}n_{i}/w_{i}}.$
+Assuming that the central limit theorem is applicable, one can derive
+$\mathbb{P}\left\{\big{|}\widehat{L}_{n}(h,w)-L(h,w)\big{|}\geq\varepsilon\right\}\lesssim\exp\bigg{(}-\frac{\varepsilon^{2}}{2\mathsf{Var}\big{(}\widehat{L}_{n}(h,w)\big{)}}\bigg{)}\lesssim\exp\bigg{(}-\frac{\varepsilon^{2}}{2}\min_{i}\frac{n_{i}}{w_{i}}\bigg{)}.$
+Armed with this result, we can extend it to accommodate all $h\in\mathcal{H}$ through the union bound.
+For a VC class with $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})=d$, the celebrated Sauer–Shelah lemma (wainwright2019high, Proposition 4.18) tells us that the set of hypotheses can be effectively compressed into a subset with cardinality no larger than $\exp\big{(}\widetilde{O}(d)\big{)}$.
+Taking the union bound then yields
+$\mathbb{P}\left(\max_{h\in\mathcal{H}}\big{|}\widehat{L}_{n}(h,w)-L(h,w)\big{|}\geq\varepsilon\right)\lesssim\exp\left(\widetilde{O}(d)-\frac{\varepsilon^{2}}{2}\min_{i}\frac{n_{i}}{w_{i}}\right).$
+Step 2: uniform concentration.
+Next, we would like to extend the above result to establish uniform concentration over all $n$ and $w$ of interest.
+Towards this, we shall invoke the union bound as well as the standard epsilon-net arguments.
+Let the set $\mathcal{X}\subseteq\Delta(k)$ be a proper discretization of $\Delta(k)$, with cardinality $\exp\big{(}\widetilde{O}(k)\big{)}$.
+In addition, given the trivial upper bound $n_{i}\leq T_{1}$ for all $i\in[k]$, we know that there exist at most $T_{1}^{k}=\exp\big{(}\widetilde{O}(k)\big{)}$ possible combinations of $\{n_{i}\}_{i\in[k]}$.
+We can then apply the union bound to show that
+$\mathbb{P}\left\{\exists w\in\mathcal{X}\text{ and feasible }n\text{ s.t.}\big{|}\widehat{L}_{n}(h,w)-L(h,w)\big{|}\geq\varepsilon\right\}$ $\lesssim\exp\bigg{(}\widetilde{O}(k)+\widetilde{O}(d)-\frac{\varepsilon^{2}}{2}\min_{i}\frac{n_{i}}{w_{i}}\bigg{)}.$ (13)
+When the discretized set $\mathcal{X}$ is chosen to have sufficient resolution, we can straightforwardly employ the standard covering argument to extend the above inequality to accommodate all $w\in\Delta(k)$ of interest.
+Key takeaways.
+The above arguments reveal the following high-probability property: whenever we collect $n=\{n_{i}\}_{i=1}^{k}$ samples in the learning process, we could obtain $\varepsilon$-approximation $\widehat{L}_{n}(h,w)$ (see (12)) of $L(h,w)$ for all $h\in\mathcal{H}$ and all $w\in\Delta(k)$ with high probability, provided that
+$\min_{i}\frac{n_{i}}{w_{i}}\gtrsim\widetilde{O}\bigg{(}\frac{k+d}{\varepsilon^{2}}\bigg{)}.$ (14)
+This makes apparent the pivotal role of the quantity $\min_{i}n_{i}/w_{i}$. In our algorithm, we design the update rule (cf. line 1 of Algorithm 1) to guaranteed that
+$\min_{i}\frac{n_{i}^{t}}{w_{i}^{t}}\gtrsim T_{1}\geq\widetilde{\Omega}\bigg{(}\frac{k+d}{\varepsilon^{2}}\bigg{)}$ (15)
+for all $1\leq t\leq T$.
+In fact, this explains our choice of $T_{1}$ in Algorithm 1.
+Crucially, the aforementioned uniform concentration result allows us to reuse samples throughout the learning process instead of drawing fresh samples to estimate $L(h,w^{t})$ in each round $t$ (note that the latter approach clearly falls short of data efficiency).
+To conclude, to guarantee $\varepsilon$-uniform convergence for all rounds, it suffices to choose $T_{1}=\widetilde{\Omega}\big{(}\frac{k+d}{\varepsilon^{2}}\big{)}$.
+Finally, recall that $n_{i}^{t}\asymp T_{1}\overline{w}_{i}^{t}$ for each $i\in[k]$ and $t\leq T$, with $\overline{w}^{t}_{i}\coloneqq\max_{1\leq\tau\leq t}w_{i}^{\tau}$; taking $n_{i}^{t}\asymp T_{1}\overline{w}_{i}^{t}$ (as opposed to $n_{i}^{t}\asymp T_{1}w_{i}^{t}$) ensures that the sample size $n_{i}^{t}$ is monotonically non-decreasing in $t$.
+With (15) in mind, the total number of samples collected within $T$ rounds in Algorithm 1 obeys
+$\frac{1}{T_{1}}\sum\nolimits_{i=1}^{k}n_{i}^{T}\asymp\sum\nolimits_{i=1}^{k}\overline{w}_{i}^{T}\eqqcolon\|\overline{w}^{T}\|_{1}.$ (16)
+This threshold $\|\overline{w}^{T}\|_{1}$ — or equivalently, the $\|\cdot\|_{1,\infty}$ norm of $\{w^{t}\}_{1\leq t\leq T}$ — is a critical quantity that we wish to control.
+In particular, in the desirable scenario where $\|\overline{w}^{T}\|_{1}\leq\widetilde{O}(1)$, the total sample size obeys $\sum\nolimits_{i=1}^{k}n_{i}^{T}\asymp T_{1}\|\overline{w}^{T}\|_{1}=\widetilde{O}\big{(}\frac{k+d}{\varepsilon^{2}}\big{)}$.
+4.2 Bounding the key quantity $\|\overline{w}^{T}\|_{1}$ by tracking the Hedge trajectory
+Perhaps the most innovative (and most challenging) part of our analysis lies in controlling the $\|\cdot\|_{1,\infty}$ norm of $\{w_{i}^{t}\}_{1\leq t\leq T}$, whose critical importance has been pointed out in Section 4.1.
+Towards this end, the key lies in carefully tracking the dynamics of the Hedge algorithm.
+To elucidate the high-level idea, let us look at a simpler minimax optimization problem w.r.t.
+the set of loss vectors in the convex hull of a set $\mathcal{Y}\subseteq\mathbb{R}^{k}$:
+$\min_{y\in\mathsf{conv}(\mathcal{Y})}\max_{w\in\Delta(k)}w^{\top}y\qquad(\text{or equivalently, }\max_{w\in\Delta(k)}\min_{y\in\mathsf{conv}(\mathcal{Y})}w^{\top}y),$ (17)
+where the equivalence arises from von Neumann’s minimax theorem (v1928theorie).
+Consider the following algorithm (cf.
+Algorithm 2) tailored to this minimax problem, assuming perfect knowledge about the loss vectors.
+[^6] Note that in Algorithm 1, we can only estimate the loss vector using the collected samples.
+Additional efforts are needed to reduce the variability (see line 1 in Algorithm 1).
 1 Input: $\mathcal{Y}\subseteq[-1,1]^{k}$, target accuracy level $\varepsilon\in(0,1)$.
-2 Initialization: $T=\frac{100\log(k)}{\varepsilon^{2}}$, $\eta=\frac{1}{10}\varepsilon$, $W^{1}_{i}=1$ for all $1\leq i\leq k$.
+2 Initialization: $T=\frac{100\log(k)}{\varepsilon^{2}}$, $\eta=\frac{1}{10}\varepsilon$, and $W^{1}_{i}=1$ for all $1\leq i\leq k$.
 3 for $t=1,2,\ldots,T$ do
-4       $w_{i}^{t}\leftarrow\frac{W_{i}^{t}}{\sum_{j}W_{j}^{t}}$ for all $1\leq i\leq k$.
-5       $y^{t}\leftarrow\arg\min_{y\in\mathcal{Y}}(w^{t})^{\top}y$.
-6       Update $W_{i}^{t+1}\leftarrow W_{i}^{t}\cdot\exp(\eta y_{i}^{t})$.
+4       compute $w_{i}^{t}\leftarrow\frac{W_{i}^{t}}{\sum_{j}W_{j}^{t}}$ for every $1\leq i\leq k$.
+5       compute $y^{t}\leftarrow\arg\min_{y\in\mathcal{Y}}~{}\langle w^{t},y\rangle$.
+6       update $W_{i}^{t+1}\leftarrow W_{i}^{t}\exp(\eta y_{i}^{t})$ for every $1\leq i\leq k$.
+7
 Algorithm 2 The Hedge algorithm for bilinear games.
-As presented in Algorithm 2, there is a set of loss vectors $\mathcal{Y}$, and we want to solve the minimax problem
-$\min_{y\in\mathsf{conv}(\mathcal{Y})}\max_{w\in\Delta(k)}w^{\top}y\qquad(\text{or equivalently, }\max_{w\in\Delta(k)}\min_{y\in\mathsf{conv}(\mathcal{Y})}w^{\top}y),$ (17)
-where the equivalence arises from von Neumann’s minimax theorem (v.
-Neumann,, 1928).
-By Hedge algorithm, we can obtain an $O(\varepsilon)$-minimax solution in $\frac{100\log(k)}{\varepsilon^{2}}$ rounds.
-The question is that:
-  • Can we bound $\|\overline{w}^{T}\|_{1}=\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ by polylogarithmic terms?
-To explain our two key techniques, we assume that
-$\min_{y\in\mathsf{conv}(\mathcal{Y})}\max_{w\in\Delta(k)}w^{\top}y=\max_{w\in\Delta(k)}\min_{y\in\mathsf{conv}(\mathcal{Y})}w^{\top}y=0,$
-Given this assumption, we then have $(w^{t})^{\top}y^{t}=\min_{y\in\mathcal{Y}}(w^{t})^{\top}y\leq 0,\forall t\in[T]$ and $\frac{1}{T}\sum_{t=1}^{T}(w^{t})^{\top}y^{t}\geq-C\epsilon$ for some constant $C$.
-In the following analysis, we assume $-(w^{t})^{\top}y^{t}=O(\epsilon)$ holds for any $t\in[T]$.
-While this assumption is not true in most cases, we can divide $[T]$ into $j_{\mathrm{max}}:=O\left(\log_{2}(1/\epsilon)\right)$ disjoint subsets as $\{\{t\in[T]:-(w^{t})^{\top}y^{t}\in(2^{-(j+1)},2^{j}]\}\}_{j=0}^{j_{\mathrm{max}}-2}$ and $\{t\in[T]:-(w^{t})^{\top}y^{t}\leq 2^{-(j_{\mathrm{max}}-1)}=O(\epsilon)\}$ and then work on each subset.
-As a trade-off, this operation leads to some additional logarithmic factors (see Lemma 29).
-4.2.1 Doubling $w_{j}$ needs $\widetilde{\Omega}(1/\varepsilon^{2})$ steps
-Instead of bounding $\|\overline{w}^{T}\|_{1}$ directly, our first attempt is to show that there are at most $\widetilde{O}(1)$ dimensions $i$ such that $\max_{1\leq t\leq T}w^{t}_{i}\geq c$, with $c>0$ some universal constant.
-For simplicity, we take $c=1/4$.
-Let
-$\mathcal{W}\coloneqq\Big{\{}i\in[k]\mid\max_{1\leq t\leq T}w_{i}^{t}\geq 1/4\big{\}}.$ (18)
-Given that the stepsize $\eta$ is small enough, one can find, for each $i\in[k]$, an interval $[s_{i},e_{i}]$ such that
-$\frac{1}{16}\leq w_{i}^{s_{i}}\leq\frac{1}{8},\qquad w_{i}^{e_{i}}\geq\frac{1}{4}\qquad\text{and}\qquad w_{i}^{t}\geq\frac{1}{8}$ (19)
-for any $t$ obeying $s_{i}<t\leq e_{i}$. Then we would like to show that $e_{i}-s_{i}\geq\widetilde{\Omega}(1/\varepsilon^{2})$.
-Since the stepsize $\eta=O(\varepsilon)$, we can easily get that $e_{i}-s_{i}=\widetilde{\Omega}(1/\varepsilon)$.
-Meanwhile, when running the Hedge algorithm for adversarial bandits, the adversarial could choose the loss properly so that $e_{i}-s_{i}=\widetilde{\Theta}(1/\varepsilon)$.
-Fortunately, we can focus on a bilinear game where the opponent plays best response.
-By definition, for any $1\leq t\leq\tau\leq T$,
-$(w^{t})^{\top}y^{t}\leq 0,\quad(w^{t})^{\top}y^{\tau}\geq(w^{t})^{\top}y^{t}.$
-For any $1\leq t_{1}\leq t_{2}\leq T$, we then have
-$(w^{t_{1}})^{\top}\sum_{\tau=t_{1}}^{t_{2}-1}y^{\tau}\geq(t_{2}-t_{1})(w^{t_{1}})^{\top}y^{t_{1}}.$ (20)
-By observing that $\sum_{\tau=t_{1}}^{t_{2}-1}y_{i}^{\tau}=\frac{1}{\eta}\log\left(\frac{W^{t_{2}}_{i}}{W^{t_{1}}_{i}}\right)$, we have
-$\sum_{i=1}^{k}w_{i}^{t_{1}}\log\left(\frac{W_{i}^{t_{2}}}{W_{i}^{t_{1}}}\right)\geq\eta(t_{2}-t_{1})(w^{t_{1}})^{\top}y^{t_{1}}.$ (21)
-On the other hand, since the optimal minimax value is 00, we can show that
-$\log\left(\frac{\sum_{i=1}^{k}W_{i}^{t_{2}}}{\sum_{i=1}^{k}W_{i}^{t_{1}}}\right)=\sum_{\tau=t_{1}}^{t_{2}-1}\log\left(\frac{\sum_{i=1}^{k}W_{i}^{\tau+1}}{\sum_{i=1}^{k}W_{i}^{\tau}}\right)\leq O\big{(}(t_{2}-t_{1})\eta^{2}\big{)}.$ (22)
-Putting (21) and (22) together, we arrive at
-$-\sum_{i=1}^{k}w_{i}^{t_{1}}\log\left(\frac{w_{i}^{t_{2}}}{w_{i}^{t_{1}}}\right)\leq O\left(\eta^{2}(t_{2}-t_{1})-\eta(t_{2}-t_{1})(w^{t_{1}})^{\top}y^{t_{1}}\right)=O(\eta^{2}(t_{2}-t_{1})),$ (23)
-which is exactly
-$\mathsf{KL}\big{(}w^{t_{1}}\,\|\,w^{t_{2}}\big{)}\leq O\big{(}\eta^{2}(t_{2}-t_{1})\big{)}.$ (24)
-Finally, we obtain that
-$t_{2}-t_{1}\geq\Omega\left(\frac{\mathsf{KL}\big{(}w^{t_{1}}\,\|\,w^{t_{2}}\big{)}}{\varepsilon^{2}}\right).$ (25)
-By setting $t_{1}=s_{i}$ and $t_{2}=e_{i}$, we can deduce that
-$e_{i}-s_{i}\geq\Omega\left(\frac{\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}}{\varepsilon^{2}}\right)=\Omega\left(\frac{1}{\varepsilon^{2}}\right).$ (26)
-Also noting that for any $t$, there are at most $8$ dimensions $i$ such that $s_{i}\leq t\leq e_{i}$, it holds that
-$8T\geq\sum_{i\in\mathcal{W}}(e_{i}-s_{i})\geq|\mathcal{W}|\cdot\Omega\left(\frac{1}{\varepsilon^{2}}\right),$ (27)
-which implies that $|\mathcal{W}|$ is at most $O(T\varepsilon^{2})=O(\log(k))$.
-In the analysis above, we analyze the property of the trajectory of Hedge algorithm with exact loss vector.
-However, in Algorithm 1, we need to sample to estimate the loss vector.
-In this case, (24) would be replaced by
-$\mathsf{KL}(w^{t_{1}}\,\|\,w^{t_{2}})\leq O(\eta^{2}(t_{2}-t_{1}))+O\left(\eta\sum_{\tau=t_{1}}^{t_{2}-1}(w^{t_{1}}+w^{\tau})^{\top}\Delta^{\tau}\right),$ (28)
-where $\Delta^{\tau}:=\left(\widehat{y}^{\tau}-y^{\tau}\right)$ with $\widehat{y}^{\tau}$ as the stochastic estimation of $y^{\tau}$.
-A naive bound for the second term in right hand side of (28) would be $\widetilde{O}\left(\eta\sqrt{(t_{2}-t_{1})}\right)$ by Azuma’s inequality.
-However, this bound is not sufficient because when $(t_{2}-t_{1})\varepsilon^{2}\ll 1$, the term $\widetilde{O}\left(\eta\sqrt{(t_{2}-t_{1})}\right)$would be the leading term.
-To address this problem, we consider weighted sampling to construct the empirical loss vector.
-As presented in Line 1 Algorithm 1, we manage to reduce the variance of $\Delta^{\tau}$ properly so that $\left(\eta\sum_{\tau=t_{1}}^{t_{2}-1}(w^{t_{1}}+w^{\tau})^{\top}\Delta^{\tau}\right)$ is roughly bounded by $O(\eta^{2}(t_{2}-t_{1}))$.
-4.2.2 Dealing with the segments
-In Section 4.2.1, we show that for any $1\leq t_{1}\leq t_{2}\leq T$
-$t_{2}-t_{1}\geq\Omega\left(\frac{\mathsf{KL}\big{(}w^{t_{1}}\,\|\,w^{t_{2}}\big{)}}{\varepsilon^{2}}\right).$ (29)
-By this inequality, we bound the size of $\mathcal{W}$ by $O(\log(k))$. A natural idea is to generalize the arguments to bound the size of
-$\mathcal{W}(p):=\left\{i\in[k]|\max_{1\leq t\leq T}w_{i}^{t}\in[2p,4p]\right\}$
-for $p\in[0,1]$.
-However, the arguments above can not provide a desired bound for $\mathcal{W}(p)$ when $p$ is small.
-For each $i\in\mathcal{W}(p)$, let $[s_{i},e_{i}]$ be the interval such that $\frac{p}{2}\leq w_{i}^{s_{i}}\leq p$, $w_{i}^{e_{i}}\geq 2p$ and $w_{i}^{t}\geq p$ for any $s_{i}<t\leq e_{i}$.
-By (29), we learn that
-$e_{i}-s_{i}\geq\Omega\left(\frac{\mathsf{KL}\big{(}\mathrm{Ber}(p)\,\|\,\mathrm{Ber}(2p)\big{)}}{\varepsilon^{2}}\right)\geq\Omega\left(\frac{p}{\varepsilon^{2}}\right),$ (30)
-where $\mathrm{Ber}(q)$ denote the Bernoulli distribution with mean $q$. Following the arguments in (27), we hope to bound
-$\sum_{i\in\mathcal{W}(p)}(e_{i}-s_{i})=O(T).$ (31)
-However, this might be not true because for each $t\in[T]$, there are possibly more than $\frac{1}{2p}$ coordinates $i\in\mathcal{W}(p)$ such that $t\in[s_{i},e_{i}]$, where we can at most derive $|\mathcal{W}(p)|\leq O(T\varepsilon^{2}/p^{2})$.
-As a result, the contribution of coordinates in $\mathcal{W}(p)$ to the target $\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ is $\widetilde{O}(1/p)$, which could be much larger than $\widetilde{O}(1)$ when $p$ is small.
-To address this problem, we note that if there are some coordinates $i\in\mathcal{W}(p)$ sharing similar $[s_{i},e_{i}]$, we can obtain a better bound.
-For example, suppose that for any $i\in\overline{\mathcal{W}}(p)\subset\mathcal{W}(p)$, $s_{i}=s$ and $e_{i}=e$, we then have that
-$e-s\geq\Omega\left(\frac{|\overline{\mathcal{W}}(p)|p}{\varepsilon^{2}}\right),$ (32)
-which is much better compared to the original bound $\Omega\left(\frac{p}{\varepsilon^{2}}\right)$ if $|\overline{\mathcal{W}}(p)|$ is large.
-By this observation, we consider to merge the dimensions in $\mathcal{W}(p)$ with similar $[s_{i},e_{i}]$. To facilitate analysis, we introduce the notation of segment.
+This algorithm is often referred to as the Hedge algorithm, which is known to yield an $\varepsilon$-minimax solution within $O\big{(}\frac{\log(k)}{\varepsilon^{2}}\big{)}$ iterations.
+A challenging question relevant to our analysis is:
+  • Question: can we bound $\|\overline{w}^{T}\|_{1}\coloneqq\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ in Algorithm 2 by poly-logarithmic terms?
+As it turns out, we can answer this question affirmatively (see Lemma 3), and the key ideas will be elucidated in the remainder of this section.
+4.2.1 First attempt: bounding the number of distributions with $\max_{1\leq t\leq T}w^{t}_{i}=\Omega(1)$
+Instead of bounding $\|\overline{w}^{T}\|_{1}$ directly, our first attempt is to look at those $i\in[k]$ with large $\max_{1\leq t\leq T}w^{t}_{i}$ (more specifically, $\max_{1\leq t\leq T}w^{t}_{i}\geq 1/4$) and show that:
+  • there exist at most $\widetilde{O}(1)$ coordinates $i\in[k]$ obeying $\max_{1\leq t\leq T}w^{t}_{i}\geq 1/4$ (or some other universal constant).
+In other words, we would like to demonstrate that the cardinality of the following set is small:
+$\mathcal{W}_{\mathsf{large}}\coloneqq\big{\{}i\in[k]\mid\max\nolimits_{1\leq t\leq T}w_{i}^{t}\geq 1/4\big{\}}.$ (18)
+To do so, note that some standard “continuity”-type argument tells us that: for a sufficiently small stepsize $\eta$, one can find, for each $i\in\mathcal{W}_{\mathsf{large}}$, a time interval $[s_{i},e_{i}]\subseteq[0,T]$ obeying
+$1/16\leq w_{i}^{s_{i}}\leq 1/8,\qquad w_{i}^{e_{i}}\geq 1/4,\qquad\text{and}\qquad w_{i}^{t}\geq 1/8~{}~{}\forall t\in(s_{i},e_{i}].$ (19)
+In words, $w_{i}^{t}$ at least doubles from $t=s_{i}$ to $t=e_{i}$. We claim for the moment that
+$e_{i}-s_{i}\geq{\Omega}(1/\eta^{2})={\Omega}(1/\varepsilon^{2})\qquad\forall i\in\mathcal{W}_{\mathsf{large}}.$ (20)
+Additionally, observe that for any $t$, there exist at most $8$ coordinates $i\in\mathcal{W}_{\mathsf{large}}$ such that $s_{i}\leq t\leq e_{i}$ (since $w_{i}^{t}\geq 1/8$ for every $t\in[s_{i},e_{i}]$).
+This reveals that
+$8T\geq\sum_{i\in\mathcal{W}_{\mathsf{large}}}(e_{i}-s_{i})\geq|\mathcal{W}_{\mathsf{large}}|\cdot\Omega\left(1/\varepsilon^{2}\right),$ (21)
+which combined with our choice of $T=\widetilde{O}(1/\varepsilon^{2})$ (cf. line 1 of Algorithm 1) yields
+$|\mathcal{W}_{\mathsf{large}}|\leq O(T\varepsilon^{2})=\widetilde{O}(1).$
+In words, the number of distributions with large $\max_{1\leq t\leq T}w^{t}_{i}$ (i.e., $\max_{1\leq t\leq T}w^{t}_{i}\geq 1/4$) is fairly small.
+Proof sketch for (20).
+Let us briefly discuss the high-level proof ideas.
+Following standard analysis for the Hedge algorithm (e.g., shalev2012online; lattimore2020bandit), we can often obtain (under certain mild conditions)
+$\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}\leq O\big{(}\eta^{2}(e_{i}-s_{i})\big{)}.$
+Combine this relation with basic properties about the KL divergence (see, e.g., Lemmas 10 and 11 in Appendix A) and the choice $\eta=\varepsilon/10$ to obtain
+$e_{i}-s_{i}$ $\geq\Omega\big{(}\eta^{-2}\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}\big{)}\geq\Omega\Big{(}\eta^{-2}\mathsf{KL}\big{(}\mathsf{Ber}(w_{i}^{s_{i}})\,\|\,\mathsf{Ber}(w_{i}^{e_{i}})\big{)}\Big{)}$
+$\geq\Omega\bigg{(}\eta^{-2}\frac{\frac{1}{16}\cdot 2^{2}}{4}\bigg{)}=\Omega\left(1/\eta^{2}\right)=\Omega\left(1/\varepsilon^{2}\right).$
+4.2.2 More general cases: issues and solutions
+Naturally, one would hope to generalize the arguments in Section 4.2.1 to cope with more general cases. More specifically, let us look at the following set
+$\mathcal{W}(p)\coloneqq\left\{i\in[k]\mid\max\nolimits_{1\leq t\leq T}w_{i}^{t}\in[2p,4p]\right\},$ (22)
+defined for each $p\in[0,1]$. If one could show that
+$|\mathcal{W}(p)|=\widetilde{O}(1/p)\qquad\text{for each }p,$ (23)
+then a standard doubling argument would immediately lead to
+$\|\overline{w}^{T}\|_{1}$ $=\sum_{i\in[k]}\max_{1\leq t\leq T}w_{i}^{t}\approx\sum_{j=1}^{\log_{2}k}\sum_{i\in\mathcal{W}(2^{-j})}\max_{1\leq t\leq T}w_{i}^{t}\leq\sum_{j=1}^{\log_{2}k}4\cdot 2^{-j}\cdot\big{|}\mathcal{W}(2^{-j})\big{|}$
+$\leq\sum_{j=1}^{\log_{2}k}4\cdot 2^{-j}\cdot\widetilde{O}\bigg{(}\frac{1}{2^{-j}}\bigg{)}=\widetilde{O}(1).$
+A technical issue.
+Nevertheless, simply repeating the arguments in Section 4.2.1 fails to deliver the desirable bound (23) on $|\mathcal{W}(p)|$ when $p$ is small.
+Briefly speaking, for each $i\in\mathcal{W}(p)$, let $[s_{i},e_{i}]$ represent a time interval (akin to (19)) such that
+$p/2\leq w_{i}^{s_{i}}\leq p,\qquad w_{i}^{e_{i}}\geq 2p\qquad\text{and}\qquad w_{i}^{t}\geq p\quad\text{for any }s_{i}<t\leq e_{i}.$ (24)
+Repeating the heuristic arguments in Section 4.2.1 leads to
+$e_{i}-s_{i}\geq\Omega\big{(}\eta^{-2}\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}\big{)}\geq\Omega\Big{(}\eta^{-2}\mathsf{KL}\big{(}\mathsf{Ber}(w_{i}^{s_{i}})\,\|\,\mathsf{Ber}(w_{i}^{e_{i}})\big{)}\Big{)}\geq\Omega(\eta^{-2}p)=\Omega\left(p/\varepsilon^{2}\right).$ (25)
+Given that each $t$ is contained within at most $1/(p/2)$ intervals associated with $\mathcal{W}(p)$, repeat the arguments for (21) to derive
+$\frac{1}{p/2}\cdot T\geq\sum_{i\in\mathcal{W}(p)}(e_{i}-s_{i})\geq|\mathcal{W}(p)|\cdot\Omega\left(p/\varepsilon^{2}\right)\qquad\Longrightarrow\qquad|\mathcal{W}(p)|\leq\widetilde{O}\bigg{(}\frac{1}{p^{2}}\bigg{)}.$
+This bound, however, is clearly loose compared to the desirable one in (23).
+Our solution.
+To address this issue, we make two key observations below that inspire our approach:
+  • Shared intervals.
+For the interval $[s_{i},e_{i}]$ associated with $i$ as defined above, if there exist other indices sharing the same interval $[s_{i},e_{i}]$ (in the sense that relations analogous to (24) are satisfied for other indices), then it is plausible to improve the bound.
+For instance, if a set $\mathcal{M}_{i}\subseteq[k]$ of indices share the same interval $[s_{i},e_{i}]$, then one can follow the heuristic argument in (25) to obtain
+    $e_{i}-s_{i}\geq\Omega\big{(}\eta^{-2}\mathsf{KL}\big{(}w^{s_{i}}\,\|\,w^{e_{i}}\big{)}\big{)}\geq\Omega\left(\eta^{-2}\mathsf{KL}\bigg{(}\mathsf{Ber}\Big{(}\sum_{j\in\mathcal{M}_{i}}w_{j}^{s_{i}}\Big{)}\,\|\,\mathsf{Ber}\Big{(}\sum_{j\in\mathcal{M}_{i}}w_{j}^{e_{i}}\Big{)}\bigg{)}\right)\geq\Omega\left(p|\mathcal{M}_{i}|/\varepsilon^{2}\right),$ (26)
+    which clearly strengthens the original bound (25) if $|\mathcal{M}_{i}|$ is large.
+  • Disjoint intervals. Consider the special case where $\mathcal{W}(p)$ can be divided into subsets $\{\mathcal{V}_{n}\}_{n=1}^{N}$ obeying
+    • (i)
+      for each $n\in[N]$, all indices in $\mathcal{V}_{n}$ share the same interval $[s_{n},e_{n}]$ (defined analogously as (24));
+    • (ii)
+      the intervals $\{[s_{n},e_{n}]\}_{n=1}^{N}$ are disjoint.
+    Then one can derive the desired bound on $|\mathcal{W}(p)|$. More precisely, it follows from (26) that
+    $e_{n}-s_{n}\geq\Omega\left(p|\mathcal{V}_{n}|/\varepsilon^{2}\right),$ (27)
+    which together with the disjointness property yields
+    $|\mathcal{W}_{j}|=\sum_{n=1}^{N}|\mathcal{V}_{n}|\leq\sum_{n=1}^{N}O\bigg{(}\frac{(e_{n}-s_{n})\varepsilon^{2}}{p}\bigg{)}\leq O\bigg{(}\frac{T\varepsilon^{2}}{p}\bigg{)}=\widetilde{O}\bigg{(}\frac{1}{p}\bigg{)}.$ (28)
+In light of the above discussion, it is helpful to (a) merge those indices that share similar intervals, and (b) identify disjoint intervals whose associated indices can cover a good fraction of $\mathcal{W}_{j}$.
+Motivated by the aforementioned observation about “shared intervals,” we introduce the notion of “segments” to facilitate analysis.
 Definition 1 (Segment).
-For $p,x>0$ and $i\in[k]$, we say $(t_{1},t_{2})$ is a $(p,q,x)$-segment iff there exists $\mathcal{I}\subset[K]$ such that $\sum_{i\in\mathcal{I}}w^{t_{1}}_{i}\in\left[\frac{p}{2},p\right]$, $\sum_{i\in\mathcal{I}}w^{t_{2}}_{i}\geq p\exp(x)$ and $\sum_{i\in\mathcal{I}}w_{i}^{t}\geq q$ for any $t_{1}\leq t\leq t_{2}$.
-We name $t_{1}$ as the starting point and $t_{2}$ as the ending point.
-And we also call $\mathcal{I}$ as the index set.
-By choosing $p/4\leq q\leq p/2$, a $(p,q,x)$-segment captures the variation of probabilities from $p$ to $p\exp(x)$.
-We say two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$ are disjoint iff $s_{1}<e_{1}\leq s_{2}<e_{2}$ or $s_{2}<e_{2}\leq s_{1}<e_{1}$.
-Using (29), it is easy to derive $t_{2}-t_{1}\geq\Omega(px^{2}/\varepsilon^{2})$ for a $(p,p/2,x)$-segment $(t_{1},t_{2})$.
-For each $i\in\mathcal{W}(p)$, there exists a $\left(\frac{p}{4},\frac{p}{8},\log(2)\right)$-segment $(s_{i},e_{i})$ with index set =$\mathcal{I}=\{i\}$.
-It then follows that $e_{i}-s_{i}\geq\Omega\left(\frac{p}{\varepsilon^{2}}\right)$ for each $i\in\mathcal{W}(p)$.
-As a result, if the segments $(s_{i},e_{i})$ are mutually disjoint, $|\mathcal{W}(p)|$ is at most $O(\log(k)/p)$.
-More generally, if we can divide the $|\mathcal{W}(p)|$ segments into disjoint blocks such that the segments in the same block share the same starting and ending points, we can also derive that $|\mathcal{W}(p)|\leq O(\log(k)/p)$.
-Taking Figure 1 as an example.
-Suppose we have $\ell$ blocks and the $i$-th block (with starting-ending points as $(\widetilde{s}^{i},\widetilde{e}^{i})$) contains $m_{i}$ coordinates satisfying that $\tilde{s}^{1}<\tilde{e}^{1}\leq\tilde{s}^{2}<\tilde{e}^{2}\leq\dots\leq\tilde{s}^{\ell}<\tilde{e}^{\ell}$.
-Then $(\widetilde{s}^{i},\widetilde{e}^{i})$ constructs a $\left(m_{i}\cdot\frac{p}{4},m_{i}\cdot\frac{p}{8},\log(2)\right)$-segment with index set as the $m_{i}$ coordinates in the $i$-th block, which means that
-$\widetilde{e}^{i}-\widetilde{s}^{i}\geq\Omega\left(\frac{m_{i}p}{\varepsilon^{2}}\right).$
-Taking sum over $1\leq i\leq\ell$, we obtain
-$|\mathcal{W}(p)|\leq\sum_{i=1}^{\ell}m_{i}\leq O\left(\sum_{i=1}^{\ell}\frac{(\widetilde{e}^{i}-\widetilde{s}^{i})\varepsilon^{2}}{p}\right)\leq O\left(\frac{T\varepsilon^{2}}{p}\right)=O\left(\frac{\log(k)}{p}\right),$ (33)
-which implies that $\sum_{i\in\mathcal{W}(p)}\max_{1\leq t\leq T}w_{i}^{t}\leq O(\log(k))$.
-With standard doubling arguments, it follows that $\sum_{i\in[k]}\max_{1\leq t\leq T}w_{i}^{t}=O(\log^{2}(k))$.
-By this observation, we introduce the concept of a regular configuration (see an example in Figure 1).
-Definition 3 (Configuration).
-A configuration $\mathsf{Conf}$ is a set of intervals $\mathsf{Conf}=\{[a_{i},b_{i}]\}_{i=1}^{m}$ with $b_{i}>a_{i}$ for each $i\in[m]$ (repeatable elements are allowed).
-We say the configuration $\mathsf{Conf}$ is regular iff for any $i,j\in[m]$, one of the three following claims holds.
-  • $a_{i}<b_{i}\leq a_{j}<b_{j}$;
-  • $a_{j}<b_{j}\leq a_{i}<b_{i}$;
-  • $a_{i}=a_{j}$, $b_{i}=b_{j}$.
-In words, (33) states that if $\{[s_{i},e_{i}]\}_{i\in\mathcal{W}(p)}$ forms a regular configuration, we then have $|\mathcal{W}(p)|=O\left(\frac{\log(k)}{p}\right)$ However, the general configuration of the segments might be irregular because it is possible that two segments are intersecting (see Figure 2).
-That is, for two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$, it holds that $s_{1}<s_{2}<e_{1}<e_{2}$.
-In this case, the approach above can at most bound $e_{2}-s_{1}$ by $\max\{e_{2}-s_{2},e_{1}-s_{1}\}$ instead of the desired bound $(e_{1}-s_{1})+(e_{2}-s_{2})$.
-To address this problem, we consider to construct a regular configuration with sub-segments [^5] A sub-segment is equivalent to a sub-interval of a segment.
-We refer to Figure 6 for more details.
-of the original segments.
-Our first step is to align one side of the segments.
-As presented in Figure 3, we divide the whole learning process into disjoint blocks such that the segments in each block have a common inner point.
-In the meantime, a segment is discarded if it intersects with more than one blocks.
-We show that there exists a regular configuration of the blocks such that at most $(1-\frac{1}{3(\log_{2}(T)+1)})\cdot|\mathcal{W}(p)|$ segments are discarded.
-In other words, at least $\frac{1}{3(\log_{2}(T)+1)}\cdot|\mathcal{W}(p)|$ segments are contained by these blocks.
-Since the blocks are disjoint to each other, it suffices to operate on one block.
-Without loss of generality, we assume that there is only one block, which also means that all segments have a common inner point $t_{\mathrm{mid}}$.
-Then we divide this block into two parts according to $t_{\mathrm{mid}}$ (see Figure 4).
-Since
-$\log\left(\frac{w_{i}^{e_{i}}}{w_{i}^{t_{\mathrm{mid}}}}\right)+\log\left(\frac{w_{i}^{t_{\mathrm{mid}}}}{w_{i}^{s_{i}}}\right)\geq\log(2),$
-it then holds either $(s_{i},t_{\mathrm{mid}})$ is a $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment with index ${i}$ or $(t_{\mathrm{mid}},e_{i})$ is a $\left(w_{i}^{t_{\mathrm{mid}}},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment with index set $\{i\}$.
-Since $w_{i}^{t_{\mathrm{mid}}}\in[p/4,p/2]$, we can roughly view $(t_{\mathrm{mid}},e_{i})$ as a $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment in the latter case.
-Therefore, we can find at least $\frac{|\mathcal{W}(p)|}{3(\log_{2}(T)+1)}\cdot\frac{1}{2}$ different $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$ segments with the same starting (ending) points.
-Without loss of generality, we assume that these $\frac{|\mathcal{W}(p)|}{6(\log_{2}(T)+1)}\cdot$ $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segments share a common starting point.
-As presented in Figure 5, we denote the common starting point as $e_{0}$, and re-order the dimensions so that $e_{i}$ is non-decreasing in $i$.
-Then we construct $O(\log_{2}(k))$ regular configurations using a recursion (see Lemma 28) , and show that at least one of these configurations contains $\widetilde{O}\left(|\mathcal{W}(p)|\right)$ $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2(\log_{2}(k)+2)}\right)$-segments.
-In Figure 6, we present a toy example in the case $k=8$.
-In this toy example We in total have $5$ configurations with 5 different colors.
-For each $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2}\right)$-segment in Figure 5, it contributes an $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2(\log_{2}(k)+2)}\right)$-segment to at least one of these configurations.
-By the pigeonhole princeton, there must be one regular configuration with $\frac{|\mathcal{W}(p)|}{6(\log_{2}(T)+1)(\log_{2}(k)+2)}$ $\left(\frac{p}{4},\frac{p}{8},\frac{\log(2)}{2(\log_{2}(k)+2)}\right)$-segments (see Lemma 28).
-Figure 1: Regular configuration.
-Figure 2: General irregular configuration.
-Figure 3: Partition of blocks.
-Figure 4: Aligning one side of the configuration.
-The unfilled part of the segments mean that the variation over $w_{i}^{s_{i}}\to w_{i}^{t_{\mathrm{mid}}}$ (i.e., $\log\left(w_{i}^{t_{\mathrm{mid}}}/w_{i}^{s_{i}}\right)$) is not significant enough.
-Figure 5: A group of segments with common starting point.
-Figure 6: Construction of the regular configurations. Each segment is cut into at most $\log_{2}(k)+1$ sub-segments with different colors.
-5 Analysis (proof of Theorem 1)
-Theorem 1 follows by combining Lemma 4, Lemma 5 and Lemma 6.
-Recall that $\varepsilon_{1}=\frac{\varepsilon}{100}$. In Lemma 4, we show that, with high-probability,
-$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (34)
-holds for any $1\leq t\leq T$.
-Assuming (34) holds for any $1\leq t\leq T$, we continue by proving that $h^{\mathsf{final}}$ is $\varepsilon$-optimal in Lemma 5, and providing a bound for the sample complexity in Lemma 6.
+For any $p,x>0$ and $i\in[k]$, we say that $(t_{1},t_{2})$ is a $(p,q,x)$-segment if there exists a subset $\mathcal{I}\subseteq[k]$ such that
+  • $\mathrm{(i)}$
+    $\sum_{i\in\mathcal{I}}w^{t_{1}}_{i}\in\left[p/2,p\right]$,
+  • $\mathrm{(ii)}$
+    $\sum_{i\in\mathcal{I}}w^{t_{2}}_{i}\geq p\exp(x)$,
+  • $\mathrm{(iii)}$
+    $\sum_{i\in\mathcal{I}}w_{i}^{t}\geq q$ for any $t_{1}\leq t\leq t_{2}$.
+We shall refer to $t_{1}$ as the starting point and $t_{2}$ as the end point, and call $\mathcal{I}$ the associated index set.
+Moreover, two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$ are said to be disjoint if $s_{1}<e_{1}\leq s_{2}<e_{2}$ or $s_{2}<e_{2}\leq s_{1}<e_{1}$.
+This definition allows one to pool indices with similar intervals together.
+In general, however, it is common for two segments to be overlapping (see Figure 2 in Appendix C.5), which precludes us from directly invoking our aforementioned observation about “disjoint intervals.” To address this issue, our strategy is to extract out shared sub-segments [^7] A sub-segment refers to a sub-interval of a segment, as illustrated in Figure 6.
+of (a subset of) these segments in a meticulous manner.
+Encouragingly, it is possible to find such sub-segments that taken collectively cover a good fraction (i.e., $\frac{1}{\mathrm{poly}\log(k,T)}$) of all segments, meaning that we do not have to discard too many segments.
+The construction is built upon careful analysis of these segments, and will be elucidated in Appendix C.
+5 Analysis for VC classes (proof of Theorem 1)
+The main steps for establishing Theorem 1 lie in proving three key lemmas, as stated below.
+The first lemma is concerned with the hypothesis $h^{t}=\arg\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})$ (cf.
+line 1 of Algorithm 1); in words, $h^{t}$ is the minimizer of the empirical loss function $\widehat{L}^{t}(\cdot,w^{t})$, computed using samples obtained up to the $t$-th round.
+The following lemma tells us that: even though $h^{t}$ is an empirical minimizer, it almost optimizes the weighted population loss $L(\cdot,w^{t})$.
+In other words, this lemma justifies that the adaptive sampling scheme proposed in Algorithm 1 ensures faithfulness of the empirical loss and its minimizer; here, we recall that $\varepsilon_{1}=\varepsilon/100$ (cf.
+line 1 of Algorithm 1).
 Lemma 1.
@@
 $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (29)
-holds for any $1\leq t\leq T$.
+holds for all $1\leq t\leq T$, where $h^{t}$ (resp. $w^{t}$) is the hypothesis (resp. weight vector) computed in round $t$ of Algorithm 1.
+Proof.
+See Appendix B.1. ∎
+Next, assuming that (29) holds, we can resort to standard analysis for the Hedge algorithm to demonstrate the quality of the final output $h^{\mathsf{final}}$.
 Lemma 2.
-Assume lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$.
-With probability exceeding $41-\delta/$ 4, the output policy $h^{\mathsf{final}}$ is $\varepsilon$-optimal.
+Suppose that lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$.
+With probability exceeding $41-\delta/$ 4, the hypothesis $h^{\mathsf{final}}$ output by Algorithm 1 is $\varepsilon$-optimal in the sense that
+$\max_{1\leq i\leq k}L(h^{\mathsf{final}},\mathrm{e}^{\mathsf{basis}}_{i})\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}L(h,\mathrm{e}^{\mathsf{basis}}_{i})+\varepsilon.$ (30)
+Here, we recall that $\mathrm{e}^{\mathsf{basis}}_{i}$ indicates the $i$-th standard basis vector.
+Proof.
+See Appendix B.2. ∎
+Taking Lemma 1 and Lemma 2 together, one can readily see that Algorithm 1 returns an $\varepsilon$-optimal randomized hypothesis $h^{\mathsf{final}}$ with probability at least $21-\delta/$ 2.
+The next step then lies in bounding the total number of samples that has been collected in Algorithm 1.
+Towards this end, recall that $\overline{w}^{T}_{i}=\max_{1\leq t\leq T}w_{i}^{t}$ for each $i\in[k]$.
+Recognizing that $\widehat{w}^{t}_{i}\leq\overline{w}^{t}_{i}$ for each $t\in[T]$ and $i\in[k]$, we can bound the total sample size by
+(sample $(\text{sample size})\qquad T_{1}\sum_{i=1}^{k}\widehat{w}_{i}^{T}+k+T\bigg{(}k\sum_{i=1}^{k}\overline{w}_{i}^{T}+k\bigg{)}$ $\leq\left(T_{1}\|\overline{w}^{T}\|_{1}+kT\|\overline{w}^{T}\|_{1}\right)+k(T+1)$
+$\lesssim\frac{d\log\left(\frac{kd}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\|\overline{w}^{T}\|_{1},$ (31)
+where the last relation follows from our choices $T\asymp\frac{\log(k/\delta)}{\varepsilon^{2}}$ and $T_{1}\asymp\frac{k\log(k/\varepsilon)+d\log(kd/\varepsilon)+\log(1/\delta)}{\varepsilon^{2}}$ (cf.
+line 1 of Algorithm 1) and the basic property that $\|\overline{w}^{T}\|_{1}\geq\sum_{i}w_{i}^{1}=1$.
+Consequently, everything then comes down to bounding $\|\overline{w}^{T}\|_{1}$, for which we resort to the following lemma.
 Lemma 3.
-Assume lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$.
-With probability at least $21-\delta/$ 2, the sample complexity of Algorithm 1 is bounded by
-$O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\left(\log^{5}(k)\log(1/\varepsilon)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)\right).$
-5.1 Optimality (proof of Lemma 5)
-Let $\mathrm{OPT}:=\min_{h\in\mathcal{H}}\max_{i}L(h,\mathbf{e}_{i})$.
-Let $v^{t}=L(h^{t},w^{t})-\mathrm{OPT}$ and $f^{t}=\min_{h\in\mathcal{H}}L(h,w^{t})-\mathrm{OPT}\leq 0$.
-By the assumption that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$, we have $v^{t}\leq f^{t}+\varepsilon_{1}\leq\varepsilon_{1}$ for any $1\leq t\leq T$.
-Let $\delta^{\prime}=\frac{\delta}{T+k+1}$. Note that $\eta\leq 1$ and $|\hat{r}_{i}^{t}|\leq 1$ for any proper $(i,t)$. Direct computation gives that
-$\log\left(\frac{\sum_{i=1}^{k}W_{i}^{t+1}}{\sum_{i=1}^{k}W_{i}^{t}}\right)$ $=\log\left(\sum_{i=1}^{k}w_{i}^{t}\exp(\eta\hat{r}^{t}_{i})\right)\leq\log\left(\sum_{i=1}^{k}w_{i}^{t}(1+\eta\hat{r}_{i}^{t}+\eta^{2}(\hat{r}_{i}^{t})^{2})\right)\leq\eta\sum_{i=1}^{k}w_{i}^{t}\hat{r}_{i}^{t}+\eta^{2}.$ (36)
-As a result, we have that
-$\eta\sum_{t=1}^{T}w^{t}\cdot\hat{r}^{t}$ $\geq\log\left(\sum_{i=1}^{k}W_{i}^{T+1}\right)-T\eta^{2}-\log(k)$
-$\geq\max_{i}\log(W_{i}^{T})-T\eta^{2}-\log(k)$
-$\geq\eta\max_{i}\sum_{t=1}^{T}\hat{r}_{i}^{t}-T\eta^{2}-\log(k).$ (37)
-By dividing both side with $\eta$, we have
-$\sum_{t=1}^{T}w^{t}\cdot\hat{r}^{t}\geq\max_{i}\sum_{t=1}^{T}\hat{r}^{t}_{i}-\left(\frac{\log(k)}{\eta}+\eta T\right).$ (38)
-By Azuma’s inequality, with probability $1-(k+1)\delta^{\prime}$, we have that
-$\left|\sum_{t=1}^{T}w^{t}\cdot\hat{r}^{t}-\sum_{t=1}^{T}L(h^{t},w^{t})\right|\leq 2\sqrt{T\log(1/\delta^{\prime})};$
-$\left|\sum_{t=1}^{T}\hat{r}_{i}^{t}-\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})\right|\leq 2\sqrt{T\log(1/\delta^{\prime})}.$
-As a result, with probability $1-(k+1)\delta^{\prime}$, it holds that
-$\sum_{t=1}^{T}L(h^{t},w^{t})$ $\geq\max_{i}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})-\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)$
-$\geq-\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right).$ (39)
-By the fact that $\varepsilon_{1}=\eta=\frac{1}{100}\varepsilon$ and $T=\frac{20000\log(\frac{k}{\delta^{\prime}\varepsilon})}{\varepsilon^{2}}$, we have that
-$\max_{i}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})$ $\leq T\mathrm{OPT}+\sum_{t=1}^{T}v^{t}+\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)$
-$\leq T\mathrm{OPT}+T\varepsilon_{1}+\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)$
-$\leq T\mathrm{OPT}+T\varepsilon.$ (40)
-As a result, we obtain that
-$\max_{i}L(h^{\mathsf{final}},\mathbf{e}_{i})=\max_{i}\frac{1}{T}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})\leq\mathrm{OPT}+\varepsilon.$ (41)
-The proof is finished by noting that $\delta^{\prime}=\frac{\delta}{(T+k+1)}$.
-5.2 Sample complexity (proof of Lemma 6)
-Recall that $\{w^{t}\}_{t=1}^{T+1}$ is the trajectory by running Algorithm 1.
-Also recall that $\overline{w}^{T}_{i}=\max_{1\leq t\leq T}w_{i}^{t}$ for all $i\in[k]$.
-We start with the lemma below to bound the sample complexity by $\|\overline{w}^{T}\|_{1}$.
+Assume that lines 1-1 in Algorithm 1 are replaced with some oracle which returns a hypothesis $h^{t}$ satisfies that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$.
+With probability at least $41-\delta/$ 4, the quantity $\|\overline{w}^{T}\|_{1}$ is bounded above by
+$\|\overline{w}^{T}\|_{1}\leq O\left(\log^{8}\left(\frac{k}{\delta\varepsilon}\right)\right).$
+It is noteworthy that the proof of Lemma 3 is the most technically challenging part of the analysis; we postpone this proof to Appendix B.3.
+Combining Lemma 3 with (31) immediately reveals that, with probability at least $1-\delta$, the sample complexity of Algorithm 1 is bounded by
+$O\left(\frac{d\log\left(\frac{kd}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\log^{8}\left(\frac{k}{\delta\varepsilon}\right)\right),$
+as claimed in Theorem 1. It remains to prove the above key lemmas, which we postpone to Appendix B.
+6 Necessity of randomization
+Given that the best-known sample complexities prior to our work were derived for algorithms that either output randomized hypotheses or invoke majority votes, awasthi2023sample raised the question about how the sample complexity is impacted if only deterministic (or “proper”) hypotheses are permitted as the output of the learning algorithms.
+As it turns out, the restriction to deterministic hypotheses substantially worsens the sample efficiency, as revealed by the following theorem.
+Theorem 2.
+Assume that $d\geq 2\log(8k)$. Consider any $\varepsilon\in(0,1/100)$, and let $N_{0}=\frac{2^{d}-1}{k}$. One can find
+  • a hypothesis class $\mathcal{H}$ containing at most $kN_{0}+1$ hypothesis,
+  • a collection of $k$ distributions $\mathcal{D}\coloneqq\{\mathcal{D}_{i}\}_{i=1}^{k}$,
+  • a loss function :×$\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$,
+such that: for any algorithm $\mathcal{G}$ that finds $h\in\mathcal{H}$ obeying
+$\max_{1\leq i\leq k}\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}\leq\min_{h^{\prime}\in\mathcal{H}}\max_{i\in[k]}\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h^{\prime},(x,y)\big{)}\big{]}+\varepsilon$ (32)
+with probability exceeding $3/4$, its sample size — denoted by $M(\mathcal{G})$ — must exceed $\mathbb{E}[M(\mathcal{G})]\geq\frac{dk}{240000\varepsilon^{2}}$.
+In words, the sample complexity when a deterministic output is required scales at least with $\Omega(\frac{dk}{\varepsilon^{2}}\big{)}$, which is considerably larger than the sample complexity $\widetilde{O}\big{(}\frac{d+k}{\varepsilon^{2}}\big{)}$ achievable via Algorithm 1 (which outputs a randomized hypothesis).
+The proof of this theorem is deferred to Appendix D.
+7 Extension: learning Rademacher classes
+In this section, we adapt our algorithm and theory to accommodate MDL for Rademacher classes.
+Note that when learning Rademacher classes, the hypotheses in $\mathcal{H}$ do not need to be binary-valued.
+7.1 Preliminaries: Rademacher complexity
+Let us first introduce the formal definition of the Rademacher complexity; more detailed introduction can be found in, e.g., shalev2014understanding.
+Definition 2 (Rademacher complexity).
+Given a distribution $\mathcal{D}$ supported on $\mathcal{Z}\coloneqq\mathcal{X}\times\mathcal{Y}$ and a positive integer $n$, the Rademacher complexity is defined as
+$\mathsf{Rad}_{n}(\mathcal{D})\coloneqq\mathop{\mathbb{E}}_{\{z_{i}\}_{i=1}^{n}}\left[\mathop{\mathbb{E}}_{\{\sigma_{i}\}_{i=1}^{n}}\left[\max_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\ell(h,z_{i})\right]\right],$ (33)
+where $\{z_{i}\}_{i=1}^{n}$ are drawn independently from $\mathcal{D}$, and $\{\sigma_{i}\}_{i=1}^{n}$ are i.i.d.
+Rademacher random variables obeying $\mathbb{P}\{\sigma_{i}=1\}=\mathbb{P}\{\sigma_{i}=-1\}=1/2$ for each $1\leq i\leq n$.
+Next, we would like to make an assumption concerning the Rademacher complexity of mixtures of distributions. Denoting by $\mathcal{D}(w)$ the mixed distribution
+$\mathcal{D}(w)\coloneqq\sum_{i=1}^{k}w_{i}\mathcal{D}_{i}$ (34)
+for any probability vector $w=[w_{i}]_{1\leq i\leq k}\in\Delta(k)$, we can state our assumption as follows.
+Assumption 1.
+For each $n\geq 1$, there exists a quantity $C_{n}>0$ (known to the learner) such that
+$C_{n}\geq\sup_{w\in\Delta(k)}\mathsf{Rad}_{n}\big{(}\mathcal{D}(w)\big{)}.$ (35)
+For instance, if the hypothesis class $\mathcal{H}$ obeys $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$, then it is well-known that Assumption 1 holds with the choice (see, e.g., mohri2018foundations)
+$C_{n}=\sqrt{\frac{2d\log(en/d)}{n}}.$
+Remark 3.
+One might raise a natural question about Assumption 1: can we use $\widetilde{C}_{n}\coloneqq\max_{i\in[k]}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ instead of ${C}_{n}$ without incuring a worse sample complexity?
+The answer is, however, negative.
+In fact, the Rademacher complexity $\mathsf{Rad}_{n}(\mathcal{D}(w))$ is not convex in $w$, and hence we fail to use $\max_{i}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ to bound $\max_{w\in\Delta(k)}\mathsf{Rad}_{n}(\mathcal{D}(w))$.
+The interested reader is referred to Appendix E.4 for more details.
+To facilitate analysis, we find it helpful to introduce another notion called weighted Rademacher complexity.
+Definition 3 (Weighted Rademacher complexity).
+Given a collection of distributions $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a set of positive integers $\{n_{i}\}_{i=1}^{k}$, the weighted (average) Rademacher complexity is defined as
+$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}\coloneqq\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right],$ (36)
+where $\big{\{}\{z_{i}^{j}\}_{j=1}^{n_{k}}\big{\}}_{i=1}^{k}$ are independently generated with each $z_{i}^{j}$ drawn from $\mathcal{D}_{i}$, and $\big{\{}\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\big{\}}_{i=1}^{k}$ are independent Rademacher random variables obeying $\mathbb{P}\{\sigma_{i}^{j}=1\}=\mathbb{P}\{\sigma_{i}^{j}=-1\}=1/2$.
+Throughout the rest of this paper, we shall often abbreviate $\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}=\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\mathcal{D})$.
+The following two lemmas provide useful properties about the weighted Rademacher complexity.
 Lemma 4.
-The sample complexity of Algorithm 1 is bounded by $O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\|\overline{w}^{T}\|_{1}\right)$.
+For any two groups of positive integers $\{n_{i}\}_{i=1}^{k}$ and $\{m_{i}\}_{i=1}^{k}$, it holds that
+$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$ $\leq\left(\sum_{i=1}^{k}(m_{i}+n_{i})\right)\widetilde{\mathsf{Rad}}_{\{m_{i}+n_{i}\}_{i=1}^{k}}$
+$\leq\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$ (37)
 Proof.
-Recall the definition of $\check{w}^{t}$ and $\overline{w}^{t}$ in Algorithm 1. It is easy to observe that $\check{w}^{t}\leq\overline{w}^{t}.$ Then the total number of samples is bounded by
-$O\left(T_{1}\cdot\|\overline{w}^{T}\|_{1}+T(k\|\overline{w}^{T}\|_{1}+k)\right)=O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\|\overline{w}^{T}\|_{1}\right).$
-On the other hand, we have the lemma below to bound $\|\overline{w}^{T}\|_{1}$.
+See Appendix E.2.∎
 Lemma 5.
-Assume Line 1-1 in Algorithm 1 is replaced by some oracle which returns a hypothesis $h^{t}$ satisfies that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$.
-With probability $41-\delta/$ 4, the sample complexity of Algorithm 1 is bounded by
-$O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\left(\log^{5}(k)\log(1/\varepsilon)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)\right).$
-The proof of Lemma 8 is rather involved. We postpone the proof to Appendix C.1 due to space limitation.
-By Lemma 7 and Lemma 8, with probability $21-\delta/$ 2, the sample complexity of Algorithm 1 is bounded by
-$O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\left(\log^{5}(k)\log(1/\varepsilon)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)\right).$
-The proof is completed.
-6 Extension: learning Rademacher classes
-In this section, we study how to adapt our algorithm and theory to accommodate MDL for Rademacher classes.
-1 input: $k$ data distributions $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, hypothesis class $\mathcal{H}$, target accuracy level $\varepsilon$, target success rate $1-\delta$, the constant $\{C_{n}\}_{n\geq 1}$ in Assumption 10.
-2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}=\min\left\{t\geq\frac{400\left(k\log(\frac{k}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}\Big{|}C_{t}\leq\frac{\epsilon_{1}}{1200}\right\}$.
+Consider any $\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for each $i\in[k]$. By taking $w\in\Delta(k)$ with $w_{i}=\frac{n_{i}}{\sum_{l=1}^{k}n_{l}}$, one has
+$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}\leq 72\mathsf{Rad}_{\sum_{i=1}^{k}n_{i}}\big{(}\mathcal{D}(w)\big{)}.$
+Proof.
+See Appendix E.3.∎
+7.2 Algorithm and sample complexity
+We are now positioned to introduce our algorithm that learns a Rademacher class in the presence of multiple distributions, which is also based on a Hedge-type strategy to learn a convex-concave game; see Algorithm 3 for full details.
+Its main distinction from Algorithm 1 lies in the subroutine to learn $h^{t}$ (see lines 3-3 in Algorithm 3) as well as the choice of $T_{1}$ (see line 3 in Algorithm 3).
+More precisely, to compute the estimator $\widehat{L}^{t}(h,w^{t})$ for $L(h,w^{t})$, instead of using the first $n_{i}^{t}$ samples from $\mathcal{D}_{i}$ for each $i\in[k]$, we choose to use the first
+$n_{i}^{t,\mathsf{rad}}=\min\big{\{}\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\big{\}}$
+samples from $\mathcal{D}_{i}$ for each $i$, where $T_{1}$ is taken to be
+$T_{1}=\min\Big{\{}t\geq\frac{4000(k\log(k/\varepsilon_{1})+\log(1/\delta))}{\varepsilon_{1}^{2}}\,\Big{|}\,C_{t}\leq\frac{\varepsilon_{1}}{4800}\Big{\}}.$ (38)
+Here, $T_{1}$ needs to take advantage of the quantities $\{C_{n}\}$ (cf. Assumption 1) that upper bound the associated Rademacher complexity.
+1 input: $k$ data distributions $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, hypothesis class $\mathcal{H}$, target accuracy level $\varepsilon$, target success rate $1-\delta$, quantities $\{C_{n}\}_{n\geq 1}$ as in Assumption 1.
+2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}=\min\Big{\{}t\geq\frac{4000(k\log(k/\varepsilon_{1})+\log(1/\delta))}{\varepsilon_{1}^{2}}\,\Big{|}\,C_{t}\leq\frac{\varepsilon_{1}}{4800}\Big{\}}$.
 3 initialization: for all $i\in[k]$, set $W^{1}_{i}=1$, $\widehat{w}^{0}_{i}=0$ and $n_{i}^{0}=0$; $\mathcal{S}=\emptyset$.
@@
 5 for $t=1,2,\ldots,T$ do
-6       set $w^{t}=[w^{t}_{i}]_{1\leq i\leq k}$ and $\widehat{w}^{t}=[\widehat{w}^{t}_{i}]_{1\leq i\leq k}$, where $w^{t}_{i}\leftarrow\frac{W^{t}_{i}}{\sum_{j}W^{t}_{j}}$ and $\widehat{w}^{t}_{i}\leftarrow\widehat{w}^{t-1}_{i}$ for all $i\in[k]$.
+6       set $w^{t}=[w^{t}_{i}]_{1\leq i\leq k}$ and $\widehat{w}^{t}=[\widehat{w}^{t}_{i}]_{1\leq i\leq k}$, with $w^{t}_{i}\leftarrow\frac{W^{t}_{i}}{\sum_{j}W^{t}_{j}}$ and $\widehat{w}^{t}_{i}\leftarrow\widehat{w}^{t-1}_{i}$ for all $i\in[k]$.
        /* recompute $\widehat{w}^{t}$ draw new samples for $\mathcal{S}_{\mathrm{w}}$ only if $w^{t}$ changes sufficiently. */
@@
 13       compute $h^{t}\leftarrow\arg\min_{h\in\mathcal{H}}\widehat{L}(h,w^{t})$, where
-$\widehat{L}^{t}(h,w^{t}):=\sum_{i=1}^{k}\frac{w^{t}_{i}}{\check{n}_{i}^{t}}\cdot\sum_{j=1}^{\check{n}_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (42)
-with $\check{n}_{i}^{t}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$, and $(x_{i,j},y_{i,j})$ being the $j$-th datapoint from $\mathcal{D}_{i}$ in $\mathcal{S}$.
+$\widehat{L}^{t}(h,w^{t}):=\sum_{i=1}^{k}\frac{w^{t}_{i}}{n_{i}^{t,\mathsf{rad}}}\cdot\sum_{j=1}^{n_{i}^{t,\mathsf{rad}}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (39)
+with $n_{i}^{t,\mathsf{rad}}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ and $(x_{i,j},y_{i,j})$ being the $j$-th datapoint from $\mathcal{D}_{i}$ in $\mathcal{S}$.
        /* estimate the loss vector and execute weighted updates. */
@@
 15       for $i=1,\ldots,k$ do
-             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
+16             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
 $\widehat{r}^{t}_{i}=\frac{1}{\lceil k\overline{w}^{t}_{i}\rceil}\sum_{j=1}^{\lceil k\overline{w}^{t}_{i}\rceil}\ell\big{(}h^{t},(x_{i,j}^{t},y_{i,j}^{t})\big{)};$
             update the weight as $W^{t+1}_{i}=W^{t}_{i}\exp(\eta\widehat{r}_{i}^{t})$. // Hedge updates.
-16
 17
+18
 output: a randomized hypothesis $h^{\mathsf{final}}$ as a uniform distribution over $\{h^{t}\}_{t=1}^{T}$.
-Algorithm 3 Hedge for Multi-distribution Learning on Rademacher Class ($\mathtt{MDL-Hedge-Rademacher}$)
-6.1 Preliminaries: Rademacher complexity
-Let us first introduce the formal definition of the Rademacher complexity; more detailed introduction can be found in Shalev-Shwartz and Ben-David, (2014).
-Definition 9 (Rademacher complexity).
-Given a distribution $\mathcal{D}$ supported on $\mathcal{Z}\coloneqq\mathcal{X}\times\mathcal{Y}$ and a positive integer $n$, the (average) Rademacher complexity is defined as
-$\mathsf{Rad}_{n}(\mathcal{D})\coloneqq\mathop{\mathbb{E}}_{\{z_{i}\}_{i=1}^{n}}\left[\mathop{\mathbb{E}}_{\{\sigma_{i}\}_{i=1}^{n}}\left[\max_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\ell(h,z_{i})\right]\right],$ (43)
-where $\{z_{i}\}_{i=1}^{n}$ are drawn independently from $\mathcal{D}$, and $\{\sigma_{i}\}_{i=1}^{n}$ are i.i.d.
-Rademacher random variables obeying $\mathbb{P}\{\sigma_{i}=1\}=\mathbb{P}\{\sigma_{i}=-1\}=1/2$ for each $1\leq i\leq n$.
-Next, we would like to make an assumption on the Rademacher complexity of mixtures of distributions. Denoting by $\mathcal{D}(w)$ the mixed distribution
-$\mathcal{D}(w)\coloneqq\sum_{i=1}^{k}w_{i}\mathcal{D}_{i}$ (44)
-for any probability vector $w\in\Delta(k)$, we can state our assumption as follows.
-Assumption 10.
-For each $n\geq 1$, there exists a universal constant $C_{n}>0$ such that
-$\mathsf{Rad}_{n}\big{(}\mathcal{D}(w)\big{)}\leq C_{n}$ (45)
-holds for all $w\in\Delta^{k}$.
-Remark 1.
-One might raise a question about Assumption 10: Can we only assume $\mathrm{Rad}_{n}(\mathcal{D}_{i})\leq C_{n}$ for $i\in[k]$ without incuring a worse sample complexity?
-The answer is negative.
-Actually, the Rademacher complexity $\mathsf{Rad}_{n}(\mathcal{D}(w))$ is not convex in $w$, where we fail to use $\max_{i}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ to bound $\max_{w\in\Delta(k)}\mathsf{Rad}_{n}(\mathcal{D}(w))$.
-We refer to Appendix D.4 for more details.
-It is well known that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ implies Assumption 10 holds with $C_{n}=\sqrt{\frac{2d\log(en/d)}{n}}$ (Mohri et al.,, 2018).
-To facilitate our analysis in this section, we find it helpful to introduce the notion of the weighted Rademacher complexity as follows.
-Definition 11 (Weighted Rademacher complexity).
-Given a collection of distributions $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a set of positive integers $\{n_{i}\}_{i=1}^{k}$, the weighted (average) Rademacher complexity is defined as
-$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\mathcal{D})\coloneqq\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right],$ (46)
-where $\big{\{}\{z_{i}^{j}\}_{j=1}^{n_{k}}\big{\}}_{i=1}^{k}$ are independently generated with each $z_{i}^{j}$ drawn from $\mathcal{D}_{i}$, and $\big{\{}\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\big{\}}_{i=1}^{k}$ are independent Rademacher random variables obeying $\mathbb{P}\{\sigma_{i}^{j}=1\}=\mathbb{P}\{\sigma_{i}^{j}=-1\}=1/2$.
-Throughout the rest of this paper, we shall often abbreviate $\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}=\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\mathcal{D})$.
-The weighted Rademacher complexity defined above satisfies an important property below.
+Algorithm 3 Hedge for multi-distribution learning on Rademacher Classes ($\mathtt{MDL}\text{-}\mathtt{Hedge}\text{-}\mathtt{Rad}$)
+Equipped with this algorithm, we are ready to establish the following theoretical guarantees.
+Theorem 3.
+Suppose Assumption 1 holds. With probability at least $1-\delta$, the output $h^{\mathsf{final}}$ returned by Algorithm 3 satisfies
+$\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},h^{\mathsf{final}}}\left[\ell\big{(}h^{\mathsf{final}},(x,y)\big{)}\right]\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}+\varepsilon.$ (40)
+In particular, the total number of samples collected by Algorithm 3 is bounded by
+$\left(\frac{k}{\varepsilon^{2}}+\min\big{\{}n\mid C_{n}\leq c_{1}\varepsilon\big{\}}\right)\,\mathrm{poly}\log\Big{(}k,\frac{1}{\varepsilon},\frac{1}{\delta}\Big{)}$
+with probability exceeding $1-\delta$, where $c_{1}>0$ is some sufficiently small constant, and $T_{1}$ is defined in (38).
+In contrast to the VC classes, the sample complexity for learning Rademacher classes entails a term related to the Rademacher complexity — namely, $\min\big{\{}n\mid C_{n}\leq c_{1}\varepsilon\big{\}}$ — as opposed to the term $d/\varepsilon^{2}$ concerning the VC-dimension.
+When it comes to the special case where $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$, taking $C_{n}\leq\sqrt{\frac{2d\log(en/d)}{n}}$ in (38) leads to $T_{1}=\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$, which recovers the sample complexity bound of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ derived in Theorem 1 for VC classes.
+Proof of Theorem 3.
+In view of Lemma 2 and Lemma 3, it boils down to showing that running Algorithm 3 results in $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ for any $1\leq t\leq T$, a property that holds with probability at least $41-\delta/$ 4.
+To accomplish this, we have the lemma below.
 Lemma 6.
-For any two groups of positive integers $\{n_{i}\}_{i=1}^{k}$ and $\{m_{i}\}_{i=1}^{k}$, it holds that
-$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$ $\leq\left(\sum_{i=1}^{k}(m_{i}+n_{i})\right)\widetilde{\mathsf{Rad}}_{\{m_{i}+n_{i}\}_{i=1}^{k}}$
-$\leq\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$ (47)
-In addition, the following allows us to bound the weighted Rademacher complexity under Assumption 10.
-Lemma 13.
-Consider any $\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for each $i\in[k]$. By taking $w\in\Delta^{k}$ with $w_{i}=\frac{n_{i}}{\sum_{l=1}^{k}n_{l}}$, one has
-$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}\leq 72\mathsf{Rad}_{\sum_{i=1}^{k}n_{i}}\big{(}\mathcal{D}(w)\big{)}.$
-6.2 Algorithm and sample complexity
-We present in Algorithm 3 to learn a Rademacher class under multi-distribution tasks.
-Algorithm 3 is also a Hedge-like algorithm to learn a convex (concave) game.
-The major difference from Algorithm 1 lies in the subroutine to learn $h^{t}$ (See lines 3-3 in Algorithm 3).
-More precisely, to compute the estimator $\widehat{L}^{t}(h,w^{t})$ for $L(h,w^{t})$, instead of using the first $n_{i}^{t}$ samples from $\mathcal{D}_{i}$ for each $i\in[k]$, we choose to use the first $\check{n}_{i}^{t}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ samples from $\mathcal{D}_{i}$ for each $i$.
-Formally, we have the result below.
-Theorem 14.
-Suppose Assumption 10 holds. By running Algorithm 3, with probability $1-\delta$, the output $h^{\mathsf{final}}$ satisfies that
-$\max_{i}L(h^{\mathsf{final}},\mathbf{e}_{i})=\max_{i}\frac{1}{T}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})\leq\min_{h\in\mathcal{H}}\max_{i}L(h,\mathbf{e}_{i})+\varepsilon.$
-Meanwhile, the sample complexity of Algorithm 3 is bounded by
-$O\left(T_{\epsilon}\cdot\log^{5}(k)\log\left(\frac{1}{\varepsilon}\right)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)$
-where $T_{\epsilon}$ is defined as
-$T_{\epsilon}=\min\left\{t\geq\frac{400\left(k\log(\frac{k}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}\Big{|}C_{t}\leq\frac{\epsilon_{1}}{1200}\right\}.$
-In the case $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$, we have $C_{n}\leq\sqrt{\frac{2d\log(en/d)}{n}}$, which implies that $T_{\epsilon}=\tilde{O}\left(\frac{d+k}{\epsilon^{2}}\right)$ and a sample complexity bound of $\tilde{O}\left(\frac{d+k}{\epsilon^{2}}\right)$.
-Proof of Theorem 14.
-By Lemma 5 and Lemma 6, it suffices to show that by running Algorithm 3 with probability $41-\delta/$ 4, $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ for any $1\leq t\leq T$.
-Formally, we have the lemma below.
-Lemma 15.
-Suppose Assumption 10 holds. By running Algorithm 3, with probability $41-\delta/$ 4, it holds that
+Suppose Assumption 1 holds. With probability at least $41-\delta/$ 4, the iterates of Algorithm 3 satisfy
 $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (41)
 for any $1\leq t\leq T$.
-The proof of Lemma 15 is postponed to Appendix D
-7 Extension: oracle-efficient multi-group learning
-Given a hypothesis set $\mathcal{H}$ where each $h\in\mathcal{H}:\mathcal{X}\to\mathcal{Y}$, a set of groups $\mathcal{G}$ such that each $g\in\mathcal{G}$ is a subset of $\mathcal{X}$, and a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ and a distribution $\mathcal{D}$ with support $\mathcal{X}\times\mathcal{Y}$, define $L_{\mathcal{D}}(h|g):=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(h(x),y)|x\in g]$.
-Define $P_{g}=\mathrm{Pr}_{(x,y)\sim\mathcal{D}}[x\in g]$.
-Let $\gamma=\min_{g\in\mathcal{G}}P_{g}$.
-The goal of the learning algorithm is to find a (possibly randomized) hypothesis $h$ to minimize $L_{\mathcal{D}}(h|g)$ for all $g\in\mathcal{G}$ up to some threshold $\epsilon\in(0,1]$.
-That is,
-$L_{\mathcal{D}}(h|g)\leq\min_{h^{\prime}\in\mathcal{H}}L_{\mathcal{D}}(h^{\prime}|g)+\epsilon,\forall g\in\mathcal{G}.$
-We continue with the compatibility assumption as below, which ensures that such near-optimal hypothesis $h$ exists.
-Assumption 16.
-There exists $h^{*}\in\mathcal{H}$, such that
-$L_{\mathcal{D}}(h^{*}|g)\leq\min_{h\in\mathcal{G}}L_{\mathcal{D}}(h|g)+\frac{\epsilon}{4},\quad\forall g\in\mathcal{G}.$
-The learning algorithm
-As presented in Algorithm 4, we first sample $N=O\left(\frac{\log(|\mathcal{G}|/\delta)+d\log(d/\epsilon)}{\gamma\epsilon^{2}}\right)$ datapoints from $\mathcal{D}$ and then estimate the optimal value $f_{g}$ for each group $g\in\mathcal{G}$.
-By taking $L_{\mathcal{D}}(h|g)-f_{g}$ as the loss function, we then invoke a Hedge algorithm over $\mathcal{G}$ to learn the problem
-$\min_{h\in\Delta(\mathcal{H})}\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h|g)-f_{g}).$ (49)
-By Assumption 16, it suffices to find some (possibly randomized) $h$ such that
-$\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h|g)-f_{g}).\leq\min_{h^{\prime}\in\mathcal{H}}\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h^{\prime}|g)-f_{g})\leq\max_{g\in\mathcal{G}}(L_{\mathcal{D}}(h^{*}|g)-f_{g})=O(\epsilon).$
-Formally, we have the following theorem.
-Theorem 17.
-Suppose Assumption 16 holds. By running Algorithm 4, with probability, the output $h^{\mathsf{final}}$ satisfies that
-$L_{\mathcal{D}}(h^{\mathsf{final}}|g)=\frac{1}{T}\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)\leq\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)+\epsilon$ (50)
-for any $g\in\mathcal{G}$. Meanwhile, the sample complexity of Algorithm 4 is bounded by
-$O\left(\frac{d\log(d/\epsilon)+\log(|\mathcal{G}|/\delta)}{\gamma\epsilon^{2}}\right).$
-Moreover, Algorithm 4 only access $\mathcal{H}$ with an $\mathrm{ERM}$ oracle.
-Under Assumption 16, Theorem 17 recovers the result in Tosh and Hsu, (2022) without direct access to the hypothesis set $\mathcal{H}$.
-In high-level idea, the proof of Theorem 17 is based on the uniform convergence argument, which is a natural extension of our main results.
-We postpone the proof to Appendix E due to space limitation.
-8 Necessity of randomization
-Given that the best-known sample complexities prior to our work were derived for algorithms that either output randomized hypotheses or invoke majority votes, Awasthi et al., (2023) raised the question about how the sample complexity is impacted if only deterministic (or “proper”) hypotheses are permitted as the output of the learning algorithms.
-As it turns out, the restriction to deterministic hypotheses substantially worsens the sample efficiency, as revealed by the following theorem.
-Theorem 18.
-Assume that $d\geq 2\log(8k)$. Consider any $\varepsilon\in(0,1/100)$, and let $N_{0}=\frac{2^{d}-1}{k}$. One can find
-  • a hypothesis class $\mathcal{H}$ containing at most $kN_{0}+1$ hypothesis,
-  • a collection of $k$ distributions $\mathcal{D}\coloneqq\{\mathcal{D}_{i}\}_{i=1}^{k}$,
-  • a loss function :×$\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$,
-such that it takes at least $\frac{dk}{240000\varepsilon^{2}}$ samples to find $h\in\mathcal{H}$ obeying
-$\max_{1\leq i\leq k}\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}\leq\min_{h^{\prime}\in\mathcal{H}}\max_{i\in[k]}\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h^{\prime},(x,y)\big{)}\big{]}+\varepsilon$ (51)
-with probability exceeding $3/4$.
-Let us breifly desribe the high-level strategy for our construction of the hard instance: for each $i\in[k]$, we build a hypothesis set $\mathcal{H}_{i}$ that performs poorly solely on $\mathcal{D}_{i}$.
-To discriminate the optimal hypothesis — denoted by $h^{\star}$ — from some $\mathcal{H}_{i}$, the learner has to call $\Omega(\log(|\mathcal{H}_{i}|)/\varepsilon^{2})=\widetilde{\Omega}(d/\varepsilon^{2})$ times to $\mathrm{Query}(\mathcal{D}_{i})$ in expectation.
-The result follows by taking sum over $i\in[k]$.
-Proof of Theorem 18.
-Note that $N_{0}=\frac{2^{d}-1}{k}$.
-Set $N=kN_{0}+1=2^{d}$.
-Set $\mathcal{X}=\{-1,0,1\}^{kN}$.
-We set $\mathcal{Y}=\{1\}$ to be a set with only one element.
-Without loss of generality we write $l(h,(x,y))=l(h,x)$.
-We now introduce our construction.
-There are $N$ hypotheses in $\mathcal{H}$, where each hypothesis corresponds to $k$ dimensions of the ground set $\mathcal{X}$.
-Without loss of generality, for $h\in\mathcal{H}$, we let $\mathcal{I}_{h}=\{j_{h,i}\}_{i=1}^{k}$ be the $k$ dimensions related to $h$.
-Note that $\mathcal{I}_{h}\cap\mathcal{I}_{h^{\prime}}=\emptyset$ for $h\neq h^{\prime}$.
-We construct $\mathcal{H}$ as $\mathcal{H}=(\cup_{i=1}^{k}\mathcal{H}_{i})\cup\{h^{\star}\}$.
-Then we define $l(h,x)$ as $l(h,x)=\min_{i\in\mathcal{I}_{h},x_{i}\neq 0}x_{i}$.
-Now we design the $k$ distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$.
-Fix $i\in[k]$, we let $\mathrm{Pr}_{\mathcal{D}_{i}}[x]=\Pi_{\ell=1}^{kN}\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]$, where $\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]=\mathbb{I}[x_{\ell}=0]$ for $\ell\notin\{j_{h,i}|h\in\mathcal{H}\}$, $\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]=\frac{1}{2}\mathbb{I}[x_{\ell}=1]+\frac{1}{2}\mathbb{I}[x_{\ell}=-1]$ for $\ell\in\{j_{h,i}|h\notin\mathcal{H}_{i}\}$, and $\mathrm{Pr}_{\mathcal{D}_{i,\ell}}[x_{\ell}]=\left(\frac{1}{2}+4\varepsilon\right)\mathbb{I}[x_{\ell}=1]+\left(\frac{1}{2}-4\varepsilon\right)\mathbb{I}[x_{\ell}=-1]$ for $\ell\in\{j_{h,i}|h\in\mathcal{H}_{i}\}$.
-In words, for each $i\in[k]$, we draw $x\sim\mathcal{D}_{i}$ by independently draw each dimension of $x$, and there are $|\mathcal{H}|=N$ non-zero dimensions.
-Furthermore, for $x\sim\mathcal{D}_{i}$ and $h\in\mathcal{H}$, we have $l(h,x)=x_{j_{h,i}}$.
-By this construction, we have the following properties:
-  1. (i).
-    $\ell(h,x)\in[-1,1]$ for any $h\in\mathcal{H}$ and $x\in\mathcal{X}$;
-  2. (ii).
-    $\mathbb{E}_{x\sim\mathcal{D}_{i}}[\ell(h,x)]=\mathbb{E}_{x\sim\mathcal{D}_{i,j_{h,i}}}[x_{j_{h,i}}]=2\varepsilon\cdot\mathbb{I}[h\in\mathcal{H}_{i}]$ for any $i\in[k]$ and $h\in\mathcal{H}$;
-  3. (iii).
-    the only $\varepsilon$-optimal hypothesis is $h^{\star}$ because for any $h\neq h^{\star}$, there exists some $i$ such that $h\in\mathcal{H}_{i}$;
-  4. (iv).
-    $|\mathcal{H}|=N=kN_{0}+1=2^{d}$, which implies $\mathrm{VC}(\mathcal{H})\leq\log_{2}(N)\leq d$.
-For each call to $\mathrm{Query}(\mathcal{D}_{i})$, we can get independent observations $\{x_{j_{h,i}}\}_{h\in\mathcal{H}}$ where $x_{j_{h,i}}\sim D_{i,j_{h,i}}$ for each $h\in\mathcal{H}$.
-Now we let the number of calls to $\mathrm{Query}(\mathcal{D}_{i})$ be $M_{i}$ for $i\in[k]$.
-Our target is to show to distinguish $h^{\star}$ from $\mathcal{H}_{i}$, $M_{i}$ has to be at least $\Omega(d/\varepsilon^{2})$.
-Now suppose there is an algorithm $\mathcal{G}$ with numbers of samples $\{M_{i}\}_{i=1}^{k}$ such that the output is $h^{\star}$ with probability $\frac{3}{4}$.
-Let $\mathrm{Pr}_{\mathcal{G}}[\cdot]$ and $\mathbb{E}_{\mathcal{G}}[\cdot]$ denote respectively the probability and expectation under running the algorithm $\mathcal{G}$.
-Let $h_{\mathrm{out}}$ be the output hypothesis.
-It then follows that
-$\mathrm{Pr}_{\mathcal{G}}[h_{\mathrm{out}}=h^{\star}]\geq\frac{3}{4}.$
-Let $\Pi_{\mathcal{H}}$ be the set of permutations over $\mathcal{H}$.
-Let $\mathrm{U}(\Pi_{\mathcal{H}})$ be the uniform distribution over $\Pi(\mathcal{H})$.
-With a slight abuse of notations, for $x\in\{-1,0,1\}^{kN}$ and $\sigma\in\Pi_{\mathcal{H}}$, we define $\sigma(x)$ to be the vector $y$ such that $y_{j_{h,i}}=x_{j_{\sigma(h),i}}$ for all $h\in\mathcal{H}$ and $i\in[k]$.
-Let $\mathcal{G}^{\prime}$ be the algorithm with $\mathcal{H}$ replaced by $\sigma(\mathcal{H})$ in the input where $\sigma\sim\mathrm{U}(\Pi_{\mathcal{H}})$.
-Because $\mathcal{G}$ returns the optimal hypothesis with probability at least $3/4$ for all problem instances, we have that
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star}]\geq\frac{3}{4}.$
-We note that $\mathcal{G}^{\prime}$ is a symmetric algorithm with respect to the hypothesis set. Formally, we have the lemma below to bound the probability of outputting a non-optimal hypothesis.
-Lemma 19.
-Fix $m\geq 0$ and $\widetilde{i}\in[k]$.
-Suppose $\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{i}\leq m]\geq\frac{1}{2}$.
-It then holds that for any $h\in\mathcal{H}_{\widetilde{i}}$
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m]\geq\frac{1}{2}\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m]\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}).$
-Moreover, it holds that $m\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}$.
-By Lemma 19, we have that
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right]\leq\frac{1}{2}.$ (52)
-Noting that $\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star}]\geq\frac{3}{4}$, we learn that
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}]\geq\frac{1}{4},$ (53)
-which implies that $\mathbb{E}_{\mathcal{G}^{\prime}}[M_{\widetilde{i}}]\geq\frac{\log(N_{0}/4)}{120000\varepsilon^{2}}\geq\frac{d-\log_{2}(8k)}{120000\varepsilon^{2}}\geq\frac{d}{240000\varepsilon^{2}}$.
-Taking sum over $i\in[k]$, we have that
-$\mathbb{E}_{\mathcal{G}^{\prime}}\left[\sum_{i=1}^{k}M_{\widetilde{i}}\right]\geq\frac{dk}{240000\varepsilon^{2}}.$ (54)
-The proof is completed.
+The proof of Lemma 6 is postponed to Appendix E. Combine this with Lemma 2 and Lemma 3 to show that: the total number of samples collected in Algorithm 3 is upper bounded by
+$T_{1}\,\mathrm{poly}\log\Big{(}k,\frac{1}{\varepsilon},\frac{1}{\delta}\Big{)}\leq\left(\frac{k}{\varepsilon^{2}}+\min\bigg{\{}C_{n}\mid C_{n}\leq\frac{1}{4800}\varepsilon\bigg{\}}\right)\,\mathrm{poly}\log\Big{(}k,\frac{1}{\varepsilon},\frac{1}{\delta}\Big{)}$
+as claimed. ∎
 8 Discussion
-In this paper, we have settled the problem of achieving optimal sample complexity in multi-distribution learning, assuming the availability of adaptive (or on-demand) sampling.
+In this paper, we have settled the problem of achieving optimal sample complexity in multi-distribution learning, assuming availability of adaptive (or on-demand) sampling.
 We have put forward a novel oracle-efficient algorithm that provably attains a sample complexity of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ for VC classes, which matches the best-known lower bound up to some logarithmic factor.
-From the technical perspective, the key novelty of our analysis lies in carefully bounding the trajectory of the Hedge algorithm on a convex (concave) optimization problem.
+From the technical perspective, the key novelty of our analysis lies in carefully bounding the trajectory of the Hedge algorithm on a convex-concave optimization problem.
 We have further unveiled the necessesity of randomization, revealing that a considerably larger sample size is necessary if the learning algorithm is constrained to return deterministic hypotheses.
-Notably, our work manages to solve three open problems presented in COLT 2023 (namely, Awasthi et al., (2023, Problems 1, 3 and 4)).
+Notably, our work manages to solve three open problems presented in COLT 2023 (namely, awasthi2023sample).
 Our work not only addresses existing challenges but also opens up several directions for future exploration.
@@
 Furthermore, can we extend our current analysis (which bounds the dynamics of the Hedge algorithm) to control the trajectory of more general first-order/second-order algorithms, in the context of robust online learning?
-Another venue for exploration is the extension of our multi-distribution learning framework to tackle other related tasks like multi-calibration (Hébert-Johnson et al.,, 2018; Haghtalab et al.,, 2023).
+Another venue for exploration is the extension of our multi-distribution learning framework to tackle other related tasks like multi-calibration (hebert2018multicalibration; haghtalab2023unifying).
 We believe that our algorithmic and analysis framework can shed light on making progress in all of these directions.
@@
 We begin by introducing three handy concentrations inequalities.
-The first result is the well-renowned Freedman inequality (Freedman,, 1975), which assists in deriving variance-aware concentration inequalities for martingales.
-Lemma 20 (Freedman’s inequality (Freedman,, 1975)).
+The first result is the well-renowned Freedman inequality (freedman1975tail), which assists in deriving variance-aware concentration inequalities for martingales.
+Lemma 7 (Freedman’s inequality (freedman1975tail)).
 Let $(M_{n})_{n\geq 0}$ be a martingale obeying $M_{0}=0$.
@@
 The second concentration result bounds the difference between the sum of a sequence of random variables and the sum of their respective conditional means (w.r.t. the associated $\sigma$-algebra).
-Lemma 21 (Lemma 10 in Zhang et al., (2022)).
+Lemma 8 (Lemma 10 in zhang2022horizon).
 Let $X_{1},X_{2},\ldots$ be a sequence of random variables taking value in the interval $[0,l]$.
@@
 where the penultimate inequality uses $x\in[0,\log(2)]$, and the last inequality holds since $y\geq q$. ∎
-Appendix B Missing proofs for optimality
+Finally, let us present a basic property related to Rademacher random variables, which will play a useful role in understanding the Rademacher complexity.
+Lemma 12.
+Let $\mathcal{L}$ be a set of vectors in $\mathbb{R}^{n}$. Let $w^{1},w^{2}\in\mathbb{R}^{n}$ be two vectors obeying $|w^{1}_{i}|\leq|w_{i}^{2}|$ for all $i\in[n]$. Then it holds that
+$\mathop{\mathbb{E}}_{\{\sigma_{i}\}_{i=1}^{n}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{1}_{i}f_{i}\right]\leq\mathop{\mathbb{E}}_{\{\sigma_{i}\}_{i=1}^{n}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{2}_{i}f_{i}\right],$ (44)
+where $\{\sigma_{i}\}$ is a collection of independent Rademacher random variables obeying $\mathbb{P}(\sigma_{i}=1)=\mathbb{P}(\sigma_{i}=-1)=1/2$.
+Proof.
+Clearly, it suffices to prove (44) for the special case where $w_{i}^{1}=w^{2}_{i}$ for $11\leq i\leq n-$ 1, and $|w^{1}_{n}|\leq|w^{2}_{n}|$.
+Fixing $\sigma_{i}$ for $11\leq i\leq n-$ 1, we can deduce that
+$\mathop{\mathbb{E}}_{\sigma_{n}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w_{i}f_{i}\right]$ $=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}f_{i}+w_{n}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}f_{i}-w_{n}f_{n}\right)$
+$=\frac{1}{2}\max_{f,\widetilde{f}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}(f_{i}+\widetilde{f}_{i})+w_{n}(f_{n}-\widetilde{f}_{n})\right)$
+$=\frac{1}{2}\max_{f,\widetilde{f}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}(f_{i}+\widetilde{f}_{i})+w_{n}(f_{n}-\widetilde{f}_{n})\right)$
+$\leq\frac{1}{2}\max_{f,\widetilde{f}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}(f_{i}+\widetilde{f}_{i})+\big{|}\widetilde{w}_{n}(f_{n}-\widetilde{f}_{n})\big{|}\right)$
+$=\frac{1}{2}\max_{f,\widetilde{f}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}(f_{i}+\widetilde{f}_{i})+\widetilde{w}_{n}(f_{n}-\widetilde{f}_{n})\right)$
+$=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}f_{i}+\widetilde{w}_{n}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}\widetilde{w}_{i}f_{i}-\widetilde{w}_{n}f_{n}\right)$
+$=\mathop{\mathbb{E}}_{\sigma_{n}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}\widetilde{w}_{i}f_{i}\right].$
+The proof is thus completed by taking expectation over $\{\sigma_{i}\}_{i=1}^{n-1}$. ∎
+Appendix B Proofs of auxiliary lemmas for VC classes
 B.1 Proof of Lemma 1
@@
 Step 1: concentration bounds for any fixed $n=\{n_{i}\}_{i=1}^{k}$ and $w\in\Delta(k)$.
-Consider any fixed $n=\{n_{i}\}_{i=1}^{k}$ obeying $0\leq n_{i}\leq T_{1}$ for all $i\in[k]$, and any fixed $w\in\Delta(k)$.
+Consider first any fixed $n=\{n_{i}\}_{i=1}^{k}$ obeying $0\leq n_{i}\leq T_{1}$ for all $i\in[k]$, and any fixed $w\in\Delta(k)$.
 For any quantity $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$, if we take
-$E(\lambda)\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right\}\right)\right]$ (57)
-with the expectation taken over the randomness of $\widetilde{\mathcal{S}}(n)$, then we can perform a standard “symmetrization” trick to bound $E(\lambda)$ as follows:
-$E(\lambda)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right\}\right)\right]$
+$E(\lambda,n,w)\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w)\right\}\right)\right]$ (45)
+with the expectation taken over the randomness of $\widetilde{\mathcal{S}}(n)$, then we can apply a standard “symmetrization” trick to bound $E(\lambda,n,w)$ as follows:
+$E(\lambda,n,w)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w)\right\}\right)\right]$
 $=\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}^{+}(n)}\left[\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\right]\right\}\right)\right]$
@@
 $=\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n),\widetilde{\mathcal{S}}^{+}(n)}\left[\mathop{\mathbb{E}}\limits_{\sigma(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i,j}\Big{(}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\Big{)}\right\}\right)\,\Big{|}\,\mathcal{C}\right]\right].$ (47)
-Note that for any dataset $\mathcal{C}$ with cardinality $|\mathcal{C}|$, the Sauer–Shelah lemma (Wainwright,, 2019, Proposition 4.18) together with our assumption that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ tells us that the cardinality of the following set obeys
-$\big{|}\mathcal{H}(\mathcal{C})\big{|}\leq(|\mathcal{C}|+1)^{d}\leq(2kT_{1}+1)^{d},$ (60)
+Note that for any dataset $\mathcal{C}$ with cardinality $|\mathcal{C}|$, the Sauer–Shelah lemma (wainwright2019high, Proposition 4.18) together with our assumption that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ tells us that the cardinality of the following set obeys
+$\big{|}\mathcal{H}(\mathcal{C})\big{|}\leq(|\mathcal{C}|+1)^{d}\leq\big{(}|\widetilde{\mathcal{S}}|+|\widetilde{\mathcal{S}}^{+}|+1\big{)}^{d}\leq(2kT_{1}+1)^{d},$ (48)
 where $\mathcal{H}(\mathcal{C})$ denotes the set obtained by applying all $h\in\mathcal{H}$ to the data points in $\mathcal{C}$, namely,
@@
 Here, the last inequality makes use of fact $\big{|}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}|\leq 2$ as well as the following elementary inequality
-$\mathop{\mathbb{E}}\limits_{\sigma_{i,j}}\big{[}\exp(\sigma_{i,j}x)\big{]}=\frac{1}{2}\big{(}\exp(x)+\exp(-x)\big{)}\leq\exp\big{(}0.5x^{2}\big{)}.$
-Taking (58) and (62) together and applying the Markov inequality reveal that, for any $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$,
-$E(\lambda)\leq(2\|n\|_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right).$ (63)
-Repeating the same arguments also yields
-$\overline{E}(\lambda)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{\sum_{i=1}^{k}w_{i}L(h,e_{i})-\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\right\}\right)\right]$
+$\mathop{\mathbb{E}}\limits_{\sigma_{i,j}}\big{[}\exp(\sigma_{i,j}x)\big{]}=\frac{1}{2}\big{(}\exp(x)+\exp(-x)\big{)}\leq\exp\big{(}x^{2}\big{)}.$
+Taking (46), (47) and (50) together reveals that
+$E(\lambda)\leq(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right).$ (51)
+Repeating the same arguments also yields an upper bound on the following quantity:
+$\overline{E}(\lambda)$ $\coloneqq\mathop{\mathbb{E}}\limits_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\exp\left(\lambda\left\{L(h,w)-\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\right\}\right)\right]$
 $\leq(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right)$
-for any $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$. Taking the above two inequalities reveals that, for any $0<\varepsilon^{\prime}\leq 1$,
-$\mathbb{P}\left(\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}\right)$
+for any $\lambda\in\big{[}0,\min_{i\in[k]}\frac{n_{i}}{w_{i}}\big{]}$. Taking the above two inequalities and applying the Markov inequality reveal that, for any $0<\varepsilon^{\prime}\leq 1$,
+$\mathbb{P}\left(\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w)\right|\geq\varepsilon^{\prime}\right)$
 $\qquad\leq\min_{0\leq\lambda\leq\min_{i}\frac{n_{i}}{w_{i}}}\frac{E(\lambda)+\overline{E}(\lambda)}{\exp(\lambda\varepsilon^{\prime})}$
 $\qquad\leq\min_{0\leq\lambda\leq\min_{i}\frac{n_{i}}{w_{i}}}2\cdot(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}-\lambda\varepsilon^{\prime}\right).$ (52)
-Step 2: uniform concentration bounds over all $n$ and $w$.
-Next, we move on to extend the above result to uniform concentration bounds. Towards this, let us first introduce a couple of notation.
+Step 2: uniform concentration bounds over epsilon-nets w.r.t. $n$ and $w$.
+Next, we move on to extend the above result to uniform concentration bounds over all possible $n$ and $w$. Towards this, let us first introduce a couple of notation.
   • Let us use $\Delta_{\varepsilon_{2}}(k)\subseteq\Delta(k)$ to denote an $\varepsilon_{2}$-net of $\Delta(k)$ — namely, for any $x\in\Delta(k)$, there exists a vector $x_{0}\in\Delta_{\varepsilon_{2}}(k)$ obeying $\|x-x_{0}\|_{\infty}\leq\varepsilon_{2}$.
@@
   • Define the following set
-    =$\mathcal{B}=\Big{\{}n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\mid T_{1}w_{i}\leq 2n_{i},0\leq n_{i}\leq T_{1},\forall i\in[k],w\in\Delta_{\varepsilon_{1}/(8k)}(k)\Big{\}}.$
-Applying the union bound yields
-$\mathbb{P}\left(\exists(n,w)\in\mathcal{B},\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}\right)$
+    =$\mathcal{B}=\bigg{\{}n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\,\Big{|}\,\frac{n_{i}}{w_{i}}\geq\frac{T_{1}}{2},0\leq n_{i}\leq T_{1},\forall i\in[k],w\in\Delta_{\varepsilon_{1}/(8k)}(k)\bigg{\}},$
+    which clearly satisfies
+    $|\mathcal{B}|\leq T_{1}^{k}\cdot\left(\frac{8k}{\varepsilon_{1}}\right)^{k}.$
+Applying the union bound yields that, for any $0<\varepsilon^{\prime}\leq 1$,
+$\mathbb{P}\left(\exists(n,w)\in\mathcal{B},\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w_{i}\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w)\right|\geq\varepsilon^{\prime}\right)$
 $\leq\sum_{(n,w)\in\mathcal{B}}\min_{0\leq\lambda\leq\min_{i}\frac{n_{i}}{w_{i}}}2\cdot(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}-\lambda\varepsilon^{\prime}\right)$
-$\leq\sum_{(n,w)\in\mathcal{B}}\min_{0\leq\lambda\leq\frac{T_{1}}{2}}2\cdot(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\frac{2}{T_{1}}-\lambda\varepsilon^{\prime}\right)$
+$\leq\sum_{(n,w)\in\mathcal{B}}\min_{0\leq\lambda\leq\frac{T_{1}}{2}}2\cdot(2kT_{1}+1)^{d}\exp\left(2\lambda^{2}\cdot\frac{2}{T_{1}}-\lambda\varepsilon^{\prime}\right)$
 $\leq\sum_{(n,w)\in\mathcal{B}}2\cdot(2kT_{1}+1)^{d}\exp\left(-\frac{T_{1}(\varepsilon^{\prime})^{2}}{16}\right)$
-$\leq|\mathcal{B}|\cdot 2\cdot(2kT_{1}+1)^{d}\exp\left(\frac{-T_{1}(\varepsilon^{\prime})^{2}}{16}\right)$
-$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(\frac{-T_{1}(\varepsilon^{\prime})^{2}}{16}\right),$
+$\leq|\mathcal{B}|\cdot 2\cdot(2kT_{1}+1)^{d}\exp\left(-\frac{T_{1}(\varepsilon^{\prime})^{2}}{16}\right)$
+$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(-\frac{T_{1}(\varepsilon^{\prime})^{2}}{16}\right),$
 where the second inequality holds since $\sum_{i=1}^{k}\frac{w_{i}^{2}}{n_{i}}\leq\frac{2}{T_{1}}\sum_{i=1}^{k}w_{i}=\frac{2}{T_{1}}$ (according to the definition of $\mathcal{B}$).
+Step 3: concentration bounds w.r.t. $n^{t}$ and $w^{t}$.
 Let $\mathcal{S}^{t}$ denote the value of $\mathcal{S}$ after line 1 of Algorithm 1 in the $t$-th round.
-Recall that $n^{t}=[n^{t}_{i}]_{1\leq i\leq k}$ denotes the number of samples in $\mathcal{S}^{t}$.
-Consequently, we can take $\mathcal{S}^{t}=\widetilde{\mathcal{S}}(n^{t})$.
+Recall that $n^{t}=[n^{t}_{i}]_{1\leq i\leq k}$ denotes the number of samples for all $k$ distributions in $\mathcal{S}^{t}$, and let $w^{t}=[w^{t}_{i}]_{1\leq i\leq k}$ represent the weight iterates in the $t$-th round.
+It is easily seen from lines 1 and 1 of Algorithm 1 that $n_{i}^{t}\leq T_{1}$ and $n_{i}^{t}/w_{i}^{t}\geq n_{i}^{t}/(2\widehat{w}_{i}^{t})\geq T_{1}/2$.
+For analysis purposes, it suffices to take $\mathcal{S}^{t}=\widetilde{\mathcal{S}}(n^{t})$.
 In view of the update rule in Algorithm 1, one can always find $(n^{t},\widetilde{w}^{t})\in\mathcal{B}$ satisfying $\|\widetilde{w}^{t}-w^{t}\|_{1}\leq k\|\widetilde{w}^{t}-w^{t}\|_{\infty}\leq\varepsilon_{1}/8$.
 As a result, for any $0<\varepsilon^{\prime}\leq 1$, we can deduce that
-$\mathbb{P}\left(\exists t\in[T],\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w^{t}_{i}\frac{1}{n^{t}_{i}}\sum_{i=1}^{n^{t}_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w^{t}_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}+\frac{\varepsilon_{1}}{4}\right)$
-$\qquad\leq\mathbb{P}\left(\exists t\in[T],\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\widetilde{w}^{t}_{i}\frac{1}{n^{t}_{i}}\sum_{i=1}^{n^{t}_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}\widetilde{w}^{t}_{i}L(h,e_{i})\right|\geq\varepsilon^{\prime}\right)$
-$\qquad\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(\frac{-T_{1}(\varepsilon^{\prime})^{2}}{16}\right),$ (65)
-where the second inequality arises from the fact that $\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\in[-1,1]$ and $L(h,e_{i})\in[-1,1]$.
-Substituting $\varepsilon^{\prime}=\frac{\varepsilon_{1}}{4}$ and $T_{1}=\frac{400\left(k\log(k/\varepsilon_{1})+d\log(kd/\varepsilon_{1})+\log(1/\delta)\right)}{\varepsilon_{1}^{2}}$ into (65), we can obtain
-$\mathbb{P}\left(\exists t\in[T],\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w^{t}_{i}\cdot\frac{1}{n^{t}_{i}}\sum_{i=1}^{n^{t}_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w^{t}_{i}L(h,e_{i})\right|\geq\frac{\varepsilon_{1}}{2}\right)$
-$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(\frac{-T_{1}\varepsilon_{1}^{2}}{16}\right)$
+$\mathbb{P}\left(\exists t\in[T],\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w^{t}_{i}\frac{1}{n^{t}_{i}}\sum_{i=1}^{n^{t}_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w^{t})\right|\geq\varepsilon^{\prime}+\frac{\varepsilon_{1}}{4}\right)$
+$\qquad\leq\mathbb{P}\left(\exists t\in[T],\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\widetilde{w}^{t}_{i}\frac{1}{n^{t}_{i}}\sum_{i=1}^{n^{t}_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,\widetilde{w}^{t})\right|\geq\varepsilon^{\prime}\right)$
+$\qquad\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(-\frac{T_{1}(\varepsilon^{\prime})^{2}}{16}\right),$ (53)
+where the second inequality arises from the fact that $\frac{1}{n_{i}}\sum_{i=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\in[-1,1]$ and $L(h,\widetilde{w}^{t})\in[-1,1]$.
+Taking $\varepsilon^{\prime}=\varepsilon_{1}/4$ and substituting $T_{1}=\frac{4000\left(k\log(k/\varepsilon_{1})+d\log(kd/\varepsilon_{1})+\log(1/\delta)\right)}{\varepsilon_{1}^{2}}$ into (53), we can obtain
+$\mathbb{P}\left(\exists t\in[T],\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}w^{t}_{i}\cdot\frac{1}{n^{t}_{i}}\sum_{i=1}^{n^{t}_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-L(h,w^{t})\right|\geq\frac{\varepsilon_{1}}{2}\right)$
+$\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\left(-\frac{T_{1}\varepsilon_{1}^{2}}{16}\right)$
 $\leq 2\cdot(8kT_{1}/\varepsilon_{1})^{k}(2kT_{1}+1)^{d}\cdot\exp\Big{(}-10\big{(}k\log(k/\varepsilon_{1})+d\log(kd/\varepsilon_{1})+\log(1/\delta)\big{)}\Big{)}$
@@
 $\leq\delta/4.$ (54)
-Step 3: putting all this together.
+Step 4: putting all this together.
 Recalling that
@@
 we can immediately deduce that
-$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}.$ (69)
+$L(h^{t},w^{t})\leq\widehat{L}(h^{t},w^{t})+\frac{\varepsilon_{1}}{2}=\min_{h\in\mathcal{H}}\widehat{L}(h,w^{t})+\frac{\varepsilon_{1}}{2}\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}.$ (57)
 This concludes the proof of Lemma 1.
-Appendix C Missing proofs for sample complexity
-We let $\delta^{\prime}=\delta/(32T^{4}k^{2})$ throughout this section.
-C.1 Proof of Lemma 8
-Define $\mathcal{W}_{j}=\{i\in[k]|\max_{1\leq t\leq T}w_{i}^{t}\in(2^{-j},2^{-(j-1)}]\}$ for $11\leq j\leq\left\lfloor\log_{2}(k)\right\rfloor+$ 1 and $\overline{\mathcal{W}}=[K]/\cup_{j}\mathcal{W}_{j}$.
-We consider to bound the size of each $\mathcal{W}_{j}$.
+B.2 Proof of Lemma 2
+Before proceeding, let us introduce some additional notation. Let $\delta^{\prime}\coloneqq\frac{\delta}{4(T+k+1)}$, and define
+$\mathsf{OPT}\coloneqq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}L(h,\mathrm{e}^{\mathsf{basis}}_{i})$
+to be the optimal objective value. Additionally, set
+$v^{t}\coloneqq L(h^{t},w^{t})-\mathsf{OPT}.$ (58)
+It follows from the assumption of this lemma (i.e., $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$) that
+$v^{t}\leq\min_{h\in\mathcal{H}}L(h,w^{t})-\mathsf{OPT}+\varepsilon_{1}=\min_{h\in\mathcal{H}}L(h,w^{t})-\min_{h\in\mathcal{H}}\max_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})+\varepsilon_{1}\leq\varepsilon_{1},\qquad\forall 1\leq t\leq T.$ (59)
+We now begin to present the proof. In view of the Azuma-Hoeffding inequality and the union bound, we see that with probability at least $1-(k+1)\delta^{\prime}$,
+$\left|\sum_{t=1}^{T}\big{\langle}w^{t},\widehat{r}^{t}\big{\rangle}-\sum_{t=1}^{T}L(h^{t},w^{t})\right|$ $\leq 2\sqrt{T\log(1/\delta^{\prime})},$ (60a)
+$\left|\sum_{t=1}^{T}\widehat{r}_{i}^{t}-\sum_{t=1}^{T}L(h^{t},\mathrm{e}^{\mathsf{basis}}_{i})\right|$ $\leq 2\sqrt{T\log(1/\delta^{\prime})}.$ (60b)
+These motivate us to look at $\sum_{t=1}^{T}\big{\langle}w^{t},\widehat{r}^{t}\big{\rangle}$ (resp.
+$\sum_{t=1}^{T}\widehat{r}_{i}^{t}$) as a surrogate for $\sum_{t=1}^{T}L(h^{t},w^{t})$ (resp.
+$\sum_{t=1}^{T}L(h^{t},\mathrm{e}^{\mathsf{basis}}_{i})$).
+We then resort to standard analysis for the Hedge algorithm. Specifically, direct computation gives
+[…]
+$\leq\log\left(1+\eta\sum_{i=1}^{k}w_{i}^{t}\widehat{r}_{i}^{t}+\eta^{2}\sum_{i=1}^{k}w_{i}^{t}(\widehat{r}_{i}^{t})^{2}\big{)}\right)\leq\eta\sum_{i=1}^{k}w_{i}^{t}\widehat{r}_{i}^{t}+\eta^{2}.$ (61)
+Here, (i) is valid since $w_{i}^{t}=\frac{W_{i}^{t}}{\sum_{j}W_{j}^{t}}$ and $W_{i}^{t+1}=W_{i}^{t}\exp(\eta\widehat{r}_{i}^{t})$ (cf.
+lines 1 and 1 of Algorithm 1); (ii) arises from the elementary inequality $e^{x}\leq 1+x+x^{2}$ for $x\in[0,1]$ as well as the facts that $\eta\leq 1$ and $|\widehat{r}_{i}^{t}|\leq 1$.
+Summing the inequality (61) over all $t$ and rearranging terms, we are left with
+$\eta\sum_{t=1}^{T}\big{\langle}w^{t},\widehat{r}^{t}\big{\rangle}$ $\geq\sum_{t=1}^{T}\left\{\log\left(\frac{\sum_{i=1}^{k}W_{i}^{t+1}}{\sum_{i=1}^{k}W_{i}^{t}}\right)-\eta^{2}\right\}$
+$=\log\left(\sum_{i=1}^{k}W_{i}^{T+1}\right)-\log\left(\sum_{i=1}^{k}W_{i}^{1}\right)-T\eta^{2}$
+$\geq\max_{1\leq i\leq k}\log(W_{i}^{T+1})-\log(k)-T\eta^{2}$
+$\geq\eta\max_{1\leq i\leq k}\sum_{t=1}^{T}\widehat{r}_{i}^{t}-\log(k)-T\eta^{2},$ (62)
+where the penultimate lines makes use of $W_{i}^{1}=1$ for all $i\in[k]$, and the last line holds since $\log\big{(}W_{i}^{T+1}\big{)}=\log\big{(}W_{i}^{T}\exp(\eta\widehat{r}_{i}^{t})\big{)}\geq\eta\widehat{r}_{i}^{t}$.
+Dividing both sides by $\eta$ yields
+$\sum_{t=1}^{T}\big{\langle}w^{t},\widehat{r}^{t}\big{\rangle}\geq\max_{i}\sum_{t=1}^{T}\widehat{r}^{t}_{i}-\left(\frac{\log(k)}{\eta}+\eta T\right).$ (63)
+Combine the above inequality with (60) to show that, with probability at least $1-(k+1)\delta^{\prime}$,
+$\sum_{t=1}^{T}L(h^{t},w^{t})$ $\geq\max_{1\leq i\leq k}\sum_{t=1}^{T}L(h^{t},\mathrm{e}^{\mathsf{basis}}_{i})-\left(\frac{\log(k)}{\eta}+\eta T+4\sqrt{T\log(1/\delta^{\prime})}\right).$ (64)
+Recalling that $\varepsilon_{1}=\eta=\frac{1}{100}\varepsilon$ and $T=\frac{20000\log(\frac{k}{\delta^{\prime}\varepsilon})}{\varepsilon^{2}}$, we can derive
+$\max_{1\leq i\leq k}\sum_{t=1}^{T}L(h^{t},\mathrm{e}^{\mathsf{basis}}_{i})$ $\leq T\mathsf{OPT}+\sum_{t=1}^{T}v^{t}+\left(\frac{\log(k)}{\eta}+\eta T+4\sqrt{T\log(1/\delta^{\prime})}\right)$
+$\leq T\mathsf{OPT}+T\varepsilon_{1}+\left(\frac{\log(k)}{\eta}+\eta T+4\sqrt{T\log(1/\delta^{\prime})}\right)$
+$\leq T\mathsf{OPT}+T\varepsilon,$ (65)
+where the penultimate line results from (59). Given that $h^{\mathsf{final}}$ is taken to be uniformly distributed over $\{h^{t}\}_{1\leq t\leq T}$, we arrive at
+$\max_{1\leq i\leq k}L(h^{\mathsf{final}},\mathrm{e}^{\mathsf{basis}}_{i})=\max_{1\leq i\leq k}\frac{1}{T}\sum_{t=1}^{T}L(h^{t},\mathrm{e}^{\mathsf{basis}}_{i})\leq\mathsf{OPT}+\varepsilon$ (66)
+with probability at least $1-(k+1)\delta^{\prime}$. This concludes the proof by recalling that $\delta^{\prime}=\frac{\delta}{4(T+k+1)}$.
+Remark 4.
+Note that the proof of this lemma works as long as $\widehat{r}_{i}^{t}\in[0,1]$ is an unbiased estimate of $L(h^{t},\mathrm{e}^{\mathsf{basis}}_{i})$ for each $i\in[k]$, regardless of how many samples are used to construct $\widehat{r}_{i}^{t}$.
+B.3 Proof of Lemma 3
+Set $\delta^{\prime}=\delta/(32T^{4}k^{2})$, and let
+$\overline{j}=\left\lfloor\log_{2}\left(\frac{k\log^{2}(2)}{50(\log_{2}(1/\eta)+1)^{2}\log_{2}^{2}(k)}\right)\right\rfloor-2.$ (67)
+Let us define
+$\mathcal{W}_{j}$$\coloneqq\big{\{}i\in[k]\mid\max_{1\leq t\leq T}w_{i}^{t}\in(2^{-j},2^{-(j-1)}]\big{\}},\qquad 1\leq j\leq\overline{j}$ (68a)
+$\overline{\mathcal{W}}$ $\coloneqq[k]/\cup_{j}\mathcal{W}_{j}.$ (68b)
+In other words, we divide the $k$ distributions into a logarithmic number of groups $\{\mathcal{W}_{j}\}$, where each $\mathcal{W}_{j}$ consists of those distributions whose corresponding $\max_{t}w_{i}^{t}$ are on the same order.
+The main step in establishing Lemma 3 lies in bounding the size of each $\mathcal{W}_{j}$, as summarized below.
 Lemma 13.
-Assume the conditions in Lemma 8 hold. For any $21\leq j\leq\log_{2}(k)-$ 2, with probability $1-8T^{4}k\delta^{\prime}$,
-$|\mathcal{W}_{j}|\leq 8\cdot 10^{7}\cdot\left((\log_{2}(k)+1)^{4}\left(\log(k)+\log(1/\delta^{\prime})\right)^{2}(\log_{2}(T)+1)\right)\cdot 2^{j}.$
-Now we prove Lemma 8. With probability $41-8T^{4}k^{2}\delta^{\prime}=1-\delta/$ 4, we have
-$\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ $\leq k\cdot\frac{16}{k}+\sum_{j=1}^{\left\lfloor\log_{2}(k)]\right\rfloor-2}|\mathcal{W}_{j}|2^{-j+1}$
-$\leq 2\cdot 10^{8}\cdot\left((\log_{2}(k)+1)^{5}\left(\log(Tk)+\log(1/\delta)\right)^{2}(\log_{2}(T)+1)\right).$
-The proof of Lemma 8 is finished.
+Suppose that the assumptions of Lemma 3 hold. Then with probability exceeding $1-8T^{4}k\delta^{\prime}$,
+$|\mathcal{W}_{j}|\leq 8\cdot 10^{7}\cdot\left(\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\log_{2}^{2}(k)\left(\log(k)+\log(1/\delta^{\prime})\right)^{3}\big{(}\log_{2}(T)+1\big{)}\right)\cdot 2^{j}$ (69)
+holds all $1\leq j\leq\overline{j}$, with $\overline{j}$ defined in (67).
+In words, Lemma 13 asserts that the cardinality of each $\mathcal{W}_{j}$ is upper bounded by
+$|\mathcal{W}_{j}|\leq\widetilde{O}(2^{j}),\qquad 1\leq j\leq\overline{j}.$
+Importantly, this lemma tells us that, with probability at least $41-8T^{4}k^{2}\delta^{\prime}=1-\delta/$ 4, one has
+$\|\overline{w}^{T}\|_{1}=\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$ $\leq k\cdot 2^{-(\overline{j}-1)}+\sum_{j=1}^{\overline{j}}|\mathcal{W}_{j}|2^{-(j-1)}$
+$\leq k\cdot\frac{800\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\log_{2}^{2}(k)}{k\log^{2}(2)}+\sum_{j=1}^{\overline{j}}|\mathcal{W}_{j}|2^{-(j-1)}$
+$\leq 2\cdot 10^{8}\cdot\left(\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\log_{2}^{2}(k)\big{(}\log(Tk)+\log(1/\delta)\big{)}^{3}\big{(}\log_{2}(T)+1\big{)}\right),$
+where the first inequality is valid since $\max_{1\leq t\leq T}w_{i}^{t}\leq 2^{-(j-1)}$ holds for any $i\in\mathcal{W}_{j}$.
+This immediately concludes the proof of Lemma 3, as long as Lemma 13 can be established.
+Noteworthily, proving Lemma 13 is the most challenging part of our analysis, and we dedicate the next section (Appendix C) to the proof of Lemma 13.
+Appendix C Controlling the Hedge trajectory (proof of Lemma 13)
+This section is devoted to proving Lemma 13. The proof relies heavily on the concepts of “segments” introduced in Section 4.2. For convenience, we restate the definition below.
+Definition 4 (Segment (restated)).
+For any $p,x>0$ and $i\in[k]$, we say that $(t_{1},t_{2})$ is a $(p,q,x)$-segment if there exists a subset $\mathcal{I}\subseteq[k]$ such that
+  • $\mathrm{(i)}$
+    $\sum_{i\in\mathcal{I}}w^{t_{1}}_{i}\in\left[p/2,p\right]$,
+  • $\mathrm{(ii)}$
+    $\sum_{i\in\mathcal{I}}w^{t_{2}}_{i}\geq p\exp(x)$,
+  • $\mathrm{(iii)}$
+    $\sum_{i\in\mathcal{I}}w_{i}^{t}\geq q$ for any $t_{1}\leq t\leq t_{2}$.
+We shall refer to $t_{1}$ as the starting point and $t_{2}$ as the end point, and call $\mathcal{I}$ the index set.
+Moreover, two segments $(s_{1},e_{1})$ and $(s_{2},e_{2})$ are said to be disjoint if either $s_{1}<e_{1}\leq s_{2}<e_{2}$ or $s_{2}<e_{2}\leq s_{1}<e_{1}$.
+In addition, throughout this section, we shall take $\delta^{\prime}=\delta/(32T^{4}k^{2})$, and focus on any $j$ obeying
+$2\displaystyle 1\leq j\leq\log_{2}(k)-$ 2 . (70)
+C.1 Main steps of the proof
+In this subsection, we present the main steps of the proof.
+Before continuing, we find it helpful to underscore a high-level idea: if $|\mathcal{W}_{j}|$ is large, then there exist many disjoint segments, thereby requiring the total length $T$ to be large enough in order to contain these segments.
+The key steps to construct such disjoint segments of interest are as follows:
+  1. 1.
+    Construct a suitable segment for each $i\in\mathcal{W}_{j}$ (see Lemma 14);
+  2. 2.
+    Identify sufficiently many disjoint blocks such that the segments within each block have nonempty intersection (see Lemma 15 and Figure 3);
+  3. 3.
+    From the above disjoint blocks, identify sufficiently many disjoint subsets such that the distributions associated with each subset can be linked with a common (sub)-segment, and that each of these (sub)-segments experiences sufficient changes between its starting and end points (see Lemma 16, Figure 4 and Figure 6).
+In the sequel, we shall present the details of our proof, which consist of multiple steps.
+C.1.1 Step 1: showing existence of a segment for each distribution in $\mathcal{W}_{j}$
+Recall that $\mathcal{W}_{j}$ contains those distributions whose corresponding weight iterates obey $\max_{1\leq t\leq T}w_{i}^{t}\in(2^{-j},2^{-j+1}]$ (cf.
+(68a)).
+As it turns out, for any $i\in\mathcal{W}_{j}$, one can find an $\big{(}\frac{1}{2^{j+1}},\frac{1}{2^{j+2}},\log(2)\big{)}$-segment, as stated in the lemma below.
+This basic fact allows one to link each distribution in $\mathcal{W}_{j}$ with a segment of suitable parameters.
+Lemma 14.
+For each $i\in\mathcal{W}_{j}$, there exists $1\leq s_{i}<e_{i}\leq T$, such that
+$\frac{1}{2^{j+2}}<w^{s_{i}}_{i}\leq\frac{1}{2^{j+1}},\qquad w^{e_{i}}_{i}>\frac{1}{2^{j}},\qquad\text{and}\qquad w^{t}_{i}>2^{-(j+2)}\quad\forall t\in[s_{i},e_{i}].$ (71)
+In other words, there exists a $\left(\frac{1}{2^{j+1}},\frac{1}{2^{j+2}},\log(2)\right)$-segment $(s_{i},e_{i})$ with the index set as $\{i\}$ (see Definition 4).
+Proof.
+From the definition (68a) of $\mathcal{W}_{j}$, it is straightforward to find a time point $e_{i}$ obeying $w^{e_{i}}_{i}>\frac{1}{2^{j}}$.
+It then remains to identify a valid point $s_{i}$.
+To this end, let us define
+$\tau=\max\big{\{}t\mid t\leq e_{i},w^{t}_{i}\leq 2^{-(j+2)}\big{\}},$
+which is properly defined since $w^{1}_{i}=1/k\leq 2^{-(j+2)}$ (see (70)). With this choice in mind, we have
+$w^{t}_{i}>2^{-(j+2)},\qquad\forall t\text{ obeying }\tau+1\leq t\leq e_{i}.$
+In addition, it follows from the update rule (cf. lines 1 and 1 of Algorithm 1) that
+$\log(w_{i}^{t+1}/w_{i}^{t})$ $=\log(W_{i}^{t+1}/W_{i}^{t})-\log\Big{(}\sum_{j}W_{j}^{t+1}/\sum_{j}W_{j}^{t}\Big{)}$
+$\leq\eta-\log\Big{(}\sum_{j}W_{j}^{t+1}/\sum_{j}W_{j}^{t}\Big{)}\leq 2\eta\leq 1/10,$
+where the last inequality results from our choice of $\eta$. This in turn allows us to show that
+$w^{\tau+1}_{i}\leq w^{\tau}_{i}\exp(1/10)\leq\frac{1}{2^{j+2}}\cdot\exp(1/10)\leq\frac{1}{2^{j+1}}.$ (72)
+As a result, it suffices to choose $s_{i}=\tau+1$, thus concluding the proof. ∎
+C.1.2 Step 2: constructing disjoint segments with good coverage
+While Lemma 14 justifies the existence of suitable segments $\{(s_{i},e_{i})\}$ associated with each distribution in $\mathcal{W}_{j}$, we need to divide (a nontrivial subset of) them into certain disjoint blocks, where the segments in each block have at least one common inner points.
+This is accomplished in the following lemma.
+Lemma 15.
+Recall the definition of $\mathcal{W}_{j}$ in (68a).
+For each $i\in\mathcal{W}_{j}$, denote by $(s_{i},e_{i})$ the segment identified in Lemma 14.
+Then there exist a group of disjoint subsets $\{\mathcal{W}_{j}^{p}\}_{p=1}^{P}$ of $\mathcal{W}_{j}$ obeying
+  1. $\mathrm{(i)}$
+    $\mathcal{W}_{j}^{p}\subseteq\mathcal{W}_{j}$, $\mathcal{W}_{j}^{p}\cap\mathcal{W}_{j}^{p^{\prime}}=\emptyset$, $\forall p\neq p^{\prime}$;
+  2. $\mathrm{(ii)}$
+    $\sum_{p=1}^{P}|\mathcal{W}_{j}^{p}|\geq\frac{|\mathcal{W}_{j}|}{3(\log_{2}(T)+1)}$;
+  3. $\mathrm{(iii)}$
+    Let $\widetilde{s}_{p}=\min_{i\in\mathcal{W}_{j}^{p}}s_{i}$ and $\widetilde{e}_{p}=\max_{i\in\mathcal{W}_{j}^{p}}e_{i}$ for each $1\leq p\leq P$.
+One has $1\leq\widetilde{s}_{1}<\widetilde{e}_{1}\leq\widetilde{s}_{2}<\widetilde{e}_{2}\leq\dots\leq\widetilde{s}_{P}<\widetilde{e}_{P}\leq T$ and $\max_{i\in\mathcal{W}_{j}^{p}}s_{i}\leq\min_{i\in\mathcal{W}_{j}^{p}}e_{i}$ for each $1\leq p\leq P$.
+Proof.
+See Appendix C.2. ∎
+In words, Lemma 15 reveals the existence a collection of disjoint subsets of $\mathcal{W}_{j}$ such that (a) they account for a sufficiently large fraction of the indices contained in $\mathcal{W}_{j}$, and (b) the segments in each subset $\mathcal{W}_{j}^{p}$ share at least one common inner point.
+Thus far, each of the segments constructed above is associated with a single distribution in $\mathcal{W}_{j}$.
+Clearly, it is likely that many of these segments might have non-trivial overlap; in other words, many of them might have shared sub-segments.
+What we intend to do next is to further group the indices in $\{\mathcal{W}_{j}^{p}\}$ into disjoint subgroups, and identify a common (sub)-segment for each of these subgroups.
+What remains unclear, however, is whether each of these (sub)-segments experiences sufficient weight changes between its starting and end points.
+We address these in the following lemma.
+Lemma 16.
+Recall the definition of $\mathcal{W}_{j}$ in (68a).
+For each $i\in\mathcal{W}_{j}$, denote by $(s_{i},e_{i})$ the segment identified in Lemma 14.
+Then there exists a group of subsets $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ of $\mathcal{W}_{j}$ satisfying the following properties:
+  1. $\mathrm{(i)}$
+    $\mathcal{V}_{j}^{n}\subseteq\mathcal{W}_{j}$, $\mathcal{V}_{j}^{n}\cap\mathcal{V}_{j}^{n^{\prime}}=\emptyset$, $\forall n\neq n^{\prime}$;
+  2. $\mathrm{(ii)}$
+    $\sum_{n=1}^{N}|\mathcal{V}_{j}^{n}|\geq\frac{|\mathcal{W}_{j}|}{24\log_{2}(k)(\log_{2}(T)+1)}$;
+  3. $\mathrm{(iii)}$
+    There exist $1\leq\widehat{s}_{1}<\widehat{e}_{1}\leq\widehat{s}_{2}<\widehat{e}_{2}\leq\dots\leq\widehat{s}_{N}<\widehat{e}_{N}\leq T$, and $\{g_{n}\}_{n=1}^{N}\in\left[1,\infty\right)^{N}$, such that for each $1\leq n\leq N$, $(\widehat{s}_{n},\widehat{e}_{n})$ is a $\left(2^{-(j+1)}g_{n}|\mathcal{V}_{j}^{n}|,2^{-(j+2)}|\mathcal{V}_{j}^{n}|,\frac{\log(2)}{2\log_{2}(k)}\right)$-segment with index set as $\mathcal{V}_{j}^{n}$.
+That is, the following properties hold for each $1\leq n\leq N$:
+    • $\mathrm{(a)}$
+      $\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j+2}}\leq\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{\widehat{s}_{n}}\leq\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j+1}}$;
+    • $\mathrm{(b)}$
+      $\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j}}\cdot\exp\left(\frac{\log(2)}{2\log_{2}(k)}\right)\leq\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{\widehat{e}_{n}}$;
+    • $\mathrm{(c)}$
+      $\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{t}\geq\frac{|\mathcal{V}_{j}^{n}|}{2^{j+2}}$ for any $t$ obeying $\widehat{s}_{n}\leq t\leq\widehat{e}_{n}$.
+Proof.
+See Appendix C.3. ∎
+In brief, the subsets $\{\mathcal{V}_{j}^{n}\}$ identified in Lemma 16 enjoy the following useful properties: (a) they are disjoint; (b) these subsets taken collectively also cover a reasonably large fraction of the elements in $\mathcal{W}_{j}$; (c) each subset $\mathcal{V}_{j}^{n}$ is linked with a suitable segment, whose associated weights have increased sufficiently from its starting point to its end point.
+C.1.3 Step 3: bounding the length of segments
+In this step, we turn attention to the length of segments, unveiling the interplay between the segment length and certain sub-optimality gaps.
+Recall the definition (58) of $v^{t}$ as follows
+$v^{t}\coloneqq L(h^{t},w^{t})-\mathsf{OPT}\qquad\text{with }\mathsf{OPT}\coloneqq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}L(h,\mathrm{e}^{\mathsf{basis}}_{i}),$ (73)
+with $\mathrm{e}^{\mathsf{basis}}_{i}$ the $i$-th standard basis vector. The following lemma assists in bounding the length of the segments defined in Definition 4.
+Lemma 17.
+Let $j_{\mathrm{max}}=\left\lfloor\log_{2}(1/\eta)\right\rfloor+1$.
+Assume the conditions in Lemma 3 hold.
+Suppose $(t_{1},t_{2})$ is a $\left(p,q,x\right)$-segment satisfying $p\geq 2q>0$.
+Then one has
+$t_{2}-t_{1}\geq\frac{x}{2\eta}.$ (74)
+Moreover, if
+$\frac{qx^{2}}{50\big{(}\log_{2}(1/\eta)+1\big{)}^{2}}\geq\frac{1}{k}$ (75)
+holds, then with probability exceeding $1-6T^{4}k\delta^{\prime}$, at least one of the following two claims holds:
+  • (a)
+    the length of the segment satisfies
+    $t_{2}-t_{1}\geq\frac{qx^{2}}{200\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\eta^{2}}.$ (76)
+  • (b)
+    the quantities $\{v^{t}\}$ obey
+    $4\sum_{\tau=t_{1}}^{t_{2}-1}(-v^{\tau}+\varepsilon_{1})\geq\frac{qx^{2}}{100\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\eta}.$ (77)
+Proof.
+See Appendix C.4. ∎
+In words, Lemma 17 shows that (i) in general the length of a segment scales at least linearly in $1/\eta$; (ii) if the parameter $q$ of a segment (i.e., some lower bound on the weights of interest within this segment) is sufficiently large, then either the length of this segment scales at least quadratically in $1/\eta$, or the sum of certain sub-optimality gaps needs to be large enough.
+C.1.4 Step 4: putting all this together
+With the above lemmas in place, we are positioned to establish Lemma 13.
+In what follows, we denote by $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ and $\{(\widehat{s}_{n},\widehat{e}_{n})\}_{n=1}^{N}$ the construction in Lemma 16.
+To begin with, it is observed that: for any $1\leq j\leq\overline{j}$ (with $\overline{j}$ defined in (67)), one has
+$2^{-(j+2)}|\mathcal{V}_{j}^{n}|\cdot\frac{\log^{2}(2)}{50(\log_{2}(1/\eta)+1)^{2}\log_{2}^{2}(k)}\geq 2^{-(\overline{j}+2)}\cdot\frac{\log^{2}(2)}{50(\log_{2}(1/\eta)+1)^{2}\log_{2}^{2}(k)}\geq\frac{1}{k}.$ (78)
+Recall that $v^{\tau}\leq\varepsilon_{1}$ (see (59)).
+Combining this fact with Lemma 17 (by setting $q=2^{-(j+2)}|{\mathcal{V}}_{j}^{n}|$ and $x=\frac{\log(2)}{\log_{2}(k)}$) reveals that, for each $1\leq n\leq N$,
+$T\eta+\bigg{\{}4T\varepsilon_{1}+4\sum_{t=1}^{T}(-v^{t})\bigg{\}}$ $\geq\sum_{n=1}^{{N}}(\widehat{e}_{n}-\widehat{s}_{n})\eta+4\sum_{n=1}^{{N}}\sum_{\tau=\widehat{s}_{n}}^{\widehat{e}_{n}-1}(-v^{\tau}+\varepsilon_{1})$ (79)
+$\geq\frac{2^{-(j+2)}\sum_{n=1}^{{N}}|{\mathcal{V}}_{j}^{n}|\log^{2}(2)}{800\log_{2}^{2}(k)\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\eta},$ (80)
+where the first inequality results from the disjoint nature of the segments $\{[\widehat{s}_{n},\widehat{e}_{n}]\}_{1\leq n\leq N}$, and the second inequality comes from Lemma 17.
+Moreover, it follows from (64) that
+$\sum_{t=1}^{T}(-v^{t})\leq\frac{\log(k)}{\eta}+\eta T+4\sqrt{T\log(1/\delta^{\prime})},$ (81)
+which taken together with (80) and the choice $\varepsilon_{1}=\eta$ gives
+$\sum_{n=1}^{{N}}|{\mathcal{V}}_{j}^{n}|$ $\leq\frac{3200\eta\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\log_{2}^{2}(k)\cdot 2^{j+2}}{\log^{2}(2)}\cdot\left(\frac{\log(k)}{\eta}+\eta T+4\sqrt{T\log(1/\delta^{\prime})}\right)$
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{4000T\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\log_{2}^{2}(k)\cdot 2^{j+2}\eta^{2}}{\log^{2}(2)}.$ (82)
+To finish up, it follows from Property (ii) of Lemma 16 that
+$|\mathcal{W}_{j}|$ $\leq 24\log_{2}(k)\big{(}\log_{2}(T)+1\big{)}\left(\sum_{n=1}^{N}|\mathcal{V}_{j}^{n}|\right)$
+$\leq 8\cdot 10^{7}\cdot\left(\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\log_{2}^{2}(k)\left(\log_{2}(k)+\log(1/\delta^{\prime})\right)^{3}\big{(}\log_{2}(T)+1\big{)}\right)\cdot 2^{j}$
+for any $1\leq j\leq\overline{j}$. The proof is completed by recalling that $\delta^{\prime}=\frac{\delta}{4(T+k+1)}$.
 C.2 Proof of Lemma 15
-Fix $j\in[1,\left\lfloor\log_{2}(k)\right\rfloor]-2$.
-Before step into the proof, we first introduce the high-level idea.
-In Lemma 29, we provide a lower bound of the length of a $(p,q,x)$ segment.
-We then want to prove that if $|\mathcal{W}_{j}|$ is large, there will be many disjoint segments, which forces the total length $T$ to be large enough to contain these segments.
-As discussed in Section 4.2, we will extract a regular configuration from a general configuration. In high-level intuition, we have the following steps.
-  1. 1.
-    Prove the existence of segments for all $i\in\mathcal{W}_{j}$ (See Lemma 26);
-  2. 2.
-    Find some disjoint blocks such that (See Lemma 27 and Figure 3 )
-    • The segments in the same blocks have a common inner point;
-    • The number of segments in these blocks is at least $\frac{1}{3(\log_{2}(T)+1)}$ of the number of all the segments, i.e., $\frac{1}{3(\log_{2}(T)+1)}|\mathcal{W}_{j}|$ ;
-    • Continue the analysis on a single block on benefit of the fact that these blocks are disjoint;
-  3. 3.
-    For each single block, use the common inner point to align either the starting points or the ending points of at least half of these segments (See Lemma 28 and Figure 4 );
-  4. 4.
-    Design a group of regular configurations (at most $O(\log_{2}(k))$ groups) so that at least one of the regular configurations contains enough segments with significant variation (See Lemma 28 and Figure 6 ).
-C.2.1 The existence of segments
-For $i\in\mathcal{W}_{j}$, by noting that $w^{1}_{i}=\frac{1}{k}\leq\frac{1}{2^{j+1}}$ and the fact that the trajectory $\{w^{t}_{i}\}$ changes slowly in $t$, we have the lemma below to show that for each $i\in\mathcal{W}_{j}$, there is an $(\frac{1}{2^{j+1}},\frac{1}{2^{j+2}},\log(2))$-segment $(s_{i},e_{i})$ for $i$.
+For any integer $11\leq x\leq\log_{2}(T)+$ 1, define
+$\mathcal{W}_{j}(x)\coloneqq\big{\{}i\in[k]\mid 2^{x-1}\leq e_{i}-s_{i}\leq 2^{x}\big{\}},$
+so that the length of each segment associated with $\mathcal{W}_{j}(x)$ lies within $[2^{x-1},2^{x}]$. Let $x^{\star}$ indicate the one that maximizes the cardinality of $\mathcal{W}_{j}(x)$:
+$x^{\star}=\arg\max_{1\leq x\leq\log_{2}(T)+1}|\mathcal{W}_{j}(x)|.$
+Given that there are at most $\log_{2}(T)+1$ choices of $x$, the pigeonhole principle gives
+$|\mathcal{W}_{j}(x^{\star})|\geq\frac{|\mathcal{W}_{j}|}{\log_{2}(T)+1}.$ (83)
+In the sequel, we intend to choose the subsets $\{\mathcal{W}_{j}^{m}\}_{m=1}^{M}$ from $\mathcal{W}_{j}(x^{\star})$.
+To proceed, let us set
+$\kappa_{1}\coloneqq\min_{i\in\mathcal{W}_{j}(x^{\star})}e_{i},\qquad\mathcal{U}_{j}^{1}\coloneqq\big{\{}i\in\mathcal{W}_{j}(x^{\star})\mid s_{i}\leq\kappa_{1}\big{\}},$ (84a)
+and then for each $o\geq 1$, take
+$\kappa_{o+1}$ $\coloneqq\min_{i\in\mathcal{W}_{j}(x^{\star})/\cup_{o^{\prime}=1}^{o}\mathcal{U}_{j}^{o^{\prime}}}e_{i},$ (84b)
+$\mathcal{U}_{j}^{o+1}$$\coloneqq\big{\{}i\in\mathcal{W}_{j}(x^{\star})/\cup_{o^{\prime}=1}^{o}\mathcal{U}_{j}^{o^{\prime}}\mid s_{i}\leq\kappa_{o+1}\big{\}}.$ (84c)
+We terminate such constructions until $\cup_{o\geq 1}\mathcal{U}_{j}^{o}=\mathcal{W}_{j}(x^{\star})$. By construction, for each $o$, we have
+$s_{i_{2}}\leq\kappa_{o}\leq e_{i_{1}},\quad\forall i_{1},i_{2}\in\mathcal{U}_{j}^{o}\qquad\Longleftrightarrow\qquad\max_{i\in\mathcal{U}_{j}^{o}}s_{i}\leq\min_{i\in\mathcal{U}_{j}^{o}}e_{i}.$ (85)
+Let us look at the three groups of subsets of $\mathcal{W}_{j}(x^{\star})$: $\{\mathcal{U}_{j}^{3o-2}\}_{o\geq 1}$, $\{\mathcal{U}_{j}^{3o-1}\}_{o\geq 1}$ and $\{\mathcal{U}_{j}^{3o}\}_{o\geq 1}$.
+Clearly, there exists $l\in\{0,1,2\}$ such that $\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-l}|\geq\frac{1}{3}\sum_{o\geq 1}|\mathcal{U}_{j}^{o}|$; without loss of generality, assume that
+$\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-2}|\geq\frac{1}{3}\sum_{o\geq 1}|\mathcal{U}_{j}^{o}|=\frac{1}{3}|\mathcal{W}_{j}(x^{\star})|.$ (86)
+With the above construction in place, we would like to verify that $\{\mathcal{U}_{j}^{3o-2}\}_{o\geq 1}$ forms the desired group of subsets.
+First of all, Condition (i) holds directly from the definition of $\{\mathcal{U}_{j}^{o}\}_{o\geq 1}$.
+When it comes to Condition (ii), it follows from (86) and (83) that
+$\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-2}|\geq\frac{1}{3}|\mathcal{W}_{j}(x^{\star})|\geq\frac{|\mathcal{W}_{j}|}{3(\log_{2}(T)+1)}.$
+Regarding Condition (iii), it suffices to verify that
+$\max_{i\in\mathcal{U}_{j}^{3o-2}}e_{i}\leq\min_{i\in\mathcal{U}_{j}^{3o+1}}s_{i}$ (87)
+for any $o$. To do so, note that for each $o\geq 1$, there exists $i\in\mathcal{W}_{j}(x^{\star})$ such that $s_{i}\geq\kappa_{o}$ and $\kappa_{o+1}=e_{i}$. We can then deduce that
+$\kappa_{o+1}=e_{i}\geq s_{i}+2^{x^{\star}-1}\geq\kappa_{o}+2^{x^{\star}-1}.$ (88)
+It then follows that, for any $i\in\mathcal{U}_{j}^{3o+1}$, one has
+$s_{i}\geq\kappa_{3o}\geq\kappa_{3o-1}+2^{x^{\star}-1}\geq\kappa_{3o-2}+2^{x^{\star}}.$
+In addition, for any $l\in\mathcal{U}_{j}^{3o-2}$, it is seen that
+$e_{l}\leq s_{l}+2^{x^{\star}}\leq\kappa_{3o-2}+2^{x^{\star}}.$
+Putting all this together yields
+$\max_{i\in\mathcal{U}_{j}^{3o-2}}e_{i}\leq\kappa_{3o-2}+2^{x^{\star}}\leq\min_{i\in\mathcal{U}_{j}^{3o+1}}s_{i}.$
+The proof is thus complete.
+C.3 Proof of Lemma 16
+We shall begin by presenting our construction of the subsets, followed by justification of the advertised properties. In what follows, we set $x=\log(2)$.
+Our construction.
+Let $\{\mathcal{W}_{j}^{p}\}_{p=1}^{P}$ and $\{(\widetilde{s}_{p},\widetilde{e}_{p})\}_{p=1}^{P}$ be the construction in Lemma 15.
+Step a): constructing $\widehat{\mathcal{W}}_{j}^{p}$. Consider any $1\leq p\leq P$. Set
+$t_{\mathrm{mid}}^{p}\coloneqq\min_{i\in\mathcal{W}_{j}^{p}}e_{i},$
+which, in view of Lemma 15, is a common inner point of the segments in $\mathcal{W}_{j}^{p}$. We can derive, for each $i\in\mathcal{W}_{j}^{p}$,
+$\max\left\{\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right),\log\left(\frac{w^{t_{\mathrm{mid}}^{p}}_{i}}{w^{s_{i}}_{i}}\right)\right\}\geq\frac{1}{2}\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)+\frac{1}{2}\log\left(\frac{w^{t_{\mathrm{mid}}^{p}}_{i}}{w^{s_{i}}_{i}}\right)=\frac{1}{2}\log\left(\frac{w^{e_{i}}_{i}}{w^{s_{i}}_{i}}\right)\geq\frac{x}{2},$
+where the last inequality holds since $(s_{i},e_{i})$ is constructed to be a $\big{(}\frac{1}{2^{j+1}},\frac{1}{2^{j+2}},x\big{)}$-segment (see Lemma 14). It then follows that
+$\sum_{i\in\mathcal{W}_{j}^{p}}\left(\mathds{1}\left\{\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)\geq\frac{x}{2}\right\}+\mathds{1}\left\{\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)\geq\frac{x}{2}\right\}\right)\geq|\mathcal{W}_{j}^{p}|.$
+Without loss of generality, we assume that
+$\sum_{i\in\mathcal{W}_{j}^{p}}\mathds{1}\left\{\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)\geq\frac{x}{2}\right\}\geq\frac{|\mathcal{W}_{j}^{p}|}{2}.$ (89)
+This means that the set define below
+$\widehat{\mathcal{W}}_{j}^{p}\coloneqq\left\{i\in\mathcal{W}_{j}^{p}\mid\log\big{(}w^{e_{i}}_{i}/w^{t_{\mathrm{mid}}^{p}}_{i}\big{)}\geq\frac{x}{2}\right\}$ (90)
+satisfies
+$|\widehat{\mathcal{W}}_{j}^{p}|\geq\frac{|\mathcal{W}_{j}^{p}|}{2}.$ (91)
+In what follows, we take [^8] We assume $\widetilde{Q}\geq 2$ without loss of generality.
+In the case $\widetilde{Q}=1$, we simply choose an arbitrary element in $\widehat{\mathcal{W}}_{j}^{p}$ as a single subset.
+In this way, we can collect at least $\frac{1}{4}|\widehat{\mathcal{W}_{j}^{p}}|$ segments.
+$Q(p)\coloneqq|\widehat{\mathcal{W}}_{j}^{p}|,\qquad\widetilde{l}(p)\coloneqq\max\big{\{}l\geq 0\mid 2^{l}\leq Q(p)\big{\}}\qquad\text{and}\qquad\widetilde{Q}(p)\coloneqq 2^{\widetilde{l}(p)}.$
+Without loss of generality, we assume
+$\widehat{\mathcal{W}}_{j}^{p}=\big{\{}1,2,\ldots,Q(p)\big{\}}\qquad\text{and}\qquad e_{1}\leq e_{2}\leq\dots\leq e_{Q(p)}.$ (92)
+In the sequel, we shall often abbreviate $Q(p)$, $\widetilde{l}(p)$ and $\widetilde{Q}(p)$ as $Q$, $\widetilde{l}$ and $\widetilde{Q}$, respectively, as long as it is clear from the context.
+Step b): constructing $\mathcal{K}_{l}$ and $\widetilde{\mathcal{W}}_{j}^{p}(l)$.
+Let us take $e_{0}=t_{\mathrm{mid}}^{p}$, and employ $[e_{0},e_{k}]\oplus a$ as a shorthand notation for $[e_{a},e_{k+a}]$.
+We can then define a group of disjoint intervals of $[T]$ as follows:
+$\mathcal{K}_{1}$ =$=\left\{[e_{0},e_{2^{\widetilde{l}-1}}]\right\};$ (93a)
+$\mathcal{K}_{2}$ =$=\left\{[e_{0},e_{2^{\widetilde{l}-2}}],[e_{0},e_{2^{\widetilde{l}-2}}]\oplus 2^{\widetilde{l}-1}\right\};$ (93b)
+$\mathcal{K}_{3}$ =$=\left\{[e_{0},e_{2^{\widetilde{l}-3}}],[e_{0},e_{2^{\widetilde{l}-3}}]\oplus 2^{\widetilde{l}-2},[e_{0},e_{2^{\widetilde{l}-3}}]\oplus 2\cdot 2^{\widetilde{l}-2},[e_{0},e_{2^{\widetilde{l}-3}}]\oplus 3\cdot 2^{\widetilde{l}-2}\right\};$ (93c)
+$\dots$
+$\mathcal{K}_{l}$ =$=\left\{[e_{0},e_{2^{\widetilde{l}-l}}],[e_{0},e_{2^{\widetilde{l}-l}}]\oplus 2^{\widetilde{l}-l+1},[e_{0},e_{2^{\widetilde{l}-l}}]\oplus 2\cdot 2^{\widetilde{l}-l+1},\dots,[e_{0},e_{2^{\widetilde{l}-l}}]\oplus(2^{l-1}-1)2^{\widetilde{l}-l+1}\right\};$ (93d)
+$\dots$
+$\mathcal{K}_{\widetilde{l}}$ =$=\left\{[e_{2i},e_{2i+1}]\mid i=0,1,2,\dots,2^{\widetilde{l}-1}-1\right\};$ (93e)
+$\mathcal{K}_{\widetilde{l}+1}$ =$=\left\{[e_{2i+1},e_{2i+2}]\mid i=0,1,2,\dots,2^{\widetilde{l}-1}-1\right\}.$ (93f)
+For each $i\in[\widetilde{Q}-1]$ with binary form $\{i_{l}\}_{l=1}^{\widetilde{l}}$ and $0\leq l\leq\widetilde{l}$, we define $\mathsf{trunc}(i,l)$ to be the number with binary form $\{i_{1},i_{2},\ldots,i_{l},0,0,\ldots,0\}$.
+For example, $\mathsf{trunc}(i,0)=0$ and $\mathsf{trunc}(i,\widetilde{l})=i$.
+From the definition (90) of $\widehat{\mathcal{W}}_{j}^{p}$, we know that for each $i\in[\widetilde{Q}-1]$,
+$\frac{x}{2}\leq\log\left(\frac{w_{i}^{e_{i}}}{w_{i}^{e_{0}}}\right)=\sum_{l=1}^{\widetilde{l}}\log\left(\frac{w_{i}^{e_{\mathsf{trunc}(i,l)}}}{w_{i}^{e_{\mathsf{trunc}(i,l-1)}}}\right)=\sum_{l=1}^{\widetilde{l}}\log\left(\frac{w_{i}^{e_{\mathsf{trunc}(i,l)}}}{w_{i}^{e_{\mathsf{trunc}(i,l-1)}}}\right)\mathds{1}\left\{e_{\mathsf{trunc}(i,l)}\neq e_{\mathsf{trunc}(i,l-1)}\right\},$ (94)
+which in turn implies that
+$\max_{1\leq l\leq\widetilde{l}}\log\left(\frac{w_{i}^{e_{\mathsf{trunc}(i,l)}}}{w_{i}^{e_{\mathsf{trunc}(i,l-1)}}}\right)\geq\frac{x}{2\widetilde{l}}.$ (95)
+By defining
+$\widetilde{\mathcal{W}}_{j}^{p}(l)\coloneqq\left\{i\in\widehat{\mathcal{W}}_{j}^{p}:\arg\max_{1\leq l^{\prime}\leq\widetilde{l}}\log\left(\frac{w_{i}^{e_{\mathsf{trunc}(i,l^{\prime})}}}{w_{i}^{e_{\mathsf{trunc}(i,l^{\prime}-1)}}}\right)=l\right\}$
+for each [^9] Without loss of generality, we assume the $\arg\max$ function is a single-valued function. $1\leq l\leq\widetilde{l}$, we can demonstrate that
+$\sum_{l=1}^{\widetilde{l}}\Big{|}\widetilde{\mathcal{W}}_{j}^{p}(l)\Big{|}\geq\widetilde{Q}-1,$ (96)
+thus implying the existence of some $1\leq l^{\star}\leq\widetilde{l}$ obeying
+$\Big{|}\widetilde{W}_{j}^{p}(l^{\star})\Big{|}\geq\frac{\widetilde{Q}-1}{\widetilde{l}}\geq\frac{\widetilde{Q}}{2\widetilde{l}}.$ (97)
+Step c): constructing $\widetilde{\mathcal{W}}_{j}^{p}(l,o)$, $\widehat{s}(p,o)$ and $\widehat{e}(p,o)$.
+By definition, for any $i$, if $\mathsf{trunc}(i,l^{\star})\neq\mathsf{trunc}(i,l^{\star}-1)$, then one has
+$[e_{\mathsf{trunc}(i,l^{\star}-1)},e_{\mathsf{trunc}(i,l^{\star})}]\in\mathcal{K}_{l^{\star}},$
+where the set $\mathcal{K}_{l}$ has been defined in (93).
+In addition, from the construction of $\widetilde{\mathcal{W}}_{j}^{p}(l^{\star})$ (see (97)), we know that $\mathsf{trunc}(i,l^{\star})\neq\mathsf{trunc}(i,l^{\star}-1)$ for any $i\in\widetilde{\mathcal{W}}_{j}^{p}(l^{\star})$.
+For each $1\leq o\leq 2^{l^{\star}-1}$, define
+$\widetilde{\mathcal{W}}_{j}^{p}(l^{\star},o)\coloneqq\left\{i\in\widetilde{\mathcal{W}}_{j}^{p}(l^{\star})\mid[e_{\mathsf{trunc}(i,l^{\star}-1)},e_{\mathsf{trunc}(i,l^{\star})}]=[e_{0},e_{2^{\widetilde{l}-l^{\star}}}]\oplus(o-1)2^{\widetilde{l}-l^{\star}+1}\right\},$ (98)
+where we employ the notation $l^{\star}$ and $\widetilde{l}$ to abbreviate $l^{\star}(p)$ and $\widetilde{l}(p)$, respectively.
+In addition, for any $1\leq p\leq P$ and $1\leq o\leq 2^{l^{\star}(p)-1}$, we set
+$\widehat{s}(p,o)$ $=e_{(o-1)2^{\widetilde{l}(p)-l^{\star}(p)+1}},$ (99a)
+$\widehat{e}(p,o)$ $=e_{2^{\widetilde{l}(p)-l^{\star}(p)}+(o-1)2^{\widetilde{l}(p)-l^{\star}(p)+1}}.$ (99b)
+In words, $[\widehat{s}(p,o),\widehat{e}(p,o)]$ can be understood as the $o$-th interval in the set $\mathcal{K}_{l^{\star}(p)}$.
+Step d): construction output. With the above construction in mind, we would like to select
+$\left\{\big{\{}\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p),o\big{)}\big{\}}_{o=1}^{2^{l^{\star}(p)-1}}\right\}_{p=1}^{P}\qquad\text{with intervals}\qquad\left\{\big{\{}\widehat{s}(p,o),\widehat{e}(p,o)\big{\}}_{o=1}^{2^{l^{\star}(p)-1}}\right\}_{p=1}^{P}$
+as the group of subsets we construct. With slight abuse of notation, we use $(p,o)$ as the index of the segments instead of $n$. In what follows, we validate this construction.
+Verification of the advertised properties.
+We now proceed to justify the claimed properties.
+Property (i).
+By construction, it is clearly seen that
+$\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p),o\big{)}\subseteq\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p)\big{)}\subseteq\widehat{\mathcal{W}}_{j}^{p}\subseteq\mathcal{W}_{j}^{p}\subseteq\mathcal{W}_{j}.$
+In addition, if
+$\widetilde{W}_{j}^{p_{1}}\big{(}l^{\star}(p_{1}),o_{1}\big{)}\cap\widetilde{W}_{j}^{p_{2}}\big{(}l^{\star}(p_{2}),o_{2}\big{)}\neq\emptyset,$
+then one has $\mathcal{W}_{j}^{p_{2}}\cap\mathcal{W}_{j}^{p_{2}}\neq\emptyset$, and as a result, $p_{1}=p_{2}$ (otherwise it violates the condition that $\mathcal{W}_{j}^{p_{2}}\cap\mathcal{W}_{j}^{p_{2}}=\emptyset$ for $p_{1}\neq p_{2}$).
+It also follows from the definition in (98) that $o_{1}=o_{2}$.
+Therefore, for any $(p_{1},o_{1})$ that does not equal $(p_{2},o_{2})$, we have $\widetilde{W}_{j}^{p_{1}}\big{(}l^{\star}(p_{1}),o_{1}\big{)}\cap\widetilde{W}_{j}^{p_{2}}\big{(}l^{\star}(p_{2}),o_{2}\big{)}=\emptyset$.
+Property (ii).
+By construction, we have
+$\sum_{o=1}^{2^{l^{\star}(p)-1}}\Big{|}\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p),o\big{)}\Big{|}=\Big{|}\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p)\big{)}\Big{|}\geq\frac{|\widehat{\mathcal{W}}_{j}^{p}|}{4\log_{2}\big{(}|\widehat{\mathcal{W}}_{j}^{p}|\big{)}}\geq\frac{|\mathcal{W}_{j}^{p}|}{8\log_{2}\big{(}|\widehat{\mathcal{W}}_{j}^{p}|\big{)}},$ (100)
+where we have made use of (97) and (91). Summing over $p$ and applying Lemma 15 yield
+$\sum_{p=1}^{P}\sum_{o=1}^{2^{l^{\star}(p)-1}}\Big{|}\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p),o\big{)}\Big{|}\geq\sum_{p=1}^{P}\frac{|\mathcal{W}_{j}^{p}|}{8\log_{2}(k)}\geq\frac{|\mathcal{W}_{j}|}{24\log_{2}(k)\big{(}\log_{2}(T)+1\big{)}}.$ (101)
+Property (iii)(a).
+Let us set the parameters $\left\{\{g(p,o)\}_{o=1}^{2^{l^{\star}(p)}}\right\}_{p=1}^{P}$ as follows:
+$g(p,o)=\frac{\sum_{i\in\widetilde{\mathcal{W}}_{j}(l^{\star}(p),o)}w_{i}^{\widehat{s}(p,o)}}{2^{-(j+2)}\cdot\big{|}\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p),o\big{)}\big{|}}\geq 1,$
+where the last inequality holds since, by construction, $w_{i}^{\widehat{s}(p,o)}\geq 2^{-(j+2)}$ (see Lemma 14). Then Property (iii)(a) is satisfied since
+$\sum_{i\in\widetilde{\mathcal{W}}_{j}(l^{\star}(p),o)}w_{i}^{\widehat{s}(p,o)}=\frac{g(p,o)\big{|}\widetilde{\mathcal{W}}_{j}(l^{\star}(p),o)\big{|}}{2^{j+2}}.$
+Property (iii)(b).
+For any $i\in\widehat{\mathcal{W}}_{j}^{p}\subseteq\mathcal{W}_{j}^{p}$, we have
+$s_{i}\leq e_{\mathsf{trunc}(i,l-1)}\leq e_{i}\qquad\text{for any }1\leq l\leq\widetilde{l}(p),$
+which is valid since $\max_{i\in\mathcal{W}_{j}^{p}}s_{i}\leq\min_{i\in\mathcal{W}_{j}^{p}}e_{i}$ (see Lemma 15) and (92). It then holds that
+$s_{i}\leq\widehat{s}(p,o)\leq e_{i}\qquad\text{for any }i\in\widehat{\mathcal{W}}^{p}_{j}.$
+Also, the construction of $(s_{i},e_{i})$ (see Lemma 14) tells us that $w_{i}^{\widehat{s}(p,o)}\geq 2^{-(j+2)}$.
+Moreover, by construction, we know that for any $i\in\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p),o\big{)}$,
+$\log\left(\frac{w_{i}^{\widehat{e}(p,o)}}{w_{i}^{\widehat{s}(p,o)}}\right)$ $\geq\frac{x}{2\widetilde{l}(p)}\qquad\text{and}\qquad w_{i}^{\widehat{s}(p,o)}\geq 2^{-(j+2)}.$
+Recalling that $x=\log(2)$, one can further derive
+$\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(l^{\star}(p),o)}w^{\widehat{s}(p,o)}_{i}$ $\geq 2^{-(j+2)}\cdot\big{|}\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p),o\big{)}\big{|}$
+$\log\left(\frac{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(l^{\star}(p),o)}w^{\widehat{e}(p,o)}_{i}}{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(l^{\star}(p),o)}w^{\widehat{s}(p,o)}_{i}}\right)$ $\geq\log\left(\frac{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(l^{\star}(p),o)}w^{\widehat{s}(p,o)}_{i}\cdot\exp\left(\frac{x}{2\widetilde{l}(p)}\right)}{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(l^{\star}(p),o)}w^{\widehat{s}(p,o)}_{i}}\right)=\frac{x}{2\widetilde{l}(p)}\geq\frac{\log(2)}{2\log_{2}(k)}.$
+Property (iii)(c).
+Note that for any $t$ obeying $\widehat{s}(p,o)\leq t\leq\widehat{e}(p,o)$, and any $i\in\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p),o\big{)}$, it holds that $s_{i}\leq\widehat{s}(p,o)\leq t\leq\widehat{e}(p,o)\leq e_{i}$.
+Recall that $w_{i}^{t}\geq 2^{-(j+2)}$ for any $t\in[s_{i},e_{i}]$ (see Lemma 14).
+As a result,
+$\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(l^{\star}(p),o)}w^{t}_{i}\geq\big{|}\widetilde{\mathcal{W}}_{j}^{p}\big{(}l^{\star}(p),o\big{)}\big{|}\cdot 2^{-(j+2)}.$
+Proper ordering.
+To finish up, it remains to verify that the intersection of $[\widehat{s}(p_{1},o_{1}),\widehat{e}(p_{1},o_{1})]$ and $[\widehat{s}(p_{2},o_{2}),\widehat{e}(p_{2},o_{2})]$ is either empty or contains only the boundary points, unless $(p_{1},o_{1})=(p_{2},o_{2})$.
+To show this, note that in the case where $p_{1}\neq p_{2}$ (assuming $p_{1}<p_{2}$), we have
+$\widetilde{s}_{p_{1}}\leq\widehat{s}(p_{1},o_{1})<\widehat{e}(p_{1},o_{1})\leq\widetilde{e}_{p_{1}}\leq\widetilde{s}_{p_{2}}\leq\widehat{s}(p_{2},o_{2})<\widehat{e}(p_{2},o_{2}),$
+which arises from Lemma 15. Also, in the case where $p_{1}=p_{2}=p$ and $o_{1}\neq o_{2}$ (assuming $o_{1}<o_{2}$), we have
+$\widehat{s}(p,o_{1})<\widehat{e}(p,o_{1})<\widehat{s}(p,o_{2})<\widehat{e}(p,o_{2}),$
+which comes from the construction (99).
+We have thus completed the proof of this lemma.
+C.4 Proof of Lemma 17
+Throughout this proof, we find it convenient to denote $Z^{t}=\sum_{i=1}^{k}W_{i}^{t}$.
+Part 1.
+We start by proving the first claim (74). Recall that $[t_{1},t_{2}]$ is assumed to be a $(p,q,x)$-segment. From the definition of the segment (see Definition 4), there exists $i\in[k]$ such that
+$\log\left(\frac{w^{t_{2}}_{i}}{w^{t_{1}}_{i}}\right)\geq x.$
+Given that $W^{t_{2}}_{i}=W^{t_{1}}_{i}\exp\big{(}\eta\sum_{\tau=t_{1}}^{t_{2}-1}\widehat{r}_{i}^{\tau}\big{)}$ and $w_{t}=W_{t}/Z_{t}$ (see lines 1 and 1 of Algorithm 1), the above inequality can be equivalently expressed as
+$\eta\sum_{\tau=t_{1}}^{t_{2}-1}\widehat{r}_{i}^{\tau}-\log(Z^{t_{2}}/Z^{t_{1}})\geq x.$ (102)
+Moreover, recognizing that
+$\log(Z^{t_{2}}/Z^{t_{1}})$ $=\log\left(\frac{\sum_{i\in[k]}W_{i}^{t_{1}}\exp\big{(}\eta\sum_{\tau=t_{1}}^{t_{2}-1}\widehat{r}_{i}^{\tau}\big{)}}{\sum_{i\in[k]}W_{i}^{t_{1}}}\right)\geq-\eta(t_{2}-t_{1})$
+and $\widehat{r}_{i}^{\tau}\leq 1$ for any $1\leq\tau\leq T$, we can invoke (102) to show that
+$x\leq 2(t_{2}-t_{1})\eta,$ (103)
+from which the claimed inequality (74) follows.
+Part 2.
+We now turn to the remaining claims of Lemma 17. For each hypothesis $h\in\mathcal{H}$, let us introduce the following vector $v_{h}\in\mathbb{R}^{k}$:
+$v_{h}=[v_{h,i}]_{i\in[k]}\qquad\text{with}~{}~{}v_{h,i}=L(h,\mathrm{e}^{\mathsf{basis}}_{i})-\mathsf{OPT}.$ (104)
+Given the $\varepsilon_{1}$-optimality of $h^{t}$ (see Lemma 3), we have the following property that holds for any $1\leq\tau,t\leq T$:
+$\langle v_{h^{\tau}},w^{t}\rangle\geq\min_{h\in\mathcal{H}}\langle v_{h},w^{t}\rangle\geq\langle v_{h^{t}},w^{t}\rangle-\varepsilon_{1}=v^{t}-\varepsilon_{1},$ (105)
+where we recall the definition of $v^{t}$ in (73). In the sequel, we divide the proof into a couple of steps.
+Step 1: decomposing the KL divergence between $w^{t}$ and $w^{t_{2}}$.
+Let us write
+$W^{t}_{i}=\exp\left(\eta\sum_{\tau=1}^{t}\widehat{r}_{i}^{\tau}\right)=\exp\left(\eta\sum_{\tau=1}^{t}\big{(}v_{h^{\tau},i}+\mathsf{OPT}+\xi^{\tau}_{i}\big{)}\right)\qquad\text{with }\xi_{i}^{\tau}=\widehat{r}_{i}^{\tau}-v_{h^{\tau},i}-\mathsf{OPT},$
+where $\xi_{i}^{\tau}=\widehat{r}_{i}^{\tau}-L(h^{\tau},\mathrm{e}^{\mathsf{basis}}_{i})$ is clearly a zero-mean random variable. Define
+$\Delta_{t_{1},t_{2}}=\sum_{\tau=t_{1}}^{t_{2}-1}\xi^{\tau}.$
+Taking $W^{t}=[W^{t}_{i}]_{i\in[k]}$ and denoting by $\log(x/y)$ the vector $\{\log(x_{i}/y_{i})\}_{i\in[k]}$ for two $k$-dimensional vectors $(x,y)$, one can then deduce that
+$\left\langle\frac{1}{\eta}\log\left(\frac{W^{t_{2}}}{W^{t_{1}}}\right)-\Delta_{t_{1},t_{2}},\,w^{t}\right\rangle-(t_{2}-t_{1})\mathsf{OPT}=\sum_{\tau=t_{1}}^{t_{2}-1}\langle v_{h^{\tau}},w^{t}\rangle\geq(t_{2}-t_{1})(v^{t}-\varepsilon_{1}),$ (106)
+where the last inequality results from (105).
+Recall that $Z^{t}=\sum_{i=1}^{k}W^{t}_{i}$ and $w^{t}_{i}=\frac{W^{t}_{i}}{Z^{t}}$. By taking $t_{1}=t$, we can derive from (106) that
+$\left\langle\log\left(\frac{w^{t_{2}}}{w^{t}}\right)-\eta\Delta_{t,t_{2}},w^{t}\right\rangle+\log\left(\frac{Z^{t_{2}}}{Z^{t}}\right)-\eta(t_{2}-t)\mathsf{OPT}\geq\eta(t_{2}-t)(v^{t}-\varepsilon_{1}).$ (107)
+As it turns out, this inequality allows us to bound the KL divergence between $w^{t}$ and $w^{t_{2}}$ as follows:
+$\mathsf{KL}(w^{t}\,\|\,w^{t_{2}})$ $\coloneqq\left\langle w^{t},\log\left(\frac{w^{t}}{w^{t_{2}}}\right)\right\rangle$
+$\leq\log(Z^{t_{2}}/Z^{t})-\eta(t_{2}-t)\mathsf{OPT}-\eta\langle w^{t},\Delta_{t,t_{2}}\rangle+\eta(t_{2}-t)(\varepsilon_{1}-v^{t}).$ (108)
+In what follows, we shall cope with the right-hand side of (108).
+Step 2: bounding the term $\log(Z^{t_{2}}/Z^{t})$.
+With probability exceeding $1-2T^{2}k\delta^{\prime}$, it holds that
+$\log(Z^{t_{2}}/Z^{t})$ $=\sum_{\tau=t}^{t_{2}-1}\log(Z^{\tau+1}/Z^{\tau})=\sum_{\tau=t}^{t_{2}-1}\log\left(\sum_{i\in[k]}\frac{W_{i}^{\tau}\exp(\eta\widehat{r}_{i}^{\tau})}{\sum_{j\in[k]}W_{i}^{\tau}}\right)$
+=$\overset{\mathrm{(i)}}{=}\sum_{\tau=t}^{t_{2}-1}\log\left(\sum_{i=1}^{k}w_{i}^{\tau}\exp(\eta\widehat{r}_{i}^{\tau})\right)\overset{\mathrm{(ii)}}{\leq}\sum_{\tau=t}^{t_{2}-1}\log\left(\sum_{i=1}^{k}w_{i}^{\tau}+\sum_{i=1}^{k}w_{i}^{\tau}(\eta\widehat{r}_{i}^{\tau})+2\sum_{i=1}^{k}w_{i}^{\tau}\eta^{2}(\widehat{r}_{i}^{\tau})^{2}\right)$
+$\overset{\mathrm{(iii)}}{\leq}\sum_{\tau=t}^{t_{2}-1}\log\left(1+\eta\sum_{i=1}^{k}w_{i}^{\tau}\widehat{r}_{i}^{\tau}+2\eta^{2}\right)\leq\sum_{\tau=t}^{t_{2}-1}\left(\eta\sum_{i=1}^{k}w_{i}^{\tau}\widehat{r}_{i}^{\tau}+2\eta^{2}\right)$
+=$\overset{\mathrm{(iv)}}{=}\eta\sum_{\tau=t}^{t_{2}-1}v^{\tau}+\eta(t_{2}-t)\mathsf{OPT}+\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}\big{\langle}w_{i}^{\tau},\widehat{r}_{i}^{\tau}-v_{h^{\tau},i}-\mathsf{OPT}\big{\rangle}+2(t_{2}-t)\eta^{2}$
+$\overset{\mathrm{(v)}}{\leq}\eta(t_{2}-t)\varepsilon_{1}+\eta(t_{2}-t)\mathsf{OPT}+\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}\big{\langle}w_{i}^{\tau},\widehat{r}_{i}^{\tau}-v_{h^{\tau},i}-\mathsf{OPT}\big{\rangle}+2(t_{2}-t)\eta^{2}.$ (109)
+Here, (i) comes from line 1 of Algorithm 1, (ii) follows from the elementary inequality $\exp(x)\leq 1+x+2x^{2}$ for any $x\leq 1$, (iii) is valid since $\sum_{i}w_{i}^{\tau}=1$ and $|\widehat{r}_{i}^{\tau}|\leq 1$, (iv) holds due to the fact that $v^{t}=\langle w^{t},v_{h^{t}}\rangle$, and (v) arises from the fact that $v^{\tau}\leq\varepsilon_{1}$ (see (59)).
+Step 3: bounding the weighted sum of $\{\xi_{i}^{\tau}\}$.
+Next, we intend to control the two random terms below:
+$\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}\big{\langle}w_{i}^{\tau},\widehat{r}_{i}^{\tau}-v_{h^{\tau},i}-\mathsf{OPT}\big{\rangle}$ $=\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{\tau}\xi_{i}^{\tau},$ (110a)
+$\eta\big{\langle}w^{t},\Delta_{t,t_{2}}\big{\rangle}$ $=\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{t}\xi_{i}^{\tau}.$ (110b)
+Let $\mathcal{F}^{\tau}$ denote what happens before the $\tau$-th round in Algorithm 1. Two properties are worth noting.
+  • The variance of $\xi_{i}^{\tau}$ is at most $O\big{(}\frac{1}{k\overline{w}^{\tau}_{i}}\big{)}$, according to the update rule (see line 1 in Algorithm 1);
+  • $\{\xi_{i}^{\tau}\}_{i\in[k]}$ are independent conditioned on $\mathcal{F}^{\tau}$.
+Let us develop bounds on the two quantities in (110) below.
+  • Letting $q^{\tau}=\sum_{i=1}^{k}w_{i}^{t}\xi_{i}^{\tau}$, one sees that
+    $|q^{\tau}|\leq 1,\qquad\mathbb{E}[q^{\tau}|\mathcal{F}^{\tau}]=0\qquad\text{and}\qquad\mathsf{Var}[q^{\tau}|\mathcal{F}^{\tau}]\leq\sum_{i=1}^{k}\frac{(w_{i}^{t})^{2}}{k\overline{w}_{i}^{\tau}}\leq\sum_{i=1}^{k}\frac{w_{i}^{t}}{k}=\frac{1}{k}.$ (111)
+    By virtue of Freedman’s inequality (cf. Lemma 7), with probability at least $1-\delta^{\prime}$ one has
+    $\left|\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{t}\xi_{i}^{\tau}\right|\leq 2\sqrt{\frac{t_{2}-t}{k}\log(2/\delta^{\prime})}+2\log(2/\delta^{\prime});$ (112)
+  • Regarding the other term, by letting $\widehat{q}^{\tau}=\sum_{i=1}^{k}w_{i}^{\tau}\xi_{i}^{\tau}$, we have
+    $|\widehat{q}^{\tau}|\leq 1,\qquad\mathbb{E}[\widehat{q}^{\tau}|\mathcal{F}^{\tau}]=0\qquad\text{and}\qquad\mathsf{Var}[\widehat{q}^{\tau}|\mathcal{F}^{\tau}]\leq\sum_{i=1}^{k}\frac{(w_{i}^{\tau})^{2}}{k\overline{w}_{i}^{\tau}}\leq\sum_{i=1}^{k}\frac{w_{i}^{\tau}}{k}=\frac{1}{k}.$
+    Invoke Freedman’s inequality (cf. Lemma 7) once again to show that, with probability exceeding $1-\delta^{\prime}$,
+    $\left|\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{\tau}\xi_{i}^{\tau}\right|\leq 2\sqrt{\frac{t_{2}-t}{k}\log(2/\delta^{\prime})}+2\log(2/\delta^{\prime}).$ (113)
+Step 4: bounding the KL divergence between $w^{t}$ and $w^{t_{2}}$.
+Combining (108), (109), (112) and (113), and applying the union bound over $(t,t_{2})$, we can demonstrate that with probability at least $1-6T^{4}k\delta^{\prime}$,
+$\mathsf{KL}(w^{t}\,\|\,w^{t_{2}})$ $\leq 2(t_{2}-t)\eta\varepsilon_{1}-(t_{2}-t)\eta v^{t}$
+$\quad\quad\quad+4\eta\sqrt{\frac{(t_{2}-t)\log(2/\delta^{\prime})}{k}}+2(t_{2}-t)\eta^{2}+4\eta\log(2/\delta^{\prime})$ (114)
+[…]
+Step 5: connecting the KL divergence with the advertised properties.
+Set
+$\tau_{\widehat{j}}$ $\coloneqq\min\{\tau\mid t_{1}\leq\tau\leq t_{2}-1,\,-v^{\tau}\leq 2^{-(\widehat{j}-1)}\},\qquad 1\leq\widehat{j}\leq j_{\mathrm{max}};$ (115a)
+$\tau_{j_{\mathrm{max}+1}}$ $\coloneqq t_{2}.$ (115b)
+By definition, we know that $\tau_{1}=t_{1}$ and $\tau_{j_{2}}\geq\tau_{j_{1}}$ for $j_{2}\geq j_{1}$.
+Let $\mathcal{I}$ be the index set associated with this segment $[t_{1},t_{2}]$, and set $y_{j}\coloneqq\sum_{i\in\mathcal{I}}w_{i}^{\tau_{j}}$.
+We then claim that there exists $1\leq\widetilde{j}\leq j_{\mathrm{max}}$ such that
+$\log\left(\frac{y_{\widetilde{j}+1}}{y_{\widetilde{j}}}\right)$ $\geq\frac{x}{\log_{2}(1/\eta)+1}.$ (116)
+Proof of (116).
+Suppose that none of $1\leq\widetilde{j}\leq j_{\mathrm{max}}$ satisfies (116).
+Then for any $1\leq\widehat{j}\leq j_{\mathrm{max}}$, it holds that $\log\left(\frac{y_{\widehat{j}+1}}{y_{\widehat{j}}}\right)<\frac{x}{\log_{2}(1/\eta)+1}$, which implies that $y_{\widehat{j}}>y_{\widehat{j}+1}\exp\left(-\frac{x}{\log_{2}(1/\eta)+1}\right)$.
+As a result, we have
+$y_{1}$ $>y_{j_{\mathrm{max}}+1}\cdot\exp\left(-j_{\mathrm{max}}\cdot\frac{x}{\log_{2}(1/\eta)+1}\right)=\bigg{(}\sum_{i\in\mathcal{I}}w_{i}^{t_{2}}\bigg{)}\cdot\exp\left(-j_{\mathrm{max}}\cdot\frac{x}{\log_{2}(1/\eta)+1}\right)$
+$\geq p\exp(x)\cdot\exp(-x)=p,$
+thus leading to contradiction (as according to the definition of the $(p,q,x)$-segment, one has $y_{1}\leq p$). ∎
+Now, assume that $\widetilde{j}$ satisfies (116). From the definition of the $(p,q,x)$-segment, we have $y_{\widetilde{j}}\geq q$. It follows from (114) that
+$\mathsf{KL}(w^{\tau_{\widetilde{j}}}\,\|\,w^{\tau_{\widetilde{j}+1}})$ $\leq 2(\tau_{\widetilde{j}+1}-\tau_{\widetilde{j}})\eta\varepsilon_{1}+(\tau_{{\widetilde{j}}+1}-\tau_{\widetilde{j}})\eta 2^{-(\widetilde{j}-1)}$
+$\quad+4\eta\sqrt{\frac{(\tau_{\widetilde{j}+1}-\tau_{\widetilde{j}})\log(2/\delta^{\prime})}{k}}+2(\tau_{\widetilde{j}+1}-\tau_{\widetilde{j}})\eta^{2}+4\eta\log(2/\delta^{\prime}).$ (117)
+Since $\log\left(\frac{y_{\widetilde{j}+1}}{y_{\widetilde{j}}}\right)\geq\frac{x}{\log_{2}(1/\eta)+1}$ and $y_{\widetilde{j}}\geq q$, we can invoke Lemma 10 and Lemma 11 to show that
+$\mathsf{KL}(w^{\tau_{\widetilde{j}}}\,\|\,w^{\tau_{\widetilde{j}+1}})\geq\mathsf{KL}\left(\mathsf{Ber}\big{(}y_{\widetilde{j}}\big{)}\,\|\,\mathsf{Ber}\big{(}y_{\widetilde{j}+1}\big{)}\right)\geq\frac{qx^{2}}{4\big{(}\log_{2}(1/\eta)+1\big{)}^{2}},$
+where $\mathsf{Ber}(x)$ denote the Bernoulli distribution with mean $x\in[0,1]$. As a result, we can obtain
+$\frac{qx^{2}}{4\big{(}\log_{2}(1/\eta)+1\big{)}^{2}}$ $\leq 2(\tau_{\widetilde{j}+1}-\tau_{\widetilde{j}})\eta\varepsilon_{1}+(\tau_{{\widetilde{j}}+1}-\tau_{\widetilde{j}})\eta 2^{-({\widetilde{j}}-1)}$
+$\quad+4\eta\sqrt{\frac{(\tau_{\widetilde{j}+1}-\tau_{\widetilde{j}})\log(2/\delta^{\prime})}{k}}+2(\tau_{\widetilde{j}+1}-\tau_{\widetilde{j}})\eta^{2}+4\eta\log(2/\delta^{\prime}),$ (118)
+which in turn results in
+$\tau_{\widetilde{j}+1}-\tau_{\widetilde{j}}$ $\geq\min\left\{\frac{qx^{2}}{100\big{(}\log_{2}(1/\eta)+1\big{)}^{2}}\min\left\{\frac{1}{\eta\varepsilon_{1}},\frac{2^{\widetilde{j}-1}}{\eta},\frac{1}{\eta^{2}}\right\},\frac{kq^{2}x^{4}}{10000\eta^{2}\log(1/\delta^{\prime})\big{(}\log_{2}(1/\eta)+1\big{)}^{4}}\right\}$
+$=\frac{qx^{2}}{100\big{(}\log_{2}(1/\eta)+1\big{)}^{2}}\min\left\{\frac{2^{\widetilde{j}-1}}{\eta},\frac{1}{\eta^{2}}\right\}$ (119)
+$=\frac{qx^{2}}{100\big{(}\log_{2}(1/\eta)+1\big{)}^{2}}\frac{2^{\widetilde{j}-1}}{\eta}.$ (120)
+Here, to see why (119) and (120) hold, it suffices to note that
+$\frac{qx^{2}}{100}\cdot\frac{2^{\widetilde{j}-1}}{\eta}\leq\frac{qx^{2}}{100}\cdot\frac{1}{\eta^{2}}\leq\frac{kq^{2}x^{4}}{10000\eta^{2}\log(1/\delta^{\prime})\big{(}\log_{2}(1/\eta)+1\big{)}^{2}},$
+a property that arises from the fact that $2^{\widetilde{j}-1}\leq 1/\eta$ and the assumption that $\frac{qx^{2}}{50(\log_{2}(1/\eta)+1)^{2}}\geq\frac{1}{k}$.
+With (120) in mind, we are ready to finish the proof by dividing into two cases.
+  • Case 1: $\widetilde{j}=j_{\mathrm{max}}$. It follows from (120) that
+    $t_{2}-t_{1}\geq\tau_{\widetilde{j}+1}-\tau_{\widetilde{j}}$ $\geq\frac{qx^{2}}{100\big{(}\log_{2}(1/\eta)+1\big{)}^{2}}\frac{2^{j_{\mathrm{max}}-1}}{\eta}\geq\frac{qx^{2}}{200\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\eta^{2}}.$
+  • Case 2: $11\leq\widetilde{j}\leq j_{\mathrm{max}}-$ 1. It comes from the definition (115) that
+    $\sum_{\tau=t_{1}}^{t_{2}-1}\mathds{1}\big{\{}-v^{\tau}\geq 2^{-\widehat{j}}\big{\}}\geq\sum_{\tau=t_{1}}^{t_{2}-1}\mathds{1}\big{\{}-v^{\tau}>2^{-\widehat{j}}\big{\}}\geq\tau_{\widehat{j}+1}-t_{1}\geq\tau_{\widehat{j}+1}-\tau_{\widehat{j}}\qquad\text{for any }1\leq\widehat{j}\leq j_{\mathrm{max}}-1.$
+    When $11\leq\widetilde{j}\leq j_{\mathrm{max}}-$ 1, the above display taken collectively with (120) gives
+    $\sum_{\tau=t_{1}}^{t_{2}-1}\mathds{1}\big{\{}-v^{\tau}\geq 2^{-\widetilde{j}}\big{\}}\geq\tau_{\widetilde{j}+1}-\tau_{\widetilde{j}}\geq\frac{qx^{2}\cdot 2^{\widetilde{j}-1}}{100\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\eta},$ (121)
+    and as a result,
+    $\sum_{\tau=t_{1}}^{t_{2}-1}\sum_{\widehat{j}=1}^{j_{\mathrm{max}}-1}\mathds{1}\big{\{}\max\{-v^{\tau},0\}\geq 2^{-\widehat{j}}\big{\}}$ $2^{-(\widehat{j}-1)}\geq\sum_{\tau=t_{1}}^{t_{2}-1}\mathds{1}\big{\{}-v^{\tau}\geq 2^{-\widetilde{j}}\big{\}}2^{-(\widetilde{j}-1)}\geq\frac{qx^{2}}{100\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\eta}.$ (122)
+    By observing that
+    $\sum_{\widehat{j}=1}^{\infty}\mathds{1}\big{\{}x\geq 2^{-\widehat{j}}\big{\}}\cdot 2^{-(\widehat{j}-1)}\leq 4x$
+    holds for any $x\geq 0$, we can combine this fact with (122) to derive
+    $4\sum_{\tau=t_{1}}^{t_{2}-1}\max\{-v^{\tau},0\}$ $\geq\frac{qx^{2}}{100\big{(}\log_{2}(1/\eta))+1\big{)}^{2}\eta}.$ (123)
+    Furthermore, recalling that $v^{\tau}\leq\varepsilon_{1}$ (cf. (59)), one can deduce that
+    $\sum_{\tau=t_{1}}^{t_{2}-1}\max\{-v^{\tau},0\}=\sum_{\tau=t_{1}}^{t_{2}-1}(-v^{\tau})-\sum_{\tau=t_{1}}^{t_{2}-1}\min\{-v^{\tau},0\}\leq\sum_{\tau=t_{1}}^{t_{2}-1}(-v^{\tau})+(t_{2}-t_{1})\varepsilon_{1},$
+    which combined with (123) yields
+    $4(t_{2}-t_{1})\varepsilon_{1}+4\sum_{\tau=t_{1}}^{t_{2}-1}(-v^{\tau})\geq\frac{qx^{2}}{100\big{(}\log_{2}(1/\eta)+1\big{)}^{2}\eta}.$
+The proof is thus complete.
+C.5 Additional illustrative figures for segment construction
+In this section, we provide several illustrative figures to help the readers understand the concept of “segments” and certain key properties, which play an important role in the segment construction described in Section 4.2.2 and the proof of Lemma 16.
+  • Figure 1 illustrates an example in which any two segments either coincide or are disjoint.
+In this case, our heuristic arguments in Section 4.2.2 about “shared intervals” and “disjoint intervals” become applicable.
+  • Figure 2 gives an example where two non-identical segments might overlap.
+Due to the presence of non-disjoint segments, our heuristic arguments in Section 4.2.2 about “shared intervals” and “disjoint intervals” are not readily applicable.
+  • In Figure 3, we provide an example of the partition of blocks (as in the proof of Lemma 15), whereas in Figure 4, we illustrate how to align one side of the segments using a common inner point (as in the proof of Lemma 16).
+  • In Figure 5 and Figure 6, we illustrate how to construct disjoint segments from a group of segments with common starting points in the case with $k=8$.
+In this particular example, we have in total $5$ groups of disjoint segments, marked with different colors.
+Figure 1: An example where any two segments either coincide or are disjoint.
+Figure 2: An example where two non-identical segments might overlap.
+Figure 3: Partition of blocks as in the proof of Lemma 16.
+Figure 4: An illustration of how we use a common inner point (i.e., $t_{\mathsf{mid}}$ to align one side of the segments (as in the proof of Lemma 16).
+The unfilled part of the segments means that the weight changes from $w_{i}^{s_{i}}$ to $w_{i}^{t_{\mathrm{mid}}}$ (i.e., $\log\left(w_{i}^{t_{\mathrm{mid}}}/w_{i}^{s_{i}}\right)$) are not significant enough.
+Figure 5: A group of segments with common starting points.
+Figure 6: Construction of groups of disjoint segments, where each of the original segments is divided into at most $\log_{2}(k)+1$ sub-segments marked with different colors.
+Appendix D Proofs of the lower bound in Theorem 2
+D.1 Proof of Theorem 2
+Note that $N_{0}=\frac{2^{d}-1}{k}$.
+Set $N=kN_{0}+1=2^{d}$.
+Set $\mathcal{X}=\{-1,0,1\}^{kN}$.
+We set $\mathcal{Y}=\{1\}$ to be a set with only a single element.
+Without loss of generality, we write $\ell(h,(x,y))=\ell(h,x)$.
+Our construction.
+We now introduce our construction, with the key components described below.
+  • Hypothesis class and loss function.
+There are $N$ hypotheses in $\mathcal{H}$, where each hypothesis is assigned $k$ unique dimensions (recall that $\mathcal{X}$ consists of $(KN)$-dimensional vectors).
+For each $h\in\mathcal{H}$, we let $\mathcal{I}_{h}=\{j_{h,i}\}_{i=1}^{k}$ denote the $k$ dimensions assigned to $h$, so that $\mathcal{I}_{h}\cap\mathcal{I}_{h^{\prime}}=\emptyset$ for $h\neq h^{\prime}$.
+We then define $h(x)$ and $\ell(h,x)$ as follows:
+    $h(x)=\ell(h,x)=\begin{cases}0,&\text{if }x_{i}=0\text{ for all }i\in\mathcal{I}_{h};\\ x_{i^{\prime}}\text{ with }i^{\prime}=\arg\min_{i\in\mathcal{I}_{h},x_{i}\neq 0}x_{i},&\text{else}.\end{cases}$
+    Next, divide $\mathcal{H}$ arbitrarily into $(k+1)$ disjoint subsets as
+    $\mathcal{H}=(\cup_{i=1}^{k}\mathcal{H}_{i})\cup\{h^{\star}\},$
+    where each $\mathcal{H}_{i}$ contains $N_{0}$ hypotheses.
+In our construction, we intend to make $h^{\star}$ the unique optimal policy, and will design properly so that the hypothesis $h\in\mathcal{H}_{i}$ performs poorly on the $i$-th distribution $\mathcal{D}_{i}$ ($i\in[k]$).
+  • Data distributions. We design the $k$ data distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$. For any given $i\in[k]$ and any $x\in\mathcal{X}$, we let
+    $\mathbb{P}_{\mathcal{D}_{i}}\{x\}=\prod_{l=1}^{kN}\mathbb{P}_{\mathcal{D}_{i,l}}\{x_{l}\}$
+    be a product distribution, where
+    $\mathbb{P}_{\mathcal{D}_{i,l}}\{x_{l}\}$ $=\mathds{1}\{x_{l}=0\},\qquad$ $l\notin\{j_{h,i}\mid h\in\mathcal{H}\},$
+    $\mathbb{P}_{\mathcal{D}_{i,l}}\{x_{l}\}$ $=\frac{1}{2}\mathds{1}\{x_{l}=1\}+\frac{1}{2}\mathds{1}\{x_{l}=-1\},\qquad$ $l\in\{j_{h,i}\mid h\notin\mathcal{H}_{i}\},$
+    $\mathbb{P}_{\mathcal{D}_{i,l}}\{x_{l}\}$ $=\left(\frac{1}{2}+4\varepsilon\right)\mathds{1}\{x_{l}=1\}+\left(\frac{1}{2}-4\varepsilon\right)\mathds{1}\{x_{l}=-1\},\qquad$ $l\in\{j_{h,i}\mid h\in\mathcal{H}_{i}\}.$
+The above construction enjoys the following properties:
+  1. (i)
+    $\ell(h,x)\in[-1,1]$ for any $h\in\mathcal{H}$ and $x\in\mathcal{X}$;
+  2. (ii)
+    $\mathbb{E}_{x\sim\mathcal{D}_{i}}[\ell(h,x)]=\mathbb{E}_{x\sim\mathcal{D}_{i,j_{h,i}}}[x_{j_{h,i}}]=8\varepsilon\cdot\mathds{1}\{h\in\mathcal{H}_{i}\}$ for any $i\in[k]$ and $h\in\mathcal{H}$;
+  3. (iii)
+    the only $\varepsilon$-optimal hypothesis is $h^{\star}$, because for any $h\neq h^{\star}$, there exists some $i$ such that $h\in\mathcal{H}_{i}$;
+  4. (iv)
+    $h(x)\in\{-1,1\}$ for $x\in\cup_{i=1}^{k}\mathrm{supp}(\mathcal{D}_{i})$, [^10] We denote by $\mathrm{supp}(\mathcal{D})$ the support of the distribution $\mathcal{D}$.
+and $|\mathcal{H}|=N=kN_{0}+1=2^{d}$, which imply that
+    $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq\log_{2}(N)\leq d$
+    over $\cup_{i=1}^{k}\mathrm{supp}(\mathcal{D}_{i})$;
+  5. (v)
+    $\ell(h,x)$ could be regarded as a function of $h(x)$ because $\ell(h,x)=h(x)$.
+Sample complexity lower bound.
+Before proceeding, let us introduce the notation $\mathsf{Query}(\mathcal{D}_{i})$ such that: for each call to $\mathsf{Query}(\mathcal{D}_{i})$, we can obtain independent observations $\{x_{j_{h,i}}\}_{h\in\mathcal{H}}$ where $x_{j_{h,i}}\sim\mathcal{D}_{i,j_{h,i}}$ for each $h\in\mathcal{H}$.
+Now for $i\in[k]$, denote by $M_{i}$ the number of calls to $\mathsf{Query}(\mathcal{D}_{i})$.
+Our aim is to show that: in order to distinguish $h^{\star}$ from $\mathcal{H}_{i}$, the quantity $M_{i}$ has to be at least $\Omega(d/\varepsilon^{2})$.
+Suppose now that there is an algorithm $\mathcal{G}$ with numbers of samples $\{M_{i}\}_{i=1}^{k}$ such that the output is $h^{\star}$ with probability at least $3/4$.
+Let $\mathbb{P}_{\mathcal{G}}\{\cdot\}$ and $\mathbb{E}_{\mathcal{G}}[\cdot]$ denote respectively the probability and expectation when Algorithm $\mathcal{G}$ is executed, and let $h_{\mathrm{out}}$ be the output hypothesis.
+It then holds that
+$\mathbb{P}_{\mathcal{G}}\{h_{\mathrm{out}}=h^{\star}\}\geq\frac{3}{4}.$
+Let $\Pi_{\mathcal{H}}$ be the set of permutations over $\mathcal{H}$, and let $\mathsf{Unif}(\Pi_{\mathcal{H}})$ be the uniform distribution over $\Pi(\mathcal{H})$.
+With slight abuse of notation, for $x\in\{-1,0,1\}^{kN}$ and $\sigma\in\Pi_{\mathcal{H}}$, we define $\sigma(x)$ to be the vector $y$ such that $y_{j_{h,i}}=x_{j_{\sigma(h),i}}$ for all $h\in\mathcal{H}$ and $i\in[k]$.
+Let $\mathcal{G}^{\prime}$ be the algorithm with $\mathcal{H}$ replaced by $\sigma(\mathcal{H})$ in the input where $\sigma\sim\mathsf{Unif}(\Pi_{\mathcal{H}})$.
+Recognizing that $\mathcal{G}$ returns the optimal hypothesis with probability at least $3/4$ for all problem instances, we can see that
+$\mathbb{P}_{\mathcal{G}^{\prime}}\{h_{\mathrm{out}}=h^{\star}\}\geq\frac{3}{4}.$
+The lemma below then assists in bounding the probability of returning a sub-optimal hypothesis.
 Lemma 18.
-For each $i\in\mathcal{W}^{j}$, there exists $1\leq s_{i}<e_{i}\leq T$, such that $\frac{1}{2^{j+2}}<w^{s_{i}}_{i}\leq\frac{1}{2^{j+1}}$, $\frac{1}{2^{j}}<w^{e_{i}}_{i}$ and $w^{t}_{i}>2^{-(j+2)}$ for any $s_{i}\leq t\leq e_{i}$.
-In words, there exists a $\left(\frac{1}{2^{j+1}},\frac{1}{2^{j+2}},\log(2)\right)$-segment $(s_{i},e_{i})$ with the index set as $\{i\}$.
+Consider $\widetilde{i}\in[k]$. If $\mathbb{P}_{\mathcal{G}^{\prime}}\{h_{\mathrm{out}}=h^{\star},M_{i}\leq m\}\geq 1/2$ for some $m\geq 0$, then for any $h\in\mathcal{H}_{\widetilde{i}}$, one has
+$\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\big{\}}\geq\frac{1}{2}\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\big{\}}\exp\big{(}-80\sqrt{m}\varepsilon-40m\varepsilon^{2}\big{)},$
+and moreover, $m$ necessarily exceeds $m\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}$.
 Proof.
-By definition, it is easy to find $e_{i}$.
-To find $s_{i}$, define $\tau=\max\{t\leq e_{i}|w^{t}_{i}\leq 2^{-(j+2)}\}$.
-$\tau$ is properly defined because $w^{1}_{i}=\frac{1}{k}\leq 2^{-(j+2)}$.
-By definition, we have that for any $\tau+1\leq t\leq e_{i}$, $w^{t}_{i}>2^{-(j+2)}$.
-Noting that $\log_{2}(w^{t+1}_{i}/w^{t}_{i})\leq\frac{\eta-\log_{2}(Z_{t+1}/Z_{t})}{\log(2)}\leq\frac{2\eta}{\log(2)}\leq\frac{1}{4}$, we have that $w^{\tau+1}_{i}\leq\frac{1}{2^{j+2}}\cdot 2^{\frac{1}{4}}\leq\frac{1}{2^{j+1}}$.
-So it suffices to choose $s_{i}=\tau+1$.
-∎
-C.2.2 Extracting regular segments from irregular segments
+See Appendix D.2. ∎
+In view of Lemma 18, we have
+$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}<\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}<\frac{1}{2}.$ (124)
+Observing that $\mathbb{P}_{\mathcal{G}^{\prime}}\{h_{\mathrm{out}}=h^{\star}\}\geq 3/4$, we can derive
+$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}$ $=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star}\right\}-\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}<\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}$
+$>\frac{3}{4}-\frac{1}{2}=\frac{1}{4},$
+which implies that
+$\mathbb{E}_{\mathcal{G}^{\prime}}[M_{\widetilde{i}}]$ $\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\cdot\mathbb{P}_{\mathcal{G}^{\prime}}\left\{M_{\widetilde{i}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\cdot\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}$
+$\geq\frac{\log(N_{0}/4)}{120000\varepsilon^{2}}\geq\frac{d-\log_{2}(8k)}{120000\varepsilon^{2}}\geq\frac{d}{240000\varepsilon^{2}}.$
+Summing over $i\in[k]$ gives
+$\mathbb{E}_{\mathcal{G}^{\prime}}\left[\sum_{i=1}^{k}M_{\widetilde{i}}\right]\geq\frac{dk}{240000\varepsilon^{2}},$ (125)
+thereby concluding the proof.
+D.2 Proof of Lemma 18
+Consider any $h\in\mathcal{H}_{\widetilde{i}}$ (and hence $h\neq h^{\star}$). For $v=[v_{p}]_{1\leq p\leq m}\in\{-1,1\}^{m}$, we let
+$n^{+}(v)=\sum_{p=1}^{m}\mathds{1}\{v_{p}=1\}$
+denote the number of $1$’s in the coordinates of $v$. Let $\mathcal{V}$ be a subset of $\{-1,1\}^{2m}$ defined as
+$\mathcal{V}\coloneqq\big{\{}v^{1},v^{2}\in\{-1,1\}^{m}\mid n^{+}(v^{1})-n^{+}(v^{2})\leq 4\sqrt{m}+2m\varepsilon\big{\}}.$
+Let $\mathbb{P}_{\mathcal{C}}\{\cdot\}$ denote the probability distribution of $\left\{\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}\right\}$, and $\mathbb{P}_{\mathcal{C}^{\prime}}\{\cdot\}$ the probability distribution of $\left\{\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}\right\}$.
+Hoeffding’s inequality then tells us that
+$\mathbb{P}_{\mathcal{C}}\{\mathcal{V}\}\geq\frac{3}{4}.$
+Also, observe that
+$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\left\{\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}\right\}\in\mathcal{V}\right\}$
+$\qquad\geq\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\right\}-\left(1-\mathbb{P}_{\mathcal{C}}\left\{\mathcal{V}\right\}\right)\geq\frac{3}{4}-\left(1-\frac{3}{4}\right)=\frac{1}{2},$ (126)
+namely,
+$\sum_{v=\{v^{1},v^{2}\}\in\mathcal{V}}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}\geq\frac{1}{2}.$
+In addition, for any $v=\{v^{1},v^{2}\}\in\mathcal{V}$, it is readily seen that
+$\mathbb{P}_{\mathcal{C}^{\prime}}\{v\}$ $=\mathbb{P}_{\mathcal{C}}\{v\}\cdot\left(1-8\varepsilon\right)^{n^{+}(v^{1})-n^{+}(v^{2})}(1+8\varepsilon)^{n^{+}(v^{2})-n^{+}(v^{1})}$
+$=\mathbb{P}_{\mathcal{C}}\{v\}\left(\frac{1-8\varepsilon}{1+8\varepsilon}\right)^{n^{+}(v^{1})-n^{+}(v^{2})}$
+$\geq\mathbb{P}_{\mathcal{C}}\{v\}\exp\Big{(}-20\big{(}n^{+}(v^{1})-n^{+}(v^{2})\big{)}\varepsilon\Big{)}$
+$\geq\mathbb{P}_{\mathcal{C}}\{v\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}),$
+where the last line follows from the definition of $\mathcal{V}$. As a result, we can demonstrate that, for any $v=\{v^{1},v^{2}\}\in\mathcal{V}$,
+$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m,\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
+$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}\mathbb{P}_{\mathcal{C^{\prime}}}\{v\}$
+$\geq\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}\mathbb{P}_{\mathcal{C}}\{v\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2})$
+$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\mid\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{2}\right\}\mathbb{P}_{\mathcal{C}}\{v\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2})$ (127)
+$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{2}\right\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}).$
+To see why (127) holds, observe that (here, for any $1\leq l\leq m$, let $v^{1}_{l}$ (resp. $v^{2}_{l}$) be the $l$-th coordinate of $v^{1}$ (resp. $v^{2}$)):
+$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
+$=\sum_{m^{\prime}=1}^{m}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}=m^{\prime}\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
+$=\sum_{m^{\prime}=1}^{m}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}=m^{\prime}\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m^{\prime}}=\{v^{1}_{l}\}_{l=1}^{m^{\prime}},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m^{\prime}}=\{v^{2}_{l}\}_{l=1}^{m^{\prime}}\right\}$ (128)
+$=\sum_{m^{\prime}=1}^{m}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}=m^{\prime}\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m^{\prime}}=\{v^{1}_{l}\}_{l=1}^{m^{\prime}},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m^{\prime}}=\{v^{2}_{l}\}_{l=1}^{m^{\prime}}\right\}$ (129)
+$=\sum_{m^{\prime}=1}^{m}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}=m^{\prime}\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$ (130)
+$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}.$
+where (129) results from Lemma 19 in Appendix D.3, and (128) and (130) hold since for $h^{\prime}=h,h^{\star}$, the event $\{h_{\mathrm{out}}=h^{\prime},M_{\widetilde{i}}=m^{\prime}\}$ is independent of $\{x^{l}(\widetilde{i})\}_{l\geq m^{\prime}+1}$.
+Taking the sum over $v=\{v^{1},v^{2}\}\in\mathcal{V}$, we obtain
+$\mathbb{P}_{\mathcal{G}^{\prime}}\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\}$
+$\geq\sum_{v\in\mathcal{V}}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m,\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{2}\right\}$
+$\geq\sum_{v\in\mathcal{V}}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{2}\right\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2})$
+$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\left\{\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m},\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}\right\}\in\mathcal{V}\right\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2})$
+$\geq\frac{1}{2}\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\big{\}}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}),$
+where the last line arises from (126).
+Summing over all $h\in\mathcal{H}_{\widetilde{i}}$, we reach
+$1\geq\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}\in\mathcal{H}_{\widetilde{i}},M_{\widetilde{i}}\leq m\big{\}}\geq\frac{N_{0}}{2}\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\big{\}}\exp(-40m\varepsilon^{2}-80\sqrt{m}\varepsilon),$ (131)
+given that each $\mathcal{H}_{\widetilde{i}}$ contains $N_{0}$ hypotheses. This in turn reveals that
+$\frac{1}{2}\leq\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\right\}\leq\frac{2}{N_{0}}\exp(40m\varepsilon^{2}+80\sqrt{m}\varepsilon).$ (132)
+Consequently, we arrive at
+$40m\varepsilon^{2}+80\sqrt{m}\varepsilon\geq\log(N_{0}/4),$
+which implies that
+$m\geq\min\bigg{\{}\frac{\log(N_{0}/4)}{80\varepsilon^{2}},\frac{\log^{2}(N_{0}/4)}{30000\varepsilon^{2}}\bigg{\}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}.$
+D.3 Statement and proof of Lemma 19
 Lemma 19.
-Recall the definition of $\mathcal{W}_{j}$ and $(s_{i},e_{i})$ for $i\in\mathcal{W}_{j}$ in Lemma 26.
-Then there exists a group of subsets $\{\mathcal{W}_{j}^{p}\}_{p=1}^{P}$ of $\mathcal{W}_{j}$ such that
-  1. (i).
-    $\mathcal{W}_{j}^{p}\subset\mathcal{W}_{j}$, $\mathcal{W}_{j}^{p}\cap\mathcal{W}_{j}^{p^{\prime}}=\emptyset$, $\forall p\neq p^{\prime}$;
-  2. (ii).
-    $\sum_{p=1}^{P}|\mathcal{W}_{j}^{p}|\geq\frac{|\mathcal{W}_{j}|}{3(\log_{2}(T)+1)}$;
-  3. (iii).
-    There exists $1\leq\tilde{s}_{1}<\tilde{e}_{1}\leq\tilde{s}_{2}<\tilde{e}_{2}\leq\dots\leq\tilde{s}_{P}<\tilde{e}_{P}\leq T$, such that for each $1\leq p\leq P$, $\tilde{s}_{p}=\min_{i\in\mathcal{W}_{j}^{p}}s_{i},\tilde{e}_{p}=\max_{i\in\mathcal{W}_{j}^{p}}e_{i}$ and $\max_{i\in\mathcal{W}_{j}^{p}}s_{i}\leq\min_{i\in\mathcal{W}_{j}^{p}}e_{i}$.
+For any $i\in[k]$ and $l\geq 1$, let $x^{l}(i)$ denote the $l$-th sample from $\mathcal{D}_{i}$.
+For any $\widetilde{i}\in[k]$, $h\in\mathcal{H}_{\widetilde{i}}$, $m>0$, and $v^{1},v^{2}\in\{-1,1\}^{2m}$, one has
+$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
+$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1}\right\}.$
 Proof.
-For $11\leq x\leq\log_{2}(T)+$ 1, we define $\mathcal{W}_{j}(x):=\{i|2^{x-1}\leq e_{i}-s_{i}\leq 2^{x}\}$ and let $x^{*}=\arg\max_{1\leq x\leq\log_{2}(T)+1}|\mathcal{W}_{j}(x^{*})|$.
-Then we have that $|\mathcal{W}_{j}(x^{*})|\geq\frac{|\mathcal{W}_{j}|}{\log_{2}(T)+1}$.
-Then we consider to choose the subsets $\{\mathcal{W}_{j}^{m}\}_{m=1}^{M}$ from $\mathcal{W}_{j}(x^{*})$.
-Let $\kappa_{1}=\min_{i\in\mathcal{W}_{j}(x^{*})}e_{i}$. We also define $\mathcal{U}_{j}^{1}:=\{i\in\mathcal{W}_{j}(x^{*})|s_{i}\leq\kappa_{1}\}$. For $o\geq 1$, we let
-$\kappa_{o+1}=\min_{i\in\mathcal{W}_{j}(x^{*})/\cup_{o^{\prime}=1}^{o}\mathcal{U}_{j}^{o^{\prime}}}e_{i}$
-$\mathcal{U}_{j}^{o+1}:=\{i\in\mathcal{W}_{j}(x^{*})/\cup_{o^{\prime}=1}^{o}\mathcal{U}_{j}^{o^{\prime}}|s_{i}\leq\kappa_{o+1}\}.$
-And we stop until $\cup_{o\geq 1}\mathcal{U}_{j}^{o}=\mathcal{W}_{j}(x^{*})$.
-Clearly, for each $o$ and $i_{1},i_{2}\in\mathcal{U}_{j}^{o}$, we have that $e_{i_{1}}\geq\kappa_{o}\geq s_{i_{2}}$.
-That is, $\max_{i\in\mathcal{U}_{j}^{o}}s_{i}\leq\min_{i\in\mathcal{U}_{j}^{o}}e_{i}$.
-We then consider the three groups of subsets of $\mathcal{W}_{j}(x^{*})$: $\{\mathcal{U}_{j}^{3o-2}\}_{o\geq 1}$, $\{\mathcal{U}_{j}^{3o-1}\}_{o\geq 1}$ and $\{\mathcal{U}_{j}^{3o}\}_{o\geq 1}$.
-Then there exists $i\in\{0,1,2\}$ such that $\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-i}|\geq\frac{1}{3}\sum_{o\geq 1}\mathcal{U}_{j}^{o}$.
-Without loss of generality, we assume that $\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-2}|\geq\frac{1}{3}\sum_{o\geq 1}\mathcal{U}_{j}^{o}=|\mathcal{W}_{j}(x^{*})|$.
-Now we verify that $\{\mathcal{U}_{j}^{3o-2}\}_{o\geq 1}$ is the desired group of sets.
-The first condition holds because of the definition of $\{\mathcal{U}_{j}^{o}\}_{o\geq 1}$.
-As for the second condition, we have that
-$\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-2}|\geq\frac{1}{3}\sum_{o\geq 1}\mathcal{U}_{j}^{o}=\frac{1}{3}|\mathcal{W}_{j}(x^{*})|\geq\frac{|\mathcal{W}_{j}|}{3(\log_{2}(T)+1)}.$
-As for the last condition, it suffices to verify that $\max_{i\in\mathcal{U}_{j}^{3o-2}}e_{i}\leq\min_{i\in\mathcal{U}_{j}^{3o+1}}s_{i}$ for any proper $o$.
-By definition, for each proper $o\geq 1$, there exists $i\in\mathcal{W}_{j}(x^{*})$ such that $s_{i}\geq\kappa_{o}$ and $\kappa_{o+1}=e_{i}$.
-Then we have that $\kappa_{o+1}=e_{i}\geq s_{i}+2^{x^{*}-1}\geq\kappa_{o}+2^{x^{*}-1}$.
-It then follows that, for any $i\in\mathcal{U}_{j}^{3o+1}$, $s_{i}\geq\kappa_{3o}\geq\kappa_{3o-2}+2^{x^{*}}$.
-On the other hand, for any $i\in\mathcal{U}_{j}^{3o-2}$, we have that $e_{i}\leq s_{i}+2^{x^{*}}\leq\kappa_{3o-2}+2^{x^{*}}$.
-Putting all together, we have that
-$\max_{i\in\mathcal{U}_{j}^{3o-2}}e_{i}\leq\kappa_{3o-2}+2^{x^{*}}\leq\min_{i\in\mathcal{U}_{j}^{3o+1}}s_{i}.$
-The proof is finished.
+Let $\overline{\sigma}$ be the permutation over $\mathcal{H}$ with $\overline{\sigma}(h^{\star})=h,\overline{\sigma}(h)=h^{\star}$ and $\overline{\sigma}(h^{\prime})=h^{\prime}$ for all $h^{\prime}\notin\{h,h^{\star}\}$.
+It then holds that $\overline{\sigma}^{-1}=\overline{\sigma}$.
+Consider a given sequence $\{m_{i}\}_{i=1}^{k}$.
+Let $X(i)=\{X^{l}(i)\}_{l=1}^{m_{i}}\in\{-1,0,1\}^{kNm_{i}}$ for $i\in[k]$, and let $x(i)=\{x^{l}(i)\}_{l=1}^{m_{i}}$ be the datapoints of the first $m_{i}$ calls to $\mathsf{Query}(D_{i})$.
+With slight abuse of notation, we take $\sigma(x(i))=\{\sigma(x^{l}(i))\}_{l=1}^{m_{i}}$ for each $i\in[k]$.
+It then follows from Lemma 20 in Appendix D.4 that
+$\mathbb{P}_{\mathcal{G},\mathcal{H}}\left\{h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right\}$
+$=\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=\sigma^{-1}(h),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right\},$
+which implies that
+$\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h^{\star},\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\sigma(x(i))=X(i),\forall i\in[k]\right\}$
+$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\overline{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right\}\cdot\frac{\mathbb{P}\left\{\sigma(x(i))=X(i),\forall i\in[k]\right\}}{\mathbb{P}\left\{\overline{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right\}}$
+$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\overline{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right\}\cdot\frac{\mathbb{P}\left\{\sigma(x(\widetilde{i}))=X(\widetilde{i})\right\}}{\mathbb{P}\left\{\overline{\sigma}\sigma(x(\widetilde{i}))=X(\widetilde{i})\right\}}$
+$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\overline{\sigma}\sigma(x(i)))=X(i),\forall i\in[k]\right\}$
+$\quad\qquad\qquad\qquad\qquad\cdot\frac{\mathbb{P}\left\{\{x_{j_{h,i}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X_{j_{\sigma(h),\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h^{\star},\widetilde{i}}}^{l}({\widetilde{i}})\}_{l=1}^{m_{\widetilde{i}}}=\{X_{j_{\sigma(h^{\star}),{\widetilde{i}}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\right\}}{\mathbb{P}\left\{\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X_{j_{\sigma(h),\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X_{j_{\sigma(h^{\star}),\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\right\}}.$ (133)
+Rearrange the equation to arrive at
+$\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\Bigg{\{}h_{\mathrm{out}}=h^{\star},\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\sigma(x(i))=X(i),\forall i\in[k]$
+|$\qquad\qquad\qquad\qquad\Big{|}\,\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\Bigg{\}}$
+$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\Bigg{\{}h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\overline{\sigma}\sigma(x(i))=X(i),\forall i\in[k]$
+|$\qquad\qquad\qquad\quad\Big{|}\,\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\Bigg{\}}.$
+Taking the sum over all possible choices of $\{X(i)\}_{i\neq\widetilde{i}}$, $\left\{\{X^{l}_{j_{h^{\prime},i^{\prime}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\right\}_{i^{\prime}\in[k],h^{\prime}\notin\{h,h^{\star}\}}$,
+$\left\{\{X^{l}_{j_{h^{\prime},i^{\prime}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\right\}_{h^{\prime}\in\{h,h^{\star}\},i^{\prime}\neq\widetilde{i}}$ and $\{m_{i}\}_{i\neq\widetilde{i}}$, we reach
+$\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\Bigg{\{}h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}=m_{\widetilde{i}}$
+|$\qquad\qquad\qquad\qquad\Big{|}\,\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\Bigg{\}}$
+$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\Bigg{\{}h_{\mathrm{out}}=h,M_{\widetilde{i}}=m_{\widetilde{i}}$
+|$\qquad\qquad\qquad\qquad\Big{|}\,\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\Bigg{\}}$
+for any $X(\widetilde{i})\in\{-1,0,1\}^{kNm_{\widetilde{i}}}$.
+Fix $m_{\widetilde{i}}=m$, and choose $\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=v_{1}$, $\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=v_{2}$.
+We then have
+$\mathbb{P}_{\mathcal{G}^{\prime},\mathcal{H}}\left\{h_{\mathrm{out}}=h^{\star},M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
+$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h^{\star},M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
+$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=\overline{\sigma}^{-1}(h^{\star}),M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1}\right\}$
+$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h,M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1}\right\}$
+$=\mathbb{P}_{\mathcal{G}^{\prime},\mathcal{H}}\left\{h_{\mathrm{out}}=h,M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1}\right\}.$
+This completes the proof. ∎
+D.4 Statement and proof of Lemma 20
 Lemma 20.
-Recall the definition of $\mathcal{W}_{j}$ and $(s_{i},e_{i})$ for $i\in\mathcal{W}_{j}$. Then there exists a group of subsets $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ such that the conditions below hold
-  1. (i).
-    $\mathcal{V}_{j}^{n}\subset\mathcal{W}_{j}$, $\mathcal{V}_{j}^{n}\cap\mathcal{V}_{j}^{n^{\prime}}=\emptyset$, $\forall n\neq n^{\prime}$;
-  2. (ii).
-    $\sum_{n=1}^{N}|\mathcal{V}_{j}^{n}|\geq\frac{|\mathcal{W}_{j}|}{24\log_{2}(k)(\log_{2}(T)+1)}$;
-  3. (iii).
-    There exists $1\leq\hat{s}_{1}<\hat{e}_{1}\leq\hat{s}_{2}<\hat{e}_{2}\leq\dots\leq\hat{s}_{N}<\hat{e}_{N}\leq T$, and $\{g_{n}\}_{n=1}^{N}\in\left[1,\infty\right)^{N}$ such that for each $1\leq n\leq N$, $(\hat{s}_{n},\hat{e}_{n})$ is a $\left(2^{-(j+1)}g_{n}|\mathcal{V}_{j}^{n}|,2^{-(j+2)}|\mathcal{V}_{j}^{n}|,\frac{\log(2)}{2\log_{2}(k)}\right)$-segment with index set as $\mathcal{V}_{j}^{n}$.
-That is, the following hold for each $1\leq n\leq N$:
-    • a.
-      $\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j+2}}\leq\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{\hat{s}_{n}}\leq\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j+1}}$ ;
-    • b.
-      $\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j}}\cdot\exp\left(\frac{\log(2)}{2\log_{2}(k)}\right)\leq\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{\hat{e}_{n}}$;
-    • c.
-      $\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{t}\geq\frac{|\mathcal{V}_{j}^{n}|}{2^{j+2}}$ for any $\hat{s}_{n}\leq t\leq\hat{e}_{n}$.
+Consider any $\{m_{i}\}_{i\in[k]}$, $\sigma\in\Pi_{\mathcal{H}}$ and $X\in\{-1,0,1\}^{kN\sum_{i=1}^{k}m_{i}}$. Let $\{X(i)\}_{i\in[k]}$ and $\{x(i)\}$ be defined as in Lemma 19. It then holds that
+$\mathbb{P}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right]$
+$=\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=\sigma^{-1}(h),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right\}.$ (134)
 Proof.
-Let $\{\mathcal{W}_{j}^{p}\}_{p=1}^{P}$ and $\{(\tilde{s}_{p},\tilde{e}_{p})\}_{p=1}^{P}$ be the construction in Lemma 27. Let $x=\log(2)$ be the parameter for the segment.
-Fix $1\leq p\leq P$. Let $t_{\mathrm{mid}}^{p}=\min_{i\in\mathcal{W}_{j}^{p}}e_{i}$. For each $i\in\mathcal{W}_{j}^{p}$, we have
-$\max\left\{\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right),\log\left(\frac{w^{t_{\mathrm{mid}}^{p}}_{i}}{w^{s_{i}}_{i}}\right)\right\}\geq\frac{1}{2}\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)+\frac{1}{2}\log\left(\frac{w^{t_{\mathrm{mid}}^{p}}_{i}}{w^{s_{i}}_{i}}\right)\geq\frac{x}{2}.$
-Then we have
-$\sum_{i\in\mathcal{W}_{j}^{p}}\left(\mathbb{I}\left[\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)\geq\frac{x}{2}\right]+\mathbb{I}\left[\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)\geq\frac{x}{2}\right]\right)\geq|\mathcal{W}_{j}^{p}|.$
-Without loss of generality, we assume that
-$\sum_{i\in\mathcal{W}_{j}^{p}}\left(\mathbb{I}\left[\log\left(\frac{w^{e_{i}}_{i}}{w^{t_{\mathrm{mid}}^{p}}_{i}}\right)\geq\frac{x}{2}\right]\right)\geq\frac{|\mathcal{W}_{j}^{p}|}{2},$ (70)
-Let $\widehat{\mathcal{W}}_{j}^{p}:=\{i\in\mathcal{W}_{j}^{p}|\log(w^{e_{i}}_{i}/w^{t_{\mathrm{mid}}^{p}}_{i})\geq\frac{x}{2}\}$.
-Then $|\widehat{\mathcal{W}}_{j}^{p}|\geq\frac{|\mathcal{W}_{j}^{p}|}{2}$.
-Let $Q$ denote $|\widehat{\mathcal{W}}_{j}^{p}|$.
-Without loss of generality, we assume $\widehat{\mathcal{W}}_{j}^{p}=\{1,2,\ldots,Q\}$ and $e_{1}\leq e_{2}\leq\ldots\leq e_{Q}$.
-Let ~$\tilde{\ell}=\max\{\ell\geq 0|2^{\ell}\leq Q\}$ and $\tilde{Q}=2^{\tilde{\ell}}$.
-[^6] We assume $\tilde{Q}\geq 2$ without loss of generality.
-In the case $\tilde{Q}=1$, we simply choose arbitrary element in $\widehat{\mathcal{W}}_{j}^{p}$ as a single subset.
-In this way, we can collect at least $\frac{1}{4}|\widehat{\mathcal{W}_{j}^{p}}|$ segments.
-Define $e_{0}=t_{\mathrm{mid}}^{p}$. We use $[e_{0},e_{k}]\oplus a$ as a shorthand of $[e_{0+a},e_{k+a}]$. Then we define a group of disjoint intervals of $[T]$ as below.
-$\mathcal{K}_{1}=\left\{[e_{0},e_{2^{\tilde{\ell}-1}}]\right\};$
-$\mathcal{K}_{2}=\left\{[e_{0},e_{2^{\tilde{\ell}-2}}],[e_{0},e_{2^{\tilde{\ell}-2}}]\oplus 2^{\tilde{\ell}-1}\right\};$
-$\mathcal{K}_{3}=\left\{[e_{0},e_{2^{\tilde{\ell}-3}}],[e_{0},e_{2^{\tilde{\ell}-3}}]\oplus 2^{\tilde{\ell}-2},[e_{0},e_{2^{\tilde{\ell}-3}}]\oplus 2\cdot 2^{\tilde{\ell}-2},[e_{0},e_{2^{\tilde{\ell}-3}}]\oplus 3\cdot 2^{\tilde{\ell}-2}\right\};$
-$\dots$
-$\mathcal{K}_{\ell}=\left\{[e_{0},e_{2^{\tilde{\ell}-\ell}}],[e_{0},e_{2^{\tilde{\ell}-\ell}}]\oplus 2^{\tilde{\ell}-\ell+1},[e_{0},e_{2^{\tilde{\ell}-\ell}}]\oplus 2\cdot 2^{\tilde{\ell}-\ell+1},\dots,[e_{0},e_{2^{\tilde{\ell}-\ell}}]\oplus(2^{\ell-1}-1)2^{\tilde{\ell}-\ell+1}\right\};$
-$\dots$
-$\mathcal{K}_{\tilde{\ell}}=\left\{[e_{2i},e_{2i+1}]|i=0,1,2,\ldots,2^{\tilde{\ell}-1}-1\right\}$
-$\mathcal{K}_{\tilde{\ell}+1}=\left\{[e_{2i+1},e_{2i+2}]|i=0,1,2,\ldots,2^{\tilde{\ell}-1}-1\right\}.$
-For each $i\in[\tilde{Q}-1]$ with binary form $\{i_{\ell}\}_{\ell=1}^{\tilde{\ell}}$ and $0\leq\ell\leq\tilde{\ell}$, we define $\mathrm{cut}(i,\ell)$ to be the number with binary form $\{i_{1},i_{2},\ldots,i_{\ell},0,0,\ldots,0\}$.
-For example, $\mathrm{cut}(i,0)=0$ and $\mathrm{cut}(i,\tilde{\ell})=i$.
-$\mathcal{J}(i)=\left\{\left[e_{\mathrm{cut}(i,\ell-1)},e_{\mathrm{cut}(i,\ell)}\right]\right\}_{\ell=1}^{\tilde{\ell}}.$
-Noting that for each $i\in[\tilde{Q}-1]$, we have that
-$\frac{x}{2}\leq\log\left(\frac{w_{i}^{e_{i}}}{w_{i}^{e_{0}}}\right)=\sum_{\ell=1}^{\tilde{\ell}}\log\left(\frac{w_{i}^{e_{\mathrm{cut}(i,\ell)}}}{w_{i}^{e_{\mathrm{cut}(i,\ell-1)}}}\right)=\sum_{\ell=1}^{\tilde{\ell}}\log\left(\frac{w_{i}^{e_{\mathrm{cut}(i,\ell)}}}{w_{i}^{e_{\mathrm{cut}(i,\ell-1)}}}\right)\mathbb{I}\left[e_{\mathrm{cut}(i,\ell)}\neq e_{\mathrm{cut}(i,\ell-1)}\right].$ (71)
-Then
-$\max_{1\leq\ell\leq\tilde{\ell}}\log\left(\frac{w_{i}^{e_{\mathrm{cut}(i,\ell)}}}{w_{i}^{e_{\mathrm{cut}(i,\ell-1)}}}\right)\geq\frac{x}{2\tilde{\ell}}.$ (72)
-By defining $\widetilde{\mathcal{W}}_{j}^{p}(\ell):=\left\{i\in\widehat{\mathcal{W}}_{j}^{p}:\ell=\arg\max_{1\leq\ell^{\prime}\leq\tilde{\ell}}\log\left(\frac{w_{i}^{e_{\mathrm{cut}(i,\ell^{\prime})}}}{w_{i}^{e_{\mathrm{cut}(i,\ell^{\prime}-1)}}}\right)\right\}$ for each $1\leq\ell\leq\tilde{\ell}$ [^7] Without loss of generality, we assume the $\arg\max$ function is a single value function., we then have that
-$\sum_{\ell=1}^{\tilde{\ell}}\left|\widetilde{\mathcal{W}}_{j}^{p}(\ell)\right|\geq\tilde{Q}-1.$ (73)
-As a result, there exists some $1\leq\ell^{*}\leq\tilde{\ell}$ such that $|\widetilde{W}_{j}^{p}(\ell^{*})|\geq\frac{\tilde{Q}-1}{\tilde{\ell}}\geq\frac{\tilde{Q}}{2\tilde{\ell}}$.
-By definition, for any $i$, if $\mathrm{cut}(i,\ell^{*})\neq\mathrm{cut}(i,\ell^{*}-1)$, then $[e_{\mathrm{cut}(i,\ell^{*}-1)},e_{\mathrm{cut}(i,\ell^{*})}]\in\mathcal{K}_{\ell^{*}}$.
-On the other hand, by definition of $\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*})$, we know that $\mathrm{cut}(i,\ell^{*})\neq\mathrm{cut}(i,\ell^{*}-1)$ for any $i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*})$.
-For each $1\leq o\leq 2^{\ell^{*}-1}$, we define
-$\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*},o):=\left\{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*})|[e_{\mathrm{cut}(i,\ell^{*}-1)},e_{\mathrm{cut}(i,\ell^{*})}]=[e_{0},e_{2^{\tilde{\ell}-\ell^{*}}}]\oplus(o-1)2^{\tilde{\ell}-\ell^{*}+1}\right\}.$ (74)
-Since we fix $p$ in the analysis above, we use *$\ell^{*}(p)$ and ~$\tilde{\ell}(p)$ to denote $\ell^{*}$ and $\tilde{\ell}$ respectively in the following analysis.
-Fix $1\leq p\leq P$ and $1\leq o\leq 2^{\ell^{*}(p)-1}$.
-We let $\hat{s}(p,o)=e_{(o-1)2^{\tilde{\ell}(p)-\ell^{*}(p)+1}}$ and $\hat{e}(p,o)=e_{2^{\tilde{\ell}(p)-\ell^{*}(p)}+(o-1)2^{\tilde{\ell}(p)-\ell^{*}(p)+1}}$.
-In words, $[\hat{s}(p,o),\hat{e}(p,o)]$ is the $o$-th interval in the set $\mathcal{K}_{\ell^{*}(p)}$.
-Then we claim that $\left\{\{\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)\}_{o=1}^{2^{\ell^{*}(p)-1}}\right\}_{p=1}^{P}$ with intervals $\left\{\{\hat{s}(p,o),\hat{e}(p,o)\}_{o=1}^{2^{\ell^{*}(p)-1}}\right\}_{p=1}^{P}$ is the desired sub-partition for $\mathcal{W}_{j}^{p}$.
-With a slight abuse of notations, we use $(p,o)$ as the index of the segments instead of $n$.
-We then verify the three conditions as below.
-The condition (i).
-It is clear that $\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)\subset\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p))\subset\widehat{\mathcal{W}}_{j}^{p}\subset\mathcal{W}_{j}^{p}\subset\mathcal{W}_{j}$.
-On the other hand, $\tilde{W}_{j}^{p_{1}}(\ell^{*}(p_{1}),o_{1})\cap\tilde{W}_{j}^{p_{2}}(\ell^{*}(p_{2}),o_{2})\neq\emptyset$ implies that $\mathcal{W}_{j}^{p_{2}}\cap\mathcal{W}_{j}^{p_{2}}\neq\emptyset$.
-Then we have $p_{1}=p_{2}$.
-Again by definition in (74), we have $o_{1}=o_{2}$.
-The condition (ii).
-Because
-$\sum_{o=1}^{2^{\ell^{*}(p)-1}}|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|=|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p))|\geq\frac{|\widehat{\mathcal{W}}_{j}^{p}|}{4\log_{2}(|\widehat{\mathcal{W}}_{j}^{p}|)}\geq\frac{|\mathcal{W}_{j}^{p}|}{8\log_{2}(|\mathcal{W}_{j}^{p}|)}.$ (75)
-Taking sum over $p$, and using Lemma 27, we have that
-$\sum_{p=1}^{P}\sum_{o=1}^{2^{\ell^{*}(p)-1}}|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|=|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p))|\geq\sum_{p=1}^{P}\frac{|\mathcal{W}_{j}^{p}|}{8\log_{2}(k)}\geq\frac{|\mathcal{W}_{j}|}{24\log_{2}(k)(\log_{2}(T)+1)}.$ (76)
-The condition (iii).
-We set the parameter $\left\{\{g(p,o)\}_{o=1}^{2^{\ell^{*}(p)}}\right\}_{p=1}^{P}$ as $g(p,o)=\frac{\sum_{i\in\widetilde{\mathcal{W}}_{j}(\ell^{*}(p),o)}w_{i}^{\hat{s}(p,o)}}{2^{-(j+2)}\cdot|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|}\geq 1$.
-Then we can verify $(a)$ as
-$\sum_{i\in\widetilde{\mathcal{W}}_{j}(\ell^{*}(p),o)}w_{i}^{\hat{s}(p,o)}=\frac{g(p,o)|\widetilde{\mathcal{W}}_{j}(\ell^{*}(p),o)|}{2^{j+2}}.$
-Next we verify $(b)$.
-For any $i\in\widehat{\mathcal{W}}_{j}^{p}$, we have that $s_{i}\leq e_{\mathrm{cut}(i,\ell-1)}\leq e_{i}$ for any $1\leq\ell\leq\tilde{\ell}(p)$.
-It then holds that $s_{i}\leq\hat{s}(p,o)\leq e_{i}$ for any $i\in\widehat{\mathcal{W}}^{p}_{j}$.By definition of $(s_{i},e_{i})$, we have that $w_{i}^{\hat{s}(p,o)}\geq 2^{-(j+2)}$.
-By definition, for any $i\in\tilde{\mathcal{w}}_{j}^{p}(\ell^{*}(p),o)$, it then holds that
-$\log\left(\frac{w_{i}^{\hat{e}(p,o)}}{w_{i}^{\hat{s}(p,o)}}\right)\geq\frac{x}{2\tilde{\ell}(p)};$
-$w_{i}^{\hat{s}(p,o)}\geq 2^{-(j+2)}.$
-Recall that $x=\log(2)$. It then follow that
-$\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{s}(p,o)}_{i}\geq 2^{-(j+2)}\cdot|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|$
-$\log\left(\frac{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{e}(p,o)}_{i}}{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{s}(p,o)}_{i}}\right)\geq\log\left(\frac{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{s}(p,o)}_{i}\cdot\exp\left(\frac{x}{2\tilde{\ell}(p)}\right)}{\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{\hat{s}(p,o)}_{i}}\right)=\frac{x}{2\tilde{\ell}(p)}\geq\frac{\log(2)}{2\log_{2}(k)}.$
-As for $(c)$, we note that for any $\hat{s}(p,o)\leq t\leq\hat{e}(p,o)$, and $i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)$, it holds that $s_{i}\leq t\leq e_{i}$. As a result, we have that
-$\sum_{i\in\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)}w^{t}_{i}\geq|\widetilde{\mathcal{W}}_{j}^{p}(\ell^{*}(p),o)|\cdot 2^{-(j+2)}.$
-Finally, we need to verify that $[\hat{s}(p_{1},o_{1}),\hat{o}(p_{1},o_{1})]$ and $[\hat{s}(p_{2},o_{2}),\hat{e}(p_{2},o_{2})]$ have no common inner point.
-In the case $p_{1}\neq p_{2}$ (assuming $p_{1}<p_{2}$), we have that $\tilde{s}_{p_{1}}\leq\hat{s}(p_{1},o_{1})<\hat{e}(p_{1},o_{1})\leq\tilde{e}_{p_{1}}\leq\tilde{s}_{p_{2}}\leq\hat{s}(p_{2},o_{2})<\hat{e}(p_{2},o_{2})$.
-In the case $p_{1}=p_{2}$ and $o_{1}\neq o_{2}$ (assuming $o_{1}<o_{2}$), we have that $\hat{s}(p_{1},o_{1})<\hat{e}(p_{1},o_{1})<\hat{s}(p_{1},o_{2})<\hat{e}(p_{1},o_{2})$.
-The proof is finished.
-C.2.3 Putting all this together
-Let $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ and $\{(\hat{s}_{n},\hat{e}_{n})\}_{n=1}^{N}$ be the construction in Lemma 28.
-We then divide $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ into two parts.
-The first part consists of $\mathcal{V}_{j}^{n}$ such that
-$2^{-(j+2)}|\mathcal{V}_{j}^{n}|\cdot\frac{\log^{2}(2)}{50(\log_{2}(k)+1)^{2}\log^{2}_{2}(k)}\geq\max\left\{2\log(1/\delta^{\prime})\eta,1/k\right\},$ (77)
-which we denote as $\{\bar{\mathcal{V}}_{j}^{n}\}_{n=1}^{\bar{N}}$. And the left part consists of the left sets, which we denote as $\{\underline{\mathcal{V}}_{j}^{n}\}_{n=1}^{\underline{N}}$.
-By Lemma 29, for each $1\leq n\leq\bar{N}$, there exists $11\leq j^{\prime}\leq 2\log_{2}(k)+$ 1 such that
-$\sum_{\tau=\hat{s}_{n}}^{\hat{e}_{n}-1}\mathbb{I}[-v^{\tau}\geq 2^{-j^{\prime}}]\geq\frac{2^{-(j+2)}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)\cdot 2^{j^{\prime}-1}}{200\log^{2}_{2}(k)(2\log_{2}(k)+1)^{2}\eta}$
-or $\hat{e}_{n}-\hat{s}_{n}\geq\frac{2^{-(j+2)}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)}{200\log^{2}_{2}(k)(\log_{2}(k)+1)^{2}\eta^{2}}$.
-As a result, for each $1\leq n\leq\bar{N}$, we have that
-$\sum_{j^{\prime}=1}^{2\log_{2}(k)+1}\sum_{\tau=\hat{s}_{n}}^{\hat{e}_{n}-1}\mathbb{I}[-v^{\tau}\geq 2^{-j^{\prime}}]2^{-(j^{\prime}-1)}+(\hat{e}_{n}-\hat{s}_{n})\eta\geq\frac{2^{-(j+2)}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)\cdot}{200\log^{2}_{2}(k)(2\log_{2}(k)+1)^{2}\eta}.$ (78)
-By noting that $\sum_{j^{\prime}=1}^{2\log_{2}(k)+1}\mathbb{I}[x\geq 2^{-j^{\prime}}]\cdot 2^{-(j^{\prime}-1)}\leq 4x$ for any $x\geq 0$, we have that
-$\sum_{\tau=\hat{s}_{n}}^{\hat{e}_{n}-1}4\cdot(-v^{\tau})+2(\hat{e}_{n}-\hat{s}_{n})\eta\geq\frac{2^{-(j+2)}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)\cdot}{200\log^{2}_{2}(k)(2\log_{2}(k)+1)^{2}\eta}.$ (79)
-Taking sum over $n$ we have that
-$T\eta+4\sum_{t=1}^{T}(-v^{t})\geq\sum_{n=1}^{\bar{n}}\left(\sum_{\tau=\hat{s}_{n}}^{\hat{e}_{n}-1}4\cdot(-v^{\tau})+(\hat{e}_{n}-\hat{s}_{n})\eta\right)\geq\frac{2^{-(j+2)}\sum_{n=1}^{\bar{N}}|\bar{\mathcal{V}}_{j}^{n}|\log^{2}(2)\cdot}{200\log^{2}_{2}(k)(2\log_{2}(k)+1)^{2}\eta}.$ (80)
-On the other hand, by (39)
-$\sum_{t=1}^{T}(-v^{t})\leq 100\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right),$ (81)
-we have that
-$\sum_{n=1}^{\bar{N}}|\bar{\mathcal{V}}_{j}^{n}|$ $\leq\frac{3200(\log_{2}(k)+1)^{4}\eta\cdot 2^{j+2}}{\log^{2}(2)}\cdot\left(100\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)\right)+$ (82)
-$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{800T(\log_{2}(k)+1)^{4}\cdot 2^{j+2}\eta^{2}}{\log^{2}(2)}.$
-On the other hand, by the first part of Lemma 29, we have that $\hat{e}_{n}-\hat{s}_{n}\geq\frac{\log(2)}{4\eta\log_{2}(k)}$. As a result, we have that
-$\sum_{n=1}^{\underline{N}}|\underline{\mathcal{V}}_{j}^{n}|$ $\leq\frac{4T\eta\log_{2}(k)}{\log(2)}\cdot 2^{j+2}\frac{2\log_{2}(k)}{\log(2)}\cdot\max\{2\log(1/\delta^{\prime})\eta,1/k\}$
-$\leq\frac{16\cdot 2^{j}T\eta^{2}\log_{2}^{2}(k)\log(1/\delta^{\prime})}{\log^{2}(2)}+\frac{82^{j}T\eta\log_{2}^{2}(k)}{k\log^{2}(2)}.$
-Note that $2\log(1/\delta^{\prime})\eta\geq\frac{1}{k}$ by the parameter configuration. We obtain that
-$\sum_{n=1}^{N}|\mathcal{V}_{j}^{n}|\leq 3200000\left((\log_{2}(k)+1)^{3}(\log(k)+\log(1/\delta^{\prime})^{2}\right)\cdot 2^{j}.$ (83)
-It then follows that
-$|\mathcal{W}_{j}|\leq 8\cdot 10^{7}\cdot\left((\log_{2}(k)+1)^{4}\left(\log(k)+\log(1/\delta^{\prime})\right)^{2}(\log_{2}(T)+1)\right)\cdot 2^{j}.$
-The proof is thus completed.
-C.3 Statement and proof of Lemma 29
-Recall the definition of segment in Definition 2. We have the following lemma to bound the length of segments.
-Lemma 29.
-Assume the conditions in Lemma 8 hold.
-Suppose $(t_{1},t_{2})$ is a $\left(p,q,x\right)$-segment such that $p\geq 2q$.
-Then $t_{2}-t_{1}\geq\frac{x}{2\eta}$.
-Moreover, if we assume that $\frac{qx^{2}}{50(\log_{2}(k)+1)^{2}}\geq\max\{2\log(1/\delta^{\prime})\eta,1/k\}$, it then holds that: with probability $1-6T^{4}k\delta^{\prime}$, at least one of the two claims holds: (1) there exists $11\leq j\leq 2\log_{2}(k)+$ 1 such that $\sum_{\tau=t_{1}}^{t_{2}-1}\mathbb{I}[-v^{\tau}\geq 2^{-j}]\geq\frac{qx^{2}\cdot 2^{\tilde{j}-1}}{100(2\log_{2}(k)+1)^{2}\eta}$; (2) $t_{2}-t_{1}\geq\frac{qx^{2}}{100(\log_{2}(k)+1)^{2}\eta^{2}}$.
-Proof.
-Let $Z^{t}=\sum_{i=1}^{k}W_{i}^{t}$. We start by proving the first argument. By definition, there exists $i\in[k]$ such that
-$\log\left(\frac{w^{t_{2}}_{i}}{w^{t_{1}}_{i}}\right)\geq x,$
-which corresponds to
-$\eta\sum_{\tau=t_{1}}^{t_{2}-1}\hat{r}_{i}^{\tau}-\log(Z^{t_{2}}/Z^{t_{1}})\geq x.$
-Because $\log(Z^{t_{2}}/Z^{t_{1}})\geq-\eta(t_{2}-t_{1})$ and $\hat{r}_{i}^{\tau}\leq 1$ for any $1\leq\tau\leq T$, we have that
-$x\leq 2(t_{2}-t_{1})\eta.$ (84)
-Then the conclusion follows.
-With a slight abuse of notation, we also use $h$ to denote the vector $\{L(h,\mathbf{e}_{i})-\mathrm{OPT}\}_{i\in[k]}$.
-By the conditions about the $\varepsilon$-optimility of $h^{t}$ in Lemma 8, we have that for any $1\leq\tau,t\leq T$, it holds that
-$h^{\tau}\cdot w^{t}\geq h^{t}\cdot w^{t}-\varepsilon_{1}=v^{t}-\varepsilon_{1}.$ (85)
-Taking sum over $\tau$, we have that
-$\sum_{\tau=t_{1}}^{t_{2}-1}h^{\tau}\cdot w^{t}\geq(t_{2}-t_{1})(v^{t}-\varepsilon_{1}).$
-On the other hand, we note that $W^{t}_{i}=\exp(\eta\sum_{\tau=1}^{t}\left(h^{\tau}_{i}+\mathrm{OPT}+\xi^{\tau}_{i}\right))$, where $\xi_{i}^{\tau}=\hat{r}_{i}^{\tau}-h^{\tau}_{i}-\mathrm{OPT}$ is a zero-mean noise.
-Then we learn that
-$\sum_{\tau=t_{1}}^{t_{2}-1}h^{\tau}\cdot w^{t}=w^{t}\cdot\left(\frac{1}{\eta}\log\left(\frac{W^{t_{2}}}{W^{t_{1}}}\right)-\Delta_{t_{1},t_{2}}\right)-(t_{2}-t_{1})\mathrm{OPT}\geq(t_{2}-t_{1})(v^{t}-\varepsilon_{1}),$ (86)
-where $\log(x/y)$ denotes the vector $\{\log(x_{i}/y_{i})\}_{i\in[k]}$ for two $k$-dimensional vectors $(x,y)$, and $\Delta_{t_{1},t_{2}}=\sum_{\tau=t_{1}}^{t_{2}-1}\xi^{\tau}$.
-Recall that $Z^{t}=\sum_{i=1}^{k}W^{t}_{i}$ and $w^{t}_{i}=\frac{W^{t}_{i}}{Z^{t}}$. By choose $t_{1}=t$, we then have that
-$w^{t}\cdot\left(\log\left(\frac{w^{t_{2}}}{w^{t}}\right)-\eta\Delta_{t,t_{2}}\right)+\log\left(\frac{Z^{t_{2}}}{Z^{t}}\right)-\eta(t_{2}-t)\mathrm{OPT}\geq\eta(t_{2}-t)(v^{t}-\varepsilon_{1}).$
-Reversing the inequality, we have that
-$\mathsf{KL}(w^{t}\,\|\,w^{t_{2}})$
-$:=w^{t}\cdot\left(\log\left(\frac{w^{t}}{w^{t_{2}}}\right)\right)$
-$\leq\log(Z^{t_{2}}/Z^{t})-\eta(t_{2}-t)\mathrm{OPT}-\eta w^{t}\cdot\Delta_{t,t_{2}}+(t_{2}-t)\eta(\varepsilon_{1}-v^{t}).$ (87)
-(87) provides an upper bound for the KL divergence between $w^{t}$ and $w^{t_{2}}$. We then carefully deal with the RHS of (87).
-We start with bounding $\log(Z^{t_{2}}/Z^{t})$. With probability $1-2T^{2}k\delta^{\prime}$, it holds that
-$\log(Z^{t_{2}}/Z^{t})$ $:=\sum_{\tau=t}^{t_{2}-1}\log(Z^{\tau+1}/Z^{\tau})$
-$=\sum_{\tau=t}^{t_{2}-1}\log\left(\sum_{i=1}^{k}w^{\tau}_{i}\exp(\eta\hat{r}_{i}^{\tau})\right)$
-$\leq\sum_{\tau=t}^{t_{2}-1}\left(\eta\sum_{i=1}^{k}w_{i}^{\tau}\hat{r}_{i}^{\tau}+2\eta^{2}\right)$
-$\leq\eta\sum_{\tau=t}^{t_{2}-1}v^{\tau}+\eta(t_{2}-t)\mathrm{OPT}+\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}(w_{i}^{\tau}(\hat{r}_{i}^{\tau}-h_{i}^{\tau}-\mathrm{OPT}))+2(t_{2}-t)\eta^{2}$ (88)
-$\leq\eta(t_{2}-t)\varepsilon_{1}+\eta(t_{2}-t)\mathrm{OPT}+\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}(w_{i}^{\tau}(\hat{r}_{i}^{\tau}-h_{i}^{\tau}-\mathrm{OPT}))+2(t_{2}-t)\eta^{2}.$ (89)
-(88) holds by the fact that $v^{t}=w^{t}\cdot h^{t}$, and (89) is by the fact that $v^{\tau}\leq f^{\tau}+\varepsilon_{1}\leq\varepsilon_{1}$.
-We then consider to bound the two following noise terms.
-$\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}(w_{i}^{\tau}(\hat{r}_{i}^{\tau}-h_{i}^{\tau}-\mathrm{OPT}))$ $=\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{\tau}\xi_{i}^{\tau},$
-$-\eta w^{t}\Delta_{t,t_{2}}$ $=-\eta\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{t}\xi_{i}^{\tau}.$ (90)
-Let $\mathcal{F}^{\tau}$ denote the event field before the $\tau$-th round in Algorithm 1.
-Note that the variance of $\xi_{i}^{\tau}$ is at most $O\left(\frac{1}{k\bar{w}^{\tau}_{i}}\right)$ by the update rule (See Line 1 in Algorithm 1), and $\{\xi_{i}^{\tau}\}$ are independent conditioned on $\mathcal{F}^{\tau}$.
-Let $q^{\tau}=\sum_{i=1}^{k}w_{i}^{t}\xi_{i}^{\tau}$.
-Then we have that $|q^{\tau}|\leq 1$, $\mathbb{E}[q^{\tau}|\mathcal{F}^{\tau}]=0$ and $\mathbb{V}[q^{\tau}|\mathcal{F}^{\tau}]\leq\sum_{i=1}^{k}\frac{(w_{i}^{t})^{2}}{k\bar{w}_{i}^{\tau}}\leq\sum_{i=1}^{k}\frac{w_{i}^{t}}{k}=\frac{1}{k}$.
-By Freedman’s inequality (Lemma 20), with probability $1-\delta^{\prime}$,
-$\left|\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{t}\xi_{i}^{\tau}\right|\leq 2\sqrt{\frac{(t_{2}-t)}{k}\log(2/\delta^{\prime}))}+2\log(2/\delta^{\prime});$ (91)
-For the other term, by letting $\hat{q}^{\tau}=\sum_{i=1}^{k}w_{i}^{\tau}\xi_{i}^{\tau}$, we then have $|\hat{q}^{\tau}|\leq 1$, $\mathbb{E}[\hat{q}^{\tau}|\mathcal{F}^{\tau}]=0$ and $\mathbb{V}[\hat{q}^{\tau}|\mathcal{F}^{\tau}]\leq\sum_{i=1}^{k}\frac{(w_{i}^{\tau})^{2}}{k\bar{w}_{i}^{\tau}}\leq\sum_{i=1}^{k}\frac{w_{i}^{\tau}}{k}=\frac{1}{k}$.
-Again, by Freedman’s inequality (Lemma 20), with probability $1-\delta^{\prime}$
-$\left|\sum_{\tau=t}^{t_{2}-1}\sum_{i=1}^{k}w_{i}^{\tau}\xi_{i}^{\tau}\right|\leq 2\sqrt{\frac{(t_{2}-t)}{k}\log(2/\delta^{\prime}))}+2\log(2/\delta^{\prime}).$ (92)
-By (87), (89), (91) and (92), and using union bounds over $(t,t_{2})$, with probability $1-6T^{4}k\delta^{\prime}$, we have that
-$\mathsf{KL}(w^{t}\,\|\,w^{t_{2}})$ $\leq 2(t_{2}-t)\eta\varepsilon_{1}-(t_{2}-t)\eta v^{t}$
-$\quad\quad\quad+2\eta\sqrt{\frac{(t_{2}-t)\log(1/\delta^{\prime})}{k}}+2(t_{2}-t)\eta^{2}+2\log(1/\delta^{\prime})\eta$ (93)
-holds for any $1\leq t<t_{2}\leq T$.
+Let $\mathcal{H}^{\prime}=\sigma(\mathcal{H})$. Let $h_{p}(\cdot)$ denote the $p$-th hypothesis in the hypothesis set. Then one has
+$\mathbb{P}_{\mathcal{G},\mathcal{H}}\left\{h_{\mathrm{out}}=h_{p}(\mathcal{H}),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right\}$
+$=\mathbb{P}_{\mathcal{G},\mathcal{H}}\left\{h_{\mathrm{out}}=h_{p}(\mathcal{H}),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\{\{x^{l}_{j_{h_{p^{\prime}}(\mathcal{H}),i}}(i^{\prime})\}_{p^{\prime}=1,i=1}^{|\mathcal{H}|,k}\}_{l=1}^{m_{i^{\prime}}}\}_{i^{\prime}=1}^{k}=X\right\}$
+$=\mathbb{P}_{\mathcal{G},\mathcal{H}^{\prime}}\left\{h_{\mathrm{out}}=h_{p}(\mathcal{H}^{\prime}),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\{\{x^{l}_{j_{h_{p^{\prime}}(\mathcal{H}^{\prime}),i}}(i^{\prime})\}_{p^{\prime}=1,i=1}^{|\mathcal{H}^{\prime}|,k}\}_{l=1}^{m_{i^{\prime}}}\}_{i^{\prime}=1}^{k}=X\right\}$ (135)
+$=\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h_{p}(\sigma(\mathcal{H})),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right\}$
+$=\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=\sigma^{-1}(h_{p}(\mathcal{H})),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right\}.$ (136)
 […]
-For $1\leq j\leq j_{\mathrm{max}}:=\left\lfloor 2\log_{2}(k)+1\right\rfloor$, we let
-$\tau_{j}:=\min\{\tau|t_{1}\leq\tau\leq t_{2}-1,\,-v^{\tau}\leq 2^{-(j-1)}\}.$
-We also define $\tau_{j_{\mathrm{max}+1}}=t_{2}$.
-By definition, we know that $\tau_{1}=t_{1}$ and $\tau_{j_{2}}\geq\tau_{j_{1}}$ for $j_{2}\geq j_{1}$.
-Let $\mathcal{I}$ be the index set of this segment.
-Let $y_{j}:=\sum_{i\in\mathcal{I}}w_{i}^{\tau_{j}}$.
-We then claim that there exists $1\leq\tilde{j}\leq j_{\mathrm{max}}$ such that
-$\log\left(\frac{y_{\tilde{j}+1}}{y_{\tilde{j}}}\right)$ $\geq\frac{x}{2\log_{2}(k)+1}.$ (94)
-Suppose that none of $1\leq\tilde{j}\leq j_{\mathrm{max}}$ satisfies (94).
-Then for any $j$, it holds that $\log\left(\frac{y_{j+1}}{y_{j}}\right)\leq\frac{x}{2\log_{2}(k)+1}$, which implies that $y_{j}\geq y_{j+1}\exp\left(\frac{x}{2\log_{2}(k)+1}\right)$.
-As a result, we have that $y_{1}\geq y_{j_{\mathrm{max}}+1}\cdot\exp\left(-j_{\mathrm{max}}\cdot\frac{x}{2\log_{2}(k)+1}\right)\geq p$, which leads to contradiction.
-Now we assume $\tilde{j}$ satisfies (94). By the definition, we have that $y_{\tilde{j}}\geq q$.
-By (93), we have that
-$\mathsf{KL}(w^{\tilde{j}}\,\|\,w^{\tau_{\tilde{j}+1}})$ $\leq 2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta\varepsilon_{1}+(\tau_{{\tilde{j}}+1}-\tau_{\tilde{j}})\eta 2^{-(\tilde{j}-1)}$
-$\quad+2\eta\sqrt{\frac{(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\log(1/\delta^{\prime})}{k}}+2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta^{2}+2\log(1/\delta^{\prime})\eta.$ (95)
-Since $\log\left(\frac{y_{\tilde{j}+1}}{y_{\tilde{j}}}\right)\geq\frac{x}{2\log_{2}(k)+1}$ and $y_{\tilde{j}}\geq q$, by Lemma 23 and Lemma 24, we have that
-$\mathsf{KL}(w^{\tau_{\tilde{j}}}\,\|\,w^{\tau_{\tilde{j}+1}})\geq\mathsf{KL}\left(\mathrm{Ber}\left(y_{\tilde{j}}\right)\,\|\,\mathrm{Ber}\left(y_{\tilde{j}+1}\right)\right)\geq\frac{qx^{2}}{4(2\log_{2}(k)+1)^{2}},$
-where $\mathrm{Ber}(x)$ denote the Bernoulli distribution with mean $x\in[0,1]$. As a result, we obtain that
-$\frac{qx^{2}}{4(2\log_{2}(k)+1)^{2}}$ $\leq 2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta\varepsilon_{1}+(\tau_{{\tilde{j}}+1}-\tau_{\tilde{j}})\eta 2^{-({\tilde{j}}-1)}$
-$\quad+2\eta\sqrt{\frac{(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\log(1/\delta^{\prime})}{k}}+2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta^{2}+2\log(1/\delta^{\prime})\eta.$ (96)
-By noting that $\frac{qx^{2}}{50(2\log_{2}(k)+1)^{2}}\geq\max\{2\log(1/\delta^{\prime})\eta,1/k\}$ and the fact that $\frac{2^{j_{\mathrm{max}}}}{\eta}\geq\frac{1}{\eta^{2}}$, we have that
-$\tau_{\tilde{j}+1}-\tau_{\tilde{j}}$ $\geq\min\left\{\frac{qx^{2}}{100(2\log_{2}(k)+1)^{2}}\min\left\{\frac{1}{\eta\varepsilon_{1}},\frac{2^{\tilde{j}-1}}{\eta},\frac{1}{\eta^{2}}\right\},\frac{kx^{2}2^{-2l_{1}}}{10000\eta^{2}\log(1/\delta^{\prime})(\log_{2}(k)+1)^{4}}\right\}$
-$\geq\frac{qx^{2}\cdot 2^{\tilde{j}-1}}{100(2\log_{2}(k)+1)^{2}\eta}.$ (97)
-By definition, we learn that $\sum_{\tau=t_{1}}^{t_{2}-1}\mathbb{I}[-v^{\tau}\geq 2^{-j}]\geq\tau_{j+1}-t_{1}\geq\tau_{j+1}-\tau_{j}$ for any $11\leq j\leq 2\log_{2}(k)+$ 1.
-As a result we have that $\sum_{\tau=t_{1}}^{t_{2}-1}\mathbb{I}[-v^{\tau}\geq 2^{-\tilde{j}}]\geq\frac{qx^{2}\cdot 2^{\tilde{j}-1}}{100(2\log_{2}(k)+1)^{2}\eta}$.
-The proof is completed by noting that $\tau_{j_{\mathrm{max}}+1}=t_{2}$.
-Appendix D Missing proofs for learning Rademacher classes
-D.1 Restatement and Proof of Lemma 15
-Lemma 15. [Restatement] By running Algorithm 3, with probability $41-\delta/$ 4, it holds that
-$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (98)
-for any $1\leq t\leq T$.
-Proof.
-We will follow the notation in the proof of Lemma 4.
-Fix $n=\{n_{i}\}_{i=1}^{k}$ such that $n_{i}\geq 12\log(2k)$ for all $i\in[k]$ and $w\in\Delta(k)$.
+Appendix E Proofs of auxiliary lemmas for Rademacher classes
+E.1 Proof of Lemma 6
+In this subsection, we shall follow the notation adopted in the proof of Lemma 1 (e.g., the dataset $\widetilde{\mathcal{S}}$, the data subset $\widetilde{\mathcal{S}}(n)$ and its independent copy $\widetilde{\mathcal{S}}^{+}(n)$, the Rademacher random variables $\{\sigma_{i}^{j}\}$).
+Consider any given $n=\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for all $i\in[k]$, and any given $w\in\Delta(k)$.
 Let $\kappa=\min_{i}\frac{n_{i}}{w_{i}}$.
-Recall that $(x_{i,j},y_{i,j})$ is the $j$-th sample from $D_{i}$.
+Recall that $(x_{i,j},y_{i,j})$ is the $j$-th sample from $\mathcal{D}_{i}$, and $\{(x_{i,j}^{+},y_{i,j}^{+}\}$ are independent copies.
 Define
-$F(n,w):=\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right)\right].$
-Let $\{\{(\tilde{x}_{i,j},\tilde{y}_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}$ be a group of ghost samples for $\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}$. We then have that
-$F(n,w)$ (99)
-$\leq\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{(\tilde{x}_{i,j},\tilde{y}_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\left(\ell(h,(x_{i,j},y_{i,j}))-\ell(h,(\tilde{x}_{i,j},\tilde{y}_{i,j}))\right)\right)\right]$
-$=\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{(\tilde{x}_{i,j},\tilde{y}_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\left(\ell(h,(x_{i,j},y_{i,j}))-\ell(h,(\tilde{x}_{i,j},\tilde{y}_{i,j}))\right)\right)\right]$
-$\leq 2\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{1}{\kappa}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,(x_{i,j},y_{i,j}))\right)\right]$ (100)
-$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}.$ (101)
-Here (100) is by Lemma 30 as below.
-Lemma 30.
-Let $\mathcal{L}$ be a subset of $\mathbb{R}^{n}$. Let $w^{1},w^{2}\in\mathbb{R}^{n}$ be such that $|w^{1}_{i}|\leq|w_{i}^{2}|$ for all $i\in[n]$. Then it holds
-$\mathbb{E}_{\{\sigma_{i}\}_{i=1}^{n}\overset{\mathrm{iid}}{\sim}\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{1}_{i}f_{i}\right]\leq\mathbb{E}_{\{\sigma\}_{i=1}^{n}\overset{\mathrm{iid}}{\sim}\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{2}_{i}f_{i}\right].$ (102)
-Proof.
-It suffices to prove (102) under the case where $w_{i}^{1}=w^{2}_{i}$ for $11\leq i\leq n-$ 1, and $|w^{1}_{n}|\leq|w^{2}_{n}|$. Fix $\sigma_{i}$ for $11\leq i\leq n-$ 1.
-$\mathbb{E}_{\sigma_{n}\sim\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w_{i}^{1}f_{i}\right]$
-$=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{1}f_{i}+w_{n}^{1}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{1}f_{i}-w_{n}^{1}f_{n}\right)$
-$=\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{1}(f_{i}^{1}+f_{i}^{2})+w_{n}^{1}(f_{n}^{1}-f_{n}^{2})\right)$
-$=\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}(f_{i}^{1}+f_{i}^{2})+w_{n}^{1}(f_{n}^{1}-f_{n}^{2})\right)$
-$\leq\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}(f_{i}^{1}+f_{i}^{2})+|w_{n}^{2}(f_{n}^{1}-f_{n}^{2})|\right)$
-$=\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}(f_{i}^{1}+f_{i}^{2})+w_{n}^{2}(f_{n}^{1}-f_{n}^{2})\right)$
-$=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}f_{i}+w_{n}^{2}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}f_{i}-w_{n}^{2}f_{n}\right)$
-$=\mathbb{E}_{\sigma_{n}\sim\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w_{i}^{2}f_{i}\right].$
-The proof is completed by taking expectation over $\{\sigma_{i}\}_{i=1}^{n-1}$. ∎
-By Lemma 22 with $c=\frac{1}{\kappa}$, we have that
-$\mathrm{Pr}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\left|\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right)-F(n,w)\right|\geq\varepsilon\right]$
-$\leq$ $2\exp\left(-\frac{2\kappa^{2}\varepsilon^{2}}{\sum_{i=1}^{k}n_{i}}\right).$ (103)
-By (101) and (103), for any $\delta^{\prime}\in(0,1]$, with probability $1-\delta^{\prime}$, it holds that
-$\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right)$
-$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{d\log(2\sum_{i=1}^{k}n_{i})+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (104)
-By replacing $(\ell,L)$ with $(-\ell,-L)$, using similar arguments we obtain that, with probability $1-2\delta^{\prime}$,
-$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right|$
-$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{d\log(2\sum_{i=1}^{k}n_{i})+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (105)
-Now we fix $\kappa\geq 0$. Define that
-~=$\tilde{\mathcal{L}}=\left\{n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\in\Delta_{\varepsilon_{1}/(8k)}(k)\mid T_{1}w_{i}\leq 2n_{i},12\log(2k)\leq n_{i}\leq T_{1},\forall i\in[k],\sum_{i=1}^{k}n_{i}\leq 2T_{1}\right\}.$
-By (105), for any $\delta^{\prime}>0$, with probability $1-\delta^{\prime}$, for any $\{n,w\}\in\tilde{\mathcal{L}}$, we have that
-$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right|$
-$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{T_{1}}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{T_{1}}\sqrt{\frac{d\log(2\sum_{i=1}^{k}n_{i})+\log(|\tilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}$
-$\leq 4\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{d\log(4T_{1})+\log(2|\tilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{T_{1}}}$
-$\leq 4\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{d\log(4T_{1})+2k\log(16kT_{1}/\epsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}$
-$\leq 600C_{T_{1}}+4\sqrt{\frac{d\log(4T_{1})+2k\log(16kT_{1}/\epsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}.$
-Here the last inequality is by Assumption 10, Lemma 12, Lemma 13 and the fact that $\sum_{i=1}^{k}n_{i}\geq\frac{T_{1}}{2}$.
-Note that in Algorithm 3,
-$\hat{L}^{t}(h,w^{t})=\sum_{i=1}^{k}\frac{w_{i}^{t}}{\check{n}_{i}^{t}}\cdot\sum_{j=1}^{\check{n}_{i}^{t}}\ell(h,(x_{i,j},y_{i,j})).$
-By the definition that $\check{n}_{i}^{t}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ for $i\in[k]$, we have that $T_{1}w_{i}^{t}\leq\check{n}_{i}^{t}-1$ and $12\log(2k)\leq\check{n}_{i}^{t}\leq T_{1}$ for all $i\in[k]$.
-On the other hand, $\sum_{1=1}^{k}\check{n}_{i}^{t}\leq T_{1}+k+12k\log(2k)\leq 2T_{1}-2$.
-Therefore, there exists some $\tilde{w}^{t}\in\Delta(k)$ such that $\{\{\check{n}_{i}^{t}\}_{i=1}^{k},\tilde{w}^{t}\}\in\tilde{\mathcal{L}}$ and $\|w^{t}-\tilde{w}^{t}\|_{1}\leq\frac{\varepsilon_{1}}{8k}$ for each $1\leq t\leq T$.
-Choose $\delta^{\prime}=\delta/4$.
-We then obtain that: with probability $41-\delta/$ 4, it holds that
-$\max_{h\in\mathcal{H}}|\hat{L}^{t}(h,w^{t})-L(h,w^{t})|$
-$\leq 600C_{T_{1}}+4\sqrt{\frac{d\log(4T_{1})+2k\log(16kT_{1}/\epsilon_{1})+\log(8/\delta)}{T_{1}}}$
-$\leq\frac{\varepsilon_{1}}{2}.$ (106)
-for any $1\leq t\leq T$. Here the last inequality is by definition of $T_{1}$.
-Finally, by noting that $h^{t}=\arg\min_{h\in\mathcal{H}}\hat{L}^{t}(h,w^{t})$, we have that
-$L(h^{t},w^{t})\leq\hat{L}^{t}(h^{t},w^{t})+\frac{\varepsilon_{1}}{2}=\min_{h\in\mathcal{H}}\hat{L}^{t}(h,w^{t})+\frac{\varepsilon_{1}}{2}\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}.$
-The proof is completed ∎
-D.2 Restatement and Proof of Lemma 12
-Lemma 12.[Restatement] For any two groups of positive integers $\{n_{i}\}_{i=1}^{k}$ and $\{m_{i}\}_{i=1}^{k}$, it holds that
-$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$ $\leq\left(\sum_{i=1}^{k}(m_{i}+n_{i})\right)\widetilde{\mathsf{Rad}}_{\{m_{i}+n_{i}\}_{i=1}^{k}}$
-$\leq\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$ (107)
-Proof.
+$F(n,w)$ $\coloneqq\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right)\right],$
+which can be upper bounded by
+$F(n,w)$ $=\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\left(\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}^{+}(n)}\left[\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\right]\right)\right)\right]$
+$\leq\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n),\widetilde{\mathcal{S}}^{+}(n)}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\left(\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\right)\right)\right]$
+$=\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n),\widetilde{\mathcal{S}}^{+}(n),\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\left\{\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\right\}\right)\right]$
+$\leq 2\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n),\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{1}{\kappa}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\right)\right]$
+$=2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{\mathsf{Rad}}}_{\{n_{i}\}_{i=1}^{k}}.$ (137)
+Here, the first inequality arises from Jensen’s inequality, whereas the penultimate line applies Lemma 12.
+Invoking the Mcdiarmid inequality (see Lemma 9) with the choice $c=1/\kappa$, we obtain
+$\mathbb{P}\left\{\left|\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right)-F(n,w)\right|\geq\varepsilon\right\}\leq 2\exp\left(-\frac{2\kappa^{2}\varepsilon^{2}}{\sum_{i=1}^{k}n_{i}}\right).$ (138)
+This taken together with (137) reveals that: for any $\delta^{\prime}\in(0,1]$, with probability at least $1-\delta^{\prime}$ we have
+$\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right)\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (139)
+Evidently, this inequality continues to hold if we replace $(\ell,L)$ with $(-\ell,-L)$. As a consequence, with probability at least $1-2\delta^{\prime}$ one has
+$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right|\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (140)
+Now, fix $\kappa\geq 0$, and define
+~=$\widetilde{\mathcal{L}}=\left\{n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\in\Delta_{\varepsilon_{1}/(8k)}(k)\mid T_{1}w_{i}\leq 2n_{i},12\log(2k)\leq n_{i}\leq T_{1},\forall i\in[k],\sum_{i=1}^{k}n_{i}\leq 2T_{1}\right\},$
+where $\Delta_{\varepsilon_{1}/(8k)}(k)$ (i.e., an $\varepsilon_{1}/(8k)$-net of $\Delta(k)$) has been defined in Appendix B.1.
+Inequality (140) combined with the union bound tells us that: for any $\delta^{\prime}>0$, with probability at least $1-\delta^{\prime}$
+$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\frac{w_{i}}{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right|$
+$\qquad\leq\frac{\sum_{i=1}^{k}n_{i}}{T_{1}/2}\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{T_{1}/2}\sqrt{\frac{\log(|\widetilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}$
+$\qquad\leq 4\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{\log(2|\widetilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{T_{1}}}$
+$\qquad\leq 4\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{2k\log(16kT_{1}/\varepsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}$
+$\qquad\leq 600C_{T_{1}}+4\sqrt{\frac{2k\log(16kT_{1}/\varepsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}$
+holds simultaneously for all $\{n,w\}\in\widetilde{\mathcal{L}}$; see the use of the union bound in Appendix B.1 too.
+Here, we have made use of Assumption 1, Lemma 4, Lemma 5 and the fact that $\sum_{i=1}^{k}n_{i}\geq T_{1}/2$.
+Note that in Algorithm 3, we take
+$\widehat{L}^{t}(h,w^{t})=\sum_{i=1}^{k}\frac{w_{i}^{t}}{n_{i}^{t,\mathsf{rad}}}\cdot\sum_{j=1}^{n_{i}^{t,\mathsf{rad}}}\ell(h,(x_{i,j},y_{i,j})).$
+Given our choice that $n_{i}^{t,\mathsf{rad}}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ for $i\in[k]$, we see that
+$T_{1}w_{i}^{t}\leq n_{i}^{t,\mathsf{rad}}-1\qquad\text{and}\qquad 12\log(2k)\leq n_{i}^{t,\mathsf{rad}}\leq T_{1}$
+for all $i\in[k]$. In addition, it is seen from our choice of $n_{i}^{t,\mathsf{rad}}$ and $T_{1}$ that
+$\sum_{i=1}^{k}n_{i}^{t,\mathsf{rad}}\leq\sum_{i=1}^{k}\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil\leq T_{1}+k+12k\log(2k)\leq 2T_{1}-2.$
+Therefore, there exists some $\widetilde{w}^{t}\in\Delta(k)$ satisfying
+$\big{\{}\{n_{i}^{t,\mathsf{rad}}\}_{i=1}^{k},\widetilde{w}^{t}\big{\}}\in\widetilde{\mathcal{L}}\qquad\text{and}\qquad\|w^{t}-\widetilde{w}^{t}\|_{1}\leq\frac{\varepsilon_{1}}{8k}$
+for each $1\leq t\leq T$. Taking $\delta^{\prime}=\delta/4$, we obtain that with probability at least $41-\delta/$ 4,
+$\max_{h\in\mathcal{H}}\big{|}\widehat{L}^{t}(h,w^{t})-L(h,w^{t})\big{|}$ $\leq\max_{h\in\mathcal{H}}\big{|}\widehat{L}^{t}(h,\widetilde{w}^{t})-L(h,\widetilde{w}^{t})\big{|}+\max_{h\in\mathcal{H}}\big{|}\widehat{L}^{t}(h,\widetilde{w}^{t})-\widehat{L}^{t}(h,w^{t})\big{|}$
+$\qquad\qquad+\max_{h\in\mathcal{H}}\big{|}L(h,\widetilde{w}^{t})-L(h,w^{t})\big{|}$
+$\leq 600C_{T_{1}}+4\sqrt{\frac{2k\log(16kT_{1}/\varepsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}+\frac{\varepsilon_{1}}{8k}+\frac{\varepsilon_{1}}{8k}$
+$\leq\frac{\varepsilon_{1}}{2}$
+for any $1\leq t\leq T$, where the last inequality results from the definition of $T_{1}$.
+Finally, the fact that $h^{t}=\arg\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})$ allows one to derive
+$L(h^{t},w^{t})\leq\widehat{L}^{t}(h^{t},w^{t})+\frac{\varepsilon_{1}}{2}=\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})+\frac{\varepsilon_{1}}{2}\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1},$
+which concludes the proof.
+E.2 Proof of Lemma 4
 In what follows, assume that each $z_{i}^{j}$ obeys $z_{i}^{j}\sim\mathcal{D}_{i}$, and each $\sigma_{i}^{j}$ is a zero-mean Rademacher random variable. Direct computation then gives
-$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$
-$=\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
-=$\overset{\mathrm{(i)}}{=}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]\right]$
-$\overset{\mathrm{(ii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
-$=\left(\sum_{i=1}^{k}(n_{i}+m_{i})\right)\widetilde{\mathsf{Rad}}_{\{n_{i}+m_{i}\}_{i=1}^{k}}$
-$\overset{\mathrm{(iii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
-$\qquad\qquad\qquad\qquad+\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=n_{i}+1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
-$=\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$
-Here, (i) is valid due to the zero-mean property of $\{\sigma_{i}^{j}\}$, (ii) comes from Jensen’s inequality, and (iii) holds since $\max_{x}(f_{1}(x)+f_{2}(x))\leq\max_{x}f_{1}(x)+\max_{x}f_{2}(x)$.
-∎
-D.3 Restatement and Proof of Lemma 13
-Lemma 13.[Restatement] Consider any $\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for each $i\in[k]$. By taking $w\in\Delta^{k}$ with $w_{i}=\frac{n_{i}}{\sum_{l=1}^{k}n_{l}}$, one has
-$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}\leq 72\mathsf{Rad}_{\sum_{i=1}^{k}n_{i}}\big{(}\mathcal{D}(w)\big{)}.$
-Proof.
-Set $\widetilde{n}=\sum_{i=1}^{k}n_{i}$.
-Let $\{X_{j}\}_{j=1}^{n}$ be $n$ i.i.d.
-multinomial random variables with parameter $\{w_{i}\}_{i=1}^{k}$, and take $\widehat{n}_{i}=\sum_{j=1}^{n}\mathds{1}\{X_{j}=i\}$ for each $i\in[k]$.
-From (44) and Definition 9, it is easily seen that
-$\mathsf{Rad}_{\widetilde{n}}\big{(}D(w)\big{)}$ $=\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{\widetilde{n}}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{\widetilde{n}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right]\right],$
+$\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}=\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
+=$\qquad\overset{\mathrm{(i)}}{=}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]\right]$
+$\qquad\overset{\mathrm{(ii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
+$\qquad=\left(\sum_{i=1}^{k}(n_{i}+m_{i})\right)\widetilde{\mathsf{Rad}}_{\{n_{i}+m_{i}\}_{i=1}^{k}}$
+$\qquad\overset{\mathrm{(iii)}}{\leq}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
+$\qquad\qquad\qquad\qquad\qquad+\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=n_{i}+1}^{n_{i}+m_{i}},\forall i\in[k]}\left[\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=n_{i}+1}^{n_{i}+m_{i}}\sigma_{i}^{j}\ell(h,z_{i}^{j})\right]\right]$
+$\qquad=\left(\sum_{i=1}^{k}n_{i}\right)\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\left(\sum_{i=1}^{k}m_{i}\right)\widetilde{\mathsf{Rad}}_{\{m_{i}\}_{i=1}^{k}}.$
+Here, (i) is valid due to the zero-mean property of $\{\sigma_{i}^{j}\}$, (ii) comes from Jensen’s inequality, and (iii) follows since $\max_{x}\big{(}f_{1}(x)+f_{2}(x)\big{)}\leq\max_{x}f_{1}(x)+\max_{x}f_{2}(x)$.
+E.3 Proof of Lemma 5
+Set $n=\sum_{i=1}^{k}n_{i}$. Let $\{X_{j}\}_{j=1}^{n}$ be $n$ i.i.d. multinomial random variables with parameter $\{w_{i}\}_{i=1}^{k}$, and take
+$\widehat{n}_{i}=\sum_{j=1}^{n}\mathds{1}\{X_{j}=i\}$
+for each $i\in[k]$. From (34) and Definition 2, it is easily seen that
+$\mathsf{Rad}_{n}\big{(}D(w)\big{)}$ $=\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{n}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{n}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right]\right],$
 where each $z_{i}^{j}$ is independently drawn from $\mathcal{D}_{i}$, and each $\sigma_{i}^{j}$ is an independent Rademacher random variable.
@@
 $\widehat{n}_{i}\geq\frac{1}{3}n_{i}-2\log(2k)\geq\frac{1}{6}n_{i}\qquad\Longrightarrow\qquad\widehat{n}_{i}\geq\left\lceil\frac{1}{6}n_{i}\right\rceil\eqqcolon\widetilde{n}_{i}$ (141)
-with probability exceeding $1-\frac{1}{2k}$. By defining $\mathcal{E}$ to be the event where $\widehat{n}_{i}\geq n_{i}/6$ for all $i\in[k]$, we can immediately see from the union bound that
+with probability exceeding $1-1/(2k)$. Defining $\mathcal{E}$ to be the event that $\widehat{n}_{i}\geq n_{i}/6$ holds for all $i\in[k]$, we can invoke the union bound to see that
 $\mathbb{P}(\mathcal{E})\geq 1/2.$
 Consequently, we can derive
-$\mathsf{Rad}_{\widetilde{n}}\big{(}\mathcal{D}(w)\big{)}$ $\geq\mathbb{P}(\mathcal{E})\cdot\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{\widetilde{n}}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{\widetilde{n}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\mid\mathcal{E}\right]\right]\right]$
-$\geq\frac{1}{2}\cdot\frac{\sum_{i=1}^{k}\widetilde{n}_{i}}{\widetilde{n}}\widetilde{\mathsf{Rad}}_{\{\widetilde{n}_{i}\}_{i=1}^{k}}$
+$\mathsf{Rad}_{n}\big{(}\mathcal{D}(w)\big{)}$ $\geq\mathbb{P}(\mathcal{E})\cdot\mathop{\mathbb{E}}\limits_{\{X_{i}\}_{i=1}^{n}}\left[\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{n}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widehat{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\mid\mathcal{E}\right]\right]\right]$
+$\geq\frac{1}{2}\cdot\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{\widehat{n}_{i}},\forall i\in[k]}\left[\frac{1}{n}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{\widetilde{n}_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\mid\mathcal{E}\right]\right]$
+$=\frac{1}{2}\cdot\frac{\sum_{i=1}^{k}\widetilde{n}_{i}}{n}\widetilde{\mathsf{Rad}}_{\{\widetilde{n}_{i}\}_{i=1}^{k}}$
 $\geq\frac{1}{12}\cdot\frac{1}{6}\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}},$ (142)
-where the last two inequalities both follow from Lemma 39. This concludes the proof. ∎
-D.4 Necessity of Assumption 10
-In this section, we will discuss the necessity of Assumption 10 in comparison to the following weaker assumption, where we only assume that the Rademacher complexity on each $\mathcal{D}_{i}$ is well bounded.
+thus concluding the proof.
+E.4 Necessity of Assumption 1
+In this subsection, we study whether Assumption 1 can be replaced by the following weaker assumption, the latter of which only assumes that the Rademacher complexity on each $\mathcal{D}_{i}$ is well-bounded.
 Assumption 2.
-For each $n\geq 1$, there exists a universal constant $C_{n}\geq 0$ such that
-$\mathsf{Rad}_{n}(\mathcal{D}_{i})\leq C_{n}$ (110)
-for all $i\in[k]$.
+For each $n\geq 1$, there exists a quantity $\widetilde{C}_{n}>0$ (known to the learner) such that
+$\widetilde{C}_{n}\geq\max_{1\leq i\leq k}\mathsf{Rad}_{n}(\mathcal{D}_{i}).$ (143)
 Formally, we have the following results.
 Lemma 21.
-Let $w^{0}=[1/k,1/k,\ldots,1/k]^{\top}$. There exists a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that
-$\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))\geq\Omega\left(\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})\right)$ (111)
+Let $w^{0}=[1/k,1/k,\ldots,1/k]^{\top}$. There exist a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that
+$\mathsf{Rad}_{n}\big{(}\mathcal{D}(w^{0})\big{)}\geq\Omega\left(\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})\right)$ (144)
 for $n\geq 12k\log(k)$.
 Proof.
-Without loss of generality, we set $\mathcal{Y}=\{0\}$ and $\ell(h,(x,y))=h(x)-y=h(x)$.
-We can then regard $\mathcal{D}_{i}$ as a distribution over $\mathcal{X}_{i}$ because there is only one element in $\mathcal{Y}$.
-We pick $k$ subsets of $\mathcal{X}$ as $\{\mathcal{X}_{i}\}_{i=1}^{k}$.
-For each $i\in[k]$, we choose the distribution $\mathcal{D}_{i}$ to be an arbitrary distribution supported by $\mathcal{X}_{i}$.
-On the other hand, we also define $\mathcal{H}_{i}$ to be a set of hypothesis $h(x)=0$ for all $x\notin\mathcal{X}_{i}$ for each $i\in[k]$.
-For $\{h_{i}\}_{i=1}^{k}$ such that $h_{i}\in\mathcal{H}_{i}$, we define $\mathsf{joint}(\{h_{i}\}_{i=1}^{k})$ to be the hypothesis $h$ such that $h(x)=h_{i}(x),\forall x\in\mathcal{X}_{i},i\in[k]$ and $h(x)=0$ for $x\notin\cup_{i}\mathcal{X}_{i}$.
-Now we construct the hypothesis set $\mathcal{H}$ as
-=$\mathcal{H}=\{\mathsf{joint}(\{h_{i}\}_{i=1}^{k})|h_{i}\in\mathcal{H}_{i},\forall i\in[k]\}.$ (112)
-Recall the definition of $\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\{\mathcal{D}_{i}\}_{i=1}^{k})$. Then we have that
-$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\{\mathcal{D}_{i}\}_{i=1}^{k})$ $=\mathop{\mathbb{E}}\limits_{\{x_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h(x_{i}^{j})\right]\right]$
-$=\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}\max_{h_{i}\in\mathcal{H}_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h_{i}(x_{i}^{j})\right]\right]$ (113)
+Without loss of generality, consider the case where $\mathcal{Y}=\{0\}$ and $\ell(h,(x,y))=h(x)-y=h(x)$.
+We can then view $\mathcal{D}_{i}$ as a distribution over $\mathcal{X}_{i}$, as there is only one element in $\mathcal{Y}$.
+Pick $k$ subsets of $\mathcal{X}$ as $\{\mathcal{X}_{i}\}_{i=1}^{k}$.
+For each $i\in[k]$, we choose the distribution $\mathcal{D}_{i}$ to be an arbitrary distribution supported on $\mathcal{X}_{i}$.
+In addition, we define $\mathcal{H}_{i}$ to be a set of hypothesis obeying $h(x)=0$ for all $x\notin\mathcal{X}_{i}$ for each $i\in[k]$.
+For a collection of hypothesis $\{h_{i}\}_{i=1}^{k}$ such that $h_{i}\in\mathcal{H}_{i}$, we define $\mathsf{joint}(\{h_{i}\}_{i=1}^{k})$ to be the hypothesis $h$ such that
+$h(x)=\begin{cases}h_{i}(x)&\text{if }x\in\mathcal{X}_{i},~{}i\in[k];\\ 0&\text{if }x\notin\cup_{i}\mathcal{X}_{i}.\end{cases}$ (145)
+The hypothesis set $\mathcal{H}$ is then constructed as
+=$\mathcal{H}=\big{\{}\mathsf{joint}(\{h_{i}\}_{i=1}^{k})\mid h_{i}\in\mathcal{H}_{i},\forall i\in[k]\big{\}}.$ (146)
+Recalling the definition of $\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$, we see that
+$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}$ $=\mathop{\mathbb{E}}\limits_{\{x_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h(x_{i}^{j})\right]\right]$
+$=\mathop{\mathbb{E}}\limits_{\{x_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}\max_{h_{i}\in\mathcal{H}_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h_{i}(x_{i}^{j})\right]\right]$ (147)
 $=\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}n_{i}\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}}\left[\frac{1}{n_{i}}\max_{h_{i}\in\mathcal{H}_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}h_{i}(x_{i}^{j})\right]\right]$
-$=\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}n_{i}\mathsf{Rad}_{n_{i}}(\mathcal{D}_{i}).$
-Here (113) is by the definition of $\mathcal{H}$.
-By taking $n_{i}=\frac{n}{k}$ for all $i\in[k]$ and Lemma 13, we have
-$\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})=\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\{\mathcal{D}_{i}\}_{i=1}^{k})\leq 72\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))$ (114)
-By Lemma 32, if we set $C_{n}\sim\sqrt{\frac{d}{n}}$ in Assumption 31, the best upper bound for $\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))\sim\sqrt{\frac{dk}{n}}$, which implies that more samples are needed to learn the mixture distribution $w^{0}$.
-Moreover, under the construction in Lemma 32, by further assuming $\min_{h_{i}\in\mathcal{H}_{i}}\mathbb{E}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]=\frac{1}{2}$ for all $i\in[k]$, to find $h$ such that
-$\max_{i}\mathbb{E}_{x\sim\mathcal{D}_{i}}[h(x)]\leq\frac{1}{2}+\epsilon,$ (115)
-we need to find $h_{i}\in\mathcal{H}_{i}$ such that
-$\mathbb{E}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]\leq\frac{1}{2}+\epsilon$ (116)
-for all $i\in[k]$.
-Following this intuition, we can construct a counter example under Assumption 31 as below.
+$=\frac{1}{\sum_{i=1}^{k}n_{i}}\sum_{i=1}^{k}n_{i}\mathsf{Rad}_{n_{i}}(\mathcal{D}_{i}),$
+where (147) results from the definition of $\mathcal{H}$. By taking $n_{i}=\frac{n}{k}$ for all $i\in[k]$ and applying Lemma 5, we reach
+$\frac{1}{k}\sum_{i=1}^{k}\mathsf{Rad}_{n/k}(\mathcal{D}_{i})=\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}\leq 72\mathsf{Rad}_{n}\big{(}\mathcal{D}(w^{0})\big{)}$ (148)
+as claimed. ∎
+By virtue of Lemma 21, if we set $\widetilde{C}_{n}=\widetilde{\Theta}(\sqrt{d/n})$ in Assumption 2, then the best possible upper bound on $\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))$ is $\mathsf{Rad}_{n}(\mathcal{D}(w^{0}))=\widetilde{\Theta}(\sqrt{dk/n})$, which implies that more samples are needed to learn the mixed distribution $\mathcal{D}(w^{0})$ than learning each individual distribution.
+Moreover, under the construction in Lemma 21, if we further assume that $\min_{h_{i}\in\mathcal{H}_{i}}\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]=1/2$ for all $i\in[k]$, then to find $h$ such that
+$\max_{1\leq i\leq k}\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{i}}[h(x)]\leq\frac{1}{2}+\varepsilon,$ (149)
+we need to find, for each $i\in[k]$, a hypothesis $h_{i}\in\mathcal{H}_{i}$ such that
+$\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{i}}[h_{i}(x)]\leq\frac{1}{2}+\varepsilon.$ (150)
+Following this intuition, we can construct a counter example under Assumption 2, with a formal theorem stated as follows.
 Theorem 4.
-There exists a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that Assumption 31 holds with $C_{n}=O\left(\sqrt{\frac{d}{n}}\right)$, and it costs at least $\tilde{\Omega}\left(\frac{dk}{\epsilon^{2}}\right)$ samples to find some $h\in\mathcal{H}$ such that
-$\max_{i\in[k]}L(h,\mathbf{e}_{i})\leq\min_{h^{\prime}\in\mathcal{H}}\max_{i\in[k]}L(h^{\prime},\mathbf{e}_{i})+\epsilon.$
+There exist a group of distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$ and a hypothesis set $\mathcal{H}$ such that Assumption 2 holds with $\widetilde{C}_{n}=\widetilde{O}\left(\sqrt{\frac{d}{n}}\right)$, and it takes at least $\widetilde{\Omega}\left(\frac{dk}{\varepsilon^{2}}\right)$ samples to find some $h\in\mathcal{H}$ obeying
+$\max_{i\in[k]}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\leq\min_{h^{\prime}\in\mathcal{H}}\max_{i\in[k]}L(h^{\prime},\mathrm{e}^{\mathsf{basis}}_{i})+\varepsilon.$
 Proof.
-With the construction in Lemma 32, it suffices to find some $\mathcal{H}^{\prime}$ and $\mathcal{D}^{\prime}$ such that the following three conditions holds.
+With the construction in Lemma 21, it suffices to find some $\mathcal{H}^{\prime}$ and $\mathcal{D}^{\prime}$ such that the following three conditions hold:
   1. 1.
-    $\mathsf{Rad}_{n}(\mathcal{D}^{\prime},\mathcal{H}^{\prime}):=\frac{1}{n}\mathbb{E}_{\{x^{j}\}_{j=1}^{n}\overset{\mathrm{iid}}{\sim}\mathcal{D}^{\prime},\{\sigma^{j}\}_{j=1}^{n}\overset{\mathrm{iid}}{\sim}\{\pm 1\}}\left[\max_{h^{\prime}\in\mathcal{H}^{\prime}}\sum_{j=1}^{n}\sigma^{j}h^{\prime}(x^{j})\right]\leq C_{n};$ (117)
+    The following inequality holds:
+    $\mathsf{Rad}_{n}(\mathcal{D}^{\prime},\mathcal{H}^{\prime})\coloneqq\frac{1}{n}\mathop{\mathbb{E}}_{\{x^{j}\}_{j=1}^{n}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{D}^{\prime},\,\{\sigma^{j}\}_{j=1}^{n}\overset{\mathrm{i.i.d.}}{\sim}\{\pm 1\}}\left[\max_{h^{\prime}\in\mathcal{H}^{\prime}}\sum_{j=1}^{n}\sigma^{j}h^{\prime}(x^{j})\right]\leq\widetilde{C}_{n};$ (151)
   2. 2.
@@
   3. 3.
-    It cost at least $\tilde{\Omega}\left(\frac{d}{\epsilon^{2}}\right)$ samples to find some $h$ such that $\mathbb{E}_{x\sim\mathcal{D}^{\prime}}[h(x)]\leq\frac{1}{2}+\epsilon$.
-This construction is also straightfoward.
-We set $N=2^{d}$ and $\mathcal{X}^{\prime}=\{0,1\}^{N}$.
-Let $\mathcal{D}^{\prime}$ to be the distribution $\mathrm{Pr}_{\mathcal{D}^{\prime}}[x]=\Pi_{n=1}^{N}\mathrm{Pr}_{\mathcal{D}^{\prime}_{n}}[x_{n}]$, where $\mathrm{Pr}_{\mathcal{D}^{\prime}_{n^{*}}}[x_{n^{*}}]=\frac{1}{2}\mathbb{I}[x_{n^{*}}=1]+\frac{1}{2}\mathbb{I}[x_{n^{*}}=0]$ for some $n^{*}$ , $\mathrm{Pr}_{\mathcal{D}^{\prime}_{n}}[x_{n}]=(\frac{1}{2}+2\epsilon)\mathbb{I}[x_{n^{*}}=1]+(\frac{1}{2}-2\epsilon)\mathbb{I}[x_{n^{*}}=0]$ for all $n\neq n^{*}$.
-Then we set =$\mathcal{H}^{\prime}=\{h^{n}\}_{n=1}^{N}$ with $h^{n}(x)=x_{n}$ for each $n\in[N]$.
-It is then easy to verify that the first two condition holds.
-As for the third condition, ollowing the arguments in Theorem 18, we need at least $\tilde{\Omega}\left(\frac{d}{\epsilon^{2}}\right)$ i.i.d.
+    It takes at least $\widetilde{\Omega}\left(\frac{d}{\varepsilon^{2}}\right)$ samples to find some $h$ such that $\mathbb{E}_{x\sim\mathcal{D}^{\prime}}[h(x)]\leq\frac{1}{2}+\varepsilon$.
+This construction is also straightforward. Set $N=2^{d}$ and $\mathcal{X}^{\prime}=\{0,1\}^{N}$. Let $\mathcal{D}^{\prime}$ be the distribution
+$\mathbb{P}_{\mathcal{D}^{\prime}}\{x\}=\prod_{n=1}^{N}\mathbb{P}_{\mathcal{D}^{\prime}_{n}}\{x_{n}\}$
+where
+$\mathbb{P}_{\mathcal{D}^{\prime}_{n^{*}}}\{x_{n^{*}}\}$ $=\frac{1}{2}\mathds{1}\{x_{n^{*}}=1\}+\frac{1}{2}\mathds{1}\{x_{n^{*}}=0\}$ for some $\text{for some }n^{*};$ (152)
+$\mathbb{P}_{\mathcal{D}^{\prime}_{n}}\{x_{n}\}$ $=\left(\frac{1}{2}+2\varepsilon\right)\mathds{1}\{x_{n^{*}}=1\}+\left(\frac{1}{2}-2\varepsilon\right)\mathds{1}\{x_{n^{*}}=0\}$ for all $\text{for all }n\neq n^{*}.$ (153)
+We then choose =$\mathcal{H}^{\prime}=\{h^{n}\}_{n=1}^{N}$ with $h^{n}(x)=x_{n}$ for each $n\in[N]$.
+It is then easy to verify that the first two conditions hold.
+Regarding the third condition, following the arguments in Theorem 2, we need at least $\widetilde{\Omega}\left(d/{\varepsilon^{2}}\right)$ i.i.d.
 samples from $\mathcal{D}^{\prime}$ to identify $n^{*}$.
-The proof is finished.
-Appendix E Proof of Theorem 17
-1 Input: The hypothesis set $\mathcal{H}$, the group set $\mathcal{G}$, target accuracy level $\epsilon$, target success rate $1-\delta$, minimal probability $\gamma$.
-2 Initialization: $T=\frac{10000\log(|\mathcal{G}|/\delta)}{\gamma\varepsilon^{2}}$, $\eta=\frac{1}{10}\varepsilon\gamma$, $N=\frac{10000(\log(|\mathcal{G}|/\delta)+d\log(d/\varepsilon))}{\gamma\varepsilon^{2}}$, $W^{1}_{g}=1$ for all $g\in\mathcal{G}$.
-3 Call $\mathrm{Query}(\mathcal{D})$ for $N$ times to get $N$ i.i.d. samples from $\mathcal{D}$, denoted by $\{(x_{i},y_{i})\}_{i=1}^{N}$.
-4 $f_{g}\leftarrow\min_{h\in\mathcal{L}}\frac{1}{\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i}),\forall g\in\mathcal{G}$,
-5 $N_{g}\leftarrow\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g],$ $p_{g}\leftarrow\frac{N_{g}}{N}$, $\forall g\in\mathcal{G}$
-6 for $t=1,2,\ldots,T$ do
-7       $w_{g}^{t}\leftarrow\frac{W_{g}^{t}}{\sum_{g\in\mathcal{G}}W_{g}^{t}},\forall g\in\mathcal{G}$,
-8       $h^{t}\leftarrow\arg\min_{h\in\mathcal{H}}\left(\sum_{g\in\mathcal{G}}w_{g}^{t}\cdot\left(\frac{1}{N_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-f_{g}\right)\right)$,
-9       Call $\mathrm{Query}(\mathcal{D})$ to get a data point $(\hat{x}_{t},\hat{y}_{t})$,
-10       $\hat{r}^{t}_{g}\leftarrow\frac{1}{p_{g}}\mathbb{I}[\hat{x}_{t}\in g]\ell(h^{t}(\hat{x}_{t}),\hat{y}_{t})-f_{g}$, $\forall g\in\mathcal{G}$,
-11       $W_{g}^{t+1}\leftarrow W_{g}^{t}\cdot\exp(\eta\hat{r}^{t}_{g})$, $\forall g\in\mathcal{G}$;
-Return: A randomized hypothesis $h^{\mathsf{final}}$ as a uniform distribution over $\{h^{t}\}_{t=1}^{T}$.
-Algorithm 4 Hedge for Multi-group Learning
-Throughout this section, we let $\mathcal{F}_{t}$ be the event field before the $t$-th round.
-Clearly, the number of samples used in Algorithm 4 is bounded by $T+N=O\left(\frac{(\log(|\mathcal{G}|/\delta))+d\log(d/\epsilon)}{\gamma\epsilon^{2}}\right)$.
-So it suffices to prove the optimality of $h^{\mathsf{final}}$.
-Formally, we have the lemma below.
-Lemma 34.
-With probability $1-\delta$, it holds that
-$\frac{1}{T}\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)\leq\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)+\epsilon$ (118)
-for any $g\in\mathcal{G}$.
-Proof.
-Let $W^{t}=\sum_{g\in\mathcal{G}}W^{t}_{g}$. By definition and the fact that $\eta\hat{r}_{g}^{t}\leq\frac{\eta}{\gamma}\leq 1$ for any proper $(g,t)$ pair,
-$\log(W_{g}^{T+1})=\eta\sum_{t=1}^{T}\hat{r}_{g}^{t},$
-$\log\left(\frac{W^{t+1}}{W^{t}}\right)=\log\left(\sum_{g\in\mathcal{G}}w_{g}^{t}\exp(\eta\hat{r}_{g}^{t})\right)\leq\log\left(\sum_{g\in\mathcal{G}}w_{g}^{t}(1+\eta\hat{r}_{g}^{t}+\eta^{2}(\hat{r}_{g}^{t})^{2})\right)\leq\sum_{g\in\mathcal{G}}w_{g}^{t}(\eta\hat{r}_{g}^{t}+\eta^{2}(\hat{r}_{g}^{t})^{2}).$
-Noting that
-$\log(W^{T+1}_{g})\log(W^{T+1})=\sum_{t=1}^{T}\log(W^{t+1}/W^{t})+\log(|\mathcal{G}|),$
-for any $\tilde{g}\in\mathcal{G}$,
-$\eta\sum_{t=1}^{T}\hat{r}_{\tilde{g}}^{t}\leq\eta\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}+\eta^{2}\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}+\log(|\mathcal{G}|).$
-Dividing both side by $\eta$, we learn that
-$\sum_{t=1}^{T}\hat{r}_{\tilde{g}}^{t}-\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}\hat{r}_{g}^{t}\leq\frac{\log(|\mathcal{G}|)}{\eta}+\eta\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}.$ (119)
-By Lemma 36, with probability $1-\frac{\delta}{16}$,
-$\sum_{t=1}^{T}\hat{r}_{\tilde{g}}^{t}-\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}\leq\frac{\log(|\mathcal{G}|)}{\eta}+\frac{3\eta T}{\gamma}\leq\frac{T\epsilon}{32}$ (120)
-for any $\tilde{g}\in\mathcal{G}$. By (120) and Lemma 37, with probability $1-\frac{3\delta}{4}$,
-$\sum_{t=1}^{T}\left(L_{\mathcal{D}}(h^{t}|g)-f_{g}\right)\leq\sum_{t=1}^{T}\hat{r}_{g}^{t}+\frac{T\epsilon}{32}\leq\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}+\frac{5T\epsilon}{32}\leq T\epsilon$ (121)
-for any $g\in\mathcal{G}$. By Lemma 35, with probability $1-\delta$, we have that
-$\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)-T\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)\leq T\epsilon,$
-which means that
-$\frac{1}{T}\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)-\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)\leq\epsilon.$
-The proof is completed.
-Lemma 35.
-With probability $1-\frac{\delta}{16}$, it holds that
-$\left|f_{g}-\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)\right|\leq\frac{\epsilon}{8}$
-for any $g\in\mathcal{G}$.
-Proof.
-Fix $g\in\mathcal{G}$.
-Let $\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}$ to be the ghost samples, another group of $N$ i.i.d.
-datapoints from $\mathcal{D}$ which is independent of $\{(x_{i},y_{i})\}_{i=1}^{N}$.
-Let $\lambda\in(0,1/2]$ be a positive real number.
-Then we have that
-$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{N}}\left[\exp\left(\lambda\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-N\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathbb{I}[x\in g]\ell(h(x),y)]\right)\right)\right]$
-$\leq\mathbb{E}_{\{(x_{i},y_{i}),(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}}\left[\exp\left(\lambda\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{N}\left(\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-\mathbb{I}[x^{\prime}_{i}\in g]\ell(h(x^{\prime}_{i}),y^{\prime}_{i})\right)\right)\right)\right]$ (122)
-$=\mathbb{E}_{\{(x_{i},y_{i}),(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}}\mathbb{E}_{\{\sigma_{i}\}_{i=1}^{N}\overset{\mathrm{i.i.d.}}{\sim}\{\pm 1\}}\left[\exp\left(\lambda\max_{h\in\mathcal{H}}\sum_{i=1}^{N}\sigma_{i}\left(\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-\mathbb{I}[x^{\prime}_{i}\in g]\ell(h(x^{\prime}_{i}),y^{\prime}_{i})\right)\right)\right]$
-$\leq\mathbb{E}_{\{(x_{i},y_{i}),(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}}\left[\exp\left(\sum_{i=1}^{N}2\lambda^{2}(\mathbb{I}[x_{i}\in g]+\mathbb{I}[x^{\prime}_{i}\in g])^{2}\right)\right]$
-$\leq\mathbb{E}_{\{(x_{i},y_{i}),(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{N}}\left[\exp\left(\sum_{i=1}^{N}4\lambda^{2}(\mathbb{I}[x_{i}\in g]+\mathbb{I}[x^{\prime}_{i}\in g])\right)\right]$
-$\leq\exp\left(16N\lambda^{2}P_{g}\right).$ (123)
-Here (122) is by Jenson’s inequality.
-As a result, we obtain that
-$\mathrm{Pr}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-N\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathbb{I}[x\in g]\ell(h(x),y)]\right)>P_{g}\epsilon/32\right]$
-$\leq\min_{\lambda\in(0,1/2]}\exp\left(4N\lambda^{2}P_{g}-\lambda P_{g}\epsilon/32\right)$
-$\leq\exp\left(-\frac{P_{g}\epsilon^{2}}{8192N}\right)$ (124)
-$\leq\exp\left(-\frac{\gamma\epsilon^{2}}{8192N}\right)$
-$\leq\frac{\delta}{64|\mathcal{G}|}.$ (125)
-Using similar arguments, we can obtain that
-$\mathrm{Pr}\left[\max_{h\in\mathcal{H}}\left(-\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})+N\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathbb{I}[x\in g]\ell(h(x),y)]\right)>P_{g}\epsilon/32\right]\leq\frac{\delta}{64|\mathcal{G}|}$ (126)
-Therefore,
-$\mathrm{Pr}\left[\left|\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-L_{\mathcal{D}}(h|g)\right|\leq\epsilon/32\right]\leq\frac{\delta}{64|\mathcal{G}|}.$ (127)
-Recall that $p_{g}=\frac{\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]}{N}$. With Chernoff’s inequality, we have that
-$\mathrm{Pr}\left[\left|NP_{g}-Np_{g}\right|\geq 2\sqrt{NP_{g}\log(128|\mathcal{G}|/\delta)}\right]\leq\frac{\delta}{64|\mathcal{G}|},$ (128)
-which means that
-$\mathrm{Pr}\left[\left|\frac{1}{NP_{g}}-\frac{1}{Np_{g}}\right|\geq\frac{\epsilon}{32}\right]\leq\frac{\delta}{64|\mathcal{G}|}.$ (129)
-By (127) and (129), and taking a union bound over $\mathcal{G}$, we obtain that with probability $1-\frac{\delta}{16}$,
-$\left|\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-L_{\mathcal{D}}(h|g)\right|\leq\frac{\epsilon}{16}$ (130)
-for any $h\in\mathcal{H},g\in\mathcal{G}$. By the definition of $h_{g}$, we have that
-$\left|f_{g}-\min_{h\in\mathcal{H}}L_{\mathcal{D}}(h|g)\right|\leq 2\max_{h\in\mathcal{H}}\left|\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-L_{\mathcal{D}}(h|g)\right|\leq\frac{\epsilon}{8}.$
-The proof is completed. ∎
-Lemma 36.
-With probability $1-\frac{\delta}{16}$, it holds that
-$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}\leq\frac{3T}{\gamma}.$ (131)
-Proof.
-By definition, we have that
-$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}\leq\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\frac{\mathbb{I}[x_{t}\in g]}{p_{g}^{2}}.$ (132)
-Let $X_{t}=\sum_{g\in\mathcal{G}}w_{g}^{t}\frac{\mathbb{I}[x_{t}\in g]}{p^{2}_{g}}$ and recall $\mathcal{F}_{t}$ is the event field before the $t$-th round.
-Then $X_{t}\leq\frac{1}{\lambda^{2}}$, and $\mathbb{E}[X_{t}^{2}|\mathcal{F}_{t}]=\sum_{g,g^{\prime}}w_{g}^{t}w_{g^{\prime}}^{t}\frac{\mathbb{E}[\mathbb{I}[\hat{x}_{t}\in g,\hat{x}_{t}\in g^{\prime}]]}{p_{g}^{2}p_{g^{\prime}}^{2}}\leq\sum_{g,g^{\prime}}w_{g}^{t}w_{g^{\prime}}^{t}\frac{1}{(p_{g}p_{g^{\prime}})^{3/2}}\leq\frac{1}{\gamma^{3}}$.
-Using Freedman’s inequality (Lemma 20), with probability $1-\frac{\delta}{16}$,
-$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(\hat{r}_{g}^{t})^{2}\leq\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\frac{1}{p_{g}}+2\sqrt{\frac{T\log(16/\delta)}{\gamma^{3}}}\leq\frac{3T}{\gamma},$ (133)
-where the last inequality is by the fact that $T\geq\frac{\log(16/\delta)}{\gamma}$.
-Lemma 37.
-Assume the events in Lemma 35 hold. With probability $1-\frac{\delta}{4}$, it holds that
-$\left|\sum_{t=1}^{T}\hat{r}_{g}^{t}-\sum_{t=1}^{T}(L_{\mathcal{D}}(h^{t}|g)-f_{g})\right|\leq\frac{T\epsilon}{8},\forall g\in\mathcal{G};$ (134)
-$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}\leq\frac{13T\epsilon}{16}.$ (135)
-Proof.
-We start with the first inequality. By definition,
-$\sum_{t=1}^{T}\hat{r}_{g}^{t}-\sum_{t=1}^{T}(L_{\mathcal{D}}(h^{t}|g)-f_{g})$
-$=\sum_{t=1}^{T}\left(\frac{1}{p_{g}}\mathbb{I}[\hat{x}_{t}\in g]\ell(h(\hat{x}_{t}),\hat{y}_{t})-L_{\mathcal{D}}(h^{t}|g)\right).$ (136)
-Defining $\hat{X}_{t}=\frac{1}{p_{g}}\mathbb{I}[\hat{x}_{t}\in g]\ell(h(\hat{x}_{t}),\hat{y}_{t})$.
-We then have that $\mathbb{E}[\hat{X}_{t}|\mathcal{F}_{t}]=\frac{P_{g}}{p_{g}}L_{\mathcal{D}}(h^{t}|g)$ and $\mathbb{E}[\hat{X}_{t}^{2}|\mathcal{F}_{t}]\leq\frac{P_{g}}{p_{g}^{2}}$.
-With Freedman’s inequality (Lemma 20) , with probability $1-\frac{\delta}{64}$,
-$\left|\sum_{t=1}^{T}\hat{r}^{g}_{t}-\frac{P_{g}}{p_{g}}\sum_{t=1}^{T}L_{\mathcal{D}}(h^{t}|g)\right|\leq\frac{2}{p_{g}}\sqrt{TP_{g}\log(128|\mathcal{G}|/\delta)}\leq\frac{T\epsilon}{16}$ (137)
-for any $g\in\mathcal{G}$. Combining (137) with (129), we learn that: with probability $1-\frac{\delta}{32}$,
-$\left|\sum_{t=1}^{T}\hat{r}_{g}^{t}-\sum_{t=1}^{T}(L_{\mathcal{D}}(h^{t}|g)-f_{g})\right|\leq\frac{T\epsilon}{8},$ (138)
-holds for any $g\in\mathcal{G}$.
-As for the second inequality, we define $\tilde{X}_{t}=\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{p_{g}}\mathbb{I}[\hat{x}_{t}\in g]\ell(h^{t}(\hat{x}_{t}),\hat{y}_{t})\right)$.
-It then follows that $\mathbb{E}[\tilde{X}_{t}|\mathcal{F}_{t}]=\sum_{g\in\mathcal{G}}w_{g}^{t}L_{\mathcal{D}}(h^{t}|g)\leq\frac{1}{\gamma}$ and
-$\mathbb{E}[\tilde{X}_{t}^{2}|\mathcal{F}_{t}]\leq\sum_{g^{\prime},g^{\prime}}w_{g}^{t}w_{g^{\prime}}^{t}\frac{\mathbb{E}[\mathbb{I}[\hat{x}_{t}\in g,\hat{x}_{t}\in g^{\prime}]]}{p_{g}p_{g^{\prime}}}\leq\sum_{g^{\prime},g^{\prime}}w_{g}^{t}w_{g^{\prime}}^{t}\frac{1}{\sqrt{p_{g}p_{g^{\prime}}}}\leq\frac{1}{\gamma}.$
-Using Freedman’s inequality (Lemma 20), we learn that, with probability $1-\frac{\delta}{64}$
-$\sum_{t=1}^{T}\left(\tilde{X}_{t}-\sum_{g\in\mathcal{G}}w_{g}^{t}L_{\mathcal{D}}(h^{t}|g)\right)\leq 2\sqrt{\frac{T\log(128/\delta)}{\gamma}}+2\frac{\log(128/\delta)}{\gamma}\leq\frac{T\epsilon}{16}.$ (139)
-As a result,
-$\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}\hat{r}_{g}^{t}\leq\sum_{t=1}^{T}\sum_{g\in\mathcal{G}}w_{g}^{t}(L_{\mathcal{D}}(h^{t}|g)-f_{g})+\frac{T\epsilon}{16}.$ (140)
-Recall the definition that
-$h^{t}=\arg\min_{h\in\mathcal{H}}\left(\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{N_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-f_{g}\right)\right).$ (141)
-By Lemma 35,
-$\max_{h\in\mathcal{H}}\left|\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h(x_{i}),y_{i})-L_{\mathcal{D}}(h|g)\right|\leq\frac{\epsilon}{16}$ (142)
-for any $g\in\mathcal{G}$.
-By the definition that $Np_{g}=N_{g}$, we have
-$\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h^{t}(x_{i}),y_{i})-f_{g}\right)$ $\leq\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h^{*}(x_{i}),y_{i})-f_{g}\right)$
-$\leq\sum_{g\in\mathcal{G}}w_{g}^{t}(L_{\mathcal{D}}(h^{*}|g)-f_{g})+\frac{\epsilon}{16}.$ (143)
-By (143), Assumption 16 and Lemma 35, we have that
-$\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h^{t}(x_{i}),y_{i})-f_{g}\right)\leq\frac{1}{2}\epsilon.$ (144)
-On the other hand, we also have
-$\sum_{g\in\mathcal{G}}w_{g}^{t}\left(L_{\mathcal{D}}(h^{t}|g)-f_{g}\right)\leq\sum_{g\in\mathcal{G}}w_{g}^{t}\left(\frac{1}{Np_{g}}\sum_{i=1}^{N}\mathbb{I}[x_{i}\in g]\ell(h^{t}(x_{i}),y_{i})-f_{g}\right)+\frac{\epsilon}{16},$
-which implies that
-$\sum_{g\in\mathcal{G}}w_{g}^{t}\left(L_{\mathcal{D}}(h^{t}|g)-f_{g}\right)\leq\frac{13\epsilon}{16}.$ (145)
-The conclusion follows by (140) and (145).
-Appendix F Missing proofs in Section 8
-F.1 Restatement and Proof of Lemma 19
-Lemma 19.
-[Restatement] Fix $m\geq 0$ and $\widetilde{i}\in[k]$.
-Suppose $\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{i}\leq m]\geq\frac{1}{2}$.
-It then holds that for any $h\in\mathcal{H}_{\widetilde{i}}$
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m]\geq\frac{1}{2}\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m]\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}).$ (146)
-Moreover, it holds that $m\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}$.
-Proof.
-For $v\in\{-1,1\}^{m}$, we let $n^{+}(v)=|\{p:v_{p}=1\}|$ be the number of $1$’s in all coordiantes of $v$.
-Let $\mathcal{V}$ be a subset of $\{-1,1\}^{2m}$ defined as $\mathcal{V}:=\{v^{1},v^{2}\in\{-1,1\}^{m}|n^{+}(v^{1})-n^{+}(v^{2})\leq 4\sqrt{m}+2m\epsilon\}$.
-Let $\mathrm{Pr}_{\mathcal{C}}[\cdot]$ denote the probability distribution of $\left\{\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}\right\}$ and $\mathrm{Pr}_{\mathcal{C}^{\prime}}[\cdot]$ denote the probability distribution of $\left\{\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}\right\}$.
-Recall the definition of $\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}$ and $\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}$.
-By Hoeffding’s inequality, we have that $\mathrm{Pr}_{\mathcal{C}}[\mathcal{V}]\geq\frac{3}{4}.$ Also noting that the distribution of $\mathcal{V}$ is independent of the algorithm $\mathcal{G}^{\prime}$, we have that
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\left\{\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}\right\}\in\mathcal{V}\right]\geq\frac{3}{4}-(1-\frac{3}{4})=\frac{1}{2}.$ (147)
-That is,
-$\sum_{v=\{v^{1},v^{2}\}\in\mathcal{V}}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\geq\frac{1}{2}.$
-On the other hand, for any $v=\{v^{1},v^{2}\}\in\mathcal{V}$, we have
-$\mathrm{Pr}_{\mathcal{C}^{\prime}}[v]$ $=\mathrm{Pr}_{\mathcal{C}}[v]\cdot\left(1-8\epsilon\right)^{n^{+}(v^{1})-n^{+}(v^{2})}(1+8\epsilon)^{n^{+}(v^{2})-n^{+}(v^{1})}$
-$=\mathrm{Pr}_{\mathcal{C}}[v]\left(\frac{1-8\epsilon}{1+8\epsilon}\right)^{n^{+}(v^{1})-n^{+}(v^{2})}$
-$\geq\mathrm{Pr}_{\mathcal{C}}[v]\exp(-20(n^{+}(v^{1})-n^{+}(v^{2}))\epsilon)$
-$\geq\mathrm{Pr}_{\mathcal{C}}[v]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2}).$
-As a result, we have that
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m,\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
-$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\mathrm{Pr}_{\mathcal{C^{\prime}}}[v]$
-$\geq\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\mathrm{Pr}_{\mathcal{C}}[v]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2})$
-$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\mathrm{Pr}_{\mathcal{C}}[v]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2})$ (148)
-$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2}).$
-Here (148) is by Lemma 38. We present the detailed computation as follows. Let $v^{o}_{\ell}$ be the $\ell$-th dimension of $v^{1}$ for $1\leq\ell\leq m$ and $o=1,2$.
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
-$=\sum_{m^{\prime}=1}^{m}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}=m^{\prime}|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
-$=\sum_{m^{\prime}=1}^{m}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}=m^{\prime}|\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m^{\prime}}=\{v^{1}_{\ell}\}_{\ell=1}^{m^{\prime}},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m^{\prime}}=\{v^{2}_{\ell}\}_{\ell=1}^{m^{\prime}}\right]$ (149)
-$=\sum_{m^{\prime}=1}^{m}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}=m^{\prime}|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m^{\prime}}=\{v^{1}_{\ell}\}_{\ell=1}^{m^{\prime}},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m^{\prime}}=\{v^{2}_{\ell}\}_{\ell=1}^{m^{\prime}}\right]$
-$=\sum_{m^{\prime}=1}^{m}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}=m^{\prime}|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$ (150)
-$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right].$
-Here (149) and (150) hold because for $h^{\prime}=h,h^{*}$ the event $\{h_{\mathrm{out}}=h^{\prime},M_{\tilde{i}}=m^{\prime}\}$ is independent of $\{x^{\ell}(\tilde{i})\}_{\ell\geq m^{\prime}+1}$.
-Taking sum over $v=\{v^{1},v^{2}\}\in\mathcal{V}$, we have that
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m]$
-$\geq\sum_{v\in\mathcal{V}}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h,M_{\tilde{i}}\leq m,\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
-$\geq\sum_{v\in\mathcal{V}}\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2})$
-$=\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m,\left\{\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}\right\}\in\mathcal{V}\right]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2})$
-$\geq\frac{1}{2}\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m]\exp(-80\sqrt{m}\epsilon-40m\epsilon^{2}).$
-(146) is proven.
-Taking sum over all $h\in\mathcal{H}_{\tilde{i}}$, we have that
-$1\geq\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}\in\mathcal{H}_{\tilde{i}},M_{\tilde{i}}\leq m]\geq\frac{N_{0}}{2}\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m]\exp(-40m\epsilon^{2}-80\sqrt{m}\epsilon)$ (151)
-which implies that
-$\frac{1}{2}\leq\mathrm{Pr}_{\mathcal{G}^{\prime}}\left[h_{\mathrm{out}}=h^{*},M_{\tilde{i}}\leq m\right]\leq\frac{2}{N_{0}}\exp(40m\epsilon^{2}+80\sqrt{m}\epsilon).$ (152)
-Therefore, we have that $40m\epsilon^{2}+80\sqrt{m}\epsilon\geq\log(N_{0}/4)$, which implies that $m\geq\min\{\frac{\log(N_{0}/4)}{80\epsilon^{2}},\frac{\log^{2}(N_{0}/4)}{30000\epsilon^{2}}\}\geq\frac{\log(N_{0}/4)}{30000\epsilon^{2}}$
-F.2 Statement and Proof of Lemma 38
-Lemma 38.
-Fix $\tilde{i}\in[k]$, and let $x^{\ell}(i)$ be the $\ell$-sample from $D_{i}$ for any $i\in[k]$ and $\ell\geq 1$. Fix $h\in\mathcal{H}_{\tilde{i}}$, $m>0$ and $v^{1},v^{2}\in\{-1,1\}^{2m}$.
-$\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h^{*},M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}]$
-$=\mathrm{Pr}_{\mathcal{G}^{\prime}}[h_{\mathrm{out}}=h,M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1}].$
-Proof.
-Let $\bar{\sigma}$ be the permutation over $\mathcal{H}$ with $\bar{\sigma}(h^{*})=h,\bar{\sigma}(h)=h^{*}$ and $\bar{\sigma}(h^{\prime})=h^{\prime}$ for all $h^{\prime}\notin\{h,h^{*}\}$.
-It then holds that $\bar{\sigma}^{-1}=\bar{\sigma}$.
-Fix $\{m_{i}\}_{i=1}^{k}$ and $X(i)=\{X^{\ell}(i)\}_{\ell=1}^{m_{i}}\in\{-1,0,1\}^{kNm_{i}}$ for $i\in[k]$.
-Let $x(i)=\{x^{\ell}(i)\}_{\ell=1}^{m_{i}}$ be the datapoints of the first $m_{i}$ calls to $\mathrm{Query}(D_{i})$.
-With a slight abuse of notations, we let $\sigma(x(i))=\{\sigma(x^{\ell}(i))\}_{\ell=1}^{m_{i}}$ for each $i\in[k]$.
-We continue with the following lemmas.
-By Lemma 39, we have that
-$\mathrm{Pr}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right]$
-$=\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=\sigma^{-1}(h),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right],$
-which means
-$\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h^{*},\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\sigma(x(i))=X(i),\forall i\in[k]\right]$
-$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\bar{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right]\cdot\frac{\mathrm{Pr}[\sigma(x(i))=X(i),\forall i\in[k]]}{\mathrm{Pr}[\bar{\sigma}\sigma(x(i))=X(i),\forall i\in[k]]}$
-$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\bar{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right]\cdot\frac{\mathrm{Pr}[\sigma(x(\tilde{i}))=X(\tilde{i})]}{\mathrm{Pr}[\bar{\sigma}\sigma(x(\tilde{i}))=X(\tilde{i})]}$
-$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\bar{\sigma}\sigma(x(i)))=X(i),\forall i\in[k]\right]$
-$\quad\qquad\qquad\qquad\qquad\cdot\frac{\mathrm{Pr}\left[\{x_{j_{h,i}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h^{*},\tilde{i}}}^{\ell}({\tilde{i}})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),{\tilde{i}}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\right]}{\mathrm{Pr}\left[\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\right]}.$ (153)
-Re-arrange the equation, we have that
-$\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\Bigg{[}h_{\mathrm{out}}=h^{*},\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\sigma(x(i))=X(i),\forall i\in[k]$
-|$\qquad\qquad\qquad\qquad\Big{|}\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\Bigg{]}$
-$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\Bigg{[}h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\bar{\sigma}\sigma(x(i))=X(i),\forall i\in[k]$
-|$\qquad\qquad\qquad\quad\Big{|}\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\Bigg{]}.$
-Taking sum over all possible choices of $\{X(i)\}_{i\neq\tilde{i}}$, $\left\{\{X^{\ell}_{j_{h^{\prime},i^{\prime}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\right\}_{i^{\prime}\in[k],h^{\prime}\notin\{h,h^{*}\}}$,
-$\left\{\{X^{\ell}_{j_{h^{\prime},i^{\prime}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\right\}_{h^{\prime}\in\{h,h^{*}\},i^{\prime}\neq\tilde{i}}$ and $\{m_{i}\}_{i\neq\tilde{i}}$, we obtain that
-$\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\Bigg{[}h_{\mathrm{out}}=h^{*},M_{\tilde{i}}=m_{\tilde{i}}$
-|$\qquad\qquad\qquad\qquad\Big{|}\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\Bigg{]}$
-$=\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\Bigg{[}h_{\mathrm{out}}=h,M_{\tilde{i}}=m_{\tilde{i}}$
-|$\qquad\qquad\qquad\qquad\Big{|}\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}},\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}\Bigg{]}$
-for any $X(\tilde{i})\in\{-1,0,1\}^{kNm_{\tilde{i}}}$.
-Fix $m_{\tilde{i}}=m$, and choose $\{X^{\ell}_{j_{\sigma(h),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=v_{1}$, $\{X^{\ell}_{j_{\sigma(h^{*}),\tilde{i}}}(\tilde{i})\}_{\ell=1}^{m_{\tilde{i}}}=v_{2}$.
-We then have that
-$\mathrm{Pr}_{\mathcal{G}^{\prime},\mathcal{H}}\left[h_{\mathrm{out}}=h^{*},M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
-$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h^{*},M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2}\right]$
-$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathrm{Pr}_{\mathcal{G},\bar{\sigma}\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=\bar{\sigma}^{-1}(h^{*}),M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1}\right]$
-$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h,M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1}\right]$
-$=\mathrm{Pr}_{\mathcal{G}^{\prime},\mathcal{H}}\left[h_{\mathrm{out}}=h,M_{i}=m|\{x_{j_{h,\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{2},\{x_{j_{h^{*},\tilde{i}}}^{\ell}(\tilde{i})\}_{\ell=1}^{m}=v^{1}\right].$
-The proof is completed. ∎
-F.3 Statement and Proof of Lemma 39
-Lemma 39.
-Fix $\{m_{i}\}_{i\in[k]}$ , $\sigma\in\Pi_{\mathcal{H}}$ and $X\in\{-1,0,1\}^{kN\sum_{i=1}^{k}m_{i}}$. Let $\{X(i)\}_{i\in[k]}$ and $\{x(i)\}$ be defined as Lemma 38. We then have that
-$\mathrm{Pr}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right]$
-$=\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=\sigma^{-1}(h),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right].$ (154)
-Proof.
-Let $\mathcal{H}^{\prime}=\sigma(\mathcal{H})$. Let $h_{p}(\cdot)$ denote the $p$-th hypothesis in the hypothesis set.
-$\mathrm{Pr}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h_{p}(\mathcal{H}),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right]$
-$=\mathrm{Pr}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h_{p}(\mathcal{H}),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\{\{x^{l}_{j_{h_{p^{\prime}}(\mathcal{H}),i}}(i^{\prime})\}_{p^{\prime}=1,i=1}^{|\mathcal{H}|,k}\}_{\ell=1}^{m_{i^{\prime}}}\}_{i^{\prime}=1}^{k}=X\right]$
-$=\mathrm{Pr}_{\mathcal{G},\mathcal{H}^{\prime}}\left[h_{\mathrm{out}}=h_{p}(\mathcal{H}^{\prime}),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\{\{x^{l}_{j_{h_{p^{\prime}}(\mathcal{H}^{\prime}),i}}(i^{\prime})\}_{p^{\prime}=1,i=1}^{|\mathcal{H}^{\prime}|,k}\}_{\ell=1}^{m_{i^{\prime}}}\}_{i^{\prime}=1}^{k}=X\right]$ (155)
-$=\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=h_{p}(\sigma(\mathcal{H})),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right]$
-$=\mathrm{Pr}_{\mathcal{G},\sigma(\mathcal{H})}\left[h_{\mathrm{out}}=\sigma^{-1}(h_{p}(\mathcal{H})),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right].$ (156)
-[…]
+The proof is thus completed.
+∎